A Complete Bibliography of Computers and Security, 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 January 2022 Version 1.01

Title word cross-reference

3 [MGYC18, NCS14, PYLL19, Woo19]. ^{2} [CDE^{+}11]. _{ABC} [GMGTdFR14b]. ERBAC [AIRC13]. C [HLB19]. D [DLH^{+}17]. E^{2} [CDE^{+}11]. G [GSZX18]. K [TFPC10, JLPJ18, NG14, XMY^{+}17, ZZC17]. l [MC18]. n [ORLS14]. O^{2} [FKSK19]. R [DLH^{+}17]. τ [ARG^{+}17].

-Anonymity [NG14]. -diversity [MC18]. -gram [ORLS14]. -means [TFPC10]. -nearest [XMY^{+}17]. -NN [ZZC17]. -safety [ARG^{+}17]. -Trustee [JLPJ18].

10-year [BB11]. 131-3 [AM18a].

2 [TH18]. 2010 [Ano10-33, Ano10-35, Ano10-34, Ano10-38, Ano10-37, Ano10-36]. 2010-2012 [Cho13b]. 2011 [Ano11y, Ano11z, Ano11-27, Ano11-29, Ano11-28]. 2012 [Ano12w, Ano12z, Ano12-27, Ano12x, Ano12y, Ano12-29, Ano12-28]. 2013 [Ano13-27, Ano13-28, Ano13y, Ano13-29, Ano13z]. 2014 [Ano14t, Ano14x, Ano14u, Ano14w, Ano14v, Ano14y]. 2015 [Ano15y, Ano15z, Ano15x, Ano15v, Ano15t, Ano15w, Ano15u]. 2016 [Ano16t, Ano16s, Ano16v, Ano16u, Ano16q, Ano16w, Ano16r, Ano16x]. 2017 [Ano17p, Ano17s, Ano17q, Ano17t, Ano17r, Ano17o, Ano17u]. 2018 [Ano18q, Ano18l, Ano18k, Ano18m, Ano18p, Ano18n, Ano18j, Ano18o]. 2019 [Ano19j, Ano19p, Ano19m, Ano19o, Ano19n, Ano19k, Ano19l, Ano19q]. 28th [Ano15i].

3-state [ZLW^{+}15]. 30/8 [Spa11d]. 30th [HF17]. 3G [XN14].

5G [FYL^{+}17, MPCK18].

6LoWPAN [MA17].

7 [Ano11-30, SCB^{+}13].

802.16m [FZZ^{+}12].

A-PANDDE [FB19]. ABE [OSNZ19]. ability [YGX^{+}19]. Abnormal [ZZ19]. Abstract [ZLLL12]. abuse [CKA^{+}17]. abuses [CWY^{+}12]. abusive [PBR11]. academic [AAMDGR18]. accelerators [KKJ^{+}16]. acceptable [RMW10]. acceptance [SYK14]. Access [ADF^{+}10, DMR19, Lop16, PG19, SRDD13, SSAK12, AAD^{+}12, ARL13, AAC^{+}19, BRH15, BJ13, BHC18, BR17, BCD^{+}10, BSL19, CHK^{+}19, CFH^{+}11, CIS17, CDS^{+}19, CZH^{+}14, GM12, GSZX18, HK17, JJ12, JJ14, KBAL13, KS14, LKJ14, LML^{+}16, LC15, MDHM18, MZZ^{+}18, MT19, PZ15, PMDM10b, PMDM10a, RLB^{+}11, RR17, SMP13, SAL12, SSVA13, SC19, SSVA19, SYGR11, UPF17, ZHZ^{+}15]. AccessMiner [FLBK15]. account [BLS^{+}16, KK16, SC19]. accountability [NFGL16]. accuracy [MMF19]. Achieving [DYL^{+}14, LZ19, XLT^{+}17]. Ack [AH13]. Acknowledgement [Ano18r]. Acknowledgment [Ano19r]. acoustic [DHZ19]. across [DPP15]. action [GHÁGGT18, KMW19, LSKF12]. actionable [CAY^{+}18]. actions [AM18b, dVM15]. Active [HWS^{+}13, VSB^{+}19, Den14, ZHZ^{+}15]. activities [BHDA14, HA14, Kha12, ZYRZ16]. activity [BFTJ18, DB13, LLK16, RR14]. actors [LCMF18]. actuator [hBCK17]. ad [BRA18, HZMH18, MJBKZ15, SRU18, Su10, XPS11, ZY11, Ano12s]. ad-hoc [BRA18, HZMH18]. adaptation [MCRB19, RNdP17]. Adaptive [PHMN16, SLP13, AAR^{+}18, BJ13, CFSF18, DRN16, EE15, MSV12, PPdCL17, RB14, TVPV19]. addition [HDK18]. address [LKPK18]. addresses [CBL10]. Addressing [LWZ^{+}17, MR11]. adjustable [BWR12]. ADLES [dLGHK18]. administration [GHL^{+}17]. administrative [JSVA14, SSVA13, SYGR11]. administrator [ZH13]. ADOPT [MJBKZ15]. adoption [HOVB17, MT17, QFG12, SWS15, UPSGF^{+}14]. ADS [HS18, LWS^{+}19]. ADS-B [HS18, LWS^{+}19]. Advanced [FB19, AWDB19, FSSF15, LCMF18, SKLD14, TAP18, XN14]. Advances [XHZ11, CBGN18, IAA^{+}19]. advantage [ABS14]. Adversarial [RSLH19, CXF^{+}18, CSS^{+}19, HM19, KKP^{+}18]. adversary [DC16, DMC19, OSNZ19]. advertisers [CKA^{+}17]. aerial [MK19]. AES [ZGZ^{+}13]. affect [DM19]. affecting [PKWP19, Ven18]. affects [KMW19]. Affordable [KN14]. aftermath [PCP^{+}18]. Against [VG14, ADAMH17, BD18, CWCC12, DCC^{+}18, DLB^{+}19, HBT^{+}19, KN14, LDC13, LLL^{+}18, MBP19, Med18, NKL^{+}18, OSNZ19, RDAS19, hRKRR13, SCLD19, SLP13, SP15, SKTO13, SSY12, TH18, WP12, YG18, YWL^{+}17, ZZ12, ZLC13]. age [OM13, TR18]. agent [HR13]. agent-based [HR13]. agents [ABR18, CP18]. aggregatable [dFGMTPL17]. aggregation [KNX18, KP18, MWAG16, RR17, ZYWZ14]. aggressive [BMMP18]. agnostic [HSPX14]. agreement [GMLZ11, HWB10, Kar12, MK11]. Aided [Gop19]. AIDIS [LJS19]. Aims [Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano10-27, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x]. air [GB19, GB19]. air-gap [GB19]. aIR-Jumper [GB19]. alarm [DVC18, LMLK16, TFPC10]. alert [LSA^{+}19, SHGH^{+}15, SK13, dABM^{+}18]. algebraic [ZGZ^{+}13]. algorithm [ASO19, BK19, CBL^{+}19, CSS^{+}19, CSXL15, DYT^{+}13, EE15, HS18, HDT^{+}19, MLM17, MZJ^{+}12, RAA15, SD17, SMC^{+}18, TFPC10, VPNP18, VDK18, YCJ12, Zin15]. algorithms [BSL19, CAM^{+}19b, LK10, SCGM14, XXLZ18]. Allocation [RW13]. allowing [JvOW11]. altering [XdZMsC18]. AMAL [MAM15]. amplification [AKK^{+}13, SMB16]. AMTRAC [SSVA13]. analyses [MA17]. analysing [ELG^{+}17]. Analysis [AM17a, CGGFGMF19, FAS^{+}17, GOB11, LMNM^{+}19, MG12, ÖTC16, SYK14, SS13, UAK16, AKFM18, AOR18, AHTS15, ABP16, AGSR^{+}18, ALF^{+}18, AG14, BHC15, BBD^{+}19, BPA^{+}11, BS17, CDKR18, CGPP11, CGPP12, COS19, CCM^{+}15, CAL15, CKT11, Coh10, CC14, CWGD15, DPPB17a, DPPB17b, DDYZ14, DJP17, FNE19, Gar13, GSA18, GAA18, GHDGRLM17, GCC18, HBBM16, HCRMBG17, JKC^{+}13, JSVA14, KGSN17, KHG17, Kha12, KPKS15, KKK18, KPD15, KSW16, LWS^{+}18, LK18, LLYD15, LXDW16, LPL18, LCZ17, LJJL15, MW11, MVD19, MJW^{+}18, MP18, MAM15, MB18, MSA11, MAC17, MBSR18, NNG^{+}18, NLC^{+}19, NAMX14, OP11, OM19, ORLS14, PF10, PK17, RM17, SS11, SJCM17, SPFB19, STCM^{+}14, SPYG15, SHGH^{+}15, SS18b, SEJ10, SKG17, SPD^{+}15, SPR18, SYGR11, SHB12, TdHRZ17, UAB19, VEPZ15, WRTA17, WZW^{+}10, WLCC17, WA19, Yad19]. analysis [YSN^{+}19, YCH18, ZMYB17, ZBBO15, ZBB17, ZTS^{+}13, ZCZ^{+}19, ZW14, ZLW^{+}15]. analytic [WLW^{+}19]. analytics [MAZAS17, PCCM16, QAR^{+}17]. analyze [JKC^{+}13, STSB17]. analyzed [WCDY18]. Analyzing [AYY^{+}17, CLS17, KLN15, SCB^{+}13, TKK15, HLFZ16, ZWNM19]. Anatomy [SCRA17]. Andro [wJKW^{+}16]. Andro-Dumpsys [wJKW^{+}16].

AndroDialysis [FAS^{+}17]. Android [AQR^{+}17, AGSR^{+}18, AMMV13, BM19, BRH15, BSTT16, BJJ^{+}17, CMV16, DLH^{+}17, DMC15, ESY^{+}15, FSZ^{+}17, FHL14, FAS^{+}17, GFKS15, HAHZ^{+}18, IRC^{+}17, JLK16, LW12a, LLCT13, LLG18, MPM15, MAC^{+}15, MTC^{+}18, NNG^{+}18, NSGF16, NVAJ19, NAMX14, OM19, PSK^{+}17, RFB17, SMM^{+}19, SLS^{+}18, TXM^{+}19, WCDY18, ZBB17, ZTC19, Zhu18]. annotations [EHBH^{+}16]. Announcement [Ano12a, Ano18a]. anomalies [CGP17, CYLL17, GTDVTSH15, MAC17]. anomalous [HS18, LJS19]. Anomaly [hBCK17, MSRF18, AYT^{+}14, AP13, CGPP11, CGPP12, CPVGTMF16, CAM^{+}18, DC11, GCD19, LWS^{+}18, MPORBR12, MLH^{+}18, PF10, PCCM16, SCGM14, SBW10, Su10, UPSGF^{+}14, VEPZ15, ZNAHK11, FB19]. anomaly-based [AYT^{+}14, SBW10]. anomaly-detection [SCGM14]. anonymisation [SWLZ11]. anonymised [Jän14]. Anonymity [NG14, BGR11, DAR14, FHF15, GH15, LLY17, PK17]. Anonymization [SVA10, GL18, JLPJ18]. Anonymizing [ANAT15, WL18]. Anonymous [APMCR13, SYWX19, Gop19, Hor16, LSQ15, YAXY19]. Anti [Ano14s, wJKW^{+}16, AAT16, JJL^{+}19, KYZ15, RB12]. anti-emulation [JJL^{+}19]. anti-forensic [RB12]. Anti-malware [wJKW^{+}16, AAT16]. Anti-virus [Ano14s, KYZ15]. anticipated [Ver18]. Apache [WJAM11]. APFS [SLP13]. API [SMM^{+}19, WA19, ZQY^{+}19]. API-related [SMM^{+}19]. app [BRH15, BJJ^{+}17, CGLP18, HSR19, NGSV16]. app-based [NGSV16]. applicability [MB18]. applicable [Ano17l, Ano17m, Ano17n, Ano18i]. Application [PGdCL16, SC19, SSK17, AAL16, CWGD15, GOGG19, RAA^{+}14, SMP13, STCM^{+}14, WA19, XNP16]. application-oriented [SMP13]. Applications [DPPB17a, APMCR13, ALRC12, BM19, BSTT16, BBD^{+}19, BHCTR12, EHBH^{+}16, FSZ^{+}17, FXC18, GJJ18, GFKS15, HZG19, HZG20, JK19b, LLL15, LLCT13, LSQ15, NSGF16, NAMX14, OSP^{+}19, PAJAP15, SBK12, SSAK12, SDM^{+}18, SMA^{+}18, SBS^{+}18, TXM^{+}19, TMP15, YCH18, ZBB17, Zhu18]. applied [CAM^{+}18, DMC19, DJP17, HDT^{+}19, MPPCFG^{+}12, OPHC16, PPdCL17]. Applying [Kar12, NML19]. Approach [Pal10, AKFM18, AYT^{+}14, ADWM18, AW18, BPA^{+}11, BOZL13, CM17, CDS^{+}19, CXF^{+}18, CAM^{+}19b, CWCC12, CLF17, DPP15, DMR19, DLLS19, GFKS15, GWWS19, GGBT16, GHÁGGT18, GWWW19, GCC18, HK13, HR13, HA14, HWX^{+}19, HPF^{+}16, HZG19, HZG20, HZH^{+}18, IGS17, JKA^{+}18, JGM^{+}13, KK17, LSC14, LWS^{+}18, MCY^{+}18, MZZ^{+}18, RDAS19, RAA^{+}14, RB11, SSTG12, SPK17, SOG15, SBE17, lTLqZqC19, TC11, WZK13, WLZ^{+}19, WLF^{+}19, YLW^{+}19, ZjHtMhBs12, ZW19, ZGD19, Zin15, SG15a]. Approaches [STSB17, BCD^{+}10, WWXZ18]. apps [HAHZ^{+}18, JLK16]. Apriori [CSXL15]. arbitrary [ARG^{+}17]. architects [JSWZ11]. Architectural [YLB^{+}17]. architecture [AGL12, CM19, CY16, CAM^{+}18, FYL^{+}17, GK14, PAP^{+}17, PDB14, RRK^{+}17, RV13, RVEK10, RM17, TVPV19, dCGÜ^{+}17]. Architectures [IH19, OS17, XPS11]. archiving [VBC^{+}15]. area [FPS11, XXLZ18]. ARES [Ano15s]. ArOMA [SBZD17]. arrangements [BTM14]. arrivals [dSFK16]. art [GJ18, NCGE15, SMP13]. Artifacts [LW12d, RAMR17]. ASMSA [Pal10]. aspect [AGSR^{+}18]. aspect-oriented [AGSR^{+}18]. Aspects [PMB^{+}14, PCP^{+}17, FC12]. Assessing [KKS13, LSSK16, NPP17, SBL19, TS13, BM17b, DM19, LMG18]. Assessment [GPGV16, TKG10, ANMN19, AAH^{+}18, BPG^{+}19, BRM^{+}18, CBB^{+}16, DE10, HBFJ19, HSW17, KBAL13, KBM16, MCT^{+}17, MMJ14, MTT^{+}16, MM17, NZGH16, RFB17, SCRA17, SSCHL14, SSABC16, SOG15, SG10b, WHZ^{+}18]. assistant [TXM^{+}19]. assisted [DHZ19, WRTA17]. association [DYT^{+}13]. assurance [CSBW18, CHRI16, HSW17, PMDM10a, SGK^{+}16]. atomic [CC14]. Attack [Kam19, NADZ18, YG18, AKFM18, AKK^{+}13, BD18, BM17b, CAM19a, DT19, DWF^{+}19, DLB^{+}19, ELG^{+}17, EM19, FNE19, GSA18, GGBT16, GSL^{+}19, HT17, JLPJ18, Kal19, KJP^{+}19, KKP^{+}18, LDB18, LXDW16, LWS^{+}19, MW11, MA17, Med18, MJW^{+}18, MLV16, NDP18, NTK^{+}19, ORLS14, PNKK17, PHMN16, RAA15, RAA^{+}14, SGLC16, SP15, SHGH^{+}15, SMB16, TC11, TP18, WZK13, XdZMsC18, ZjHtMhBs12, ZW19]. attack-resilient [JLPJ18]. attacker [KKK18, bORF^{+}15]. attackers [DLB^{+}19, KK10a]. Attacking [XN14]. Attacks [Bay12a, GFN11, HZG19, HZG20, AH13, AF14, ADAMH17, AYT^{+}14, AAL16, AHGC19, BSMT14, CMPR15, CXF^{+}18, CSS^{+}19, CWCC12, DCC^{+}18, DSNH17, EGM10, FRRP17, GJ18, GHDGRLM17, HK13, HR13, HL18, HAGTdFR13, JLGM19, JC14, JCD^{+}15, JJA13, JCD14, KPS16, KK10b, KPKS15, KN14, LDC13, LAVK19, Len12, LLL^{+}18, LLD^{+}19, MBP19, MAC^{+}15, MPCK18, MK19, MMJ14, NYE17, NT18, NKL^{+}18, PTM^{+}18, RDAS19, RB12, Ril13, hRKRR13, Rru16, SCLD19, SLP13, SSK16, SSK17, SKTO13, TTCB18, TUB16, TR18, TCNFW16, TH18, VC12, WLZ^{+}19, WP12, YWL^{+}17, ZTL^{+}19, ZGZ^{+}13, ZJK^{+}17, ZLK10, ZLC13]. Attention [YLW^{+}19, YSN^{+}19]. Attention-based [YLW^{+}19]. attestation [TTJ19]. attitude [GGV18]. attitudes [Kok17, SWS15]. attractors [SDM^{+}18]. attribute [Bay13, BTK15, LZCX18, LAHV17, SSVA19, WLWG11]. attribute-based [LZCX18, LAHV17, SSVA19, WLWG11]. attributes [WL18]. attribution [ZTC19]. auction [DJP17, SYCS11]. auctions [Len12, ML14]. Audio [SG10b, HZH^{+}18, LNK16, YWYZ12]. auditable [DMR19, PMP19]. auditing [LXH^{+}18]. auditors [JSWZ11]. audits [PWL^{+}19]. augmentation [GWWW19]. augmenting [NPP17]. August [Ano16t, Ano18q, Ano19p]. Aura [HFCR13]. Australian [BCK19, DPP15]. Austrian [ZS15]. authenticated [ABR15, HWB10, JKA^{+}18]. authenticating [Bel18, RR16]. Authentication [HFCR13, KLN15, NSBM17, SPK17, TWNC18, ARL13, APMCR13, ACC^{+}13, BOP14, BD18, CAM19a, CRS13, DRN16, FHM^{+}10, FZZ^{+}12, GEAHR11, GH15, Gop19, GMMJ11, GBC19, HW19, HHBS18, HK17, HL12, IAA^{+}19, KS19, LNNH13, LSQ15, MMLN15, MDHM18, MCRB19, NDSA17, NM18, NML19, NXS10, NMX15, OSP^{+}19, QLZ19, RR17, SCFB15, SK18, SSNS15, SYWX19, SNG^{+}17, SCR19, SMS^{+}16, TZTC16, VSB^{+}19, YI17]. authenticity [VBC^{+}15]. authority [JWB16, LML^{+}16]. Authorization [LJJL15, ZDBW11, AP14, FHM^{+}10, JMGW17, JAH^{+}11, UFF15]. authorizing [Bel18]. Authorship [PSMG18]. Auto [NNNQ18]. Auto-detection [NNNQ18]. Automated [CXF^{+}18, DF16, GAA18, MVK16, Pal10, dRMC^{+}16, BLR17, GMMJ11, Hol12, LMS14, MAM15, SCLD19, SHB12]. Automatic [AHP18, DBH17, GTDVTSH15, KCB13, VG14, WWY^{+}18, CDS^{+}19, NRM19]. Automatically [ELG^{+}17, LHI^{+}19]. Automating [AQR^{+}17]. automation [EMWE19, HT12a, PHP19, TTCB18]. automaton [WLZ^{+}19]. automotive [CSBW18]. autonomic [SBZD17, ZNAHK11]. Autonomous [HYLY19, DFC13, DFCC16, GHDGRLM17, SMS^{+}16]. Autonomously [BVWM19]. autonomy [JL10, LLY17]. AutoPrivacy [WWY^{+}18]. availability [TFE10]. available [LCMF18, UB12]. avoid [DPP15]. avoidance [AM18b, MHS17, Su10]. aware [ARL13, BAH18, BOZL13, CHK^{+}19, GWKFC19, ODB18, PHP19, RFGL18, RVEK10, SRDD13, SRU18, WLZ^{+}19]. awareness [AH10, AFF14, BBC17, BFS^{+}18, FE16, FB14, HPJ^{+}19, JvS17, KDC^{+}14, KAF17, KvS10, MKG13, ÖTC16, PMB^{+}14, TRB18, TZZG13, TKK15, WAMS14].

B [HS18, LWS^{+}19]. B3G [NXS10]. BACI [AAC^{+}19]. back [CAM^{+}18]. backbone [MPORBR12]. backup [JL10]. balance [SS11]. balancing [BGR11, MKP12]. bandwidth [AKJR14]. Bank [CWCC12]. banking [BGL18, CCM^{+}15, GWWS19, GMMJ11]. banks [BBC17]. BankSealer [CCM^{+}15]. bargaining [GCC18]. Based [GMGTdFR14b, SRDD13, AR17, ARL13, ATAAd18, AAC^{+}19, AYT^{+}14, AKP14, ABR18, ATG11, ACC^{+}13, AAL16, AW18, AGL12, AP13, BK19, BD16, BRH15, hBCK17, BD18, BHC18, BLR17, BTK15, BL15, BSL19, CPVGTMF16, CMV16, CFH^{+}11, CM17, CVS18, CIS17, CJA^{+}19, CLL^{+}13, CSXL15, CAM^{+}19b, CKJ^{+}17, DSNH17, DT19, DC11, DMR19, DAR14, DYT^{+}13, DDYZ14, DXCL18, DLLS19, DWZ19, EGM10, EE15, EM19, FB19, FSZ^{+}17, FHL14, FNE19, FLBK15, FYL^{+}17, FXC18, FHF15, FRRP17, GOB11, GWWS19, GWKFC19, GHÁGGT18, GBC19, HB18, HDK18, HR13, HW19, HHBS18, HXW^{+}19, HBFJ19, HBF13, HDMKS10, HWB10, HA15, HYZ18, HYLY19, IGS17, JKC^{+}13, wJKW^{+}16, JWZX19, JJA13, JJ12, JJ14, KTCS19, KK10a, KHG17, KMM13, KJKM11, KKK18, KP18, KTG^{+}18, KKM19, KKS13, LMM^{+}12, Len12, LHM13, LMLK16, LZCX18, LWS^{+}19]. based [LMLL19, LHI^{+}19, LLX^{+}17, LLHT13, LC15, LGS14, LAHV17, MFCMC^{+}18, MKP12, MLK14, MLKC17, MCRB19, MPPS10, MAM15, MKG16, NSFB15, NCS14, NGSV16, NM18, NYE17, NMX15, ÖB15, Par18, PWWY14, PYLL19, PK17, PMP19, QHW^{+}19, QLZ19, RG19, RP19, RD19, RB14, RB11, RSLH19, RWS^{+}19, hRKRR13, SD17, SS16, SS17a, SBZD17, SERF12, SFK15, SK18, SP15, SAL12, SSVA13, SYK14, SMC^{+}18, SVA10, SSK19, SSVA19, SPK17, SOG15, SBW10, SSY12, SKG17, SYGR11, SMA^{+}18, Sul19, SMK18b, lTLqZqC19, TUB16, TKG10, THA^{+}13, TDJB18, USB17, VDK18, VVS12, WLWG11, WN15, WW15, WLCC17, WL18, WLZ^{+}19, WHZ^{+}18, WW18, XXLZ18, XSD19, YCJ12, YYL^{+}15, YLW^{+}19, YLN^{+}19, YXD^{+}11, ZLLL12, qZLxCjY17, ZQY^{+}19, ZTL^{+}19, ZZ19, ZTS^{+}13, ZGZ^{+}13, ZY14, ZjHtMhBs12, ZWX^{+}19, ZLC13, ZGD19, ZZZY13, ZW14, ZH13, ZHB^{+}13, ZLW19]. Basics [Mic12a, Wed12]. batch [HYLY19]. Bayesian [Kha12]. be [HDT^{+}19]. begins [TMW17]. Behavior [HBF13, BRH15, DDYZ14, DXCL18, GGV18, GBC^{+}18, Guo13, HSL19, Ifi12, KLN15, LMNM^{+}19, LJS19, MAM15, MBAE^{+}11, ÖTC16, Pad12, PRS13, SS18a, STCM^{+}14, SPYG15, SWS15, Sil19, SPK17, TMW17, TRB18, VSB^{+}19, ZZ19, ZTS^{+}13, ZH13]. Behavior-based [HBF13, MAM15]. behavioral [BOP14, CAM19a, CJL^{+}13, FAE14, KKK18, MMCA11, PC12]. behaviorist [Wor10]. behaviors [AAA19, DM19, FSZ^{+}17, HNSD18, JLK16, ySTJA^{+}16, Ver18, ZMYB17]. Behaviour [NF15, AH10, FanE^{+}18, KK16, Kok17, SSS^{+}15]. behavioural [DPPB17a]. behaviours [ACS^{+}18, BGL18, ESCAM17, MLKC17]. behind [SB14]. belong [SW19]. benchmark [GDR15, SSTG12]. bending [LW19]. benefit [WZK13]. best [MM15a, CJA^{+}19]. better [BBC17, BCRAA13]. between [BLR17, GMMP18, JGLE16, JSWZ11, MCG19, SS11]. Beyond [JvS17]. bias [MM17]. bias-free [MM17]. biases [TKK15]. Bible [LW12e]. bid [ML14, SYCS11]. Big [CGGFGMF19, CBL^{+}19, LCZ17, LZ19, MCY^{+}18]. bigram [HDK18]. bilateral [HKK17]. billing [Sul19]. bimodal [LNK16]. binary [YYL^{+}15, ZLW^{+}15, ZWNM19]. binding [HZG19, HZG20, NNNQ18]. Bio [BCRAA13, ZHL^{+}11]. bio-cryptographic [ZHL^{+}11]. Bio-Hashed [BCRAA13]. biometric [AAJ12, GBC19, PPdCL17, SS17a, YST^{+}19]. biometric-based [GBC19]. biometric-enabled [YST^{+}19]. biometrics [BOP14, BD16, ICRB14, RB14, SS16, SCFB15, SSY12]. biometrics-based [BD16]. birthmark [ZLLL12]. Bitcoin [CGR18, SBL19]. BitTorrent [AKJR14, SFK15]. black [CSS^{+}19]. black-box [CSS^{+}19]. blacklistable [YAXY19]. blacklisting [DM17]. block [DWF^{+}19]. Blockchain [AAC^{+}19, CJA^{+}19, NML19, ZLW19, DMR19, HHBS18, HYLY19, KKM19, PMP19, YLN^{+}19, ZWX^{+}19]. Blockchain-based [CJA^{+}19, ZLW19, HHBS18, HYLY19]. Blockchain-enabled [NML19]. blocks [NTK^{+}19, WLW^{+}19]. Bloom [KP18, MT19, RB14, hRKRR13, SS17a]. Bluetooth [HBHA18, PO14]. Board [Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano12t, Ano12u, Ano12v, Ano13x, Ano15q, Ano15r, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano18b, Ano18c, Ano18i, Ano19a, HK13, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. boo [HNSD18]. Book [SG15a, TG15]. bootstrapping [KKJ^{+}16]. border [dRMC^{+}16]. Borg [SCLD19]. both [LAHV17]. Botnet [ZTS^{+}13, GGSZ14, RMG11, SLH14, SSK19, CKA^{+}17]. botnets [BVWM19, CFSF18, FCPA16, SG10a, WLCC17]. bots [BLR17, JHJ^{+}16]. bound [MW11]. bounding [MOV14]. bounds [CKA^{+}17]. bowtie [AKFM18]. box [CSS^{+}19]. BPNN [ZQY^{+}19]. brain [BLR17]. branching [CVS18]. Brazilian [ABC^{+}19]. BRB [WHZ^{+}18]. Breaking [AMMV13, HZG19, HZG20, ZBB17, SHB12]. Bridging [JSWZ11]. broadband [TMP15]. broadcast [TMP15]. broker [HB18]. broker-based [HB18]. browser [ACC^{+}13, JCD^{+}15, SS11, SGLC16, SWZL14, VMTG15]. browser-based [ACC^{+}13]. browsers [SGLC16, TMN^{+}17, ZY14]. browsing [TMN^{+}17]. Bubbles [HHBS18]. Budget [SRDD13]. Budget-aware [SRDD13]. Bug [SG14b, DEP13]. bugnet [FW10]. build [CAM^{+}19b]. Building [CSBW18, CAY^{+}18, JAAS19, Mic12e, RMVC^{+}12, PHP19, RG19]. bulk [Gar13]. bulk extractor [Gar13]. Bulwark [BKK19]. Burrows [Kül12]. business [AP14, CDV12]. business-driven [CDV12]. Busters [VMTG15]. buyer [CTH10]. buyer-seller [CTH10]. BYOD [DPP15, ZBBO15]. BYOD-enabled [DPP15]. Bypassing [BHCTR12]. bytecode [KTG^{+}18].

cache [JCD^{+}15, ZGZ^{+}13]. calculation [KTCS19]. calculus [DJP17, PS10]. Call [Ano10a, Ano10b, Ano10c, Ano11a, Ano11b, Ano11m, Ano11y, Ano11-30, Ano12b, Ano13-30, EMBH14, LLCT13]. caller [AM13, AM13]. Caller-REP [AM13]. calls [AM13, YXD^{+}11]. cameras [GB19]. campaigns [CGR18, KDC^{+}14]. Can [KK16, LHS^{+}18, TBDM14]. Cancelable [QLZ19, RB14, BD16, BCRAA13, DDSP17]. Candidate [CH14]. can’t [WA19]. capabilities [KMG17, bORF^{+}15]. capability [Lop16, MCG19, SS18b]. CAPPCHA [GMMP18]. CAPTCHA [BLR17, HCR10, HCRMBG17, SG10b]. CAPTCHAs [NDMM13, NCS14, SWS16, Woo19]. captured [SPK17]. card [BCRAA13, BL15, HLB19, RVEK10, MLM17]. cards [JAN19]. care [SSS^{+}15, vDBD13]. carving [YYL^{+}15]. cascade [HMFE10, MT19]. case [BMH16, BM17a, CSBW18, CKA^{+}17, DSP16, FCPA16, HCR10, HCRMBG17, KPKS15, OS17, WYS18, dVM15]. categorical [MSV12]. categorization [YtLpZD14]. causal [PBR11, HOVB17]. Causality [ZYRZ16, PK17]. CDroid [WCDY18]. cellular [XNP16]. CENDroid [BM19]. center [LMG18]. centralised [DFC13]. centred [KAF17]. centric [wJKW^{+}16, JMGW17, JAH^{+}11, JJ12]. Certificate [ATL^{+}17, KKM19, ZDBW11, KS14, MJBKZ15]. Certificate-less [ATL^{+}17]. certification [AAD^{+}18]. Certified [FGOPL10, BPA^{+}11, PW13, Tau11]. CertLedger [KKM19]. chain [BS17, Par18]. challenge [SPK17]. Challenges [AHR12, BM17a, Cho11, DLS14, LCZ17, Mic12c, PMP^{+}15, YHL^{+}15, AM17a, BTM14, CBGN18, GJ18, Hor16, RJJ15, STSB17, SSK17, SSK19, USB17, ZKM^{+}17]. changed [SBK12]. changes [BM17b, PBR11]. changing [CW19, JXK18, ROW19]. channel [CAM19a, FNE19, HZH^{+}18, JK19a, JMGW17, KJKM11, KJP^{+}19, MZJ^{+}12, PO14, ZTL^{+}19, ZYPL10, ZLL^{+}19]. channels [AG14, BEB^{+}18, BKK19]. chaos [ZZZY13]. chaos-based [ZZZY13]. characteristic [HBF13]. characteristics [BEB^{+}18, HXW^{+}19, TAP18]. Characterization [GPAD^{+}14]. charging [MJW^{+}18]. chasm [JSWZ11]. chat [MBSR18]. Chauhan [TG15]. cheating [BHC15]. Checker [BPG^{+}19]. Checking [XYY^{+}12, ABP16, AP14, BKP10, BPA^{+}11, MSA11]. checkpoint [YST^{+}19]. chess [BHC15]. children [NRM19]. China [HLFZ16]. Chinese [ZDW19]. choice [MMPC18]. chronicity [KGSN17]. CIFC [WCDY18]. CIPA [CY16]. circles [SKLD14]. circuit [CPGA13]. CISOs [AS13]. cities [VHBJ19]. City [SMA^{+}18, GWKFC19]. claimant [YI17]. class [LWZ^{+}17]. classical [YGX^{+}19]. classification [BEGW^{+}19, BFTJ18, FPS11, GPAD^{+}14, Kha12, MC18, MAM15, MZJ^{+}12, SPFB19, SKG17]. classifier [BM19, KKP^{+}18, VSO^{+}18]. classifiers [AR17, VDK18]. Classifying [HMNT10, LJS19]. ClearChart [MSMK18]. click [MLKC17]. click-draw [MLKC17]. client [JM13]. clients [NADZ18]. cloned [KL18]. closer [HT17]. Cloud [Ano12r, BTK15, CGMP18, DSNH17, HB18, HW12, IH19, KW14, Mic12d, OM13, AAT16, AMK18, ANMN19, ABR18, AAD^{+}18, BHC18, Bel18, BSBG19, CIS17, CAM^{+}19b, CBGN18, DVC18, DYL^{+}14, DYZ^{+}15, HOVB17, JWB16, LML^{+}16, LZCX18, LXH^{+}18, MTT^{+}16, NGSV16, NLC^{+}19, NFGL16, PHMN16, PDM19, RC15, RG19, RR16, RRK^{+}17, SCB^{+}13, SSSC18, SBE17, TL15, TG14, WLWG11, WW15, WHZ^{+}18, XYY^{+}12, XLS^{+}19, Yad19, YWS17, ZKM^{+}17, ZY14, ZZC17, Zin15, ZHB^{+}13, ZS15]. Cloud-based [DSNH17, CIS17, WW15, ZY14, ZHB^{+}13]. Cloud-of-Things [CGMP18, CBGN18, SSSC18]. clouds [HSW17, RMVC^{+}12]. cluster [AGL12, BM19, KJKM11, LWS^{+}18, SMK18b]. cluster-based [AGL12, KJKM11]. cluster-ensemble [BM19]. Clustering [WLCC17, CMPR15, Ngu18, PRS13, dABM^{+}18]. clusters [MKP12]. CNN [ZQY^{+}19]. Co [GMGTdFR14a, HFCR13]. Co-operative [HFCR13]. Co-owned [GMGTdFR14a]. code [AQR^{+}17, HLB19, JvOW11, KTG^{+}18, MKP11, PMDM10b, PMDM10a, SKK^{+}14, TOZJ13, YXD^{+}11]. Coded [XXLZ18]. codes [FCSD10, RB14]. Coding [LW12b]. Coercion [KNM^{+}17, HK13]. Coercion-resistant [KNM^{+}17, HK13]. Cognitive [YST^{+}19, BSBG19, DM19, LLD^{+}19, TKK15]. collaboration [DYZ^{+}15, NM19]. Collaborative [KKP10, CHK^{+}19, CY16, GM12, PAP^{+}17, PSMG18, UFF15, qZLxCjY17, ZLK10]. Collateral [SBS^{+}18]. collection [PHMN16, SKA^{+}19]. collective [AH10, SSF16]. collectivism [MWL18]. Colluding [SGLC16]. collusion [ADAMH17]. colors [MMLN15]. combat [TCNFW16]. Combating [FSSF15, FFGL18, JHJ^{+}16, NT18]. combination [CAM^{+}19b, DLLS19, VSO^{+}18]. combined [CAM^{+}18, DWF^{+}19]. Combining [CGPP11, AKFM18, BCRAA13]. coming [BMO18]. Command [PO14, SLH14]. Comment [WJ15]. commerce [BGR11]. commercial [DB13, TTCB18]. commitment [BTLT16]. Committee [Ano10x, Ano10y, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x]. commodity [JvOW11, KKJ^{+}16, ZW14]. Common [HA15, HCRMBG17, JvOW11, PRS13]. Communication [KTCS19, VG14, ATL^{+}17, ABP16, BRH15, BBD^{+}19, BJJ^{+}17, BEB^{+}18, BKK19, DLH^{+}17, Hor16, JK19a, MVD19, PDB14, RB11, YGX^{+}19, YKH10]. Communication-efficient [KTCS19]. communications [ROL13, WZW16]. communities [STSB17]. Community [JJ12, TBSR14]. Community-centric [JJ12]. companies [BMH16]. comparative [AG14, MK11, MP18, RM17, VHBJ19, WWXZ18, XPS11]. Comparing [DPP15, CZH^{+}14]. comparison [GGSZ14, MCG19, SS18b]. competition [BWMG17, BRM^{+}18]. competitions [BWMG17]. competitive [ABS14]. compile [SKK^{+}14]. compile-time [SKK^{+}14]. complementary [MMLN15]. complex [BTM14, MZJ^{+}12, SJCM17]. Complexity [TBDM14, CDOV10, LJJL15, YGX^{+}19]. Compliance [JSWZ11, BTM14, BBC17, HKK17, HSL19, Ifi12, KD13, LLK16, LSKF12, SVF16, SW19, TBDM14]. compliant [AM18a, BCRAA13, Pad12, Sil19]. comply [AM17b, RE19]. component [BR17, BCK19, DLH^{+}17]. component-oriented [BR17]. composable [MFBC11, PMDM10b]. composite [KSG17, NDSA17]. Composition [Med18]. Comprehensive [BR18, AMK18, CDS^{+}19, GGBT16, GK14, SK13, SBS^{+}18, WB16, ZGZ^{+}13, ZHB^{+}13]. compression [Kül12]. Computation [SKK^{+}14, WW15]. computations [DLOP17, LXH^{+}18]. Computer [Mic12d, SG14a, AHP18, ABR16, BLR17, Cho13a, LMNM^{+}19, SPCCMMS18, TMW17, ZHZ^{+}15]. Computers [Ano12r, Ano18b, Ano18c, Ano19a, GZ20, GZG20, HZG20, MMPC18]. Computing [BTK15, HW12, IH19, AAA19, ABR18, Ano13-30, CKJ^{+}17, DYL^{+}14, DYZ^{+}15, Gop19, JWZX19, KKK17, LZCX18, MHL^{+}17, PDM19, SCB^{+}13, SSSC18, SBE17, VPNP18, WHZ^{+}18, XLS^{+}19, Yad19, YWND14, YWS17, ZKM^{+}17, ZGD19]. concept [Hat18]. Concepts [DPPB17a, TG15, MSVA16]. Conceptual [Nie17]. concerns [MR11, MT17, PMDM10b]. concolic [BJLS17, ZWNM19]. Conditional [RW13]. Conference [HF17, Ano15i]. confidential [TOZJ13]. confidentiality [WW15]. configuration [AAD^{+}18, LDB18, ROL13, SKLD14]. configurations [BPG^{+}19, MAZAS17]. Confirmation [TP18, LH11]. conformance [ALF^{+}18, MCL^{+}13]. confront [KK10a]. confronted [MBAE^{+}11]. congruence [KDC^{+}14]. congruency [DLS14]. Connect [NB19]. connects [HZMH18]. conscious [SSS^{+}15]. consensus [YLN^{+}19]. consequences [HOVB17]. considerations [MMCA11]. considered [DLS14]. Consistency [LKJ14, CC14]. Consolidation [HW12]. Constant [MK11]. constrained [BRA18, HS13, HSPX14, HXS^{+}16]. constraints [AAD^{+}18]. construct [Bay13]. Constructing [MT19]. construction [CPGA13, EHBH^{+}16]. consumer [MSMK18, PCC^{+}18]. consumption [MMF19]. containment [CLLK10]. content [ODB18]. Contents [Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, DC16]. Context [MM17, PHP19, BOZL13, GWKFC19, HBFJ19, HZMH18, LKPK18, ODB18, PTM^{+}18, Sil19]. Context- [MM17]. context-aware [BOZL13, GWKFC19, ODB18]. contexts [AAA19, DPP15, YW11]. contextual [HBFJ19]. continuity [Nie17]. Continuous [SCFB15, ARL13, AAD^{+}18, CRS13, SKA^{+}19, SCR19]. contract [HKK17]. contrast [HB16]. Contributions [Ano11y]. Control [AKFM18, BSTT16, DMR19, DDYZ14, PG19, PO14, SRDD13, AAD^{+}12, ARL13, ALC17, AAC^{+}19, ADF^{+}10, BJ13, BHC18, BCD^{+}17, BR17, BCD^{+}10, BL15, BSL19, CHK^{+}19, CFH^{+}11, CIS17, CDS^{+}19, CLL^{+}13, CZH^{+}14, CRJ18, Coh10, CJSM17, GMGTdFR14b, GM12, HPJ^{+}19, HLB19, HT12b, JJ12, JJ14, KMM13, KBAL13, KS14, LML^{+}16, LPL18, LC15, LAR13, Lop16, MDHM18, MSR19, MMJ14, MSRF18, NNNQ18, PZ15, PMDM10b, PMDM10a, RLB^{+}11, RR17, RARL19, SLH14, SAL12, SSVA13, SSAK12, SYK14, SC19, SSVA19, SYGR11, UPF17, WW15, YYJC12, YG18, ZHZ^{+}15, ZZ19, dCGÜ^{+}17, dRMC^{+}16]. control-flow [HLB19, HT12b]. controlled [ACS19, KJP^{+}19, SM10]. controlled-channel [KJP^{+}19]. controller [AM18a]. controls [ASFI18, AM17b, KKK17, SMP13]. controversy [ASO19]. Convergence [GCC18]. convolution [HW19]. convolutional [KMG17, YLW^{+}19]. CooPeD [GMGTdFR14a]. cooperation [JJ14]. cooperative [HS13, IHED15, LKJ14]. coordinated [ZLK10]. Coordinates [SLH14]. coordination [JAAS19]. copy [YCJ12]. corporate [Wor10]. CorrCorr [GCD19]. correct [PS10]. correctly [KKP^{+}18]. correlated [LZ19]. Correlating [GBC^{+}18, YW11]. correlation [FSSF15, GCD19, HXW^{+}19, RAA15, SHGH^{+}15, TFPC10]. Corrigendum [GZ20, GZG20, HZG20]. Cortana [SS17b]. cost [COS19, SGK^{+}16, WZK13, ZS18, ZNAHK11, Zhu18, Zin15]. cost-benefit [WZK13]. cost-sensitive [ZS18, ZNAHK11]. counter [EGM10]. counterattack [AWDB19]. counterfeit [MHS17]. Countering [JvOW11]. countermeasure [KJP^{+}19, MK19, hRKRR13, VVS12, ZTL^{+}19]. countermeasures [ArMS18, AZ17, CW19, FHL14, HLB19, SHB12]. counters [DCC^{+}18]. countries [MT17, NTK^{+}19]. counts [JXK18]. coupons [FGHHR18]. courses [GMdF19]. Court [ZY11]. Cover [Ano17l, Ano17m, Ano17n, Ano18i, MW11]. coverage [Zin15]. Covert [GB19, ROL13, SKK^{+}14, AG14, KJKM11, MZJ^{+}12, ZYPL10]. CPBAC [JJ14]. CPI [LGZ^{+}19]. CPI-enabled [LGZ^{+}19]. create [ROW19]. Creating [RM17, GZG19, GZG20, RAMR17]. creation [GJJ18, YGZG19]. creator [wJKW^{+}16]. credentials [LAHV17, YAXY19]. Creditability [LLHT13]. Creditability-based [LLHT13]. CRiBAC [JJ12]. crime [LSC14, RR14]. criminals [MAC17]. Critical [Mic12c, Sil19, ASFI18, BCK19, CAL15, GSA18, MSR19, NZGH16, ZH13]. Criticality [TKG10]. cross [MWL18, VC12]. cross-cultural [MWL18]. cross-site [VC12]. crowd [XLT^{+}17]. crowdsourcing [AAA19]. crucial [WLF^{+}19]. Crypt [HHAW19]. crypto [CW19]. crypto-ransomware [CW19]. cryptographic [BJ10, GJJ18, KKJ^{+}16, Ksi12, ZHL^{+}11]. cryptographically [CBL10]. Cryptography [Ano19i, DA18, Kam19]. Cuckoo [Zin15]. cultural [MWL18, TKK15]. culture [AOR18, AS13, DE10, DSP16, FAE14, FE16, MT17, VV10, dVM15]. cultures [dVM17]. cure [BBC17]. currency [ML14]. Current [ANMN19, AAMDGR18, RARL19, TLJ14, Kok17]. curtain [SB14]. customizable [MFBC11]. CUSUM [CGPP12]. Cyber [ANMN19, AM17a, Bay12a, BJW19, Cae12, FB14, Kal19, KvS10, MK19, Mic12c, NWD^{+}12, RJJ15, TCNFW16, TG12, WMPA19, ZH13, vdWEG18, AKFM18, AAH^{+}18, ALC17, AM18a, Ano13-30, ABR16, BRM^{+}18, CHK^{+}19, CM19, CBB^{+}16, Cho11, CJSM17, Den14, DM19, EM19, Fra17, GBC^{+}18, GHDGRLM17, HBBM16, HLB^{+}18, HWS^{+}13, HLL17, LSC14, LDB18, MMJ14, MWAG16, NPP17, OS17, PWL^{+}19, PC12, PTM^{+}18, QAR^{+}17, RFW^{+}18, Rua17, RARL19, RR14, STSB17, SSF16, SKTO13, TR18, TS13, VHBJ19, YWND14, YG18, ZLL^{+}18, dLGHK18, vSvN13]. Cyber-attack [Kal19]. Cyber-attacks [MK19, GHDGRLM17, PTM^{+}18]. cyber-crime [RR14]. cyber-defense [RARL19]. cyber-exercises [dLGHK18]. Cyber-Physical [Mic12c, BJW19, ZH13, ALC17, CHK^{+}19, HLB^{+}18]. cyber-safety [DM19]. Cyber-security [vdWEG18, LDB18, MWAG16]. cyber-threat [QAR^{+}17]. Cyber-Warfare [NWD^{+}12]. cyber-wargame [HWS^{+}13]. cyberbullying [MBSR18]. Cybercrime [BS19, GT13, MR11, CW19, KSW16, OM13, TVPV19]. cybernomics [Rua17]. Cybersecurity [CDKR18, ATAAd18, ADWM18, BWMG17, BTP^{+}18, CBGN18, GMdF19, Hat18, LSSK16, Les18, MB17, SSSC18, dFGMTPL17]. cycle [JLK16]. cyclically [MHL^{+}17]. cyclostationarity [MFCMC^{+}18]. cyclostationarity-based [MFCMC^{+}18]. Cylindrical [SLH14]. CyTrONE [BTP^{+}18].

D [NCS14, Woo19, MGYC18, PYLL19]. D2D [Gop19, JLGM19]. D2D-Aided [Gop19]. damage [SBS^{+}18]. danger [AP13]. DangerNeighbor [GSL^{+}19]. Data [CGGFGMF19, DC11, FB19, GMGTdFR14a, LML^{+}17, LW19, QAR^{+}17, XYY^{+}12, AMK18, ALF^{+}18, AHGC19, BRH15, BTK15, BFTJ18, CMPR15, CBL^{+}19, DFJ^{+}13, DEP13, DMC15, DYL^{+}14, DYZ^{+}15, EE15, Gar13, GZS19, GSBY18, HMNT10, Hor18, HLL17, JKA^{+}18, JGM^{+}13, Jän14, JWB16, JAN19, LKJ14, LZCX18, LWS^{+}19, LMLL19, LCZ17, LXH^{+}18, LZ19, LMG18, MCN10, MCY^{+}18, Med18, MGYC18, MAZAS17, MKG16, MMMG13, NAS19, PKK^{+}14, PCC^{+}18, PAS^{+}18, QLZ19, RR16, RR17, Ril13, RWS^{+}19, SKA^{+}19, SPFB19, SLB^{+}13, SMK^{+}18a, SA19, SH15, SS18b, SSVA19, SPK17, SWLZ11, WLWG11, WL18, WLZ^{+}19, WX19, WLF^{+}19, YWZM18, YWS17, ZKM^{+}17, ZYWZ14, ZLL^{+}18, ZZC17]. data-access [BRH15]. data-aware [WLZ^{+}19]. Data-driven [QAR^{+}17, MAZAS17]. Data-flow [LW19]. data-hiding [EE15]. data-only [Ril13]. data-oriented [WLZ^{+}19]. database [AAD^{+}12, EHBH^{+}16, LMLL19, SVGE14, SSVA19, TFE10, VEPZ15, XMY^{+}17, YGX^{+}19]. databases [GDR15, JMGW17]. dataset [AOR18, MFCMC^{+}18, TAP18, UB12]. datasets [HMNT10, SSTG12]. day [TCNFW16, VG14]. DBMS [SERF12]. DBod [WLCC17]. DCT [TLZL16]. DDoS [DT19, MPCK18, SBZD17, SP15, SKKP12, SMB16, SSK17, SKTO13, ZGD19]. Dead [Ano14s, TBDM14]. Deadline [VPNP18]. debate [Gar16]. Debian [WN15]. Debugging [OM19]. decade [HBBM16]. Deceiving [ÖB15]. Decentralized [CMPR15, YAXY19, BAH18, HHBS18, HBB12, YYJC12]. deception [HWS^{+}13, vdWEG18]. deceptive [AW18]. decimal [DDSP17]. Decision [Jän14, NDdL15, BWMG17, BS19, CCM^{+}15, DE16, MWAG16, RMW10, ST19, SAL12, WYS18]. decision-making [BWMG17, DE16, WYS18]. decoder [HS18]. decomposition [CDV12]. decoys [VSB^{+}19]. decryptable [MBP19]. deductions [ZGZ^{+}13]. deductions-based [ZGZ^{+}13]. Deep [CAM^{+}19b, HM19, NNNQ18, NQZ18, XSD19]. defeat [VC12]. Defeating [NDMM13]. defence [CJSM17, GJ18, SKTO13]. defending [CP18]. Defense [ADAMH17, VG14, ADA^{+}18, AAL16, BRM^{+}18, DCC^{+}18, Den14, HWS^{+}13, HEK^{+}18, KN14, RARL19, SSF16, lTLqZqC19, TUB16, WP12, ZS18, qZLxCjY17]. defenses [CXF^{+}18, GFN11]. Defined [AX15, GWKFC19, TDJB18, HHK^{+}19]. Defining [dVM17]. Definition [AWDB19, DMR19]. degree [ZLC13]. Delay [DFC13, DFCC16, KUA^{+}18, MVK16]. delays [GOB11]. Delegate [MKG13]. delegation [AWZ^{+}17, AAC^{+}19, DLH^{+}17, HMFE10, PRMD10, ZDBW11]. deletion [YWZM18]. deliberately [AAL16]. delivery [KMM13]. demand [KKJ^{+}16]. demographically [AOR18]. demographically-diverse [AOR18]. demonstrator [DBH17]. deniable [CCC^{+}18, KKKP18]. Denial [BSMT14, EGM10, HWS^{+}13, SLP13, SSK16, TUB16, TH18]. Denial-of-Service [BSMT14, EGM10, SLP13, TUB16]. deobfuscation [BJLS17]. dependable [AGSR^{+}18, Hor18]. dependence [ESY^{+}15]. dependency [KKKP18]. deploying [HPF^{+}16, dLGHK18]. deployment [JJL^{+}19]. depth [ASFI18]. Deriving [PRS13]. Design [GMdF19, Hor18, JM13, JCD14, KPKS15, MSR19, THA^{+}13, WRTA17, WJ15, Woo19, ZKM^{+}17, ZJK^{+}17, ZS15, BBDP16, BR17, BOZL13, DRN16, HCR10, HCRMBG17, JSWZ11, KFGL19, PMP^{+}15, PAS^{+}18, RDAS19, SVGE14, YHL^{+}15]. Designing [KJKM11, PDM19, CM17, BBC17]. desktop [LLX^{+}17, TMN^{+}17]. destruction [GSBY18]. desynchronisation [LDC13]. detect [HR13, XXLZ18, YW11]. Detecting [AM13, HXW^{+}19, HL18, JLGM19, JC14, JLK16, KPS16, KYZ15, LJS19, MJW^{+}18, TTCB18, YYL^{+}15, BM19, BVWM19, Cho13a, DLLS19, ESCAM17, GK14, HS18, HZH^{+}18, LHI^{+}19, MZJ^{+}12, NTK^{+}19, PDM19, SCLD19, WLCC17, ZYRZ16, ZjHtMhBs12]. Detection [CYLL17, FB19, FAS^{+}17, LAVK19, MK19, NAS19, NVAJ19, NCGE15, XSD19, AR17, AHTS15, AQR^{+}17, AYT^{+}14, AH18, AG14, ATG11, AMS18, AP13, BK19, BMMP18, hBCK17, BHC15, BJJ^{+}17, BEGW^{+}19, BLS^{+}16, BRA18, CGPP11, CGPP12, CGP17, CPVGTMF16, CGGFGMF19, COS19, CMPR15, CXF^{+}18, CAM^{+}18, CCST15, CH14, CWCC12, CFSF18, DT19, DC11, DF16, DYT^{+}13, DDYZ14, DXCL18, DBH17, EMBH14, ESY^{+}15, EM19, FNE19, FSSF15, FRRP17, GGSZ14, GOB11, GWWS19, GCD19, GWWW19, GZ10, GZP^{+}13, HDK18, HX17, HS13, HSPX14, HXS^{+}16, HDT^{+}19, HB16, HS11, IRC^{+}17, Kal19, KGSN17, KK10a, KK10b, KK17, KKM11, LWS^{+}18, LDC13, LMLK16, LWS^{+}19, LSA^{+}19, LLHT13, LWZ^{+}17, LGS14, MPCK18, MCT^{+}17, MCY^{+}18, MLK14, MMF19, MPORBR12, MLH^{+}18, MSRF18, NADZ18, NDP18, NNNQ18, NRM19, NLC^{+}19, ORLS14, ÖB15, PF10, PSK^{+}17, PHP19]. detection [PAP^{+}17, PCCM16, RAA15, RW13, RP19, RAA^{+}14, RWS^{+}19, SD17, SE11, SS18a, SMM^{+}19, SG10a, SLH14, STCM^{+}14, SWZL14, SRU18, SCGM14, SMC^{+}18, SKKP12, SSTG12, SHGH^{+}15, SSK19, SK10, SBW10, SDM^{+}18, SKG17, Su10, TXM^{+}19, TC11, TH18, UPSGF^{+}14, USB17, VMS18, VSO^{+}18, VDK18, WB16, WW18, WWXZ18, XPS11, Yad19, YCJ12, YG18, YXD^{+}11, ZY11, ZNAHK11, ZQY^{+}19, ZTC19, ZZ19, ZTS^{+}13, ZLK10, ZW19, ZW14, vdWEG18, VRRK17]. detection/prevention [YG18]. detector [CMV16, VMA16]. deter [ROW19]. Determinants [TMW17, DM19]. Determining [PMB^{+}14]. deterrence [CLL^{+}13]. deterring [ESCAM17]. develop [WA19]. Developing [GPGV16, HSL19, MT17, NTK^{+}19, SSTG12]. Development [SWS16, AAD^{+}12, BCCS18, EHBH^{+}16, FT16]. deviations [STCM^{+}14]. device [CRS13, LMS14, OSP^{+}19, SSSC18, TMW17, WHSE15]. Devices [SG15b, BKK19, CCC^{+}18, DC16, DMC15, HZMH18, HYLY19, KMW19, NSBM17, NRM19, SM19a, SSNS15, SPCCMMS18, SCR19, TZTC16, YWZM18]. devour [SP15]. Dexteroid [JLK16]. DGA [WLCC17]. DGA-based [WLCC17]. DGAs [XSD19]. DHCP [HT17]. Diagrams [NDdL15]. dialogue [AH10]. Diary [SG14b]. dictionaries [ABR15]. dictionary [SS13]. difference [dSFK16]. differences [KK16, LHM13]. differencing [SH15]. different [CZH^{+}14, LK18]. Differential [WX19, CVS18, JXK18, LZ19, WS17]. Differentially [CSX^{+}15, CSXL15, WLF^{+}19]. Digital [EAML15, Gar13, GT13, LW12d, Wed12, ALZ^{+}11, DM19, HAGTdFR13, Hor18, KEE13, MMMG13, RV13, RB11, RB12, SE11]. Dilemma [Sil13]. dimensions [SSF16]. direct [XdZMsC18, YLB^{+}17]. directional [SVA10]. directions [ArMS18, AMS18, Cho11, CJL^{+}13, LCZ17, MK19, NCGE15, SH15, WMPA19]. Dirichlet [RW13]. disambiguation [PSMG18]. disaster [BMO18]. discipline [CDKR18]. disclose [PKWP19]. disclosing [PKP17]. disclosure [AF14, HLFZ16, MW11, ZZC17]. disclosures [JGLE16, Ven18]. discounts [PWL^{+}19]. Discouraging [BWOD13]. discourse [KHG17]. Discovering [GZS19]. Discovery [LC15, MCG19, RW13, SLD12, WJAM11, XLT^{+}17]. discrete [MMPC18]. disinformation [AWDB19]. disk [KJKM11]. displays [Rru16]. disruption [FKSK19]. Dissecting [MG12]. dissemination [TOZJ13]. dissimilarity [ATAAd18]. distance [KTCS19, MOV14, WN15]. distinguish [BLR17, GMMP18]. distortion [PYLL19]. Distributed [BSMT14, CZZ^{+}19, FW10, MCL^{+}13, ZY11, AKAuR19, BVWM19, DLDTGMMP15, JAAS19, KS14, Len12, NDSA17, PDB14, SLP13, TUB16, UFF12, UFF15, Uzu16, XHC^{+}10]. distribution [KYZ15, MFBC11, ZYWZ14]. diverse [AOR18, DM19]. diversity [MC18]. divided [HYZ18]. DMTP [YtLpZD14]. DNP3 [MCN10]. DNS [AKK^{+}13, HBF13, NAS19, SSK19, SPD^{+}15, WLCC17]. DNS64 [LK18]. Do [SW19]. document [MKG16]. documentation [MGYC18]. documents [ALZ^{+}11]. Does [KBM16, MOV14, ROW19]. Domain [HT12a, MJRK18, BPG^{+}19, CAY^{+}18, CWY^{+}12, HBBM16, MPPS10, SLH14, XSD19, ZW19]. DomainChroma [CAY^{+}18]. dominant [dVM17]. domRBAC [GM12]. Don’t [BWOD13]. DOPdefender [WLZ^{+}19]. DoS-resistant [TOZJ13]. DoS/DDoS [SKTO13]. Double [MCRB19, MB18]. download [GAA18]. DP [CSXL15]. DP-Apriori [CSXL15]. draw [MLKC17, NSBM17]. DRAW-A-PIN [NSBM17]. drawn [NSBM17]. driven [APMCR13, CDV12, LSQ15, MAZAS17, QAR^{+}17, ZGZ^{+}13]. driver [GBC19, OP11]. DriverAuth [GBC19]. DRM [DC16]. Droid [MPM15]. DroidInjector [FSZ^{+}17]. DroidNative [AQR^{+}17]. Dual [DLOP17, Sil13]. Dual-use [Sil13]. Dumpsys [wJKW^{+}16]. during [QFG12]. duty [UPF17]. Dynamic [HEK^{+}18, LWS^{+}18, LLG18, SAL12, TZZG13, VMS18, BMMP18, BSBG19, BM17b, COS19, DT19, DFJ^{+}13, FSZ^{+}17, HXW^{+}19, HBT^{+}19, HPF^{+}16, KTG^{+}18, LDC13, LKPK18, LPL18, MRS12, SKA^{+}19, SM10, SKG17, SPR18, SMK18b, TZMG15, ZZ12, ZLW^{+}15, ZWNM19]. dynamics [GEAHR11, GDR15, ICRB14, MCRB19, PGdCL16, RMG11, SCGM14, SSY12, TZTC16, UB12]. Dypermin [LLG18]. Dytaint [ZLW^{+}15].

E-ADOPT [MJBKZ15]. e-auction [SYCS11]. e-commerce [BGR11]. e-gates [dRMC^{+}16]. e-health [OSP^{+}19]. e-mail [BPA^{+}11, PW13]. e-voting [Gar16]. Early [RBJ18]. Early-stage [RBJ18]. ECC [CBL10]. ECG [HW19]. eclecticism [ASO19]. EclipseIoT [AAR^{+}18]. ecommerce [MT17]. economic [CGR18, KSW16, Rua17]. Economy [GT13]. ecosystem [AYY^{+}17, KNX18, KBM16]. ecosystems [NNG^{+}18, PAP^{+}17, SM19a]. edge [HZMH18, JWZX19, SSSC18, TXM^{+}19]. editing [PSMG18]. Editorial [Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano12r, Ano12k, Ano12l, Ano12t, Ano12u, Ano12v, Ano13k, Ano13l, Ano13m, Ano13n, Ano13x, Ano14i, Ano14j, Ano15i, Ano15j, Ano15q, Ano15r, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano18b, Ano18c, Ano18i, Ano19a, BCCS18, GE10, GD10, GS10, GL10, GT13, GJ11, HF17, HLL17, RCB^{+}14, Spa10a, Spa10b, Spa11a, Spa11b, Spa11c, Spa11d, Spa12a, Spa12b, Spa12c, Spa12d, Spa12e, Spa13a, Spa13b, Spa14, Spa19, XHZ11, YWND14, Ano19i, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13o, Ano13p, Ano13q, Ano13r]. Editorial [Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. eDocument [LPR10]. education [Bon19, CDKR18, HSL19, RE19]. Effect [AAT16, CGW15, AM17b, AM18b, DLS14, FAE14, ROW19]. Effective [BEGW^{+}19, SWZL14, Yad19, DYL^{+}14, KH12, KKK17, LSA^{+}19, TUB16, ZWNM19]. effectively [LMG18]. Effectiveness [BJLS17, FAS^{+}17, ACS19, AHGC19, BWMG17, HEK^{+}18, LW19, SMM^{+}19, SGK^{+}16]. effects [HOVB17, MAC^{+}15]. efficacious [GK14]. efficacy [BWMG17]. efficiency [PAS^{+}18]. Efficient [DA18, DT19, GZ10, GZP^{+}13, LNNH13, LPL18, LXH^{+}18, OSP^{+}19, WZW16, XMY^{+}17, YWS17, ZZC17, AuRMC^{+}18, ABR15, BEGW^{+}19, BRA18, CBL^{+}19, CTH10, CDE^{+}11, FZZ^{+}12, GMLZ11, HDMKS10, HMFE10, HYLY19, KTCS19, KPD15, LML^{+}16, MDHM18, MJBKZ15, MTT^{+}16, MT19, NLC^{+}19, NXS10, PDM19, SKA^{+}19, SLD12, SMK18b, THA^{+}13, VSO^{+}18, XLT^{+}17, ZLL^{+}19]. Efficiently [MHL^{+}17]. EFM [MLK14]. eID [ZS15]. electrical [BMH16]. electroencephalograms [BD16]. electromagnetic [UAK16]. electronic [AuRMC^{+}18, FGHHR18, FGOPL10, HK13, JAH^{+}11, LPR10, MPPCFG^{+}12]. Elephants [SP15]. Eliciting [NFGL16]. Elsevier [SG15a, TG15]. Email [VRRK17, ATL^{+}17, MB18, VVS12]. embedded [NZGH16, VSO^{+}18, WHZ^{+}19]. embedding [LHM13, SA19, ZQY^{+}19]. Emerging [YST^{+}19, AAMDGR18, DPPB17b]. emini [GSZX18]. eMotion [PPKK19]. Empirical [GSA18, GHDGRLM17, ZDW19, CMV16, DPPB17a, GGSZ14, HW18, HL18, KL18, RHL15, SBK12, SYX^{+}16, SHB12, YGZG19]. employee [AM17b, HSL19, MMPC18, PMB^{+}14]. employing [GCC18]. employment [SW19]. emulation [JJL^{+}19]. Enable [ZDBW11]. enabled [CJA^{+}19, DPP15, LGZ^{+}19, NML19, YST^{+}19]. enclave [CKJ^{+}17]. enclave-based [CKJ^{+}17]. enclaves [PPKK19]. encoder [HS18]. encoder-decoder [HS18]. encoding [DDSP17]. encrypted [BBDP16, BTK15, XMY^{+}17, ZZC17]. Encryption [SERF12, BTK15, CJW^{+}19, HHAW19, LK10, MBP19, OPHC16, OSNZ19, SVGE14, WLWG11, ZCZ^{+}19]. Encryption-based [SERF12]. end [Bel18, CZZ^{+}19, JFR13, RR14]. end-to-end [CZZ^{+}19, JFR13]. end-users [RR14]. endowment [ROW19]. enemy [AS13]. Energy [CDE^{+}11, BMMP18, CJA^{+}19, SMK18b, VSO^{+}18]. energy-efficient [VSO^{+}18]. Enforcement [KV15, TDJB18, AHP18, BCD^{+}10, BR18, CKT11, KvS10, LKJ14, MT19, PMDM10a]. Enforcing [AIRC13, BGR11, DFJ^{+}13, NSGF16]. engine [BHC15, HDMKS10, RP19]. engineered [JLK16]. Engineering [AKP14, CCC^{+}12, HW18, CWGD15, DF16, ELG^{+}17, FE16, Hat18, Hat19, HL18, LM10, MLM17, MMKV15, MLV16, OP11, SNG^{+}17, ZBBO15]. Enhance [KTG^{+}18, Kar12]. Enhanced [PGdCL16, YI17, DEP13, MLK14, SD17, SMA^{+}18]. enhancement [TZZG13, TZMG15]. enhancements [NCGE15]. Enhancing [EMBH14, FanE^{+}18, MLK14, SK13, HZNF15, MLKC17]. enigmatic [KK16]. ensemble [AR17, BM19, GWWW19, HWX^{+}19, IRC^{+}17, LWZ^{+}17, ZW19]. ensure [Hor18]. Ensuring [MSMK18]. Enterprise [WL12, KLN15, ODB18, PAS^{+}18, SC19, BMO18]. entity [RW13]. entropy [CMV16, ÖB15]. entry [LNK16, NKL^{+}18, YHL^{+}15]. enumeration [KPKS15]. environment [AAT16, AAMDGR18, CAM^{+}19b, JK19a, KS19, LKJ14, PAJAP15, PSMG18, RG19, RR16, SYWX19, VPNP18]. environments [ANMN19, AZ17, AIRC13, BPA^{+}11, BR17, DLDTGMMP15, Lop16, SVA10, SSSC18, VGL14, WHZ^{+}18]. EPC [LC15]. EPCglobal [SLD12]. epidemic [WZW^{+}10]. episode [RAA15]. equilibria [AAdM19]. equipment [PYLL19]. error [CJW^{+}19, EHMJ19, KPD15, SKG17]. error-related [EHMJ19]. ESPOON [AIRC13]. Establishing [DFC13]. establishment [BEB^{+}18, NM19]. Estimating [RMG11]. estimation [CLLK10, GZ18, GZ20, RDAS19, bORF^{+}15]. estimators [SKKP12]. ETGuard [JLGM19]. Ethereum [WLW^{+}19]. Ethical [RMW10]. ethics [Hat19, MMKV15, Spa10c, Spa11e]. European [RR14]. evade [KJKM11]. evaluate [DC16]. Evaluating [ACS19, AM17b, CZH^{+}14, LDB18, Len12, MB18, NAMX14, RE19, EHMJ19]. Evaluation [NDdL15, NMX19, SMB16, TBSR14, AMK18, AAJ12, CHRI16, Cho13b, HL18, HT12b, KS19, MFCMC^{+}18, MVK16, PTM^{+}18, SLG19, SCGM14, SHB12, THA^{+}13, WJ15, WLW^{+}19, WYS18, Woo19, XPS11, XNP16, YHL^{+}15, ZJK^{+}17]. evaluations [AIB^{+}14, CSBW18]. evasion [JHJ^{+}16, KKP^{+}18]. event [FSSF15]. Events [BSMT14, SJCM17, SA16, TS13, ZYRZ16]. evidence [GZS19, MMMG13, PHMN16]. evidential [SOG15]. Evil [BMO18, JLGM19, NADZ18]. EVIV [JFR13]. Evolution [CDKR18, Hat18, LWS^{+}18]. evolutionary [BSL19]. Evolvable [YWL^{+}17]. evolving [FFGL18, FC12, LSSK16]. EvoPass [YWL^{+}17]. examination [MWL18, PBR11]. Examining [HKK17, LW12d, MT17]. example [KKP^{+}18]. examples [MLV16]. exams [BGLR17]. exchange [CIS17, MVK16, SBL19]. excuses [BWOD13]. executable [UPSGF^{+}14]. execution [JvOW11, JK19a, LLL^{+}18, SMC^{+}18, SS18b, ZWNM19].

Executive [WL12]. exercises [BM17a, dLGHK18]. Exfiltrating [DMC15]. Exfiltration [FB19, GB19, NAS19]. exfiltration/infiltration [GB19]. existence [VBC^{+}15]. Existing [SG10b, CM17]. experience [PAS^{+}18, SK18]. experiences [MPORBR12, VRRK17]. experiment [HWS^{+}13, MMPC18]. Experimental [Hor18, SDM^{+}18]. experiments [WN15]. Expert [EAML15, HA15, MWAG16]. expert-based [HA15]. Expertise [GPGV16]. Explaining [GGV18, HXW^{+}19, Sil19]. explanation [KMW19]. explanatory [RE19]. Exploiting [HBF13, HL12, SMK^{+}18a, GZS19, SH15, YWYZ12]. exploits [MTC^{+}18]. exploration [LZBS13]. Exploratory [PCCM16, KDC^{+}14, WYS18]. Exploring [AM18b, HPJ^{+}19, KMG17, NGSV16, TMN^{+}17, WZK13, CKJ^{+}17]. exposure [RR14, SS11]. expressive [GMGTdFR14b]. Extended [YtLpZD14, BCD^{+}17, Kam19, MAC^{+}15]. Extending [AWZ^{+}17, CKT11]. extensible [JAAS19, SLS^{+}18, ZTC19]. Extension [WA19, PPKK19, SGLC16, Ver18, MM15a]. Extensions [HDMKS10, OBRK15, AF14, JAAS19, KTCS19, MJRK18, SWZL14]. external [BTLT16, LKPK18, WHZ^{+}18]. extraction [MZJ^{+}12, SPYG15]. extreme [ZJK^{+}17]. extrusion [MKG16].

face [SWS16]. face-recognition [SWS16]. Facebook [PZ15, SCRA17, SBS^{+}18]. Facebook-style [PZ15]. faces [DLS14]. facial [dRMC^{+}16]. factor [BD18, DRN16, GMMJ11, MDHM18, NMX15, SNG^{+}17]. Factors [PKWP19, ArMS18, CGW15, HK17, PKP17]. fails [MHS17]. Fair [ALZ^{+}11, CTH10, DAR14, PW13]. fake [BLS^{+}16, KYZ15]. False [LXDW16, LMLK16, LLHT13, SK10]. Familiarity [JvS17]. families [DWZ19]. family [DXCL18, LLL15, ZTC19]. Fast [LAHV17, ASO19, DYT^{+}13, LNNH13]. FATF [Cho13b]. fault [AGL12, DEP13]. fault-tolerant [AGL12]. FCMDT [BSBG19]. FDM [BD18]. Feature [MZJ^{+}12, YXD^{+}11, ATAAd18, BK19, BFTJ18, DF16, GCD19, GWWW19, HDK18, KK17, VDK18, ZQY^{+}19]. feature-hybrid [ZQY^{+}19]. features [LGS14, NMX15, PMDM10b, VMA16, YCH18]. February [Ano10-33, Ano12w, Ano13-27, Ano14t, Ano15y, Ano16s]. federated [BSBG19, LC15]. feedback [HBB12]. fi [NADZ18]. fidelity [MAM15]. field [ABP16, RW13]. Fields [SLN^{+}15]. File [XYY^{+}12, CWGD15, Kha12, LWS^{+}18, SFK15, VSB^{+}19, WP10]. files [GSZX18, NCGE15]. Filler [Ano12s]. filming [MJW^{+}18]. Filter [SLP13, MLK14, hRKRR13, SS17a]. filtering [TFPC10, TZMG15, UPSGF^{+}14, YKH10]. filters [KP18, MT19, RB14, TZMG15]. financial [CKA^{+}17, LSC14, WB16]. Find [ZTL^{+}19]. finding [DEP13, KPKS15]. Fine [KJP^{+}19, RLB^{+}11, BHC18]. Fine-grained [KJP^{+}19, RLB^{+}11, BHC18]. finger [NSBM17]. finger-drawn [NSBM17]. fingerprint [BCRAA13, HW19, ZHL^{+}11]. fingerprinting [BHDA14, CTH10, FXC18]. fingerprints [DSNH17]. Firefly [BK19].

Firefox [OBRK15]. firewall [GACCB^{+}13, HHK^{+}19, LGZ^{+}19, TZZG13, TZMG15]. firewalls [AGL12, MBE^{+}12]. firm [BS19]. firms [KCJB12]. firmware [HYLY19]. first [PAS^{+}18]. Fishbone [KS19]. fixing [AMMV13, SHB12]. Flash [BSMT14]. flaw [ACC^{+}13]. flaws [HCRMBG17]. flexible [Gar16]. flights [Zin15]. flood [DT19, SSK17]. flooding [DSNH17, HK13, hRKRR13]. Flow [AMMV13, FXC18, RSLH19, USB17, BSTT16, BCD^{+}17, BL15, DT19, DDYZ14, DBH17, DPP^{+}19, GWWS19, HLB19, HT12b, LW19, NNNQ18, QHW^{+}19, STA18, WW15, YYJC12, ZTS^{+}13, dCGÜ^{+}17]. Flow-based [RSLH19, USB17, DT19, DDYZ14, GWWS19]. Flow-net [FXC18]. flow-sensitive [BCD^{+}17, STA18]. fly [MRS12]. FNF [FXC18]. Focusing [LLK16]. fog [Gop19, SSSC18, ZGD19]. following [PBR11]. forecasting [GGBT16]. Forensic [BBD^{+}19, Cho13a, Coh10, EAML15, KEE13, NAMX14, RV13, RB11, RB12, SS17b, SS18b, ZBB17]. Forensics [LW12d, Wed12, HBBM16, Hor18, KJKM11, LW12a, ZLW19]. foresight [GTGB16]. forgeries [SSY12]. forgery [Cho13a, HZH^{+}18]. Formal [ABP16, DJP17, MBE^{+}12, SHB12, TdHRZ17, GKB14, NZGH16, RDAS19, RB12, THA^{+}13, WJ15, WCDY18]. formal-analyzed [WCDY18]. formalization [BR17]. formalized [NML19]. Formally [AW18, DPP^{+}19, HLB19]. format [CWGD15, SS18b]. formation [DPPB17b, SSS^{+}15, SMK18b]. formats [MP18]. formed [SLB^{+}13]. Foundations [Mic12c, Nie17]. four [CKA^{+}17]. Fragmentation [NVAJ19]. frame [HZH^{+}18]. frames [UFF15]. Framework [Den14, Hor18, NF15, AHTS15, AAL16, ABR16, BPG^{+}19, BJ13, BCRAA13, BTP^{+}18, BSMT14, BS17, CCC^{+}12, CFSF18, CRS13, CWGD15, DE10, EAML15, FNE19, HB18, HL18, HT12b, HAHZ^{+}18, JAH^{+}11, KBAL13, KS19, LZBS13, LLG18, MKP12, MB17, MAZAS17, NSGF16, PDM19, PMDM10b, PMDM10a, PMLK^{+}11, Ril13, Rua17, SBZD17, SLG19, SLS^{+}18, SSVA19, SSSC18, SKTO13, VEPZ15, ZBBO15, ZLLL12, ZTC19, ZLW^{+}15, ZWNM19]. frameworks [RE19]. framing [CGLP18]. fraud [CCM^{+}15, HX17, LMNM^{+}19, WB16]. FRED [Hor18]. free [DJP17, JMGW17, MM17]. freeware [GAA18]. Freeze [HHAW19]. frequent [CSXL15, CSX^{+}15, HB16, LMLL19]. Friend [KKP^{+}18, JJA13]. Friend-safe [KKP^{+}18]. friendly [CCC^{+}18, KKP^{+}18, ZS15]. friends [KPKS15]. Front [Ano17l, Ano17m, Ano17n, Ano18i]. FT [AGL12]. FT-FW [AGL12]. FTC [BB11]. Full [LLY17]. fully [BVWM19, MBP19]. FunCAPTCHA [HCRMBG17]. function [AHR12]. functions [HL12]. Fundamentals [Mic12a]. further [PCP^{+}17]. fusion [CM19, HW19, HXW^{+}19, HZH^{+}18, WW18]. fusion-based [WW18]. Future [CJL^{+}13, ANMN19, AMS18, BMH16, Cho11, FYL^{+}17, LCZ17, MK19, ZKM^{+}17]. fuzz [KCB13]. fuzzer [CSWX19]. fuzzing [CCM^{+}18]. fuzzy [AAH^{+}18, BSBG19, HK17, SKKP12]. FW [AGL12].

GA [CSS^{+}19, KK17]. GA-LR [KK17]. gained [RR14]. Gait [DB13, NMX15]. Gaithashing [NMX15]. gambling [HWX^{+}19]. game [AAdM19, AW18, KK10a, LAVK19, MPM15, MBSR18, SYK14, SKTO13, lTLqZqC19]. game-based [AW18, KK10a]. games [AAdM19, DLB^{+}19, GCC18, HBT^{+}19]. GAN [HM19]. gap [GB19]. GARS [MMJ14]. gate [QHW^{+}19]. gates [dRMC^{+}16]. Gateway [NADZ18]. Gaussian [AH18]. general [WP12]. generalised [LML^{+}17]. generate [SSTG12]. generated [CBL10, XSD19]. generating [BSMT14]. Generation [Mic12b, BD16, CM17, CMM^{+}10, DBH17, DDSP17, GTDVTSH15, HM19, KCB13, RSLH19, SMC^{+}18, SPK17, ZHL^{+}11]. Generative [RSLH19]. generator [ZZZY13]. generic [HK13, NXS10]. genetic [CSS^{+}19, VDK18]. genetic-algorithm-based [VDK18]. genomic [PCC^{+}18]. genuineness [TS13]. geometry [MZZ^{+}18]. gesture [SCR19]. gesture-typing [SCR19]. GET [SSK17]. Global [SYK14, SM19b]. GM [EM19]. Goalkeeper [BR18]. Golbeck [SG15a]. gold [ELG^{+}17]. governance [FAE14, HSL19]. GPGPU [CBL10]. grained [BHC18, KJP^{+}19, LLCT13, RLB^{+}11]. gram [ORLS14, SS18a]. grammar [KCB13]. Graph [BM17b, DXCL18, DLB^{+}19, KPS16, LDB18, NNNQ18, PRS13, PWWY14, WZK13, WL18]. graph-based [WL18]. graphical [MLKC17, YWL^{+}17]. graphs [CPGA13, EMBH14, GSA18, KLN15]. Gray [SG14a]. graylist [YtLpZD14]. great [JL10]. Greedy [BSL19]. Grid [HX17, PAP^{+}17, CRJ18, JK19b, KNX18, LSSK16, Les18, Sul19, WZW16, XXLZ18, XdZMsC18, ZH13]. grids [KKS13]. grooming [MMJ14]. Ground [Mic12e, SPD^{+}15]. group [MK11]. grouping [KPD15, XXLZ18]. grouping-based [XXLZ18]. growth [RHL15]. GTRBAC [MSA11]. Guest [Ano19i, GSZX18]. Guest-transparent [GSZX18]. Guide [KG13, LW12d, MG12, TG14]. guidelines [JCD14].

habitual [KKK18]. hacker [GPGV16]. Hacking [LW12c, Mic12b, SG13, SG15b, Hat19, TG15]. Hadoop [VPNP18]. HAIS [PMB^{+}14, PCP^{+}17]. HAIS-Q [PMB^{+}14, PCP^{+}17]. HAISA [Ano11m]. halved [SSF16]. hand [JAN19]. Handbook [Mic12c]. handling [RC15]. handoff [RRK^{+}17]. handover [FZZ^{+}12, LNNH13]. Hands [MG12, BTP^{+}18, SG15a, dLGHK18]. Hands-On [MG12, BTP^{+}18, SG15a, dLGHK18]. handshake [SM10]. Happy [DLS14]. harassment [Wor10]. Hardening [DLB^{+}19, WZK13]. hardware [DCC^{+}18, GZ10, QHW^{+}19, SMC^{+}18]. hash [HL12, Par18]. Hashed [BCRAA13]. hashing [NMX19, TLZL16, YCJ12]. haystack [CGGFGMF19]. health [AuRMC^{+}18, CIS17, OSP^{+}19, PKP17, PKWP19, XHC^{+}10, vDBD13]. healthcare [ADF^{+}10, JAH^{+}11, KCJB12]. Hearing [AHGC19]. HEART [EHMJ19]. HEART-IS [EHMJ19]. heavy [SMK18b]. Help [Sil13, TBDM14, dABM^{+}18]. heterogeneous [WW18]. Heuristic [BCD^{+}10, LZBS13]. Heuristic-Systematic [LZBS13]. heuristics [DRP^{+}12]. Hidden [ATG11, hBCK17, VMS18, HS11, RFB17, SMK^{+}18a, SKLD14]. Hiding [SKK^{+}14, EE15, SH15]. Hierarchical [RB12, WLWG11, GMLZ11, LWS^{+}19, dABM^{+}18]. hierarchies [BCD^{+}10].

High [MAM15, ABR18]. High-fidelity [MAM15]. high-performance [ABR18]. higher [Bon19, HSL19, RE19]. highly [ZJK^{+}17]. hinder [Sil13]. HIPAA [KCJB12, LSKF12]. history [RD19]. history-based [RD19]. HMM [CMV16]. HMMPayl [ATG11]. hoc [BRA18, HZMH18, MJBKZ15, SRU18, Su10, XPS11, ZY11]. Home [BD18, HLB^{+}18, HT12a, KMW19, KvS10, SYWX19, TTCB18, TMW17, XN14, TMW17]. homology [GL18]. homomorphic [MBP19]. honey [CJW^{+}19, GSZX18]. honeyfile [GHÁGGT18]. honeyfile-based [GHÁGGT18]. honeypot [HBT^{+}19, TP18]. honeypots [AYY^{+}17]. honeyword [CM17]. hooking [WP12]. hop [TOZJ13]. Host [THA^{+}13]. Hot [CWCC12]. hotspots [ABR16]. HTML5 [GSL^{+}19]. HTTP [DF16, GTDVTSH15, ORLS14, SSK17, TH18, WJAM11]. HTTP-GET [SSK17]. HTTP/2 [TH18]. HTTPS [JCD^{+}15]. hub [AAR^{+}18]. Human [LLD^{+}19, PMB^{+}14, PCP^{+}17, ACS^{+}18, BLR17, EHMJ19, FC12, GBC^{+}18, HW19, Hat19, HL18]. human-as-a-security-sensor [HL18]. humans [GMMP18]. Hunter [SG14b]. Hybrid [NG14, YKH10, AR17, AGSR^{+}18, HSW17, STA18, SMK18b, TMP15, ZQY^{+}19]. Hypervisor [FLBK15, AP13, BR18, GSZX18]. Hypervisor-based [FLBK15, AP13]. hypervisor-level [GSZX18].

IaaS [Kar12]. ICISSP [Ano19i]. ICS [CSWX19, RM17]. ICT [Ano15j]. Identification [dSFK16, Pal10, BOP14, CLF17, GTDVTSH15, LK18, MMJ14, NQZ18, PYLL19, SG10a, YLW^{+}19]. identifier [FCSD10]. identify [HCRMBG17, SSSC18]. Identifying [ABR16, Bel18, GBCD19, LLCT13, MGYC18, SKLD14, CAM19a, dVM17]. Identity [THA^{+}13, vdWEG18, BSBG19, HPJ^{+}19, HFCR13, HWB10, NML19, VGL14, WHZ^{+}19, ZMYB17]. identity-authentication [NML19]. identity-based [HWB10]. IDS [CAL15, CAM^{+}19b, SK13]. IDSs [MFCMC^{+}18]. IEC [MM15a, AM18a]. IEEE [FZZ^{+}12]. IEEE802.11s [AG18]. if [Ano17l, Ano17m, Ano17n, Ano18i, STA18]. If-transpiler [STA18]. IFC [Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano12t, Ano12u, Ano12v, Ano13x, Ano15q, Ano15r, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k]. IFIP [Ano15i, HF17, Ano10-32, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano12r, Ano12s, Ano17v]. IFIP/SEC [Ano12r]. IFIPTM [Ano11y]. IFTTT [BD18]. II [Spa11e]. IIoT [ZGD19]. IIS [WJAM11]. illegal [HWX^{+}19]. illustrated [dVM15]. image [Ano17l, Ano17m, Ano17n, Ano18i, HM19, KMG17, TLZL16, WHZ^{+}19, YYL^{+}15]. images [NQZ18]. imbalance [LWZ^{+}17]. imminent [NTK^{+}19]. Impact [ACC^{+}13, KCJB12, SW19, WX19, HLB^{+}18, LSC14, LXDW16, MWL18, MT17, MM17, Sil19, SA16, TRB18, WLW^{+}19, WHZ^{+}18]. imperfect [WLW^{+}19]. impersonated [RW13]. impersonation [FNE19]. implantable [BKK19]. Implementation [HL18, ZWNM19, BR17, FT16, HCR10, SVGE14, SG10b, WCDY18, ZLW^{+}15, dVM15]. implementations [FK13, MBE^{+}12, PS10]. Implementing [MM15a, SVGE14, ZYPL10]. implication [SBL19, SGLC16]. implications [BGR^{+}17]. implicit [SSNS15]. importance [MCT^{+}17]. improve [CSWX19, CWCC12]. improved [CBL10, RP19, YtLpZD14, CYLL17]. Improved-MSPCA [CYLL17]. improvements [SCRA17]. Improving [AAH^{+}18, AH10, CGPP12, RMW10, WHSE15, dVM15, IHED15, PTM^{+}18, SH15]. IMS [VGL14]. In-depth [ASFI18]. in-game [MBSR18]. in-memory [SSVA19]. in-vehicle [PDB14]. inaccurate [WX19]. Inadvertent [Ven18]. incentive [Sul19]. incentive-based [Sul19]. Incident [AHR12, BMH16, FSSF15, HDT^{+}19, MP18, PWL^{+}19, RC15, TLJ14, TVPV19]. incidents [EHMJ19, HLFZ16, LMNM^{+}19]. Incompetents [MAC17]. Incorporating [ODB18, bORF^{+}15]. incremental [CH14]. independent [LLD^{+}19, NADZ18, SCR19]. indeterministic [PK17]. individual [MT17, PKP17, Sil19]. individualized [LLY17]. individuals [DWZ19, ICRB14]. indoor [DHZ19, GHDGRLM17]. industrial [AM18a, CJSM17, DPP^{+}19, EMWE19, MSRF18, OSP^{+}19, RARL19, YG18, ZZ19, AKFM18]. inertia [Sil19]. infected [GBCD19]. infectious [AAT16]. Inference [AF14, ZLC13, ASO19, AAH^{+}18, TFE10, ZH13]. Inferring [HNSD18]. infiltration [GB19]. Influence [CGLP18, DPPB17b, TAP18]. influences [BTLT16]. Information [Ano15i, BTLT16, CM19, FAE14, FT16, GSL^{+}19, HF17, KKK17, Mic12a, Mic12e, NF15, Nie17, PMB^{+}14, PCP^{+}17, SSS^{+}15, SVF16, SGK^{+}16, TLJ14, VV10, WYS18, ACS19, AF14, AH10, AFF14, AM17b, BMH16, BBC17, BCD^{+}17, BHCTR12, Bon19, BCK19, CGP17, CIS17, CHRI16, CLS17, CGW15, CJL^{+}13, DE10, DPPB17a, DPPB17b, DSP16, DBH17, DE16, EHMJ19, FE16, FPS11, Guo13, HPJ^{+}19, HKK17, Ifi12, wJKW^{+}16, KDC^{+}14, KHG17, KK16, KCJB12, KAF17, KD13, LLK16, MMCA11, MM15a, MM15b, NM19, ÖTC16, Pad12, PKP17, PKWP19, QAR^{+}17, QFG12, QHW^{+}19, RC15, RE19, RS15, RRK12, SPFB19, SMM^{+}19, SSABC16, SW19, SWS15, SSF16, SOG15, SA16, TBSR14, TRB18, TKK15, UAK16, WHZ^{+}19, WAMS14, XHC^{+}10, YYJC12, YLB^{+}17, ZTL^{+}19, dVM15, dVM17, dFGMTPL17, vDBD13, vSvN13, Ano19i]. information-flow [BCD^{+}17]. information-theoretic [CGP17]. InfoSec [Mic12a]. infrared [GB19]. Infrastructure [HW12, Mic12c, CGW15, HNSD18, IH19, MSR19, NGSV16, PMP19, TAP18]. infrastructures [CAL15, FHM^{+}10, ZH13]. inherent [PYLL19]. initial [DFC13, SWS15]. initiated [hBCK17]. initiation [TUB16]. Injecting [SWLZ11]. injection [AAL16, CWCC12, FSZ^{+}17, JC14, JLPJ18, KPS16]. injection-based [FSZ^{+}17]. inlined [BCD^{+}17]. Inlining [STA18, MRS12]. innovation [RM17]. innovations [JKC^{+}13]. Input [SPYG15, DEP13, SBK12]. insecure [AAL16]. Insecurity [LSQ15]. Insider [AH18, BJ13, CGMP18, PBR11]. Insider-threat [AH18]. insidious [ZJK^{+}17]. insight [MAC^{+}15, SS17b]. insights [HSL19]. inspection [BKP10, XXLZ18]. installers [GAA18]. instances [GZP^{+}13]. instant [ZWX^{+}19]. Insteon [TTCB18]. Institutional [HSL19]. institutions [HSL19, KCJB12]. instrument [DE10]. insurance [Fra17, PWL^{+}19]. Integrally [ST19]. Integrated [BTP^{+}18, KEE13, BCRAA13, CLL^{+}13, DM19]. integrating [AAD^{+}12]. integration [Ifi12, RLB^{+}11]. integrative [AAMDGR18, HKK17]. Integrity [VBC^{+}15, ADA^{+}18, AYT^{+}14, AAMDGR18, HBFJ19, HLB19, JK19b, LKPK18, LW19, MSMK18, XLS^{+}19, YWS17, ZKM^{+}17, dCGÜ^{+}17]. Íntegro [BLS^{+}16]. Intelligence [TG15, CAY^{+}18, KGSN17, KKM11, QAR^{+}17, TR18, WMPA19, ZW19]. Intelligent [CAM^{+}19b, CP18, WB16, CJA^{+}19, RAA^{+}14, SMS^{+}16, ZCZ^{+}19]. intensify [HL12]. Intent [FAS^{+}17, AM17b]. intention [DPP15, FE16, KMW19, PKWP19, RE19]. intentions [GBC^{+}18, SWS15]. inter [BJJ^{+}17, DLH^{+}17, HZH^{+}18, LLD^{+}19, MPPS10]. inter-app [BJJ^{+}17]. inter-component [DLH^{+}17]. inter-domain [MPPS10]. inter-frame [HZH^{+}18]. inter-keystroke [LLD^{+}19]. interaction [BBDP16, HK17, JJ12]. interactions [CC14]. interactive [AM17b, GMLZ11, JM13, LH11]. interconnection [ALC17]. interdependencies [TKG10]. interdependent [MHL^{+}17]. interests [BWMG17]. interfaces [JvOW11]. internal [KK10a, PBR11, WHZ^{+}18]. internalization [TKK15]. International [Ano15i, HF17]. Internet [AAC^{+}19, AAR^{+}18, ABHS18, Bel18, BOZL13, CZZ^{+}19, CBGN18, GWEA18, HYLY19, JvS17, JFR13, MAZAS17, NSFB15, PCC^{+}18, SP15, Tau11, VMS18]. Internet- [CBGN18]. interpretation [Hor18, ZLLL12]. interpretation-based [ZLLL12]. Interpreting [DSP16]. interpretive [LSKF12]. interval [CC14, HYZ18]. intervals [ZTS^{+}13]. intervention [AH10]. Introducing [MB17, Rua17]. Introduction [Ano15s, SG14a, SG15a]. introspection [HNSD18]. intruder [BKP10]. intruders [YLB^{+}17]. Intrusion [SHGH^{+}15, VDK18, AR17, ATG11, BK19, BMMP18, BRA18, CGGFGMF19, CY16, CAM^{+}18, CH14, DC11, GTDVTSH15, GOB11, GTGB16, GWWW19, GZ10, GZP^{+}13, HDK18, HS13, HSPX14, HXS^{+}16, KK10a, KK17, KKM11, LMLK16, LLHT13, MLK14, PHP19, PAP^{+}17, RWS^{+}19, SD17, SS18a, SSCHL14, SRU18, SSTG12, SK10, USB17, VSO^{+}18, XPS11, YW11, ZY11, ZLK10, dABM^{+}18]. intrusions [PDM19]. invariant [ZW14]. invariants [SKG17]. inventions [Ven18]. Investigating [FAE14, LZBS13, GGV18]. Investigation [DPPB17b, SG15a, CCM^{+}15, HA15, KDC^{+}14, KJKM11, RB11, RB12, SFK15]. investment [DE16]. investments [WYS18]. Invi [PNKK17]. Invi-server [PNKK17]. Invisible [GMMP18, PNKK17]. Invitation [Ano10-32]. iOS [DC16]. IoT [APMCR13, CP18, GWKFC19, GMMP18, HHBS18, LSQ15, TTJ19]. IoT-based [GWKFC19]. IoTChecker [MAZAS17]. IP [CDE^{+}11, HDMKS10, MA17, NT18, SMK^{+}18a, SSK16, YW11]. IPv6 [LK18]. IR [GB19]. iris [DDSP17, RB14, ZZZY13]. iris-codes [RB14]. irrespective [DLS14]. Isabelle [Kam19]. ISBPF [NF15]. ISCP [ASFI18]. ISO [MM15a]. ISO/IEC [MM15a]. Isolated [MJRK18, KKJ^{+}16]. isolation [HDMKS10]. ISRA [SSABC16]. Issue [Ano12r, Ano19i, BJW19, Ano13-30, Ano15j, Ano17v, HLL17, LM10, YWND14]. Issues [FHL14, RJJ15, SSK19, Hol12, LK18, SVGE14]. item [SYCS11]. items [CC14, LGS14]. itemset [CSXL15, LMLL19].

Jail [RP19]. Jail-Phish [RP19]. jammer [IHED15]. January [Ano11z, Ano17p, Ano18l, Ano19m]. Java [BL15, CMM^{+}10, MLM17, PS10, WA19]. JavaScript [STA18]. Jennifer [SG15a]. JFCGuard [MJW^{+}18]. Johnny [WA19]. judgments [CGW15, MB18]. judicious [RFW^{+}18]. juice [MJW^{+}18]. July [Ano12-27, Ano13-28, Ano14u, Ano15x, Ano16u, Ano17q, Ano18m, Ano19n]. Jumper [GB19]. June [Ano10-35, Ano11-27, Ano12z, Ano14x, Ano15z, Ano16v, Ano17s, Ano18k, Ano19o]. justifications [VVS12].

k-modes [Ngu18]. KakaoTalk [KPKS15]. KAMAS [WRTA17]. Kano [YCH18]. keeping [JL10]. Kerberos [PMLK^{+}11]. Kernel [MJRK18, HHAW19, JvOW11, LKPK18, LLL^{+}18, LJS19, WP12]. kernels [JvOW11]. key [ALRC12, BD16, BCD^{+}10, BEB^{+}18, GMLZ11, HWB10, MFBC11, MK11, MVK16, RR16, ZHL^{+}11, ZZC17, ZCZ^{+}19]. key-disclosure [ZZC17]. keyloggers [DKG13]. keys [AAJ12, dSFK16]. keystroke [FNE19, GEAHR11, GDR15, ICRB14, LLD^{+}19, MCRB19, PGdCL16, SSY12, UB12]. keystroke-dynamics [SSY12]. keyword [DLLS19, JMGW17]. keyword-based [DLLS19]. Kid [NRM19]. kill [BS17]. kill-chain [BS17]. Kingdom [Hor16]. knowledge [ABS14, DWZ19, FAE14, MDHM18, MCY^{+}18, RAMR17, RM17, TBSR14, WRTA17, ZW19]. knowledge-assisted [WRTA17]. knowledgeable [XHC^{+}10]. Kumar [TG15].

LAAP [Gop19]. labels [YYJC12]. landscape [Cho11, CW19, SSK17]. language [Ksi12]. large [BLS^{+}16, CTH10, JJL^{+}19, SJCM17, SDM^{+}18, VSB^{+}19, ZWX^{+}19]. large-scale [JJL^{+}19, SJCM17, SDM^{+}18, ZWX^{+}19]. laser [UAK16]. latency [BMMP18, CPGA13, JK19b]. Latent [RW13]. later [HBBM16]. lateral [HBT^{+}19]. Launching [AMMV13]. layer [IHED15, SSK17, TKG10, YLB^{+}17]. lazy [NNNQ18]. lazy-binding [NNNQ18]. LDA [EM19]. LDoS [ZjHtMhBs12]. leadership [BRM^{+}18, FE16]. Leakage [YHL^{+}15, ABS14, BHCTR12, GSL^{+}19, LLYD15, MOV14]. Leakage-resilient [YHL^{+}15]. leakages [UAK16]. leaks [DBH17]. learned [OM19]. learner [ACS19]. learning [COS19, CXF^{+}18, CAM^{+}19b, HWX^{+}19, HCRMBG17, IRC^{+}17, KMG17, LWZ^{+}17, MBSR18, NNG^{+}18, NNNQ18, NQZ18, PKP17, RFB17, UAB19, WYS18, ZW19]. least [Bon19]. Legacy [OBRK15]. Legally [BB11]. legend [Ano17l, Ano17m, Ano17n, Ano18i]. LEGO [KJP^{+}19]. lens [MB18]. less [ATL^{+}17]. lessons [OM19]. level [CKJ^{+}17, GSZX18, KGSN17, Kar12, LSC14, QHW^{+}19, SE11, SOG15, WHSE15]. Leveraging [BLS^{+}16, NNG^{+}18, PC12, DSNH17, DHZ19]. Lévy [Zin15]. lexical [LGS14, ZDW19]. liabilities [HZG19, HZG20]. library [TXM^{+}19]. life [JLK16, SYX^{+}16]. lifecycle [MM15a]. Light [KP18, NWD^{+}12]. Light-weight [KP18]. Lightweight [Gop19, GZ18, GZ20, BR17, DA18, GH15, JK19b, KCB13, MJRK18, VMA16, XLS^{+}19, ZLW^{+}15, ZHB^{+}13]. likelihoods [MHL^{+}17]. limitations [CM17]. limited [LZCX18, ZZC17]. limits [BHC15]. line [CWGD15, NDMM13]. line-noise [NDMM13]. lineage [UCCJ18]. linear [LML^{+}17]. linearly [MBP19]. lines [UAK16]. Linings [Ano12r]. link [HL18]. Linking [ALF^{+}18]. Linux [HHAW19, SCB^{+}13]. list [Ano18r]. lists [Lop16]. literature [Bon19, FB14, GGV18, IAA^{+}19, IH19, SA16, TLJ14, XL19]. live [TBDM14]. Living [PCC^{+}18]. LLE [TLZL16]. load [MKP12, SMK18b, XdZMsC18]. load-balancing [MKP12]. local [CLLK10, YYL^{+}15]. Localising [PSS19]. localization [DEP13, DHZ19, GHDGRLM17]. Location [JLPJ18, MZZ^{+}18, MOV14, BHC18, CVS18, JWZX19, SVA10]. location-based [BHC18]. locations [CC14]. lock [KN14]. Locker [GHÁGGT18]. locus [HPJ^{+}19]. log [BS17, HDMKS10, LWS^{+}18]. log-based [HDMKS10]. logging [PMP19]. Logic [RB11, AAH^{+}18, LXDW16]. Logic-based [RB11]. logical [SKG17]. logo [BWR12, CCST15]. long [TBDM14, VBC^{+}15]. long-term [VBC^{+}15]. longitudinal [AYY^{+}17]. look [DDSP17, HT17, SM19a, UB12]. look-up [DDSP17]. loss [HX17]. lossless [Kül12]. Low [LW12c, SG15b, JK19b, NAS19]. lower [CKA^{+}17]. LPSE [GZ18, GZ20]. LR [KK17]. LSB [LHM13]. LSTM [HS18]. LTE [hBCK17].

M2M [ABHS18]. M2M-REP [ABHS18]. MAC [HDT^{+}19, JWB16, SMK18b]. Machine [MBSR18, ABC^{+}19, BFTJ18, COS19, CXF^{+}18, CAM^{+}19b, HCRMBG17, KTG^{+}18, NNG^{+}18, SRU18, UAB19, VDK18]. machine-learning [NNG^{+}18]. machines [ABHS18, RG19]. Macroeconomic [MAC17]. mail [BPA^{+}11, FGOPL10, PW13, Tau11]. main [HHAW19]. Maintaining [TFE10]. make [BWOD13]. making [BWMG17, DE16, MWAG16, PNKK17, RMW10, WYS18]. mal [GFN11]. mal-packets [GFN11]. maladaptive [MVD19]. Malboard [FNE19]. MalDAE [HXW^{+}19]. Malicious [KKK18, MG12, AYY^{+}17, BM19, CAY^{+}18, GPAD^{+}14, JLK16, LHI^{+}19, LLCT13, NAS19, NCGE15, OSNZ19, SWZL14, SSSC18, SPD^{+}15, XXLZ18, YXD^{+}11]. Malware [BFTJ18, DDYZ14, DRP^{+}12, FAS^{+}17, UCCJ18, MG12, NVAJ19, NQZ18, AAT16, AHTS15, AQR^{+}17, AGSR^{+}18, BGL18, CMV16, COS19, CXF^{+}18, DPP15, DYT^{+}13, DXCL18, DWZ19, EMBH14, ESY^{+}15, FLBK15, GMMP18, HXW^{+}19, HB16, HYLY19, IRC^{+}17, wJKW^{+}16, LLL15, LPL18, MPM15, MAC^{+}15, MCT^{+}17, MCY^{+}18, MMF19, MAM15, NNG^{+}18, NNNQ18, NLC^{+}19, PSK^{+}17, PRS13, RBJ18, Rru16, SS11, STCM^{+}14, SPCCMMS18, SPR18, TCNFW16, UAB19, WRTA17, Yad19, YSN^{+}19, ZYRZ16, ZQY^{+}19, ZTC19]. malware-proof [HYLY19]. Man [JCD^{+}15]. Man-in-the-browser-cache [JCD^{+}15]. manage [TBDM14]. Management [GACCB^{+}13, Mic12e, NDdL15, SYK14, ALRC12, BJ13, BMH16, BTLT16, BCRAA13, BOZL13, BSBG19, Bon19, DVC18, DAR14, GTGB16, GMGTdFR14a, HB18, HLL17, KKP10, MLM17, MB17, MPPS10, MKG16, PKK^{+}14, RV13, RRK12, SE11, SK18, THA^{+}13, TLJ14, TMP15, VV10, VHBJ19, WAMS14]. manager [JMGW17, ZY14]. manager-centric [JMGW17]. Managing [BCD^{+}10, DLDTGMMP15, QFG12, SSVA19, NM19]. MANETs [LMM^{+}12, TP18]. manual [SCLD19, PTM^{+}18]. many [MSVA16]. many-valued [MSVA16]. Mapping [TG12, DDSP17]. Maps [LMNM^{+}19, BSBG19, BFTJ18]. March [Ano10-34, Ano12x, Ano15v, Ano16q, Ano17t, Ano18p, Ano19k]. market [Fra17, KCJB12, PCP^{+}18, SA16]. marketplaces [MSMK18]. markets [HSR19]. Markov [ATG11, hBCK17, GCC18, HS11, PWWY14, RFB17, lTLqZqC19, VMS18]. mashup [PAJAP15]. MASKER [KNX18]. Masking [KNX18]. Masquerade [HS11, TC11]. mass [PCP^{+}18]. Masses [PG19]. massive [YLW^{+}19]. master [CDKR18]. matching [DA18, GZ10, LHM13, SM10]. Math [HCR10]. mathematical [MSR19, RMG11]. matrix [DLOP17]. maximal [CSX^{+}15, TFE10]. Maximizing [LMLL19]. May [Ano12y, Ano13y, Ano14w, Ano15t, Ano16w, Ano17r, Ano18n, Ano19l]. MDC [BPG^{+}19]. MDC-Checker [BPG^{+}19]. me [ZTL^{+}19, CMM^{+}10]. meaningful [GOGG19]. means [CGPP12, MMPC18, TFPC10]. Measure [GPGV16, JGLE16]. measurement [QLZ19]. measurements [Sav13]. measures [RR14, SS17a]. Measuring [CKA^{+}17, MM15b, ABR16, HEK^{+}18, Rua17]. mechanical [dSFK16]. mechanics [LPL18]. Mechanism [DVC18, Pal10, ADA^{+}18, AuRMC^{+}18, CVS18, GSL^{+}19, GMMP18, HMFE10, JXK18, KK10a, KKJ^{+}16, LSA^{+}19, MLK14, MCRB19, NXS10, SYCS11, YSN^{+}19]. mechanisms [AAL16, DAR14, GJ18, JHJ^{+}16, MBAE^{+}11, Ngu18]. Media [SG15a, Gar13, UAK16, Ven18, WS17, vdWEG18]. Median [EM19]. medical [BKK19]. meets [CGMP18, LLD^{+}19]. membership [Ano11a, Ano11b]. memorable [GZG19, GZG20]. memory [BBD^{+}19, CSWX19, HHAW19, JAN19, LWS^{+}19, MLM17, NLC^{+}19, OP11, SSVA19]. Merkle [HZG20, HZG19]. mesh [HSPX14, HXS^{+}16, LNNH13, VDK18]. message [BKP10, KDC^{+}14]. message-person [KDC^{+}14]. messages [HS18, ROL13]. messaging [BBD^{+}19, CCC^{+}12, FKSK19]. Meta [DLDTGMMP15, HBFJ19]. meta-model [HBFJ19]. Meta-Policies [DLDTGMMP15]. metamorphic [AHTS15, EMBH14]. Metasploit [KG13]. metering [Sul19, TAP18]. meters [GZ18, GZ20, XXLZ18]. method [ATAAd18, CGP17, Cho13a, Coh10, DXCL18, GWKFC19, GCD19, HYZ18, KPD15, LH11, LNK16, NZGH16, NLC^{+}19, WW18, XSD19, ZZ19].

Methodology [LK18, AAD^{+}12, CDV12, HSW17, LMG18, OP11, PAJAP15, SFK15, TKG10]. methods [AMS18, CBB^{+}16, Cho13b, GGSZ14, HZMH18, IAA^{+}19, IRC^{+}17, RNdP17, SAL12, SS13, TAP18, YXD^{+}11]. metric [ZH13]. metrics [AG14, BM17b, HEK^{+}18, NFGL16, Sav13]. mHealth [HAHZ^{+}18]. Mice [SP15]. microaggregation [ANAT15, MSV12]. microblog [YLW^{+}19]. microdata [MSV12]. Microsoft [JKC^{+}13, LW12e]. middleware [JGM^{+}13]. migrating [PPKK19]. migration [Len12]. migration-based [Len12]. mimicry [TC11]. Mind [SNG^{+}17]. mindset [PBR11]. MinHash [NLC^{+}19]. Minimal [HB16]. minimization [SA19]. Mining [GOGG19, HSR19, MSVA16, BSL19, CSXL15, CSX^{+}15, CDOV10, CDV12, DYT^{+}13, DWF^{+}19, HB16, LMLL19, LLG18, MSRF18, RNdP17, WLW^{+}19, WLF^{+}19, YYJC12, ZMYB17, dABM^{+}18]. minutiae [BCRAA13]. misbehaving [KUA^{+}18]. misconfiguration [GACCB^{+}13]. misinformation [YLW^{+}19]. missing [NTK^{+}19]. mission [MM17]. Misuse [FRRP17, GZS19]. Misuse-based [FRRP17]. mitigate [BJ13, PC12]. Mitigating [SNG^{+}17, GSA18, NB19]. mitigation [FW10, LAVK19, SBZD17, ZGD19, Zin15, bORF^{+}15]. mix [FHF15, PK17]. mix-based [FHF15]. mixed [AAdM19]. Mixing [RB14]. Mixture [AH18]. ML [Ksi12]. mnemonic [YGZG19]. Mobile [GJ18, STCM^{+}14, ABR18, BPA^{+}11, BHC18, BFS^{+}18, BRA18, CCC^{+}12, CCC^{+}18, CGLP18, CRS13, FCPA16, FHM^{+}10, GFKS15, GMMP18, IAA^{+}19, JWZX19, KKK17, KS19, LH11, MJBKZ15, MMF19, NF15, NRM19, NAMX14, RRK^{+}17, SCRA17, SSNS15, SPCCMMS18, SCR19, Su10, SMA^{+}18, TZTC16, TMW17, WWY^{+}18, WHSE15, XPS11, XNP16, XLS^{+}19, YWZM18, YCH18, YKH10, ZY11]. mobile-cloud [ABR18]. Mobility [CGMP18, KW14, THA^{+}13, TG14]. Mobility-as-a-Service [CGMP18]. modal [BOP14, GBC19, SCFB15]. Model [AKP14, AP14, hBCK17, LZBS13, NSFB15, VMS18, AWDB19, ASFI18, AAH^{+}18, ABP16, ABR18, ARG^{+}17, BKP10, BPA^{+}11, BLR17, BSBG19, CHK^{+}19, CLL^{+}13, CHRI16, DC16, DMC19, DM19, DWF^{+}19, DE16, GTGB16, GMGTdFR14b, GM12, GKB14, HKK17, HBFJ19, JSVA14, JJ12, JJ14, Kar12, KKK17, KEE13, KFGL19, KS19, KKM19, KKS13, LLY17, LGZ^{+}19, LC15, LZS^{+}16, LLL^{+}18, MSR19, MB18, MSA11, MKG16, NDSA17, NZGH16, PWWY14, RFB17, RR16, SVF16, SERF12, SA19, ST19, SK18, SSVA13, SC19, SOG15, SEJ10, TBSR14, TZZG13, TCNFW16, UPF17, WZW^{+}10, WL18, WAMS14, WHZ^{+}18, Wor10, WCDY18, XHC^{+}10, YCH18, ZZ19]. Model-based [AKP14, NSFB15]. Modeling [AAD^{+}18, MMCA11, MA17, MBAE^{+}11, WP10, WJAM11, ZMYB17, AW18, CLS17, LLD^{+}19, RNR15, VSB^{+}19, XL19]. Modelling [MWAG16, RFGL18, SPCCMMS18, BS19, DJP17, FHF15, Ksi12]. Models [AH18, ATG11, HW18, ANMN19, ABR15, HS11, JLK16, LML^{+}17, MCG19, PMDM10b, RMG11, SCLD19, SS18a, SVA10, UPF17, ZCZ^{+}19]. modern [GM12, PDB14]. modes [Ngu18]. modification [JvOW11, SH15]. modified [EE15, CM17]. modified-UI [CM17]. modifying [MKP11]. Modular [JM13, ADWM18]. module [KGSN17]. monitor [BCD^{+}17, STA18]. Monitoring [vDBD13, CPVGTMF16, CMM^{+}10, HPF^{+}16, TS13, dVM15]. monitors [CKT11, LKPK18, MRS12]. Montgomery [SLN^{+}15]. mood [DLS14]. most [Pal10]. motion [SPYG15]. motion-sensor [SPYG15]. motivated [AWDB19]. Motivation [DPP15, HSL19, Ifi12, MWL18, ySTJA^{+}16, Ver18]. motives [Wor10]. mounted [Rru16]. Mouse [MPM15, SCGM14]. move [ZS18]. movement [HBT^{+}19, PSS19]. moving [HEK^{+}18, lTLqZqC19, ZS18, qZLxCjY17]. MP3 [YWYZ12]. MSABMS [ZjHtMhBs12]. MSABMS-based [ZjHtMhBs12]. MSPCA [CYLL17]. MTRA [TTJ19]. Multi [TTJ19, WW18, BTM14, BOP14, BOZL13, CGPP11, DRN16, DLOP17, GBC19, HZH^{+}18, JWB16, LSC14, LAVK19, LMLK16, LML^{+}16, MPCK18, MLKC17, ML14, NDP18, RAA15, RB14, SE11, SCFB15, SYCS11, TOZJ13, TKG10, FYL^{+}17]. multi-authority [JWB16, LML^{+}16]. multi-biometrics [RB14]. multi-channel [HZH^{+}18]. multi-currency [ML14]. multi-factor [DRN16]. multi-hop [TOZJ13]. multi-item [SYCS11]. multi-layer [TKG10]. multi-level [LSC14, SE11]. multi-modal [BOP14, GBC19, SCFB15]. multi-party [DLOP17]. multi-RANs [FYL^{+}17]. multi-service [BOZL13]. Multi-source [WW18]. multi-stage [LAVK19]. multi-stakeholder [BTM14]. multi-step [NDP18, RAA15]. multi-tenant [MPCK18]. Multi-Tier [TTJ19]. multi-touch [MLKC17]. multi-view [LMLK16]. multifaceted [CHRI16]. multifactor [KS19]. multifold [WL18]. multilayer [LMM^{+}12]. multilevel [SERF12, TZMG15]. multimedia [GJJ18, KMM13, RM17]. multimodal [SS16]. multipath [OPHC16]. Multiple [XYY^{+}12, BPG^{+}19, GJJ18, NDMM13, RVEK10, SLH14, VDK18, WL18, WYS18, ZGZ^{+}13]. Multiple-File [XYY^{+}12]. multiplication [DLOP17, SLN^{+}15]. multisourcing [NM19]. multitenant [RMVC^{+}12]. Multivariate [CGGFGMF19, CPVGTMF16, GCD19]. mutation [qZLxCjY17]. Mutual [JJA13, Cho13b]. Mutual-friend [JJA13]. MVPSys [LMLK16]. my [ROW19].

name [GWEA18]. names [CAY^{+}18, XSD19]. narrative [BCK19]. Nash [GCC18]. NAT64 [LK18]. national [FAE14, RM17]. native [AQR^{+}17]. Navigating [SG14a]. Near [DM17, ABP16]. Near-optimal [DM17]. Nearest [EM19, XMY^{+}17]. Necessity [MMKV15]. need [KBM16]. needle [CGGFGMF19]. needs [ANMN19, HBBM16]. negatives [LLHT13]. negotiation [MKP12]. negotiations [KFGL19, RFGL18]. neighborhood [XXLZ18]. Neighbors [EM19]. neighbour [XMY^{+}17]. NeMHIP [WJ15, THA^{+}13]. net [FXC18, FCPA16]. Network [AF14, CAM^{+}19b, DPPB17b, HDK18, PF10, PKK^{+}14, SFK15, SG14a, TDJB18, qZLxCjY17, ABR16, AHGC19, BK19, BPG^{+}19, hBCK17, CGPP11, CGPP12, CGP17, CPVGTMF16, CFH^{+}11, CMPR15, CY16, CYLL17, CAM^{+}18, CH14, CLF17, CGW15, DPPB17a, DT19, DC11, EGM10, EM19, FKSK19, FSSF15, FZZ^{+}12, FRRP17, GL18, GOB11, GGBT16, GCD19, GBCD19, GZ10, HW19, Hol12, JKC^{+}13, JJA13, KLN15, KK17, LMLK16, MBE^{+}12, MFCMC^{+}18, MPCK18, MSR19, MA17, Med18, MLK14, MKG16, MCG19, PDM19, PSS19, RSLH19, RWS^{+}19, ROL13, SD17, SCLD19, SS11, SS18a, STCM^{+}14, SLG19, SLD12, TFPC10, THA^{+}13, VMS18, VDK18, WLW^{+}19, XN14, XHZ11, XSD19, ZYRZ16, ZYPL10]. network-based [JKC^{+}13, RWS^{+}19]. network-initiated [hBCK17]. networked [SM19a]. Networking [GWKFC19]. Networks [AX15, DFC13, GMGTdFR14b, RD19, RB11, RSLH19, VG14, AG18, AWZ^{+}17, AKAuR19, ASO19, ADAMH17, ACS^{+}18, hBCK17, BM17b, BRA18, CVS18, CTH10, CAM^{+}18, DFCC16, DLB^{+}19, EE15, FYL^{+}17, FFGL18, GFN11, HR13, HM19, HS13, HSPX14, HXS^{+}16, HHK^{+}19, HT17, IHED15, JJ14, KMG17, KUA^{+}18, Kha12, LAVK19, LNNH13, LLYD15, MPCK18, MK19, MJBKZ15, MVK16, MPORBR12, NXS10, OPHC16, OAO17, PZ15, PNKK17, RR17, RBJ18, SJCM17, SRU18, SSAK12, SC19, Su10, SMK18b, TOZJ13, TTJ19, UB12, WW18, XPS11, XN14, XNP16, ZY11, ZNAHK11, ZHZ^{+}15]. Neural [CAM^{+}19b, YSN^{+}19, CAM^{+}18, EE15, HW19, HM19, KMG17, Kha12, MCG19, RBJ18, TFPC10, UB12, XSD19]. Neurokey [BD16]. neutralization [BWOD13]. Next [Mic12b, CMM^{+}10]. NFD [HX17]. NIC [Rru16]. nifty [PAP^{+}17]. NN [ZZC17]. no [LHS^{+}18]. nodes [AAT16, KUA^{+}18]. noise [NDMM13, QLZ19, ZYWZ14, HM19]. Noise-GAN [HM19]. noisy [YLW^{+}19]. Non [HX17, BBC17, GMLZ11, JM13, PW13, Sil19, VBC^{+}15]. non-compliance [BBC17]. non-compliant [Sil19]. non-interactive [GMLZ11, JM13]. non-parallelizable [JM13]. non-repudiable [PW13]. non-repudiation [VBC^{+}15]. Non-technical [HX17]. Noncespaces [VC12]. nonlinear [PF10]. normalization [PPdCL17]. Norman [LSKF12]. norms [HOVB17]. note [FCSD10]. Nothing [RR14]. Novel [MTT^{+}16, TUB16, BPG^{+}19, BSBG19, BS17, CIS17, CAM^{+}18, CFSF18, CLF17, EHMJ19, FNE19, GWWW19, HR13, IRC^{+}17, LLY17, PAJAP15, SWS16, SCR19, ZZZY13, ZLW^{+}15]. November [Ano10-38, Ano11-29, Ano12-29, Ano13-29, Ano15w, Ano16r, Ano17o, Ano18j, Ano19j]. NTFS [Cho13a]. number [ZZZY13]. numbers [KPKS15]. Nutan [TG15].

OAuth [SMA^{+}18]. obfuscation [BSTT16, HT12b, KTG^{+}18, MAC^{+}15, MKP11, SKK^{+}14, Zhu18]. object [DBH17, MCN10, OAO17]. object-oriented [DBH17]. objective [DYT^{+}13]. objective-oriented [DYT^{+}13]. objectives [FK13]. objects [Bel18, MCT^{+}17, PYLL19]. oblivious [KTCS19]. Obscuring [VGL14]. observable [HBT^{+}19]. observation [AYY^{+}17]. observational [BRM^{+}18]. October [Ano11-28, Ano10-37, Ano14v]. off [CWGD15]. off-line [CWGD15]. Offline [FKSK19, ZZC17]. offs [MMPC18]. old [LSSK16]. On-demand [KKJ^{+}16]. On-the-fly [MRS12]. One [Par18, LHS^{+}18, SPK17]. One-time [Par18, SPK17]. Onion [CPGA13, ZLC13]. Online [LWS^{+}19, PKK^{+}14, AAA19, ACS^{+}18, CCM^{+}15, CKA^{+}17, ELG^{+}17, FFGL18, GMdF19, HWX^{+}19, JJ14, LLYD15, MB17, MSMK18, RFB17, RR14, SSAK12, ySTJA^{+}16, Wor10]. only [Ril13]. ontologies [NNG^{+}18]. Ontology [DWZ19, RAA^{+}14, SOG15]. Ontology-based [DWZ19]. opcode [DDYZ14, ZQY^{+}19]. opcodes [COS19]. Open [CB12, TG15, AM18a, CC14, GJ18, MGYC18, NMX19, Sil13]. open-source [MGYC18]. OpenID [NB19, SHB12]. OpenPLC [AM18a]. OpenSGX [CKJ^{+}17]. OpenSSL [WN15]. operating [CZH^{+}14, RHL15, WHSE15, ZW14]. operation [ZLL^{+}18]. Operational [MPORBR12, AFF14, LAR13, Sul19]. operations [KKJ^{+}16]. operative [HFCR13]. operators [MWAG16]. opinion [ANS13]. opp [DLH^{+}17]. OppNets [ADA^{+}18]. Opportunistic [RD19, ADAMH17]. Optimal [GJJ18, SLN^{+}15, lTLqZqC19, YYJC12, CAM^{+}18, DM17, PKK^{+}14, SMK18b, ZYWZ14]. optimality [AAdM19, HS13]. optimism [RRK12]. optimization [EE15, MZZ^{+}18, SD17, Zin15]. optimized [CSS^{+}19]. Optimizing [HBT^{+}19, AQR^{+}17, ZNAHK11]. Optiwords [GZG19, GZG20]. Order [AKAuR19]. organisation [DPPB17b, KK16]. organisational [AHR12, AS13]. organising [BFTJ18]. Organizational [KD13, ABS14, DSP16, Sil19]. organizations [CLL^{+}13, FAE14, SSS^{+}15, SVF16, Sil13]. Organizing [LMNM^{+}19]. Oriented [HW12, AGSR^{+}18, BR17, DYT^{+}13, DBH17, KKKP18, LCZ17, NML19, RRK^{+}17, SMP13, SK18, WLZ^{+}19]. OSINT [KGSN17]. OSN [PKK^{+}14, BBDP16, KGSN17]. OSNs [BLS^{+}16]. OSSF [MB17]. other [KCJB12]. OTR [FKSK19]. outbreak [CLLK10]. outsourced [AIRC13, LCZ17]. Outsourcing [OSNZ19, BTM14, DFJ^{+}13, JGM^{+}13, LXH^{+}18, XMY^{+}17]. overcome [BBC17]. overcoming [CM17]. overlay [MPCK18]. OWL [KSG17]. own [AS13]. owned [GMGTdFR14a]. owner [ZZC17]. ownership [MWL18, VHBJ19].

P2P [CMPR15, IGS17]. PaaS [RMVC^{+}12]. packed [UPSGF^{+}14]. Packet [ADA^{+}18, YW11, ADAMH17, TZMG15]. packets [GFN11]. packing [BEGW^{+}19]. Page [Ano17l, Ano17m, Ano17n, Ano18i]. Page/Cover [Ano17l, Ano17m, Ano17n, Ano18i]. Pages [Ano10-33, Ano10-35, Ano10-34, Ano10-38, Ano10-37, Ano10-36, Ano11z, Ano11-27, Ano11-29, Ano11-28, Ano12w, Ano12z, Ano12-27, Ano12x, Ano12y, Ano12-29, Ano12-28, Ano13-27, Ano13-28, Ano13y, Ano13-29, Ano13z, Ano14t, Ano14x, Ano14u, Ano14w, Ano14v, Ano14y, Ano15y, Ano15z, Ano15x, Ano15v, Ano15t, Ano15w, Ano15u, Ano16t, Ano16s, Ano16v, Ano16u, Ano16q, Ano16w, Ano16r, Ano16x, Ano17p, Ano17s, Ano17q, Ano17t, Ano17r, Ano17o, Ano17u, Ano18q, Ano18l, Ano18k, Ano18m, Ano18p, Ano18n, Ano18j, Ano18o, Ano19p, Ano19m, Ano19o, Ano19n, Ano19k, Ano19l, Ano19q, GZ20, GZG20]. pairings [HWB10]. palmprint [QLZ19]. Panda [TG15]. PANDDE [FB19]. Panning [ELG^{+}17]. Papers [Ano11-30, Ano10a, Ano10b, Ano10c, Ano11m, Ano12b, Ano13-30]. paradigm [BD16, Gop19]. paradox [GGV18, Kok17]. Parallel [BTK15, HW19, HT12b, JCD14]. parallelizable [JM13]. parameterized [SYGR11]. parameters [CAM^{+}18, SS11]. parent [FCSD10]. Part [Spa10c, Spa11e]. partially [HBT^{+}19]. participants [BWMG17, HWX^{+}19]. participation [AH10]. particle [EE15]. parties [BGLR17]. party [AF14, DLOP17, HWB10, PW13, SSAK12, SBS^{+}18, TXM^{+}19, VHBJ19]. passing [Ano14z]. passive [ZYPL10]. Password [CJW^{+}19, AOR18, GZ18, GZG19, GZ20, GZG20, LHS^{+}18, Lop16, MCRB19, NMX19, Par18, SYX^{+}16, YHL^{+}15, YGZG19, YWL^{+}17, ZY14]. password-strength [GZ18, GZ20]. passwords [GZG19, ICRB14, MLKC17, ROW19, SYX^{+}16, ZDW19]. patches [AW18]. patent [Ven18]. path [HDMKS10]. paths [GSA18, HMFE10]. Patient [JAH^{+}11, PKP17, PKWP19]. Patient-centric [JAH^{+}11]. pattern [DA18, GZ10, HB16, Kal19, KGSN17, KN14]. patterns [BRH15, GOGG19, HW18, HBF13, LHM13, NML19, SMC^{+}18, SPK17, UFF12, UFF15, WLF^{+}19, YYL^{+}15]. payment [Cho13b, ZWX^{+}19]. PCA [CPVGTMF16]. PCA-based [CPVGTMF16]. PDF [NCGE15]. Peek [HNSD18]. Peek-a-boo [HNSD18]. Peer [SFK15, DHZ19, JL10]. peer-assisted [DHZ19]. Peer-to-Peer [SFK15, JL10]. Penetration [CB12, KG13, LW12b, SG15b, Hat19, KBM16]. penetration-testing [Hat19]. people [FC12, ROW19]. per-context [LKPK18]. Perceived [SGK^{+}16, HOVB17, KMW19]. perceiver [DLS14]. perception [HSR19]. perceptions [GMMJ11, MT17, MMPC18]. perceptual [KK16]. Performance [Hol12, Kha12, SCGM14, ABR18, CGPP12, CSWX19, CWCC12, DCC^{+}18, MLK14, SK13, TZZG13, UB12, Zhu18]. performances [CBL10]. permissible [ZLW19]. Permission [FHL14, GFKS15, AAC^{+}19, DLH^{+}17, LLG18]. permissioned [PMP19]. permissions [NNG^{+}18]. perpetrators [KSW16]. persistent [AWDB19, FSSF15, GL18, LCMF18, XN14]. Persisting [JCD^{+}15]. person [KDC^{+}14]. Persona [KAF17]. Persona-centred [KAF17]. Personal [GMGTdFR14a, Pal10, CLS17, ÖTC16, WS17, WHZ^{+}19]. Personality [SWS15, MPM15]. personalized [BAH18]. perspective [ABS14, BTM14, Bay12b, CGMP18, CGR18, HKK17, SM19b, ySTJA^{+}16, VV10, Wor10]. Perspectives [RS15, ALF^{+}18, AAMDGR18, EAML15, YLB^{+}17]. Perturbation [CSS^{+}19]. Pervasive [FHM^{+}10]. phase [WLF^{+}19]. phenomenon [Kok17]. phish [VMA16, RP19]. Phishing [AZ17, KSW16, RW13, VRRK17, CCST15, DLLS19, GJ18, GK14, HA14, LSA^{+}19, LZBS13, MB18, PMP^{+}15, RP19, ZJK^{+}17]. phone [KPKS15, NF15, NRM19]. phones [DB13]. photo [WWY^{+}18, ZLW19]. PHP [LHI^{+}19, WN15]. PHP-based [LHI^{+}19]. Physical [BEB^{+}18, Mic12c, NKL^{+}18, ALC17, Ano13-30, AM17a, BBD^{+}19, BJW19, CHK^{+}19, CM19, HLB^{+}18, HZMH18, IHED15, LXDW16, YWND14, YLB^{+}17, ZH13]. pi [DJP17]. PIN [KH12, LNK16, NSBM17, NKL^{+}18, SS13]. PIN-entry [LNK16]. PIndroid [IRC^{+}17]. PINs [LLD^{+}19]. pioneer [Ano14z]. pirate [ZZ12]. pitfall [ZHL^{+}11]. Pitfalls [HCR10]. pixel [LHM13, SH15]. PKCS15 [BCRAA13]. PKI [DFC13, KKM19, MPPS10, YI17]. PKI-based [MPPS10]. places [Bon19]. Plaintext [MBP19]. planned [Ifi12]. platform [BRH15, CKJ^{+}17, CMM^{+}10, HT12a, HYLY19, HAHZ^{+}18, LLG18]. platforms [GBC19, HLFZ16, MKG13, NGSV16, NMX19, RMVC^{+}12, SLB^{+}13, SLS^{+}18, vdWEG18]. POBA [CSS^{+}19]. POBA-GA [CSS^{+}19]. pod [KK10b]. points [MZZ^{+}18]. poisoning [CXF^{+}18, IGS17, JCD^{+}15]. Policies [DLDTGMMP15, AAD^{+}12, AIRC13, BBC17, BSL19, CDS^{+}19, HPF^{+}16, Kar12, KHG17, KSG17, KH12, RE19, RLB^{+}11, RMW10, SSVA19, SBE17, TKK15, TdHRZ17]. Policy [NDdL15, AHP18, ADF^{+}10, BWOD13, CLL^{+}13, FT16, GHL^{+}17, GZG19, GZG20, HKK17, Ifi12, MBE^{+}12, MCL^{+}13, PAS^{+}18, SVF16, SW19, TVPV19]. Polite [VVS12]. poll [JCPACRV12]. poll-site [JCPACRV12]. polls [ANS13]. Polymorphic [VG14]. portals [GAA18]. positioning [MZZ^{+}18]. positives [LLHT13, SK10]. possible [FCPA16]. Post [PWL^{+}19, MFBC11, SHGH^{+}15, SK13, YGX^{+}19]. post-correlation [SHGH^{+}15]. Post-incident [PWL^{+}19]. post-processing [MFBC11, SK13, YGX^{+}19]. postMessage [GSL^{+}19]. posts [YLW^{+}19]. potential [LK18]. potentially [SDM^{+}18]. Power [FC12, MCN10, NTK^{+}19, SG15b, BMH16, CKT11, CRJ18, KD13, MMF19, RDAS19, RE19, UAK16]. poweRed [VRRK17]. PRAC [WZW16]. PRACIS [dFGMTPL17]. Practical [LCZ17, MG12, CWGD15, HXS^{+}16, LMLK16, MFBC11, ML14, SHB12]. Practical-oriented [LCZ17]. practically [DPP^{+}19, WCDY18]. Practice [KHG17, Mic12a, KL18, SPFB19, SYX^{+}16, TLJ14]. Practice-based [KHG17]. practices [DM19, MM15a, MM15b]. practitioner [LDB18, TG14]. Practitioners [Cae12, RS15]. pragmatic [PSK^{+}17]. precise [HDMKS10]. predict [BWMG17]. Predicting [SWS15]. prediction [BLS^{+}16, KKK17, MCG19, RBJ18]. predictors [AAA19]. Preface [Ano17v, YRA15]. preferences [LDB18]. preliminary [TFPC10]. PREON [HMFE10]. preparedness [BM17a]. preprocessing [DC11]. presence [TFE10]. preservation [CVS18, FZZ^{+}12, GL18, SS16]. preserved [AG18, JWZX19]. Preserving [HBB12, PMDM10b, ANS13, AMK18, AKAuR19, APMCR13, BBDP16, BHC18, CBL^{+}19, CZZ^{+}19, CRJ18, CLF17, DYL^{+}14, DDSP17, GWKFC19, GH15, HDT^{+}19, KNX18, LLY17, LML^{+}17, LMLL19, LSQ15, LCZ17, Ngu18, OSP^{+}19, QLZ19, RD19, RR16, SKA^{+}19, SMS^{+}16, Sul19, TMP15, TAP18, XLT^{+}17, XLS^{+}19, ZYWZ14, dFGMTPL17]. pressed [dSFK16]. pressures [AFF14]. prevent [HK13, JCD14, MKG16]. preventing [CAM19a, WP10]. Prevention [BBC17, CY16, DLH^{+}17, PAP^{+}17, SRU18, YG18]. PRIDE [HXS^{+}16]. Prime [SLN^{+}15]. priming [CGLP18]. principles [Den14]. printed [PYLL19]. printer [MGYC18, UAK16]. printouts [LPR10]. prioritisation [SHGH^{+}15]. Privacy [ANS13, AMK18, Ano12r, Ano15i, Ano19i, BRH15, BHC18, BJW19, CRJ18, HBFJ19, JWZX19, KSG17, Kok17, KKKP18, LLYD15, LMLL19, LSKF12, Ngu18, RD19, SS16, Sul19, WX19, ZBB17, dFGMTPL17, AG18, ARL13, APMCR13, ARG^{+}17, AuRMC^{+}18, Ano15j, BAH18, BBDP16, BGR11, BCCS18, CVS18, CBL^{+}19, CDE^{+}11, CLS17, CZZ^{+}19, CGLP18, CLF17, DYL^{+}14, DDSP17, FZZ^{+}12, GGV18, GWKFC19, HBB12, HSR19, HZNF15, HYZ18, JXK18, JLPJ18, KNX18, KKK17, KKP10, KV15, Kül12, LLY17, LML^{+}17, LSQ15, LCZ17, LZ19, MZZ^{+}18, MOV14, MT17, NSGF16, NAMX14, OSP^{+}19, PCP^{+}18, PMLK^{+}11, PAS^{+}18, QLZ19, RR16, RFGL18, RVEK10, SGLC16, SKA^{+}19, SA19, SSNS15, SYCS11, SKLD14, SMS^{+}16, SMA^{+}18, TMP15, TAP18, WZW16, WS17, WWY^{+}18, XLT^{+}17, XLS^{+}19, ZYWZ14, ZS15]. privacy-aware [ARL13, BAH18, RFGL18]. privacy-enhanced [SMA^{+}18]. privacy-friendly [ZS15]. Privacy-oriented [KKKP18]. Privacy-preserving [ANS13, AMK18, CRJ18, Ngu18, RD19, Sul19, dFGMTPL17, APMCR13, BBDP16, CZZ^{+}19, DYL^{+}14, DDSP17, GWKFC19, KNX18, LSQ15, LCZ17, QLZ19, RR16, SMS^{+}16, TMP15, XLT^{+}17]. PrivaKERB [PMLK^{+}11]. Private [HW12, KKK17, CSXL15, CSX^{+}15, DLOP17, KTCS19, NLC^{+}19, ST19, TMN^{+}17, WLF^{+}19, YGX^{+}19]. Privilege [AM17b]. privileged [SC19]. privileges [DFJ^{+}13, ZZ12]. PriviPK [ATL^{+}17]. pro [ZHZ^{+}15]. pro-active [ZHZ^{+}15]. Proactively [OPHC16]. Probabilistic [SLP13, SPR18, BPA^{+}11, MM17, SEJ10, WZK13]. probabilities [Kam19]. probability [AAD^{+}18, KPD15]. probing [BHDA14]. problem [LWZ^{+}17, LJJL15, SSF16, SPD^{+}15, TFE10]. procedure [AM17b, PSK^{+}17]. Process [dABM^{+}18, AWDB19, ALF^{+}18, BR18, CJSM17, DE16, EMWE19, FSZ^{+}17, KEE13, MSRF18, Nie17, PWWY14, PAS^{+}18, RRK^{+}17, WJAM11]. processes [AP14, LJS19, MM15a, MWAG16]. processing [MFBC11, SK13, YGX^{+}19]. processor [MJW^{+}18]. product [JGLE16, YI17]. products [PAS^{+}18]. professional [TBSR14, TRB18]. professionalisation [RS15]. profile [HS11]. Profiles [AH18, RR14]. Profiling [BWMG17, ESY^{+}15, ICRB14, NF15, HA14, PSS19]. Program [Mic12e, SS18b, SKG17, HNSD18, TL15]. programmable [CY16]. programmer [Bay12b]. programming [AGSR^{+}18, WLZ^{+}19]. programs [BBC17, CDKR18, DBH17, TKK15, ZLW^{+}15, ZWNM19]. progress [BCD^{+}17]. progress-sensitive [BCD^{+}17]. promoted [GAA18]. Proof [YLN^{+}19, HYLY19, QHW^{+}19, VBC^{+}15]. Proof-of-QoS [YLN^{+}19]. propagating [GFN11]. Propagation [PWWY14, CSWX19, CAM^{+}18, RMG11, SPCCMMS18]. Properties [FGOPL10, DPP^{+}19]. Property [JJ14, FCSD10]. Property-based [JJ14]. proposed [HWB10]. protect [LMM^{+}12]. protected [HMNT10, PNKK17]. Protecting [ABS14, DSNH17, WS17, ABR18, WW15]. Protection [DPP15, Ano15j, BCCS18, BHCTR12, DC16, FLBK15, GSBY18, HSL19, HYZ18, Ifi12, JK19b, KKK18, KvS10, Ksi12, LLL^{+}18, MWL18, NKL^{+}18, PK17, PCC^{+}18, SS17a, ySTJA^{+}16, TMN^{+}17, Ver18, WZW16, WWY^{+}18]. protective [AM18b]. Protocol [TWNC18, VMS18, AG18, BGR^{+}17, BPA^{+}11, CIS17, CSWX19, FK13, GH15, Gop19, GMLZ11, KKKP18, MVK16, NAS19, NML19, RVEK10, SKA^{+}19, Su10, SMA^{+}18, SMK18b, TOZJ13, TUB16, THA^{+}13, YtLpZD14, YLN^{+}19]. protocol-based [TUB16]. protocols [ACC^{+}13, BKP10, BJ10, DJP17, DLOP17, HWB10, JWZX19, JAAS19, JCD14, KP18, Ksi12, LDC13, LCZ17, MK11, PS10]. prototyping [Ril13]. provable [SYWX19]. Provably [PS10, SM10]. prove [HZG19, HZG20]. Provenance [FB19, AKAuR19, HK17, JKA^{+}18, RG19]. Provenance-based [FB19]. provide [Kül12]. provided [Tau11]. Provider [OM13, ANMN19, JL10, TL15]. providers [HBB12, TBDM14]. Providing [SS17a]. provision [Hor16]. provisioning [MKP12]. provisions [KCJB12]. proximity [ARL13, GKB14]. proximity-based [ARL13]. proxy [CIS17, KNM^{+}17]. pseudo [WN15]. psychological [HKK17, MWL18, Wor10]. psychophysics [BLR17]. public [GDR15, MZZ^{+}18, SPFB19, ZS15]. publication [ARG^{+}17, LZ19]. Publications [SG15a, TG15]. publicly [LCMF18, UB12]. publish [Uzu16]. publish/subscribe [Uzu16]. publishing [ANS13, AMK18, Spa10c, Spa11e]. PUFs [KPD15]. purification [AAdM19]. purpose [SWLZ11]. puzzles [JM13].

Q [PMB^{+}14, PCP^{+}17]. QoP [Ksi12]. QoP-ML [Ksi12]. QoPI [KKK17]. QoS [YLN^{+}19]. quadratic [NDMM13]. Quality [KKK17, Ksi12, Sav13, CZH^{+}14, RRK^{+}17, RMW10]. quality-of-service-supported [RRK^{+}17]. Quantitative [BPA^{+}11, HT12b, MVD19]. quantum [Kam19, MFBC11, YGX^{+}19]. Quasi [AAdM19]. Quasi-purification [AAdM19]. queries [YGX^{+}19]. Query [CWCC12, LJJL15, JC14, XMY^{+}17, XLS^{+}19, ZZC17]. querying [AIB^{+}14]. Question [TWNC18]. Questionnaire [PMB^{+}14, PCP^{+}17]. queue [WP12]. quorum [IGS17]. quorum-based [IGS17].

R [GHÁGGT18]. R-Locker [GHÁGGT18]. R2D2 [GSBY18]. radiation [UAK16]. rail [NTK^{+}19]. Random [RW13, LHM13, QLZ19, WN15, ZZZY13]. randomised [TC11]. randomization [VC12]. randomized [TTJ19]. range [XLS^{+}19]. ranging [DHZ19]. RANs [FYL^{+}17]. Ransomware [ArMS18, CW19, CGR18, GHÁGGT18, SMM^{+}19]. rapid [LMS14, WZW^{+}10, HSPX14]. rare [GOGG19]. rate [TH18]. rates [SBL19]. rating [AIB^{+}14]. ratings [MSMK18]. RBAC [CDOV10, GKB14, JSVA14, LZS^{+}16, LJJL15, ZDBW11]. RDF [AAL16]. RDF-based [AAL16]. re [DLH^{+}17, Par18]. re-delegation [DLH^{+}17]. re-registration [Par18]. reachability [SYGR11]. Reactive [GSBY18]. readiness [EAML15, RV13]. Real [MMJ14, RAA15, SKKP12, AHTS15, GHDGRLM17, LSA^{+}19, MTT^{+}16, NZGH16, NTK^{+}19, SYX^{+}16]. real-life [SYX^{+}16]. Real-time [MMJ14, AHTS15, LSA^{+}19, MTT^{+}16, NZGH16, NTK^{+}19]. realistic [BSMT14, GH15]. Really [Ano14s, KH12, SW19]. reasonable [BB11, RFW^{+}18]. reasoned [KMW19]. reasoning [PBR11, SOG15, ZYRZ16]. rebroadcast [ZZ12]. receipt [DJP17]. receipt-free [DJP17]. receiver [MW11]. receiver-bound [MW11]. recipient [VVS12]. reciprocity [ASO19]. recognition [BMMP18, DB13, DRP^{+}12, FC12, Kal19, SS16, SWS16, dRMC^{+}16]. recognized [KKP^{+}18]. Recognizing [ACS^{+}18, HWX^{+}19]. Recommendations [TKK15, GMdF19]. Reconciling [SSNS15]. Reconnaissance [TG15]. Reconstruction [LPR10, LMLL19, UAK16]. Record [NT18, CC14]. recording [CAM19a, NKL^{+}18]. records [AuRMC^{+}18, BBD^{+}19]. recovery [MBP19, TCNFW16]. recruitment [BWMG17]. recurrent [RBJ18]. recursive [HDK18]. redirection [AYY^{+}17, GSZX18]. reduce [BWOD13]. Reducing [PNKK17, SK10, YGX^{+}19, LLHT13]. reduction [LMLK16]. redundancy [GZS19, GSBY18]. Reed [FCSD10]. reference [CHRI16]. refinement [CDS^{+}19]. reflection [AH10, WRTA17]. registration [CWY^{+}12, Par18]. regret [Ver18]. regulating [HDT^{+}19]. regulation [Hor16]. related [EHMJ19, Guo13, MMCA11, SMM^{+}19]. relational [SEJ10]. Relationship [LC15, BSL19, PWWY14]. Relationship-based [LC15, BSL19]. relaxed [YGX^{+}19]. Relay [IHED15]. releasing [Jän14]. reliability [ALZ^{+}11, ODB18]. Reliable [Hor18, GZS19, HHK^{+}19, OAO17, VSO^{+}18]. remaining [JAN19]. remediating [Hol12]. remediations [ACC^{+}13]. Remote [XYY^{+}12, BS17, HL12, TTJ19]. REP [AM13, ABHS18]. repackaged [LLCT13]. replay [JCD14]. replicated [YWS17]. report [BS19]. reported [TLJ14]. reporting [MP18]. reports [Cho13b, LCMF18]. Representation [AIB^{+}14, DWZ19, KKK17, RAMR17, XSD19, YXD^{+}11]. representative [GZP^{+}13]. repudiable [PW13]. repudiation [VBC^{+}15]. Reputation [ABHS18, BAH18, HBB12, KKS13, LMM^{+}12, TMP15, YAXY19]. reputation-based [LMM^{+}12]. requirements [AP14, BB11, KV15, Les18]. rescue [ZLC13]. Research [SSK17, ArMS18, AM18a, ADWM18, BEB^{+}18, CLS17, Cho11, CBGN18, CJL^{+}13, DPPB17a, DMC19, FPS11, GJ18, Hor18, Kok17, LCZ17, MK19, MMKV15, SPFB19, WMPA19]. researchers [PMP^{+}15]. residual [MGYC18]. resilience [CJW^{+}19, TCNFW16]. Resilient [ALC17, CAM19a, JLPJ18, YHL^{+}15]. resist [FE16]. resistant [HK13, KNM^{+}17, MLH^{+}18, TOZJ13]. resizing [YYL^{+}15]. resolution [GWEA18]. Resource [PAS^{+}18, BRA18, HS13, HSPX14, HXS^{+}16, HDT^{+}19, LZCX18, RFB17, VVS12]. resource-constrained [BRA18, HSPX14]. resource-limited [LZCX18]. resource-preserving [HDT^{+}19]. resource-saving [VVS12]. response [AHR12, GTGB16, Len12, PCP^{+}18, SSCHL14, TVPV19, ZNAHK11]. responses [MVD19]. responsibilities [VVS12]. responsibilization [RFW^{+}18]. responsive [ZH13]. restrictions [SMP13]. result [JC14]. results [VSB^{+}19]. Rethinking [JJL^{+}19]. retrieval [ZLL^{+}18]. retrospective [BB11]. return [ABC^{+}19]. Revealing [HSR19, MMPC18]. Reverse [MLM17, CWGD15, JLK16, OP11]. reverse-engineered [JLK16]. Review [SG15a, TG15, AAMDGR18, Bon19, CCM^{+}18, CBB^{+}16, Cho13b, CP18, DC11, ESCAM17, FB14, GGV18, GDR15, Guo13, IAA^{+}19, IH19, Kok17, Les18, SSK16, SA16, SDM^{+}18, WB16, XL19]. Reviewer [Ano18r, Ano19r]. revisit [RHL15]. revisited [AKK^{+}13, FGOPL10]. Revisiting [HBBM16]. revocable [LML^{+}16]. Revocation [DFCC16, HMFE10, LAHV17, LXH^{+}18, WLWG11]. revoking [ZZ12]. RFID [GH15]. ride [GBC19]. ride-sharing [GBC19]. Rights [SE11, Ven18]. RIPsec [LMM^{+}12]. rise [CW19]. Risk [Mic12e, TDJB18, AKFM18, ANMN19, AAH^{+}18, ABR16, BPG^{+}19, BRH15, BJ13, CBB^{+}16, CGW15, GBC19, HBFJ19, HOVB17, KBAL13, KL18, KMW19, MB17, MHS17, MHL^{+}17, NZGH16, NML19, PC12, RFB17, RFW^{+}18, Rua17, SK18, SAL12, SSCHL14, SSABC16, SEJ10, SKG17, VHBJ19, WAMS14, bORF^{+}15]. risk-based [GBC19, SK18, SAL12, SKG17]. risk-oriented [NML19]. risks [LSSK16, QFG12, vDBD13]. RO [KPD15]. RO-PUFs [KPD15]. roaming [XN14]. robots [GHDGRLM17]. Robust [DLH^{+}17, TLZL16, ASO19, BWR12, BEGW^{+}19, BLS^{+}16, KMG17, SS17a, lTLqZqC19, WHZ^{+}19, YCJ12]. Robustness [SSY12]. rogue [TTCB18, VMTG15]. roid [DLH^{+}17]. Role [SRDD13, CDOV10, CDV12, DMC19, HPJ^{+}19, JJ12, PKP17, SSVA13, SYGR11, TKK15, VHBJ19]. role-based [SSVA13, SYGR11]. Roles [FPS11, GOGG19, HWX^{+}19, MSVA16]. root [JM13]. rootkit [Ril13]. ROP [DCC^{+}18]. ROPSentry [DCC^{+}18]. rough [DKG13]. round [MK11]. route [NT18]. routing [BGR^{+}17, CPGA13, IGS17, MAC17, RD19, Su10, ZLC13]. Roving [FW10]. RTECA [RAA15]. Rule [LSKF12, GZS19, KD13, YWYZ12]. rules [LKJ14]. Runtime [BJ10, CMM^{+}10, DCC^{+}18, COS19, FSZ^{+}17, TS13].

S [CKJ^{+}17]. S-OpenSGX [CKJ^{+}17]. SaaS [TL15]. SADUS [YWZM18]. safe [KKP^{+}18, RMVC^{+}12, ZTL^{+}19]. Safeguarding [NML19]. safety [AKFM18, ARG^{+}17, DM19, ySTJA^{+}16]. safety/security [AKFM18]. sandbox [MJRK18]. sandboxes [SMP13]. sanitization [LMS14]. sanitizing [IGS17]. satisfaction [MM15b]. Saving [BMMP18, VVS12]. SCADA [AYT^{+}14, ADWM18, CBB^{+}16, Kal19, LXDW16, LGZ^{+}19, NPP17, NWD^{+}12, SD17]. SCADAWall [LGZ^{+}19]. Scalable [ZLL^{+}19, CBL^{+}19, DYL^{+}14, DYZ^{+}15, WLWG11, ZTC19]. scale [BLS^{+}16, CGPP11, CTH10, GPGV16, JJL^{+}19, OS17, SJCM17, SDM^{+}18, VSB^{+}19, ZWX^{+}19]. scales [JXK18]. scanning [Hol12]. scenarios [KKS13, MPPS10, MLV16, MHL^{+}17, PHMN16, RAA15]. schedule [CC14]. Scheduling [SLP13, KTG^{+}18, SP15, SMK18b, VPNP18]. scheme [ARL13, AKAuR19, BD18, BWR12, CTH10, DDSP17, FZZ^{+}12, Gar16, GJJ18, GBC19, HL12, JWB16, JK19b, LML^{+}17, NMX15, OSNZ19, PZ15, hRKRR13, SH15, SYWX19, SCR19, SMS^{+}16, VSO^{+}18, WHZ^{+}19, XLS^{+}19, ZLL^{+}19, ZZ12]. schemes [ALRC12, HK13, IHED15, MBP19, NMX19, SS17a, SP15, SSK16, Sul19, ZKM^{+}17]. science [HLL17, PC12]. Scope [Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano10-27, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x]. Score [PPdCL17, HW19]. Scoring [HA15]. scrambling [Kül12]. screen [IAA^{+}19]. Screening [LLL15, YST^{+}19]. scripting [VC12]. SDC [RNdP17]. SDN [CJA^{+}19, CY16, FYL^{+}17, PG19, SBZD17]. SDN-enabled [CJA^{+}19]. SDSS [JWB16]. SDSS-MAC [JWB16]. sealed [ML14]. SEAM [GPGV16, YYL^{+}15]. search [BTK15, JMGW17, RP19, SS17b, VMA16, Zin15]. SEC [HF17, Ano10-32, Ano12r, Ano17v]. SEC’14 [Ano15j]. SecKit [NSFB15]. Secloud [ZHB^{+}13]. SECO [DYZ^{+}15]. second [JAN19, SNG^{+}17, UB12]. secret [BEB^{+}18, GEAHR11, LH11, Par18, SM10]. secrets [BS19, LDC13]. SECRYPT [Ano19i]. sector [LSC14]. Secure [AP13, DYZ^{+}15, HM19, JKA^{+}18, JWB16, JMGW17, LNK16, LML^{+}16, LZCX18, MDHM18, MMPC18, OM19, OAO17, RR17, RRK^{+}17, YWZM18, ZDBW11, AG18, AKAuR19, ATL^{+}17, AAR^{+}18, BCRAA13, Bon19, CCC^{+}12, CJA^{+}19, CM19, DA18, FYL^{+}17, GWEA18, HW18, HZMH18, JL10, JJ14, JAAS19, KP18, LMS14, LM10, LLX^{+}17, LXH^{+}18, NKL^{+}18, OSNZ19, ODB18, PDB14, PMP19, RAMR17, SLS^{+}18, SLD12, SYCS11, SM10, THA^{+}13, WHZ^{+}19, WA19, XMY^{+}17, ZY14, ZWX^{+}19]. Secure-channel [JMGW17]. secured [NLC^{+}19]. securely [JGM^{+}13]. Securing [BKK19, DHZ19, EMWE19, JK19a, KW14, Mic12c, Mic12d, OBRK15, TXM^{+}19, TG14, UFF12, BCK19, GH15]. Security [AAJ12, AX15, AKP14, Ano12r, Ano15i, Ano18b, Ano18c, Ano19a, Ano19i, BTM14, Bay13, BJW19, BCCS18, Cae12, EHBH^{+}16, GPGV16, Guo13, GZ20, GZG20, HBHA18, HF17, HZG20, HSW17, JSVA14, LW12e, MCT^{+}17, Mic12a, Mic12d, Mic12e, MSA11, NSFB15, NF15, OM19, PMB^{+}14, PCP^{+}17, RDAS19, RAMR17, SCRA17, SLH14, Spa10c, Spa11e, SG14a, TDJB18, UFF15, VMTG15, WL12, ZCZ^{+}19, AKFM18, ACS19, AHP18, AHR12, ASFI18, AAH^{+}18, AH10, ABP16, AFF14, AAdM19, AM18a, AAMDGR18, Ano15j, AIRC13, AM17a, AM17b, AM18b, BGR^{+}17, BMH16, BM17a, BMO18, BTLT16, BKP10, BBC17, BFS^{+}18, Bon19, BB11, BS17, BRM^{+}18, BCK19, CSBW18, CLL^{+}13, CHRI16, CBB^{+}16, CGLP18, CJL^{+}13, DE10, DPPB17a, DPPB17b, DSP16, DMC19, DE16, EGM10, EHMJ19, FHL14, FAE14, FE16, FT16, FK13, FPS11]. security [FC12, FanE^{+}18, GBC^{+}18, GCC18, GMMJ11, GB19, HPJ^{+}19, HKK17, HL18, HSL19, HT12a, Hol12, HEK^{+}18, HPF^{+}16, HZMH18, HLFZ16, HLL17, HAHZ^{+}18, IHED15, IAA^{+}19, Ifi12, JKC^{+}13, JCD14, KDC^{+}14, Kar12, KHG17, KK16, KCJB12, KAF17, KMW19, KBM16, KD13, KvS10, KTG^{+}18, KKS13, LMM^{+}12, LDB18, LDC13, LAVK19, LLK16, LK18, LGZ^{+}19, LSA^{+}19, LMNM^{+}19, MBE^{+}12, MRS12, MKP12, MCN10, MMCA11, MM15a, MWAG16, MTT^{+}16, MAZAS17, MBAE^{+}11, MCL^{+}13, MM15b, NM19, NPP17, NCS14, NZGH16, NWD^{+}12, NML19, NTK^{+}19, ÖTC16, OS17, Pad12, PKP17, PDM19, PMDM10a, PC12, PCCM16, PS10, PAS^{+}18, PBR11, QFG12, QHW^{+}19, QLZ19, RC15, RE19, RAA^{+}14, RS15, RFW^{+}18, RRK12, RMVC^{+}12, RVEK10, RR14, RHL15, RM17, SS17a, SSS^{+}15, SVF16, SBL19, SCB^{+}13, SJCM17, SPFB19, Sav13, STA18, SWS16, SSABC16]. security [SLG19, SW19, SYX^{+}16, SWS15, SYWX19, Sil13, Sil19, SSF16, SOG15, SEJ10, SA16, SHB12, TBSR14, TBDM14, TMW17, THA^{+}13, TLJ14, TRB18, TKK15, Uzu16, VV10, Ver18, VHBJ19, WZK13, WJ15, WAMS14, WHZ^{+}18, WYS18, WHSE15, WW18, XNP16, XHZ11, XHC^{+}10, YYJC12, YCH18, ZBBO15, ZBB17, ZLL^{+}18, ZH13, ZHB^{+}13, dVM15, dVM17, vDBD13, vdWEG18, vSvN13, HB18, MKG13, TMW17]. Security-by-construction [EHBH^{+}16]. security-critical [NZGH16]. Security-related [Guo13]. Security-SLAs [HB18]. security-state [WHZ^{+}18]. see [LHS^{+}18]. seeders [AKJR14]. SEIQV [WZW^{+}10]. Selecting [ALRC12, OM13, TL15, ASFI18, Pal10]. Selection [CH14, BK19, CGLP18, DRN16, GCD19, IHED15, KK17, KH12, ST19, SS13, lTLqZqC19, VDK18, YXD^{+}11, ZS18]. selective [GTDVTSH15]. Self [BWMG17, LMNM^{+}19, ABR18, BFTJ18, GFN11, HDT^{+}19, MKP11, ZNAHK11]. Self-efficacy [BWMG17]. self-modifying [MKP11]. self-optimizing [ZNAHK11]. Self-Organizing [LMNM^{+}19]. self-propagating [GFN11]. self-protecting [ABR18]. self-regulating [HDT^{+}19]. Selfholding [DWF^{+}19]. selfish [DWF^{+}19, WLW^{+}19]. seller [CTH10]. Semantic [CFH^{+}11, MSV12, HL18, MBSR18, XSD19, ZLLL12, AAL16]. semi [hBCK17, PWWY14, SHB12]. semi-automated [SHB12]. semi-Markov [hBCK17, PWWY14]. sender [VVS12]. senior [BTLT16]. sensing [XLT^{+}17]. sensitive [BCD^{+}17, STA18, ZS18, ZNAHK11]. Sensitivity [AH18, BMMP18]. sensor [hBCK17, CVS18, GFN11, HR13, HL18, JWZX19, MMMG13, OAO17, RR17, SPYG15, SMK18b, TOZJ13]. sensors [BVWM19]. sentiment [ZDW19]. SENTINEL [OBRK15]. separation [PMDM10b, UPF17]. September [Ano10-36, Ano11-28, Ano12-28, Ano13z, Ano14y, Ano15u, Ano16x, Ano17u, Ano18o, Ano19q]. sequence [CSX^{+}15, YLB^{+}17]. sequences [LLCT13, WN15]. sequential [ARG^{+}17, LXDW16, LWS^{+}19]. serial [MCRB19]. series [RNR15]. server [BBDP16, BR17, PNKK17, PNKK17]. servers [PAS^{+}18, SM19a, WLWG11, WJAM11]. Service [BSMT14, CGMP18, HW12, LC15, OM13, AAD^{+}18, BOZL13, DYL^{+}14, EGM10, HK17, IH19, JL10, Kar12, MKP12, MB17, RRK^{+}17, SFK15, SLP13, SLD12, SVA10, SSK16, TL15, TUB16, TBDM14, TH18]. Service-Oriented [HW12, RRK^{+}17]. services [Bel18, CAM19a, DYZ^{+}15, FHF15, JAH^{+}11, KSG17, NDSA17, NGSV16, PKK^{+}14, RVEK10]. session [JCD14, TUB16]. sets [MCN10, RWS^{+}19, SPK17]. SGX [CKJ^{+}17, KJP^{+}19, PPKK19]. SGX-LEGO [KJP^{+}19]. Shades [SG14a]. Shadow [SB14, Sil19].

SHADuDT [AP13]. Shaping [FE16, HSL19, MT17]. share [GJJ18]. shared [GEAHR11, LDC13, LXH^{+}18, Par18, SSF16]. sharing [DYL^{+}14, FAE14, GBC19, JWB16, LKJ14, LH11, LZCX18, QAR^{+}17, SSF16, TBSR14, WMPA19, WLWG11, WHZ^{+}19, dFGMTPL17, dLGHK18]. ShellBreaker [LHI^{+}19]. shells [LHI^{+}19]. shortfalls [SMP13]. shoulder [YWL^{+}17]. shoulder-surfing [YWL^{+}17]. side [CAM19a, FNE19, NADZ18]. side-channel [FNE19]. SIEM [BS17]. SIGMA [KKKP18]. sigmoidal [RHL15]. Sign [ACC^{+}13, SCRA17]. Sign-On [ACC^{+}13, SCRA17]. signal [UAK16]. signaling [hBCK17]. signals [LNK16]. signature [GOB11, MLK14]. signature-based [GOB11, MLK14]. signatures [CIS17, GTDVTSH15, HAGTdFR13, LLL15, LPR10]. signed [ASO19, ALZ^{+}11]. significance [BCK19, CGR18]. Significantly [CBL10]. signing [ALZ^{+}11]. silence [SMK^{+}18a]. Silver [Ano12r]. similarities [BGL18]. Similarity [DVC18, BM17b, wJKW^{+}16]. SimMonitor [XNP16]. simple [UPF17]. simulated [KBM16]. Single [ACC^{+}13, GMMJ11, MCRB19, RVEK10, SCRA17]. single-factor [GMMJ11]. sinkhole [HR13]. SIP [DSNH17, EGM10, KMM13, hRKRR13]. SIP-based [EGM10, KMM13]. site [JCPACRV12, VC12]. sites [VMTG15]. situation [WAMS14]. situational [FB14]. size [JC14, SP15]. size-based [SP15]. sketches [CGPP11, CYLL17]. skip [SS18a]. skip-gram [SS18a]. SLAs [HB18]. Slow [TH18]. SlowBot [FCPA16]. slurping [KK10b]. small [OS17]. small-scale [OS17]. Smart [BD18, BJW19, HX17, LAR13, PAP^{+}17, SMA^{+}18, VHBJ19, Bel18, BL15, CRJ18, GWKFC19, HLB^{+}18, HLB19, JK19b, KNX18, KMW19, LSSK16, Les18, RVEK10, SYWX19, Sul19, WZW16, XXLZ18, XdZMsC18, ZH13]. smarter [ADF^{+}10]. smartphone [BRH15, HSR19, KN14, LLL15, MKG13, PWWY14, Ver18, AFF14, MMMG13]. smartphones [MLKC17, MJW^{+}18, SPYG15, ZHB^{+}13]. smartwatches [NM18]. smooth [DKG13]. SMS [LH11]. SMSes [SNG^{+}17]. SMT [TdHRZ17]. smudge [KN14]. Snake [BBDP16]. snakes [NDMM13]. SNAPS [RG19]. snapshot [RG19]. Snatched [BS19]. SNMP [CMPR15]. Snowden [PCP^{+}18]. Social [AF14, GMGTdFR14b, Hat18, MLV16, PKK^{+}14, SG15a, AWZ^{+}17, ASO19, AIB^{+}14, ACS^{+}18, Ano13-30, AM13, CFH^{+}11, CLL^{+}13, DPPB17a, ELG^{+}17, FE16, FFGL18, GL18, Hat19, HL18, JHJ^{+}16, JJA13, JJ14, LLYD15, Med18, MMKV15, Nie17, PZ15, PSS19, PWWY14, SS11, SSAK12, SNG^{+}17, SKLD14, Ven18, WS17, WWY^{+}18, YWND14, vdWEG18]. Socialized [GHL^{+}17]. socially [AM17b]. socioeconomic [DM19]. Sociometrics [BRM^{+}18]. Socket [WA19]. SOFIA [dCGÜ^{+}17]. Soft [ICRB14, SS16]. Software [AX15, KL18, MG12, TDJB18, dCGÜ^{+}17, AAT16, ABC^{+}19, BS17, EMWE19, HLB19, HHK^{+}19, JJL^{+}19, JGLE16, KYZ15, LM10, MM15a, RMVC^{+}12, SA19, Sil13, TBDM14, ZLLL12, Zhu18, GWKFC19]. software-defined [HHK^{+}19]. Solomon [FCSD10]. solution [MDHM18, ML14, SCRA17, SVGE14, UFF15, ZHB^{+}13]. solutions [AM17a, SG10b, Uzu16]. Solving [GCC18]. SOM [TFPC10]. SoNeUCON [GMGTdFR14b]. sophisticated [NNNQ18, TR18]. Source [CB12, PYLL19, TG15, AM18a, HDMKS10, MGYC18, NMX19, RM17, Sil13, WW18]. sources [SPFB19, SS18b]. space [CC14, YWZM18]. spaces [ANAT15, ADF^{+}10, CHK^{+}19, CC14]. spam [LWZ^{+}17, LGS14, VVS12, WWXZ18, YKH10]. spammers [FFGL18, VMS18]. SPARQL [AAL16]. SPARQL/SPARUL [AAL16]. sparse [ANAT15]. SPARUL [AAL16]. Spatial [JXK18]. spatio [MCY^{+}18]. spatio-temporal [MCY^{+}18]. Special [Ano12r, Ano13-30, Ano19i, BJW19, HLL17, LM10, YWND14, Ano15j, Ano17v]. Specification [CHK^{+}19, KFGL19, SBW10]. specification-based [SBW10]. specifications [PS10]. specifiers [ODB18]. Specifying [dLGHK18]. spectrum [YLB^{+}17]. Spherical [ANAT15]. Spi [PS10]. spies [MAC17]. SPIT [AMS18]. splay [TZMG15]. split [MPM15]. split-personality [MPM15]. splitting [CSXL15]. spoofing [MA17]. spread [YLB^{+}17]. spreading [WZW^{+}10]. SQL [CWCC12, JC14, KPS16]. SQLiGoT [KPS16]. square [JM13]. squaring [SLN^{+}15]. Stability [WZW^{+}10]. Stable [LGS14]. stack [LLL^{+}18]. Stackelberg [GCC18]. stage [CAL15, LAVK19, RBJ18, TFPC10]. Staircase [CVS18]. stakeholder [BTM14]. stakeholders [MR11, RR16]. stalkers [ESCAM17]. Stalking [ESCAM17]. standardisation [KBM16]. standardization [HB18]. standards [Les18, VHBJ19]. starvation [HT17]. State [GJ18, NCGE15, RDAS19, SMP13, WHZ^{+}18, ZTL^{+}19, ZLW^{+}15]. state-of-the [NCGE15]. stateful [AGL12, GACCB^{+}13, LK18]. stateless [BBDP16]. Static [ZW14, CKT11, DRP^{+}12, GDR15, HXW^{+}19, YXD^{+}11]. stationary [CC14]. Statistical [TZMG15, CPVGTMF16, HMNT10, MW11, TZZG13, WN15]. status [AM17b, MJBKZ15, SW19]. Stealing [AKJR14]. Stealth [MAC^{+}15]. stealthy [RDAS19, ZYRZ16]. steganalysis [LHM13]. Steganography [YWYZ12]. step [NDP18, RAA15]. stepping [YW11]. stepping-stone [YW11]. steps [CGGFGMF19]. Sticky [SBE17]. stochastic [HBT^{+}19]. stock [KCJB12, SA16]. stone [YW11]. storage [CCC^{+}18, JWB16, LML^{+}16, PAS^{+}18, XYY^{+}12]. stories [ZMYB17]. storm [AH13]. strategic [DLB^{+}19, RM17]. Strategically [AWDB19]. Strategically-motivated [AWDB19]. strategies [AAdM19, AAMDGR18, DRN16, HBT^{+}19, TUB16, ZS15]. strategy [MHS17, PTM^{+}18, lTLqZqC19, ZS18]. stream [RNdP17]. streams [CBL^{+}19, WLF^{+}19]. strength [AM13, GZ18, GZ20]. Stress [KGSN17, LLK16]. strong [GH15, GZG19, GZG20]. structural [ALZ^{+}11, CMV16, DSNH17]. structure [JKA^{+}18]. stubborn [WLW^{+}19]. student [NF15]. studies [PMP^{+}15, PCP^{+}17]. Study [JCPACRV12, AH10, AHGC19, BMH16, BM17a, BTLT16, CMV16, CKA^{+}17, DSP16, GFN11, HCR10, HLFZ16, KPKS15, KL18, LSKF12, MK11, Nie17, OS17, SBK12, SWS16, SYX^{+}16, SS18b, SBS^{+}18, VHBJ19, VSB^{+}19, WYS18, WWXZ18, YGZG19, ZDW19, ZGZ^{+}13, dVM15]. Studying [GL18]. Stuxnet [MSR19]. style [PZ15, WP10]. Styx [BRH15]. Sub [AAdM19].

Sub-optimality [AAdM19]. subcultures [dVM17]. subfields [RM17]. subjective [HOVB17]. subscribe [Uzu16]. subsequent [MVD19]. substations [LAR13]. subtle [YGX^{+}19]. success [ArMS18]. Successive [XdZMsC18]. Sudhanshu [TG15]. suggestion [WWY^{+}18]. suitable [Pal10]. supervision [JL10]. supply [CRJ18]. Support [CH14, MC18, CCM^{+}15, GZ10, HHAW19, Jän14, NGSV16, SRU18, VDK18]. supported [RRK^{+}17]. supporting [AHR12, FanE^{+}18, HZMH18]. suppression [SMK^{+}18a]. SURF [YCJ12]. surface [BM17b]. surfaces [ELG^{+}17, PNKK17]. surfing [YWL^{+}17]. surrounding [Hor16]. surveillance [FW10, PCP^{+}18]. Survey [AMS18, EGM10, HZMH18, IH19, JAN19, KS14, LCMF18, STSB17, UAB19, WMPA19, WWXZ18, AR17, ArMS18, AZ17, AX15, BGL18, BEB^{+}18, FGHHR18, FPS11, HBBM16, HAGTdFR13, JvOW11, KKM11, LK10, MTC^{+}18, MPPCFG^{+}12, NDP18, NPP17, NCGE15, RC15, RWS^{+}19, RMVC^{+}12, SMP13, SSK19, SSF16, Sul19, Tau11, TZTC16, TR18, UFF12, Uzu16, VBC^{+}15, ZKM^{+}17, ZLK10]. sustainability [LMG18]. sustainable [BCCS18, VPNP18]. SVM [CH14, GWWW19, KPS16]. Swarm [KKM11, EE15]. Sweden [Fra17]. switching [YWYZ12]. symbiosis [SBW10]. Symbolic [SMC^{+}18, SYGR11]. symmetric [ZCZ^{+}19]. Sync [SFK15]. synchronisation [SFK15]. synthesis [BJLS17, Guo13]. synthetic [SSY12]. System [BD18, HA15, ZY11, AAH^{+}18, ATG11, ABHS18, BAH18, BTLT16, BOZL13, CCM^{+}15, CJA^{+}19, CKJ^{+}17, CJSM17, CGW15, FSZ^{+}17, FKSK19, GH15, GHDGRLM17, HDK18, HHBS18, HXS^{+}16, IRC^{+}17, wJKW^{+}16, JFR13, Kha12, LMS14, Len12, LXDW16, LMLK16, LML^{+}16, LLX^{+}17, LLCT13, MCN10, MK19, MCT^{+}17, MMJ14, MB18, MLH^{+}18, NTK^{+}19, ODB18, OS17, PW13, RDAS19, RG19, RP19, RV13, RHL15, SE11, SS16, SJCM17, SMM^{+}19, SSCHL14, SRU18, SLD12, SC19, SOG15, TFPC10, VDK18, VSB^{+}19, WRTA17, WHSE15, XHZ11, YWS17, YG18, YXD^{+}11, ZZ19, ZLL^{+}18, ZWX^{+}19, ZH13, ZS15]. system-level [CKJ^{+}17, WHSE15]. Systematic [DEP13, LZBS13, Bon19, CSBW18, CCM^{+}18, FB14, GGV18, IAA^{+}19, IH19, NDP18, SSTG12, SSK16, SA16, XL19]. Systematically [SHB12]. Systems [AF14, Ano19i, AMS18, BJW19, AR17, AHP18, ALC17, AIB^{+}14, AYT^{+}14, Ano15j, AM17a, CM19, CXF^{+}18, CZH^{+}14, CBB^{+}16, CZZ^{+}19, DMR19, DLDTGMMP15, DPP^{+}19, EGM10, GOB11, GM12, Guo13, GKB14, HW18, HBB12, Ifi12, JCPACRV12, JWB16, JJA13, Kal19, KBAL13, KL18, KN14, LM10, MDHM18, MLK14, MMF19, MSRF18, NZGH16, Nie17, NFGL16, PHP19, PK17, PPdCL17, RARL19, RAMR17, SA19, SAL12, SK10, SMS^{+}16, TTCB18, Tau11, TS13, UFF12, UFF15, Uzu16, VEPZ15, VSO^{+}18, WP10, XHC^{+}10, XLT^{+}17, YG18, YLB^{+}17, ZHZ^{+}15, ZGD19, ZW14, dRMC^{+}16, AKFM18].

table [DDSP17, IGS17]. tabular [AMK18]. Tactics [Cae12, Mic12d, AWDB19]. tagging [WWY^{+}18]. taint [CWGD15, ZLW^{+}15]. Take [DKG13]. taking [AM18b]. Tallinn [PTM^{+}18]. Taming [CDOV10].

Tangible [HZMH18]. Tap [NM18]. Tap-based [NM18]. target [APMCR13, DRP^{+}12, HEK^{+}18, LSQ15, lTLqZqC19, ZS18, qZLxCjY17]. target-driven [APMCR13, LSQ15]. Taxonomising [CW19]. Taxonomy [BFS^{+}18, Pad12, SSCHL14, SSABC16, ArMS18, HLB^{+}18, HAGTdFR13, HZNF15, KUA^{+}18, MKP11, PRMD10, ZKM^{+}17]. TC [Ano15i]. TC-11 [Ano15i]. TC11 [Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano10z, Ano10-27, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x]. TCP [AH13, YW11]. TCP/IP [YW11]. TDL4 [CKA^{+}17]. TDSS [CKA^{+}17]. TDSS/TDL4 [CKA^{+}17]. teaching [AAL16]. teaming [BRM^{+}18]. teams [AHR12]. Tech [LW12c]. Technical [Ano10x, Ano10y, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, HX17, TR18]. technique [EHMJ19, HDK18, LH11, qZLxCjY17]. Techniques [Cae12, Mic12d, USB17, AZ17, BJJ^{+}17, CCM^{+}18, GCD19, HEK^{+}18, JJL^{+}19, Kal19, MTT^{+}16, NNG^{+}18, NPP17, SCLD19, SGK^{+}16, UAB19, Zin15, TG15]. technologies [HZNF15, LK18, YST^{+}19]. technology [ESCAM17, HBHA18, HOVB17, KCJB12, MT17, QFG12, Spa10c, Spa11e, TRB18]. telephone [GMMJ11]. telephony [SMK^{+}18a, SG10b]. template [BCRAA13, DDSP17, PGdCL16, SS17a]. templates [BCRAA13, MLV16, QLZ19]. temporal [ABR16, BM17b, DSNH17, JSVA14, Kal19, LWS^{+}19, MCY^{+}18, MSVA16, SSVA13]. tenant [MPCK18]. term [VBC^{+}15]. terrorist [STSB17]. test [SMC^{+}18]. test-patterns [SMC^{+}18]. testbeds [ADWM18]. Tester [CB12, KG13]. Testers [LW12b]. Testing [SG15b, CSWX19, EMWE19, Hat19, KCB13, KBM16, Ril13, WN15]. text [JKC^{+}13, NCS14, ZMYB17]. text-based [NCS14]. textual [MVD19]. TFTP [SMB16]. thanks [CBL10]. theft [BS19, ZMYB17]. Their [AS13, PSS19, ROW19, SMP13, SS18b]. Theorem [QHW^{+}19]. theoretic [CGP17, SKTO13, YLB^{+}17]. Theoretical [HSL19, Bay13, EAML15, LZBS13, RE19, Wor10]. Theory [DPP15, Mic12a, AP13, CLL^{+}13, Ifi12, LSKF12, RB12, SYK14, ySTJA^{+}16, Ver18]. there [VHBJ19]. Things [CGMP18, PCC^{+}18, CBGN18, SBK12, SSSC18, AAC^{+}19, AAR^{+}18, ABHS18, Bel18, BOZL13, HYLY19, MAZAS17, NSFB15]. thinking [Bay12c]. third [AF14, PW13, SSAK12, SBS^{+}18, TXM^{+}19, VHBJ19]. third-party [AF14, SBS^{+}18, TXM^{+}19]. Thor [LW12e]. thread [LLCT13]. thread-grained [LLCT13]. Threat [LK18, XL19, AWDB19, ArMS18, AH18, CGMP18, CAY^{+}18, Cho11, FCPA16, FW10, HLL17, KGSN17, KKS13, LCMF18, MPM15, MVD19, NTK^{+}19, QAR^{+}17, SCLD19, TR18, WMPA19, XN14, ZMYB17, ZW19]. threats [AM17b, BJ13, BJJ^{+}17, FSSF15, HBHA18, HLB^{+}18, JvS17, LAR13, NB19]. three [CAL15, HWB10, MVD19, WLF^{+}19]. three-party [HWB10]. three-phase [WLF^{+}19]. three-stage [CAL15]. thresholding [DT19]. throughput [NAS19]. Thumbs [NKL^{+}18]. thwart [Rru16]. Thwarting [GHÁGGT18, WLZ^{+}19]. ticketing [MPPCFG^{+}12]. Tier [TTJ19]. Time [CC14, JGLE16, MMF19, RNR15, AHTS15, AAD^{+}18, CGPP11, dSFK16, GHDGRLM17, HYZ18, LSA^{+}19, LPL18, MMJ14, MTT^{+}16, NZGH16, NTK^{+}19, Par18, RAA15, SKK^{+}14, SKKP12, SPK17]. time-scale [CGPP11]. timestamp [Cho13a]. timing [AG14, LLD^{+}19, MZJ^{+}12, ZYPL10]. TinyLock [KN14]. TinyVisor [SLS^{+}18]. tips [YGZG19]. TLS [SM19a]. TOCTTOU [WP10]. tokens [KPS16]. Tolerant [DFC13, AGL12, DFCC16, KUA^{+}18, MVK16]. toning [Zin15]. tool [BWMG17]. Toolkit [CB12, NSFB15]. Tools [Cae12, Gar16]. TOoltip [VRRK17]. TOoltip-poweRed [VRRK17]. topic [CLS17]. topology [CHK^{+}19]. topology-aware [CHK^{+}19]. TORPEDO [VRRK17]. touch [IAA^{+}19, MLKC17, NSBM17, TZTC16]. touchloggers [DKG13]. TR [FKSK19]. trace [ZGZ^{+}13]. traceback [AR14, HDMKS10, NT18, SSK16]. traces [SPR18]. Tracing [ZZ12]. tracking [DEP13, FSZ^{+}17, HBF13, OAO17, QHW^{+}19, SM19b]. trade [BS19, MMPC18]. trade-offs [MMPC18]. tradeoff [MMF19]. trading [CJA^{+}19]. Traffic [HSPX14, hBCK17, BSMT14, CGP17, CLF17, DT19, GPAD^{+}14, GBCD19, HBF13, RSLH19, SYK14, SPD^{+}15, SMK18b, TZZG13, WLCC17, ZTS^{+}13]. Traffic-agnostic [HSPX14]. training [ACS19, BMH16, BTP^{+}18, HM19, MLH^{+}18]. training-resistant [MLH^{+}18]. traits [GBC^{+}18]. trajectories [WS17]. Trajectory [HYZ18]. Transaction [MMLN15, CSXL15]. transactional [SPK17]. transactions [CGR18]. transfer [BL15, KTCS19]. transfers [SMK^{+}18a]. transform [Kül12]. transformation [DSP16, LZS^{+}16]. transformational [FE16]. transformations [ZLW19]. transition [Kar12, LK18]. Transparency [KKM19]. transparent [CRS13, GSZX18]. transpiler [STA18]. transport [MPPCFG^{+}12]. transportation [CJA^{+}19, LAVK19, SMS^{+}16]. transversal [MPCK18]. Tree [SLB^{+}13, AKFM18, CSWX19, TZMG15]. Tree-formed [SLB^{+}13]. trees [Kam19, ST19]. trends [CLS17, JKC^{+}13, RARL19, ZKM^{+}17]. triage [CJSM17, Gar13, ZLL^{+}18]. Triangular [RR16]. trick [SCLD19]. Trickbot [GWWS19]. trigger [ESY^{+}15]. triggered [PAS^{+}18]. Triple [DVC18]. Triple-Similarity [DVC18]. Trojan [GWWS19, SMC^{+}18]. truncation [CKT11]. Trust [BOZL13, HHBS18, HOVB17, SRU18, SYK14, ASO19, Ano13-30, BGR11, BSBG19, DFC13, DFCC16, DLOP17, KFGL19, MKP12, MPPS10, RFGL18, SWLZ11, YWND14, ZLC13]. trust-based [ZLC13]. Trusted [HT12a, IH19, PKK^{+}14, BGLR17, BHCTR12, FNE19, JK19a, PAJAP15, PW13, SLB^{+}13, TL15, YI17]. Trustee [JLPJ18]. Trustworthy [BGLR17, DLS14]. truth [SPD^{+}15, XLT^{+}17]. Ts [LZS^{+}16]. Ts-RBAC [LZS^{+}16]. tunnel [DF16]. tweet [LSA^{+}19]. Twin [NADZ18]. Twins [JLGM19, BMO18]. Twitter [LSA^{+}19, LWZ^{+}17, WWXZ18]. Two [HWB10, PCP^{+}17, BD18, GMMJ11, IHED15, MDHM18, NMX15, SWS16, TFPC10, NKL^{+}18]. two-factor [BD18, GMMJ11, MDHM18, NMX15]. two-stage [TFPC10]. Two-Thumbs-Up [NKL^{+}18]. two-way [IHED15]. type [LLK16]. types [MVD19, WL18]. typing [ICRB14, SCR19]. typographical [CJW^{+}19].

U [XNP16]. ubiquitous [LJS19]. UFace [LHS^{+}18]. UGR‘16 [MFCMC^{+}18]. UI [CM17]. UK [JAN19, OS17, RS15]. ultimate [BL15]. unaided [CAM19a]. unauthorized [JvOW11]. uncertainty [AM18b]. uncle [WLW^{+}19]. Unconstrained [GEAHR11]. Uncoupled [HDT^{+}19]. Uncovering [AAA19]. Understanding [CLL^{+}13, CWY^{+}12, Ifi12, LLK16, MR11, Mic12a, NB19, PBR11, ySTJA^{+}16, Ver18, AAMDGR18, LSC14]. Underworld [TG12]. unfair [AIB^{+}14]. Unified [HK17, SSVA19]. Uniform [AR14]. unifying [Rua17]. Unintended [AF14]. Union [RR14]. United [Hor16]. units [LLL^{+}18]. universal [KS19, LHS^{+}18, PW13]. Universally [MFBC11]. universe [Bon19]. university [DPP15]. Unix [WP10]. Unix-style [WP10]. unmanned [MK19]. Unrealistic [RRK12]. unrevealed [HSR19]. unsafe [AAA19]. Unsupervised [SS18a, AYT^{+}14, LWS^{+}18]. untrusted [MJRK18]. unwanted [AM13, SYK14, SDM^{+}18]. update [DFCC16, GOB11, HYLY19, PGdCL16]. updates [ARG^{+}17, TFE10]. URL [AYY^{+}17, KKK18]. USA [SG15a, TG15]. usability [GMMJ11, SS11, SWS16, WA19]. usable [GMMP18, MMPC18, ZY14]. Usage [KMM13, CRJ18, GOGG19, GMGTdFR14b, KGSN17, KS14, MJW^{+}18, RFB17]. USB [NYE17]. USB-based [NYE17]. use [KMW19, LDC13, RMW10, Sil13]. used [RVEK10, SPFB19]. User [BOP14, CCC^{+}18, GMMJ11, LLD^{+}19, LJJL15, SYX^{+}16, VRRK17, BBC17, BFS^{+}18, CLF17, ESY^{+}15, FNE19, FanE^{+}18, HFCR13, HL12, KLN15, LMNM^{+}19, LXH^{+}18, MM15b, MKG13, NADZ18, NM18, OS17, PMLK^{+}11, SCFB15, SGLC16, SK18, SSNS15, SPK17, VSB^{+}19, VGL14, WLWG11, YWZM18]. User-friendly [CCC^{+}18]. User-independent [LLD^{+}19]. user-side [NADZ18]. user-trigger [ESY^{+}15]. users [BLR17, Bel18, FHM^{+}10, GBCD19, HSR19, KvS10, LZCX18, LAHV17, MBAE^{+}11, NF15, PSS19, RR14, WS17]. Using [BLR17, CSWX19, HS18, HCRMBG17, JKC^{+}13, LMM^{+}12, SS11, VC12, ASO19, AAH^{+}18, ABP16, AGSR^{+}18, AH18, ADF^{+}10, AW18, AP13, BOP14, BD16, BFTJ18, COS19, CKT11, CIS17, CSBW18, CYLL17, CAM^{+}19b, CLS17, DA18, DCC^{+}18, DT19, DB13, DWF^{+}19, DLB^{+}19, DDSP17, EMBH14, ESCAM17, GWKFC19, GWWW19, GZP^{+}13, Guo13, HK17, HM19, HFCR13, HBT^{+}19, HS11, IRC^{+}17, JLGM19, JKA^{+}18, JC14, JLK16, Kal19, KPS16, KSG17, LH11, LLL15, LNK16, LML^{+}17, LSKF12, LPL18, LWZ^{+}17, LMNM^{+}19, LGS14, MMLN15, MCN10, MLK14, MLKC17, MWAG16, MSVA16, MSA11, MSRF18, NADZ18, NSBM17, NNNQ18, NQZ18, NLC^{+}19, ORLS14, PMB^{+}14, PHMN16, RW13, RFB17, RR17, RBJ18, RSLH19, hRKRR13, SCFB15, SH15, SKKP12, SHGH^{+}15, SSVA19, SPK17, TP18, TFPC10, TdHRZ17, UB12, UFF12, VMA16, VMS18, VDK18, VSB^{+}19, WLCC17, YYL^{+}15, ZQY^{+}19, Zin15]. Utilisation [CCST15]. utility [GL18, KKS13, LMLL19]. utility-based [KKS13].

Validation [FK13, Wor10, HW18, PCP^{+}17, SBK12, WRTA17]. validations [MJBKZ15]. value [KCJB12, SH15]. valued [MSVA16]. variance [ABR16]. variants [AQR^{+}17, MSA11, NT18, ZQY^{+}19]. vault [ZBB17]. vector [ANAT15, MC18, SRU18, VDK18]. Vectors [CH14]. vehicle [PDB14, WZW16]. vehicle-to-grid [WZW16]. vehicles [PDB14]. vehicular [SRU18]. ventured [RR14]. verifiable [CZZ^{+}19, JFR13]. verification [BJ10, CHK^{+}19, CDS^{+}19, GFKS15, HFCR13, HYLY19, JCPACRV12, KSG17, LKPK18, QHW^{+}19, SLB^{+}13, YWS17]. verified [HLB19]. verifiers [LAHV17]. verifying [DPP^{+}19]. version [AKFM18, BCD^{+}17]. very [ASO19]. via [CPGA13, CSS^{+}19, DRP^{+}12, EHBH^{+}16, GL18, GBCD19, GSZX18, GSL^{+}19, GB19, JCD^{+}15, KGSN17, KKKP18, MJW^{+}18, MGYC18, SPYG15, TLZL16, TXM^{+}19]. vibrations [dSFK16]. victim [AFF14, BLS^{+}16]. victimization [LZBS13]. video [DC16, HZH^{+}18, LK10, YCJ12]. view [LMLK16, SB14]. violation [BWOD13, CLL^{+}13, SE11]. virtual [KTG^{+}18, LPL18, NADZ18, PDM19, RG19, TBSR14]. virtual-machine-based [KTG^{+}18]. Virtualization [ADWM18, HW12, LLX^{+}17, WHSE15]. virtualization-based [LLX^{+}17]. virtualized [BR17, HNSD18]. Virtuous [Hat19]. virulence [CLLK10, RMG11]. virus [Ano14s, KYZ15, MSR19]. Visibility [RB12]. Visual [TWNC18, BLR17, GJJ18, WRTA17]. Visualization [SLH14, ATAAd18, NQZ18, dABM^{+}18]. visualizations [SLG19]. vmOS [LLX^{+}17]. voice [CDE^{+}11, VMS18]. VoIP [CDE^{+}11, AMS18, SG10b, SG13, VGL14]. VoIP/IMS [VGL14]. Volatile [NLC^{+}19]. Volume [GZ20, GZG20, HZG20]. voluntary [AM18b]. voting [ABC^{+}19, CZZ^{+}19, Gar16, HK13, JCPACRV12, JFR13, KNM^{+}17, LLHT13]. VQAP [TWNC18]. vs [BCCS18, SM19a, VMTG15]. Vulnerabilities [Zin15, ABC^{+}19, GSA18, KL18, RNR15, RHL15, SBK12, WP10]. Vulnerability [HA15, MCG19, Hol12, HLFZ16, JGLE16, WJAM11]. vulnerable [SWZL14].

warehousing [SSVA19]. Warfare [Cae12, NWD^{+}12, TG12, RJJ15]. wargame [HWS^{+}13]. WARP [Su10]. watchlist [YST^{+}19]. watermarking [BWR12, KMG17]. WAVE [CGPP12, BGR^{+}17, FRRP17]. WAVE-CUSUM [CGPP12]. wavelet [CGPP11, CGPP12]. way [IHED15, KvS10, ROW19]. WBANs [OSP^{+}19]. weakest [HL18]. Web [TG15, AMK18, CFH^{+}11, EHBH^{+}16, GOGG19, LHI^{+}19, LGS14, NMX19, PAJAP15, RAA^{+}14, SGLC16, SM19b, SBK12, ZY14, AAL16, GMGTdFR14b, GPAD^{+}14, HLFZ16, NDSA17, SM19a, VMTG15]. web-based [CFH^{+}11, GMGTdFR14b]. webpages [DLLS19, GK14, KYZ15]. website [CCST15]. websites [ZDW19]. weight [KP18]. weighted [ASO19, LLHT13]. Wheeler [Kül12]. while [ICRB14, JL10]. whitelist [hRKRR13]. whitelist-based [hRKRR13]. WhiteScript [SS11]. who [FT16, MBAE^{+}11]. wi [NADZ18]. wi-fi [NADZ18]. wild [UCCJ18, MTC^{+}18, SYX^{+}16]. wildcards [DA18]. window [YWYZ12]. Windows [OP11, SCB^{+}13, SS17b, SS18b]. wireless [hBCK17, BEB^{+}18, CVS18, HR13, HS13, HSPX14, HXS^{+}16, HT17, JLGM19, JWZX19, LNNH13, NADZ18, OAO17, RR17, SMK18b, TTCB18, TOZJ13, VDK18]. WISA [YRA15]. WISE [Ano11-30]. Wishful [Bay12c]. withholding [DWF^{+}19]. within [AAD^{+}12, KLN15, RMG11, Rru16, SBE17, vDBD13]. without [BGLR17, CC14, DFC13, HNSD18, Par18, PW13]. word [SCR19]. word-independent [SCR19]. work [HPJ^{+}19, MOV14]. Workbench [AKP14]. Workload [FHF15]. workplace [Guo13]. world [CP18, HSL19]. Worm [CLLK10]. wormhole [Su10, TP18]. wormhole-avoidance [Su10]. Worms [VG14, PWWY14, WZW^{+}10]. WorSE [AKP14]. worst [AS13]. WPA2 [BMO18]. wrapper [KK17]. write [DFJ^{+}13]. WSN [ALRC12].

X.509 [DAR14]. X.509-based [DAR14]. x86 [ZLW^{+}15]. XACML [DLDTGMMP15, NDdL15, TdHRZ17]. XBOX [LW12d]. XSS [ZW19].

year [BB11, CKA^{+}17].

Z [BGR^{+}17, FRRP17, ZLL^{+}19]. Z-Channel [ZLL^{+}19]. Z-Wave [BGR^{+}17, FRRP17]. Zero [VG14, MDHM18, TCNFW16]. Zero-day [VG14, TCNFW16]. zero-knowledge [MDHM18]. Zerocash [ZLL^{+}19]. zone [ZTL^{+}19].

References

Alomar:2019:UPU

[AAA19] Noura Alomar, Mansour Alsaleh, and Abdulrahman Alarifi. Uncovering the predictors of unsafe computing behaviors in online crowdsourcing contexts. Computers & Security, 85(??):300–312, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

                https://www.sciencedirect.com/science/article/pii/
                S0167404818302165.

Ali:2019:BBP

[AAC^{+}19] Gauhar Ali, Naveed Ahmad, Yue Cao, Muhammad Asif, Haitham Cruickshank, and Qazi Ejaz Ali. Blockchain based permission delegation and access control in Internet of Things (BACI). Computers & Security, 86(??):318–334, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

                1872-6208 (electronic). URL https://www.sciencedirect.
                com/science/article/pii/S0167404819301208.

Abramov:2012:MIA

[AAD^{+}12] Jenny Abramov, Omer Anson, Michal Dahan, Peretz Shoval, and Arnon Sturm. A methodology for integrating access control policies within database development. Computers & Security, 31(3):299–314, May 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000089.

Anisetti:2018:MTP

[AAD^{+}18] M. Anisetti, C. A. Ardagna, E. Damiani, N. El Ioini, and F. Gaudenzi. Modeling time, probability, and configuration constraints for continuous cloud service certification. Computers & Security, 72(??):234–254, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302018.

Alshawish:2019:QPM

[AAdM19] Ali Alshawish, Mohamed Amine Abid, and Hermann de Meer. Quasi-purification of mixed game strategies: Sub-optimality of equilibria in security games. Computers & Security, 87(??):Article 101575, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300458.

Alali:2018:IRA

[AAH^{+}18] Mansour Alali, Ahmad Almogren, Mohammad Mehedi Hassan, Iehab A. L. Rassan, and Md Zakirul Alam Bhuiyan. Improving risk assessment model of cyber security using fuzzy logic inference system. Computers & Security, 74(??):323– 339, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302006.

Al-Assam:2012:SEB

[AAJ12] Hisham Al-Assam and Sabah Jassim. Security evaluation of biometric keys. Computers & Security, 31(2):151–163, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000065.

Asghar:2016:DIR

[AAL16] Hira Asghar, Zahid Anwar, and Khalid Latif. A deliberately insecure RDF-based Semantic Web application framework for teaching SPARQL/SPARUL injection attacks and defense mechanisms. Computers & Security, 58(??):63–82, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815001807.

Amigud:2018:IRS

[AAMDGR18] Alexander Amigud, Joan Arnedo-Moreno, Thanasis Daradoumis, and Ana-Elena Guerrero-Roldan. An integrative review of security and integrity strategies in an academic environment: Current understanding and emerging perspectives. Computers & Security, 76(??):50–70, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301792.

Anthi:2018:ESA

[AAR^{+}18] Eirini Anthi, Shazaib Ahmad, Omer Rana, George Theodorakopoulos, and Pete Burnap. EclipseIoT: a secure and adaptive hub for the Internet of Things. Computers & Security, 78(??):477–490, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818309052.

Abazari:2016:EAM

[AAT16] Farzaneh Abazari, Morteza Analoui, and Hassan Takabi. Effect of anti-malware software on infectious nodes in cloud environment. Computers & Security, 58(??):139–148, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815001959.

Aranha:2019:RSV

[ABC^{+}19] Diego F. Aranha, Pedro Y. S. Barbosa, Thiago N. C. Cardoso, Caio Lüders Araújo, and Paulo Matias. The return of software vulnerabilities in the Brazilian voting machine. Computers & Security, 86(??):335–349, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301191.

Azad:2018:MRR

[ABHS18] Muhammad Ajmal Azad, Samiran Bag, Feng Hao, and Khaled Salah. M2M-REP: Reputation system for machines in the Internet of Things. Computers & Security, 79(??):1–16, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308666.

Alexiou:2016:FSA

[ABP16] Nikolaos Alexiou, Stylianos Basagiannis, and Sophia Petridou. Formal security analysis of near field communication using model checking. Computers & Security, 60(??):1–14, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300244.

Atighehchi:2015:NME

[ABR15] Kevin Atighehchi, Alexis Bonnecaze, and Gabriel Risterucci. New models for efficient authenticated dictionaries. Computers & Security, 53(??):203–214, September 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000607.

Awan:2016:ICR

[ABR16] Malik Shahzad Kaleem Awan, Pete Burnap, and Omer Rana. Identifying cyber risk hotspots: a framework for measuring temporal variance in computer network risk. Computers & Security, 57(??):31–46, March 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001674.

Angin:2018:SPA

[ABR18] Pelin Angin, Bharat Bhargava, and Rohit Ranchal. A selfprotecting agents based model for high-performance mobilecloud computing. Computers & Security, 77(??):380–396, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818303948.

Ahmad:2014:POC

[ABS14] Atif Ahmad, Rachelle Bosua, and Rens Scheepers. Protecting organizational competitive advantage: a knowledge leakage perspective. Computers & Security, 42(??):27–39, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000054.

Armando:2013:AFB

[ACC^{+}13] Alessandro Armando, Roberto Carbone, Luca Compagna, Jorge Cuéllar, Giancarlo Pellegrino, and Alessandro Sorniotti. An authentication flaw in browser-based single sign-on protocols: Impact and remediations. Computers & Security, 33(??):41–58, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001356.

Amato:2018:RHB

[ACS^{+}18] Flora Amato, Aniello Castiglione, Aniello De Santo, Vincenzo Moscato, Antonio Picariello, Fabio Persia, and Giancarlo Sperlí. Recognizing human behaviours in online social networks. Computers & Security, 74(??):355–370, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301219.

Abraham:2019:EEL

[ACS19] Sherly Abraham and InduShobha Chengalur-Smith. Evaluating the effectiveness of learner controlled information security training. Computers & Security, 87(??):Article 101586, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818308873.

Ahmad:2018:PID

[ADA^{+}18] Asma’a Ahmad, Robin Doss, Majeed Alajeely, Sarab F. Al Rubeaai, and Dua’a Ahmad. Packet integrity defense mechanism in OppNets. Computers & Security, 74(??):71–93, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818300269.

Alajeely:2017:DAP

[ADAMH17] Majeed Alajeely, Robin Doss, Asma’a Ahmad, and Vicky Mak-Hau. Defense against packet collusion attacks in opportunistic networks. Computers & Security, 65(??):269–282, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301717.

Ardagna:2010:ACS

[ADF^{+}10] Claudio A. Ardagna, Sabrina De Capitani di Vimercati, Sara Foresti, Tyrone W. Grandison, Sushil Jajodia, and Pierangela Samarati. Access control for smarter healthcare using policy spaces. Computers & Security, 29(8):848–858, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000623.

Alves:2018:VST

[ADWM18] Thiago Alves, Rishabh Das, Aaron Werth, and Thomas Morris. Virtualization of SCADA testbeds for cybersecurity research: a modular approach. Computers & Security, 77(??):531–546, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304905.

Ahmadinejad:2014:UDI

[AF14] Seyed Hossein Ahmadinejad and Philip W. L. Fong. Unintended disclosure of information: Inference attacks by third-party extensions to social network systems. Computers & Security, 44(??):75–91, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481400056X.

Allam:2014:SIS

[AFF14] Sean Allam, Stephen V. Flowerday, and Ethan Flowerday. Smartphone information security awareness: a victim of operational pressures. Computers & Security, 42(??):56–65, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000169.

Archibald:2014:CAD

[AG14] Rennie Archibald and Dipak Ghosal. A comparative analysis of detection metrics for covert timing channels. Computers & Security, 45(??):284–292, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000467.

Abouhogail:2018:NSP

[AG18] Reham Abdellatif Abouhogail and Mohammed S. Gadelrab. A new secure and privacy preserved protocol for IEEE802.11s networks. Computers & Security, 77(??):745–755, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301992.

Ayuso:2012:FFC

[AGL12] Pablo Neira Ayuso, Rafael M. Gasca, and Laurent Lefevre. FT-FW: a cluster-based fault-tolerant architecture for stateful firewalls. Computers & Security, 31(4):524–539, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000156.

Ali-Gombe:2018:TMD

[AGSR^{+}18] Aisha I. Ali-Gombe, Brendan Saltaformaggio, J. “Ram” Ramanujam, Dongyan Xu, and Golden G. Richard. Toward a more dependable hybrid analysis of Android malware using aspect-oriented programming. Computers & Security, 73(??):235–248, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302377.

Albrechtsen:2010:IIS

[AH10] Eirik Albrechtsen and Jan Hovden. Improving information security awareness and behaviour through dialogue, participation and collective reflection. An intervention study. Computers & Security, 29(4):432–445, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001436.

Abramov:2013:TAS

[AH13] Raz Abramov and Amir Herzberg. TCP Ack storm DoS attacks. Computers & Security, 33(??):12–27, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001411.

Altabash:2018:ITD

[AH18] Kholood Al tabash and Jassim Happa. Insider-threat detection using Gaussian mixture models and sensitivity profiles. Computers & Security, 77(??):838–859, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302487.

Axon:2019:HAN

[AHGC19] Louise Axon, Jassim Happa, Michael Goldsmith, and Sadie Creese. Hearing attacks in network data: an effectiveness study. Computers & Security, 83(??):367–388, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303377.

Adi:2018:ASP

[AHP18] Kamel Adi, Lamia Hamza, and Liviu Pene. Automatic security policy enforcement in computer systems. Computers & Security, 73(??):156–171, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302286.

Ahmad:2012:IRT

[AHR12] Atif Ahmad, Justin Hadgkiss, and A. B. Ruighaver. Incident response teams — challenges in supporting the organisational security function. Computers & Security, 31(5):643–652, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000624.

Alam:2015:FMM

[AHTS15] Shahid Alam, R. Nigel Horspool, Issa Traore, and Ibrahim Sogukpinar. A framework for metamorphic malware analysis and real-time detection. Computers & Security, 48(??):212–233, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001576.

Allahbakhsh:2014:RQU

[AIB^{+}14] Mohammad Allahbakhsh, Aleksandar Ignjatovic, Boualem Benatallah, Seyed-Mehdi-Reza Beheshti, Norman Foo, and Elisa Bertino. Representation and querying of unfair evaluations in social rating systems. Computers & Security, 41(??):68–88, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001375.

Asghar:2013:EES

[AIRC13] Muhammad Rizwan Asghar, Mihaela Ion, Giovanni Russello, and Bruno Crispo. ESPOON_{ERBAC} : Enforcing security policies in outsourced environments. Computers & Security, 35(??):2–24, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001824.

Ahmed:2019:OPS

[AKAuR19] Idrees Ahmed, Abid Khan, Mansoor Ahmed, and Saif ur Rehman. Order preserving secure provenance scheme for distributed networks. Computers & Security, 82(??):99–117, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818306667.

Abdo:2018:SSR

[AKFM18] H. Abdo, M. Kaouk, J.-M. Flaus, and F. Masse. A safety/ security risk analysis approach of Industrial Control Systems: a cyber bowtie — combining new version of attack tree with bowtie analysis. Computers & Security, 72(??):175–195, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301931.

Adamsky:2014:SBB

[AKJR14] Florian Adamsky, Syed Ali Khayam, Rudolf Jäger, and Muttukrishnan Rajarajan. Stealing bandwidth from BitTorrent seeders. Computers & Security, 46(??):126–140, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001205.

Anagnostopoulos:2013:DAA

[AKK^{+}13] Marios Anagnostopoulos, Georgios Kambourakis, Panagiotis Kopanos, Georgios Louloudakis, and Stefanos Gritzalis. DNS amplification attack revisited. Computers & Security, 39 (part B)(??):475–485, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001405.

Amthor:2014:WWM

[AKP14] Peter Amthor, Winfried E. Kühnhauser, and Anja Pölck. WorSE: a workbench for model-based security engineering. Computers & Security, 42(??):40–55, May 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000066.

Alcaraz:2017:RIC

[ALC17] Cristina Alcaraz, Javier Lopez, and Kim-Kwang Raymond Choo. Resilient interconnection in cyber-physical control systems. Computers & Security, 71(??):2–14, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300573.

Alizadeh:2018:LDP

[ALF^{+}18] Mahdi Alizadeh, Xixi Lu, Dirk Fahland, Nicola Zannone, and Wil M. P. van der Aalst. Linking data and process perspectives for conformance analysis. Computers & Security, 73(??):172–193, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302262.

Alcaraz:2012:SKM

[ALRC12] Cristina Alcaraz, Javier Lopez, Rodrigo Roman, and Hsiao-Hwa Chen. Selecting key management schemes for WSN applications. Computers & Security, 31(8):956–966, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001034.

Arnellos:2011:FDS

[ALZ^{+}11] Argyris Arnellos, Dimitrios Lekkas, Dimitrios Zissis, Thomas Spyrou, and John Darzentas. Fair digital signing: the structural reliability of signed documents. Computers & Security, 30(8):580–596, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481100112X.

Azad:2013:CRD

[AM13] Muhammad Ajmal Azad and Ricardo Morla. Caller-REP: Detecting unwanted calls with caller social strength. Computers & Security, 39 (part B)(??):219–236, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001090.

Ashibani:2017:CPS

[AM17a] Yosef Ashibani and Qusay H. Mahmoud. Cyber physical systems security: Analysis, challenges and solutions. Computers & Security, 68(??):81–97, July 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300809.

Aurigemma:2017:PPE

[AM17b] Salvatore Aurigemma and Thomas Mattson. Privilege or procedure: Evaluating the effect of employee status on intent to comply with socially interactive information security threats and controls. Computers & Security, 66(??):218–234, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300329.

Alves:2018:OIC

[AM18a] Thiago Alves and Thomas Morris. OpenPLC: an IEC 61,131-3 compliant open source industrial controller for cyber security research. Computers & Security, 78(??):364–379, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818305388.

Aurigemma:2018:EEU

[AM18b] Salvatore Aurigemma and Thomas Mattson. Exploring the effect of uncertainty avoidance on taking voluntary protective security actions. Computers & Security, 73(??):219–234, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302328.

Abdelhameed:2018:PPT

[AMK18] Saad A. Abdelhameed, Sherin M. Moussa, and Mohamed E. Khalifa. Privacy-preserving tabular data publishing: a comprehensive evaluation from web to cloud. Computers & Security, 72(??):74–95, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301840.

Armando:2013:BFA

[AMMV13] Alessandro Armando, Alessio Merlo, Mauro Migliardi, and Luca Verderame. Breaking and fixing the Android launching flow. Computers & Security, 39 (Part A)(??):104–115, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000540.

Azad:2018:SMS

[AMS18] Muhammad Ajmal Azad, Ricardo Morla, and Khaled Salah. Systems and methods for SPIT detection in VoIP: Survey and future directions. Computers & Security, 77(??):1–20, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818302414.

Abril:2015:SMA

[ANAT15] Daniel Abril, Guillermo Navarro-Arribas, and Vicenç Torra. Spherical microaggregation: Anonymizing sparse vector spaces. Computers & Security, 49(??):28–44, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001679.

Akinrolabu:2019:CRA

[ANMN19] Olusola Akinrolabu, Jason R. C. Nurse, Andrew Martin, and Steve New. Cyber risk assessment in cloud provider environments: Current models and future needs. Computers & Security, 87(??):Article 101600, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301543.

Anonymous:2010:CPa

[Ano10a] Anonymous. Call for papers. Computers & Security, 29(1):163–164, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001369.

Anonymous:2010:CPb

[Ano10b] Anonymous. Call for papers. Computers & Security, 29 (2):296, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404809001618.

Anonymous:2010:CPc

[Ano10c] Anonymous. Call for papers. Computers & Security, 29 (2):297, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404810000039.

Anonymous:2010:Ca

[Ano10d] Anonymous. Contents. Computers & Security, 29(1):iii, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740480900131X.

Anonymous:2010:Cb

[Ano10e] Anonymous. Contents. Computers & Security, 29(2):iii, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809001552.

Anonymous:2010:Cc

[Ano10f] Anonymous. Contents. Computers & Security, 29(3):iv, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481000012X.

Anonymous:2010:Cd

[Ano10g] Anonymous. Contents. Computers & Security, 29(4):iii, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404810000271.

Anonymous:2010:Ce

[Ano10h] Anonymous. Contents. Computers & Security, 29(5):iv, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481000043X.

Anonymous:2010:Cf

[Ano10i] Anonymous. Contents. Computers & Security, 29(6):iii, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000568.

Anonymous:2010:Cg

[Ano10j] Anonymous. Contents. Computers & Security, 29(7):iii, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000672.

Anonymous:2010:Ch

[Ano10k] Anonymous. Contents. Computers & Security, 29(8):iii, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000969.

Anonymous:2010:IEBa

[Ano10l] Anonymous. IFC — editorial board. Computers & Security, 29(1):ifc, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480900128X.

Anonymous:2010:IEBb

[Ano10m] Anonymous. IFC — editorial board. Computers & Security, 29(2):ifc, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001527.

Anonymous:2010:IEBc

[Ano10n] Anonymous. IFC — editorial board. Computers & Security, 29(3):ifc, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740481000009X.

Anonymous:2010:IEBd

[Ano10o] Anonymous. IFC — editorial board. Computers & Security, 29(4):ifc, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404810000246.

Anonymous:2010:IEBe

[Ano10p] Anonymous. IFC — editorial board. Computers & Security, 29(5):ifc, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404810000404.

Anonymous:2010:IEBf

[Ano10q] Anonymous. IFC — editorial board. Computers & Security, 29(6):ifc, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000532.

Anonymous:2010:IEBg

[Ano10r] Anonymous. IFC — editorial board. Computers & Security, 29(7):ifc, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000647.

Anonymous:2010:IEBh

[Ano10s] Anonymous. IFC — editorial board. Computers & Security, 29(8):ifc, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000933.

Anonymous:2010:ITAa

[Ano10t] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 29(1):158, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001357.

Anonymous:2010:ITAb

[Ano10u] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 29(2):291, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001606.

Anonymous:2010:ITAc

[Ano10v] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 29(3):380, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000131.

Anonymous:2010:ITAd

[Ano10w] Anonymous. IFIP TC11 — aims and scope. Computers & Security, 29(4):515, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000283.

Anonymous:2010:ITAe

[Ano10x] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 29(5):636–640, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000441.

Anonymous:2010:ITAf

[Ano10y] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 29(6):724–728, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481000057X.

Anonymous:2010:ITAg

[Ano10z] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 29(7):807–811, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000726.

Anonymous:2010:ITAh

[Ano10-27] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 29(8):903–907, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404810000970.

Anonymous:2010:ITCa

[Ano10-28] Anonymous. IFIP technical committee. Computers & Security, 29(1):159–162, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001321.

Anonymous:2010:ITCb

[Ano10-29] Anonymous. IFIP technical committee. Computers & Security, 29(2):292–295, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001564.

Anonymous:2010:ITCc

[Ano10-30] Anonymous. IFIP technical committee. Computers & Security, 29(3):i–iv, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000143.

Anonymous:2010:ITCd

[Ano10-31] Anonymous. IFIP technical committee. Computers & Security, 29(4):516–519, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000295.

Anonymous:2010:IIS

[Ano10-32] Anonymous. Invitation to IFIP sec 2011. Computers & Security, 29(8):908, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481000101X.

Anonymous:2010:PF

[Ano10-33] Anonymous. Pages 1–164 (February 2010). Computers & Security, 29(1):??, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2010:PM

[Ano10-34] Anonymous. Pages 165–298 (March 2010). Computers & Security, 29(2):??, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2010:PJ

[Ano10-35] Anonymous. Pages 381–520 (June 2010). Computers & Security, 29(4):??, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2010:PS

[Ano10-36] Anonymous. Pages 641–728 (September 2010). Computers & Security, 29(6):??, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2010:PO

[Ano10-37] Anonymous. Pages 729–812 (October 2010). Computers & Security, 29(7):??, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2010:PN

[Ano10-38] Anonymous. Pages 813–908 (November 2010). Computers & Security, 29(8):??, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2011:CMa

[Ano11a] Anonymous. Call for membership. Computers & Security, 30(1):87, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000113.

Anonymous:2011:CMb

[Ano11b] Anonymous. Call for membership. Computers & Security, 30(4):293, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404811000605.

Anonymous:2011:Ca

[Ano11c] Anonymous. Contents. Computers & Security, 30(1):iii, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000216.

Anonymous:2011:Cb

[Ano11d] Anonymous. Contents. Computers & Security, 30(2–3):iii, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000344.

Anonymous:2011:Cc

[Ano11e] Anonymous. Contents. Computers & Security, 30(2–3):iii, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000344.

Anonymous:2011:Cd

[Ano11f] Anonymous. Contents. Computers & Security, 30(2–3):iii, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000344.

Anonymous:2011:Ce

[Ano11g] Anonymous. Contents. Computers & Security, 30(4):iii, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000563.

Anonymous:2011:Cf

[Ano11h] Anonymous. Contents. Computers & Security, 30(5):iii, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000782.

Anonymous:2011:Cg

[Ano11i] Anonymous. Contents. Computers & Security, 30 (6–7):iii, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000903.

Anonymous:2011:Ch

[Ano11j] Anonymous. Contents. Computers & Security, 30 (6–7):iii, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000903.

Anonymous:2011:Ci

[Ano11k] Anonymous. Contents. Computers & Security, 30 (6–7):iii, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000903.

Anonymous:2011:Cj

[Ano11l] Anonymous. Contents. Computers & Security, 30(8):iii–iv, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811001271.

Anonymous:2011:HCP

[Ano11m] Anonymous. HAISA — call for papers. Computers & Security, 30(2–3):170, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000393.

Anonymous:2011:IEBa

[Ano11n] Anonymous. IFC — editorial board. Computers & Security, 30(1):ifc, January 2011. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000186.

Anonymous:2011:IEBb

[Ano11o] Anonymous. IFC — editorial board. Computers & Security, 30(2–3):ifc, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000319.

Anonymous:2011:IEBc

[Ano11p] Anonymous. IFC — editorial board. Computers & Security, 30(4):ifc, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404811000538.

Anonymous:2011:IEBd

[Ano11q] Anonymous. IFC — editorial board. Computers & Security, 30(5):ifc, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404811000757.

Anonymous:2011:IEBe

[Ano11r] Anonymous. IFC — editorial board. Computers & Security, 30(6–7):ifc, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000873.

Anonymous:2011:IEBf

[Ano11s] Anonymous. IFC — editorial board. Computers & Security, 30(8):ifc, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001246.

Anonymous:2011:ITAa

[Ano11t] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 30(1):81–85, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000071.

Anonymous:2011:ITAb

[Ano11u] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 30(2–3):165–169, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000356.

Anonymous:2011:ITAc

[Ano11v] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 30(4):288–292, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000575.

Anonymous:2011:ITAd

[Ano11w] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 30(5):346–350, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000794.

Anonymous:2011:ITAe

[Ano11x] Anonymous. IFIP TC11 — aims, scope and technical committee. Computers & Security, 30(6–7):547–551, September/ October 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000915.

Anonymous:2011:ICC

[Ano11y] Anonymous. IFIPTM 2011 — call for contributions. Computers & Security, 30(1):86, January 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000101.

Anonymous:2011:PJa

[Ano11z] Anonymous. Pages 1–88 (January 2011). Computers & Security, 30(1):??, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2011:PJb

[Ano11-27] Anonymous. Pages 171–294 (June 2011). Computers & Security, 30(4):??, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2011:PSO

[Ano11-28] Anonymous. Pages 351–552 (September–October 2011). Computers & Security, 30(6–7):??, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2011:PN

[Ano11-29] Anonymous. Pages 553–840 (November 2011). Computers & Security, 30(8):??, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2011:WCP

[Ano11-30] Anonymous. WISE 7 — call for papers. Computers & Security, 30(1):88, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000125.

Anonymous:2012:A

[Ano12a] Anonymous. Announcement. Computers & Security, 31(8):ii, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001575.

Anonymous:2012:CP

[Ano12b] Anonymous. Call for papers. Computers & Security, 31(8):iii–iv, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001587.

Anonymous:2012:Ca

[Ano12c] Anonymous. Contents. Computers & Security, 31(1):iii, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481100157X.

Anonymous:2012:Cb

[Ano12d] Anonymous. Contents. Computers & Security, 31(2):iii, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000302.

Anonymous:2012:Cc

[Ano12e] Anonymous. Contents. Computers & Security, 31(3):iii, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481200048X.

Anonymous:2012:Cd

[Ano12f] Anonymous. Contents. Computers & Security, 31(4):iii–iv, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000739.

Anonymous:2012:Ce

[Ano12g] Anonymous. Contents. Computers & Security, 31(5):iii, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000880.

Anonymous:2012:Cf

[Ano12h] Anonymous. Contents. Computers & Security, 31(6):iii, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001113.

Anonymous:2012:Cg

[Ano12i] Anonymous. Contents. Computers & Security, 31(7):iii, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001290.

Anonymous:2012:Ch

[Ano12j] Anonymous. Contents. Computers & Security, 31(8):iii, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481200154X.

Anonymous:2012:Ea

[Ano12k] Anonymous. Editorial. Computers & Security, 31(3):271– 272, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000570.

Anonymous:2012:Eb

[Ano12l] Anonymous. Editorial. Computers & Security, 31(4):375– 376, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000685.

Anonymous:2012:EBa

[Ano12m] Anonymous. Editorial Board. Computers & Security, 31(4): ifc, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000703.

Anonymous:2012:EBb

[Ano12n] Anonymous. Editorial Board. Computers & Security, 31(5): ifc, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000855.

Anonymous:2012:EBc

[Ano12o] Anonymous. Editorial Board. Computers & Security, 31(6):ifc, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001083.

Anonymous:2012:EBd

[Ano12p] Anonymous. Editorial Board. Computers & Security, 31 (7):ifc, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404812001265.

Anonymous:2012:EBe

[Ano12q] Anonymous. Editorial Board. Computers & Security, 31(8):ifc, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001514.

Anonymous:2012:ECS

[Ano12r] Anonymous. Editorial computers and security special issue IFIP/SEC 2010 “Security and Privacy — Silver Linings in the Cloud”. Computers & Security, 31(7):805, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481200137X.

Anonymous:2012:FAI

[Ano12s] Anonymous. Filler AD:IFIP. Computers & Security, 31(4): i, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000764.

Anonymous:2012:IEBa

[Ano12t] Anonymous. IFC — editorial board. Computers & Security, 31(1):ifc, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001544.

Anonymous:2012:IEBb

[Ano12u] Anonymous. IFC — editorial board. Computers & Security, 31(2):ifc, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000272.

Anonymous:2012:IEBc

[Ano12v] Anonymous. IFC — editorial board. Computers & Security, 31(3):ifc, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404812000454.

Anonymous:2012:PF

[Ano12w] Anonymous. Pages 1–148 (February 2012). Computers & Security, 31(1):??, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PMa

[Ano12x] Anonymous. Pages 149–270 (March 2012). Computers & Security, 31(2):??, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PMb

[Ano12y] Anonymous. Pages 271–374 (May 2012). Computers & Security, 31(3):??, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PJa

[Ano12z] Anonymous. Pages 375–640 (June 2012). Computers & Security, 31(4):??, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PJb

[Ano12-27] Anonymous. Pages 641–738 (July 2012). Computers & Security, 31(5):??, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PS

[Ano12-28] Anonymous. Pages 739–804 (September 2012). Computers & Security, 31(6):??, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2012:PN

[Ano12-29] Anonymous. Pages 857–1014 (November 2012). Computers & Security, 31(8):??, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:Ca

[Ano13a] Anonymous. Contents. Computers & Security, 32(??):iii–iv, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001873.

Anonymous:2013:Cb

[Ano13b] Anonymous. Contents. Computers & Security, 33(??):iii, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000114.

Anonymous:2013:Cc

[Ano13c] Anonymous. Contents. Computers & Security, 34(??):iii, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000412.

Anonymous:2013:Cd

[Ano13d] Anonymous. Contents. Computers & Security, 35(??):iii, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000643.

Anonymous:2013:Ce

[Ano13e] Anonymous. Contents. Computers & Security, 36(??):iii, July 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000734.

Anonymous:2013:Cf

[Ano13f] Anonymous. Contents. Computers & Security, 37(??):iii, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001041.

Anonymous:2013:Cg

[Ano13g] Anonymous. Contents. Computers & Security, 38(??):iii, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481300117X.

Anonymous:2013:Ch

[Ano13h] Anonymous. Contents. Computers & Security, 39 (Part A)(??):iii, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001508.

Anonymous:2013:Ci

[Ano13i] Anonymous. Contents. Computers & Security, 39 (part B)(??):iii–iv, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001582.

Anonymous:2013:Cj

[Ano13j] Anonymous. Contents. Computers & Security, 39 (part B)(??):iii–iv, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001582.

Anonymous:2013:Ea

[Ano13k] Anonymous. Editorial. Computers & Security, 33(??):1–2, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000072.

Anonymous:2013:Ec

[Ano13l] Anonymous. Editorial. Computers & Security, 35(??):1, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000369.

Anonymous:2013:Ed

[Ano13m] Anonymous. Editorial. Computers & Security, 37(??):iv–v, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001053.

Anonymous:2013:Ee

[Ano13n] Anonymous. Editorial. Computers & Security, 39 (Part A)(??):1, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001454.

Anonymous:2013:EBa

[Ano13o] Anonymous. Editorial Board. Computers & Security, 32(??):ifc, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001848.

Anonymous:2013:EBb

[Ano13p] Anonymous. Editorial Board. Computers & Security, 33(??): ifc, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000308.

Anonymous:2013:EBc

[Ano13q] Anonymous. Editorial Board. Computers & Security, 34(??): ifc, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000382.

Anonymous:2013:EBd

[Ano13r] Anonymous. Editorial Board. Computers & Security, 35(??): ifc, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000618.

Anonymous:2013:EBe

[Ano13s] Anonymous. Editorial Board. Computers & Security, 37(??):ifc, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000989.

Anonymous:2013:EBf

[Ano13t] Anonymous. Editorial Board. Computers & Security, 38 (??):ifc, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404813001144.

Anonymous:2013:EBg

[Ano13u] Anonymous. Editorial Board. Computers & Security, 39 (Part A)(??):ifc, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001478.

Anonymous:2013:EBh

[Ano13v] Anonymous. Editorial Board. Computers & Security, 39 (part B)(??):ifc, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001557.

Anonymous:2013:EBi

[Ano13w] Anonymous. Editorial Board. Computers & Security, 39 (part B)(??):ifc, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001557.

Anonymous:2013:IEB

[Ano13x] Anonymous. IFC — editorial board. Computers & Security, 36(??):ifc, July 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000709.

Anonymous:2013:PM

[Ano13y] Anonymous. Pages 1–140 (May 2013). Computers & Security, 34(??):1–140, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:PS

[Ano13z] Anonymous. Pages 1–228 (September 2013). Computers & Security, 37(??):1–228, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:PF

[Ano13-27] Anonymous. Pages 1–270 (February 2013). Computers & Security, 32(??):1–270, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:PJ

[Ano13-28] Anonymous. Pages 1–68 (July 2013). Computers & Security, 36(??):1–68, July 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:PN

[Ano13-29] Anonymous. Pages 117–502 (November 2013). Computers & Security, 39 (part B)(??):117–502, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2013:SIT

[Ano13-30] Anonymous. Special issue on trust in cyber, physical and social computing: Call for papers. Computers & Security, 34(??):iii, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000448.

Anonymous:2014:Ca

[Ano14a] Anonymous. Contents. Computers & Security, 40(??):iii, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001752.

Anonymous:2014:Cb

[Ano14b] Anonymous. Contents. Computers & Security, 41(??):iii, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000121.

Anonymous:2014:Cc

[Ano14c] Anonymous. Contents. Computers & Security, 42(??):iii, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481400039X.

Anonymous:2014:Cd

[Ano14d] Anonymous. Contents. Computers & Security, 43(??):iii, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000674.

Anonymous:2014:Ce

[Ano14e] Anonymous. Contents. Computers & Security, 44(??):iii, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000777.

Anonymous:2014:Cf

[Ano14f] Anonymous. Contents. Computers & Security, 45(??):iii–iv, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001126.

Anonymous:2014:Cg

[Ano14g] Anonymous. Contents. Computers & Security, 46(??):iii, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001278.

Anonymous:2014:Ch

[Ano14h] Anonymous. Contents. Computers & Security, 47(??):iii, November 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001412.

Anonymous:2014:Ea

[Ano14i] Anonymous. Editorial. Computers & Security, 40(??):iv, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000042.

Anonymous:2014:Ec

[Ano14j] Anonymous. Editorial. Computers & Security, 42(??):iv– v, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481400042X.

Anonymous:2014:EBa

[Ano14k] Anonymous. Editorial Board. Computers & Security, 40(??):ifc, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001727.

Anonymous:2014:EBb

[Ano14l] Anonymous. Editorial Board. Computers & Security, 41(??): ifc, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000091.

Anonymous:2014:EBc

[Ano14m] Anonymous. Editorial Board. Computers & Security, 42(??): ifc, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000364.

Anonymous:2014:EBd

[Ano14n] Anonymous. Editorial Board. Computers & Security, 43(??): ifc, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000649.

Anonymous:2014:EBe

[Ano14o] Anonymous. Editorial Board. Computers & Security, 44(??): ifc, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000741.

Anonymous:2014:EBf

[Ano14p] Anonymous. Editorial Board. Computers & Security, 45(??):ifc, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001096.

Anonymous:2014:EBg

[Ano14q] Anonymous. Editorial Board. Computers & Security, 46 (??):ifc, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404814001242.

Anonymous:2014:EBh

[Ano14r] Anonymous. Editorial Board. Computers & Security, 47(??):ifc, November 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001382.

Anonymous:2014:AVR

[Ano14s]  Anonymous. Is anti-virus really dead?  Computers

& Security, 44(??):iv, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000820.

Anonymous:2014:PF

[Ano14t] Anonymous. Pages 1–114 (February 2014). Computers & Security, 40(??):1–114, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PJb

[Ano14u] Anonymous. Pages 1–118 (July 2014). Computers & Security, 44(??):1–118, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PO

[Ano14v] Anonymous. Pages 1–174 (October 2014). Computers & Security, 46(??):1–174, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PM

[Ano14w] Anonymous. Pages 1–192 (May 2014). Computers & Security, 42(??):1–192, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PJa

[Ano14x] Anonymous. Pages 1–218 (June 2014). Computers & Security, 43(??):1–218, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PS

[Ano14y] Anonymous. Pages 1–306 (September 2014). Computers & Security, 45(??):1–306, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2014:PP

[Ano14z] Anonymous. The passing of a pioneer. Computers & Security, 40(??):v–vi, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001788.

Anonymous:2015:Ca

[Ano15a] Anonymous. Contents. Computers & Security, 48(??):iii–iv, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001758.

Anonymous:2015:Cb

[Ano15b] Anonymous. Contents. Computers & Security, 49(??):iii, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000103.

Anonymous:2015:Cc

[Ano15c] Anonymous. Contents. Computers & Security, 50(??):iii, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000280.

Anonymous:2015:Cd

[Ano15d] Anonymous. Contents. Computers & Security, 51(??):iii, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000498.

Anonymous:2015:Ce

[Ano15e] Anonymous. Contents. Computers & Security, 52(??):iii– iv, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000772.

Anonymous:2015:Cf

[Ano15f] Anonymous. Contents. Computers & Security, 53(??):iii–iv, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001066.

Anonymous:2015:Cg

[Ano15g] Anonymous. Contents. Computers & Security, 54(??):iii, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001352.

Anonymous:2015:Ch

[Ano15h] Anonymous. Contents. Computers & Security, 55(??):iii, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001455.

Anonymous:2015:EIT

[Ano15i] Anonymous. Editorial — 28th IFIP TC-11 International Information Security and Privacy Conference. Computers & Security, 52(??):162–163, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000711.

Anonymous:2015:ESS

[Ano15j] Anonymous. Editorial — SEC’14 special issue on ICT systems security and privacy protection. Computers & Security, 53(??):174, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001157.

Anonymous:2015:EBa

[Ano15k] Anonymous. Editorial Board. Computers & Security, 48(??):ifc, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001722.

Anonymous:2015:EBb

[Ano15l] Anonymous. Editorial Board. Computers & Security, 49(??): ifc, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000073.

Anonymous:2015:EBc

[Ano15m] Anonymous. Editorial Board. Computers & Security, 50(??): ifc, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000255.

Anonymous:2015:EBd

[Ano15n] Anonymous. Editorial Board. Computers & Security, 51(??): ifc, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000462.

Anonymous:2015:EBe

[Ano15o] Anonymous. Editorial Board. Computers & Security, 52(??): ifc, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000747.

Anonymous:2015:EBf

[Ano15p] Anonymous. Editorial Board. Computers & Security, 53(??):ifc, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001030.

Anonymous:2015:IEBa

[Ano15q] Anonymous. IFC — editorial board. Computers & Security, 54(??):ifc, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001327.

Anonymous:2015:IEBb

[Ano15r] Anonymous. IFC — editorial board. Computers & Security, 55(??):CO2, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481500142X.

Anonymous:2015:IA

[Ano15s] Anonymous. Introduction — ARES 2014. Computers & Security, 54(??):1, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001558.

Anonymous:2015:PMb

[Ano15t] Anonymous. Pages 1–106 (May 2015). Computers & Security, 50(??):1–106, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PS

[Ano15u] Anonymous. Pages 1–246 (September 2015). Computers & Security, 53(??):1–246, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PMa

[Ano15v] Anonymous. Pages 1–276 (March 2015). Computers & Security, 49(??):1–276, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PN

[Ano15w] Anonymous. Pages 1–280 (November 2015). Computers & Security, 55(??):1–280, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PJb

[Ano15x] Anonymous. Pages 1–292 (July 2015). Computers & Security, 52(??):1–292, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PF

[Ano15y] Anonymous. Pages 1–298 (February 2015). Computers & Security, 48(??):1–298, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2015:PJa

[Ano15z] Anonymous. Pages 1–62 (June 2015). Computers & Security, 51(??):1–62, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:Ca

[Ano16a] Anonymous. Contents. Computers & Security, 56(??):iii, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001728.

Anonymous:2016:Cb

[Ano16b] Anonymous. Contents. Computers & Security, 57(??):iii, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481500190X.

Anonymous:2016:Cc

[Ano16c] Anonymous. Contents. Computers & Security, 58(??):iii– iv, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300165.

Anonymous:2016:Cd

[Ano16d] Anonymous. Contents. Computers & Security, 59(??):iii, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300396.

Anonymous:2016:Ce

[Ano16e] Anonymous. Contents. Computers & Security, 60(??):iii, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481630058X.

Anonymous:2016:Cf

[Ano16f] Anonymous. Contents. Computers & Security, 61(??):iii, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300815.

Anonymous:2016:Cg

[Ano16g] Anonymous. Contents. Computers & Security, 62(??):iii, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301092.

Anonymous:2016:Ch

[Ano16h] Anonymous. Contents. Computers & Security, 63(??):iii, November 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301316.

Anonymous:2016:IEBa

[Ano16i] Anonymous. IFC — editorial board. Computers & Security, 56(??):CO2, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001698.

Anonymous:2016:IEBb

[Ano16j] Anonymous. IFC — editorial board. Computers & Security, 57(??):CO2, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481500187X.

Anonymous:2016:IEBc

[Ano16k] Anonymous. IFC — editorial board. Computers & Security, 58(??):CO2, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481630013X.

Anonymous:2016:IEBd

[Ano16l] Anonymous. IFC — editorial board. Computers & Security, 59(??):CO2, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300360.

Anonymous:2016:IEBe

[Ano16m] Anonymous. IFC — editorial board. Computers & Security, 60(??):CO2, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300554.

Anonymous:2016:IEBf

[Ano16n] Anonymous. IFC — editorial board. Computers & Security, 61(??):CO2, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300785.

Anonymous:2016:IEBg

[Ano16o] Anonymous. IFC — editorial board. Computers & Security, 62(??):CO2, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301067.

Anonymous:2016:IEBh

[Ano16p] Anonymous. IFC — editorial board. Computers & Security, 63(??):CO2, November 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301286.

Anonymous:2016:PMa

[Ano16q] Anonymous. Pages 1–106 (March 2016). Computers & Security, 57(??):1–106, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PN

[Ano16r] Anonymous. Pages 1–116 (November 2016). Computers & Security, 63(??):1–116, November 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PF

[Ano16s] Anonymous. Pages 1–162 (February 2016). Computers & Security, 56(??):1–162, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PA

[Ano16t] Anonymous. Pages 1–184 (August 2016). Computers & Security, 61(??):1–184, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PJb

[Ano16u] Anonymous. Pages 1–226 (July 2016). Computers & Security, 60(??):1–226, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PJa

[Ano16v] Anonymous. Pages 1–254 (June 2016). Computers & Security, 59(??):1–254, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PMb

[Ano16w] Anonymous. Pages 1–284 (May 2016). Computers & Security, 58(??):1–284, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2016:PS

[Ano16x] Anonymous. Pages 1–348 (September 2016). Computers & Security, 62(??):1–348, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:Ca

[Ano17a] Anonymous. Contents. Computers & Security, 64(??):iii, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301493.

Anonymous:2017:Cb

[Ano17b] Anonymous. Contents. Computers & Security, 65(??):iii–iv, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817300184.

Anonymous:2017:Cc

[Ano17c] Anonymous. Contents. Computers & Security, 66(??):iii, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300457.

Anonymous:2017:Cd

[Ano17d] Anonymous. Contents. Computers & Security, 67(??):iii– iv, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817300767.

Anonymous:2017:Ce

[Ano17e] Anonymous. Contents. Computers & Security, 68(??):iii, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301013.

Anonymous:2017:E

[Ano17f] Anonymous. Editorial. Computers & Security, 71(??):1, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301645.

Anonymous:2017:IEBa

[Ano17g] Anonymous. IFC — editorial board. Computers & Security, 64(??):CO2, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301468.

Anonymous:2017:IEBb

[Ano17h] Anonymous. IFC — editorial board. Computers & Security, 65(??):CO2, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300159.

Anonymous:2017:IEBc

[Ano17i] Anonymous. IFC — editorial board. Computers & Security, 66(??):CO2, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300421.

Anonymous:2017:IEBd

[Ano17j] Anonymous. IFC — editorial board. Computers & Security, 67(??):CO2, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300731.

Anonymous:2017:IEBe

[Ano17k] Anonymous. IFC — editorial board. Computers & Security, 68(??):CO2, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300986.

Anonymous:2017:IFCa

[Ano17l] Anonymous. Inside front cover — editorial board page/ cover image legend if applicable. Computers & Security, 69(??):CO2, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301311.

Anonymous:2017:IFCb

[Ano17m] Anonymous. Inside front cover — editorial board page/ cover image legend if applicable. Computers & Security, 70(??):CO2, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301864.

Anonymous:2017:IFCc

[Ano17n] Anonymous. Inside front cover — editorial board page/ cover image legend if applicable. Computers & Security, 71(??):CO2, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301712.

Anonymous:2017:PN

[Ano17o] Anonymous. Pages 1–132 (November 2017). Computers & Security, 71(??):1–132, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PJa

[Ano17p] Anonymous. Pages 1–134 (January 2017). Computers & Security, 64(??):1–134, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PJc

[Ano17q] Anonymous. Pages 1–196 (July 2017). Computers & Security, 68(??):1–196, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PMb

[Ano17r] Anonymous. Pages 1–234 (May 2017). Computers & Security, 66(??):1–234, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PJb

[Ano17s] Anonymous. Pages 1–368 (June 2017). Computers & Security, 67(??):1–368, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PMa

[Ano17t] Anonymous. Pages 1–432 (March 2017). Computers & Security, 65(??):1–432, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PS

[Ano17u] Anonymous. Pages 1–756 (September 2017). Computers & Security, 70(??):1–756, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2017:PIS

[Ano17v] Anonymous. Preface for the IFIP SEC special issue. Computers & Security, 71(??):87, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301657.

Anonymous:2018:A

[Ano18a] Anonymous. Announcement. Computers & Security, 79 (??):iv, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404818310848.

Anonymous:2018:CSEa

[Ano18b] Anonymous. Computers and security editorial board. Computers & Security, 78(??):ii–iii, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818310071.

Anonymous:2018:CSEb

[Ano18c] Anonymous. Computers and security editorial board. Computers & Security, 79(??):ii–iii, November 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818310812.

Anonymous:2018:EBa

[Ano18d] Anonymous. Editorial Board. Computers & Security, 73(??): ii, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818300440.

Anonymous:2018:EBb

[Ano18e] Anonymous. Editorial Board. Computers & Security, 74(??): ii, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818301032.

Anonymous:2018:EBc

[Ano18f] Anonymous. Editorial Board. Computers & Security, 75(??): ii, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481830316X.

Anonymous:2018:EBd

[Ano18g] Anonymous. Editorial Board. Computers & Security, 76(??): ii, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818305455.

Anonymous:2018:EBe

[Ano18h] Anonymous. Editorial Board. Computers & Security, 77(??):ii–iii, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308034.

Anonymous:2018:IFC

[Ano18i] Anonymous. Inside front cover — editorial board page/ cover image legend if applicable. Computers & Security, 72(??):CO2, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730216X.

Anonymous:2018:PN

[Ano18j] Anonymous. Pages 1–208 (November 2018). Computers & Security, 79(??):1–208, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PJb

[Ano18k] Anonymous. Pages 1–218 (June 2018). Computers & Security, 75(??):1–218, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PJa

[Ano18l] Anonymous. Pages 1–272 (January 2018). Computers & Security, 72(??):1–272, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PJc

[Ano18m] Anonymous. Pages 1–366 (July 2018). Computers & Security, 76(??):1–366, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PMb

[Ano18n] Anonymous. Pages 1–400 (May 2018). Computers & Security, 74(??):1–400, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PS

[Ano18o] Anonymous. Pages 1–490 (September 2018). Computers & Security, 78(??):1–490, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PMa

[Ano18p] Anonymous. Pages 1–544 (March 2018). Computers & Security, 73(??):1–544, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:PA

[Ano18q] Anonymous. Pages 1–886 (August 2018). Computers & Security, 77(??):1–886, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2018:RAL

[Ano18r] Anonymous. Reviewer acknowledgement list — 2017. Computers & Security, 72(??):i–v, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302195.

Anonymous:2019:CSE

[Ano19a] Anonymous. Computers and security editorial board. Computers & Security, 80(??):ii–iii, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818312434.

Anonymous:2019:EBa

[Ano19b] Anonymous. Editorial Board. Computers & Security, 81 (??):ii–iii, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404818314585.

Anonymous:2019:EBb

[Ano19c] Anonymous. Editorial Board. Computers & Security, 82(??): ii–iii, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819300379.

Anonymous:2019:EBc

[Ano19d] Anonymous. Editorial Board. Computers & Security, 83(??): ii–iii, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819300793.

Anonymous:2019:EBd

[Ano19e] Anonymous. Editorial Board. Computers & Security, 84(??): ii–iii, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301014.

Anonymous:2019:EBe

[Ano19f] Anonymous. Editorial Board. Computers & Security, 85(??):ii–iii, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301282.

Anonymous:2019:EBf

[Ano19g] Anonymous. Editorial Board. Computers & Security, 86(??):ii–iii, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301452.

Anonymous:2019:EB

[Ano19h] Anonymous. Editorial Board. Computers & Security, 87 (??):Article 101624, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301713.

Anonymous:2019:GES

[Ano19i] Anonymous. Guest editorial: Special issue on Information Systems Security, Privacy, Security and Cryptography (ICISSP 2017 and SECRYPT 2017). Computers & Security, 86(??):419, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301233.

Anonymous:2019:N

[Ano19j] Anonymous. November 2019. Computers & Security, 87(??): ??, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PMa

[Ano19k] Anonymous. Pages 1–184 (March 2019). Computers & Security, 81(??):1–184, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PMb

[Ano19l] Anonymous. Pages 1–328 (May 2019). Computers & Security, 82(??):1–328, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PJa

[Ano19m] Anonymous. Pages 1–334 (January 2019). Computers & Security, 80(??):1–334, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PJc

[Ano19n] Anonymous. Pages 1–402 (July 2019). Computers & Security, 84(??):1–402, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PJb

[Ano19o] Anonymous. Pages 1–406 (June 2019). Computers & Security, 83(??):1–406, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PA

[Ano19p] Anonymous. Pages 1–452 (August 2019). Computers & Security, 85(??):1–452, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:PS

[Ano19q] Anonymous. Pages 1–512 (September 2019). Computers & Security, 86(??):1–512, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2019:RA

[Ano19r] Anonymous. Reviewer acknowledgment. Computers & Security, 81(??):i–v, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818314639.

Abdalaal:2013:PPP

[ANS13] Ahmed Abdalaal, Mehmet Ercan Nergiz, and Yucel Saygin. Privacy-preserving publishing of opinion polls. Computers & Security, 37(??):143–154, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000539.

AlSabah:2018:YCY

[AOR18] Mashael AlSabah, Gabriele Oligeri, and Ryan Riley. Your culture is in your password: an analysis of a demographically-diverse password dataset. Computers & Security, 77(??):427–441, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302979.

Azmi:2013:SSH

[AP13] Reza Azmi and Boshra Pishgoo. SHADuDT: Secure hypervisor-based anomaly detection using danger theory. Computers & Security, 39 (part B)(??):268–288, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813001223.

Armando:2014:MCA

[AP14] Alessandro Armando and Serena Elisa Ponta. Model checking authorization requirements in business processes. Computers & Security, 40(??):1–22, February 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001429.

Alcaide:2013:AAP

[APMCR13] Almudena Alcaide, Esther Palomar, José Montero-Castillo, and Arturo Ribagorda. Anonymous authentication for privacy-preserving IoT target-driven applications. Computers & Security, 37(??):111–123, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000904.

Alam:2017:DAO

[AQR^{+}17] Shahid Alam, Zhengyang Qu, Ryan Riley, Yan Chen, and Vaibhav Rastogi. DroidNative: Automating and optimizing detection of Android native code malware variants. Computers & Security, 65(??):230–246, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630164X.

Alenezi:2014:UT

[AR14] Mohammed N. Alenezi and Martin J. Reed. Uniform DoS traceback. Computers & Security, 45(??):17–26, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000601.

Aburomman:2017:SID

[AR17] Abdulla Amin Aburomman and Mamun Bin Ibne Reaz. A survey of intrusion detection systems based on ensemble and hybrid classifiers. Computers & Security, 65(??):135–152, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301572.

Anjum:2017:SPM

[ARG^{+}17] Adeel Anjum, Guillaume Raschia, Marc Gelgon, Abid Khan, Saif ur Rehman Malik, Naveed Ahmad, Mansoor Ahmed, Sabah Suhail, and M. Masoom Alam. τ-safety: a privacy model for sequential publication with arbitrary updates. Computers & Security, 66(??):20–39, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300019.

Agudo:2013:PAC

[ARL13] Isaac Agudo, Ruben Rios, and Javier Lopez. A privacy-aware continuous authentication scheme for proximity-based access control. Computers & Security, 39 (part B)(??):117–126, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000874.

Al-rimy:2018:RTS

[ArMS18] Bander Ali Saleh Al-rimy, Mohd Aizaini Maarof, and Syed Zainudeen Mohd Shaid. Ransomware threat success factors, taxonomy, and countermeasures: a survey and research directions. Computers & Security, 74(??):144–166, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481830004X.

Ashenden:2013:COC

[AS13] Debi Ashenden and Angela Sasse. CISOs and organisational culture: Their own worst enemy? Computers & Security, 39 (part B)(??):396–405, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001338.

Al-Safwani:2018:IDM

[ASFI18] Nadher Al-Safwani, Yousef Fazea, and Huda Ibrahim. ISCP: In-depth model for selecting critical security controls. Computers & Security, 77(??):565–577, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818305534.

Akilal:2019:VFR

[ASO19] Karim Akilal, Hachem Slimani, and Mawloud Omar. A very fast and robust trust inference algorithm in weighted signed social networks using controversy, eclecticism, and reciprocity. Computers & Security, 83(??):68–78, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308708.

AlShboul:2018:VCM

[ATAAd18] Rabah AlShboul, Fadi Thabtah, Neda Abdelhamid, and Mofleh Al-diabat. A visualization cybersecurity method based on features’ dissimilarity. Computers & Security, 77(??):289–303, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818303572.

Ariu:2011:HID

[ATG11] Davide Ariu, Roberto Tronci, and Giorgio Giacinto. HMM-Payl: an intrusion detection system based on hidden Markov models. Computers & Security, 30(4):221–241, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000022.

AlSabah:2017:PCL

[ATL^{+}17] Mashael AlSabah, Alin Tomescu, Ilia Lebedev, Dimitrios Serpanos, and Srini Devadas. PriviPK: Certificate-less and secure email communication. Computers & Security, 70(??):1–15, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300834.

Anjum:2018:EPM

[AuRMC^{+}18] Adeel Anjum, Saif ur Rehman Malik, Kim-Kwang Raymond Choo, Abid Khan, Asma Haroon, Sangeen Khan, Samee U. Khan, Naveed Ahmad, and Basit Raza. An efficient privacy mechanism for electronic health records. Computers & Security, 72(??):196–211, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302031.

Avery:2018:FMD

[AW18] Jeffrey Avery and John Ross Wallrabenstein. Formally modeling deceptive patches using a game-based approach. Computers & Security, 75(??):182–190, June 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301330.

Ahmad:2019:SMA

[AWDB19] Atif Ahmad, Jeb Webb, Kevin C. Desouza, and James Boorman. Strategically-motivated advanced persistent threat: Definition, process, tactics and a disinformation model of counterattack. Computers & Security, 86(??):402–418, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818310988.

Ahmad:2017:ESN

[AWZ^{+}17] Adnan Ahmad, Brian Whitworth, Furkh Zeshan, Elisa Bertino, and Robert Friedman. Extending social networks with delegation. Computers & Security, 70(??):546–564, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301517.

Alsmadi:2015:SSD

[AX15] Izzat Alsmadi and Dianxiang Xu. Security of software defined networks: a survey. Computers & Security, 53(??):79–108, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481500070X.

Almalawi:2014:UAB

[AYT^{+}14] Abdulmohsen Almalawi, Xinghuo Yu, Zahir Tari, Adil Fahad, and Ibrahim Khalil. An unsupervised anomaly-based detection approach for integrity attacks on SCADA systems. Computers & Security, 46(??):94–110, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001072.

Akiyama:2017:AEM

[AYY^{+}17] Mitsuaki Akiyama, Takeshi Yagi, Takeshi Yada, Tatsuya Mori, and Youki Kadobayashi. Analyzing the ecosystem of malicious URL redirection through longitudinal observation from honeypots. Computers & Security, 69(??):155–173, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481730007X.

Aleroud:2017:PET

[AZ17] Ahmed Aleroud and Lina Zhou. Phishing environments, techniques, and countermeasures: a survey. Computers & Security, 68(??):160–196, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300810.

Bag:2018:PAD

[BAH18] Samiran Bag, Muhammad Ajmal Azad, and Feng Hao. A privacy-aware decentralized and personalized reputation system. Computers & Security, 77(??):514–530, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304930.

Bayuk:2012:CA

[Bay12a] Jennifer L. Bayuk. Cyber attacks. Computers & Security, 31(2):251, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001519.

Bayuk:2012:PP

[Bay12b] Jennifer L. Bayuk. A programmer’s perspective. Computers & Security, 31(5):736–737, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481200079X.

Bayuk:2012:WT

[Bay12c] Jennifer L. Bayuk. Wishful thinking. Computers & Security, 31(8):1011–1012, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001241.

Bayuk:2013:STA

[Bay13] Jennifer L. Bayuk. Security as a theoretical attribute construct. Computers & Security, 37(??):155–175, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000515.

Breaux:2011:LRS

[BB11] Travis D. Breaux and David L. Baumer. Legally “reasonable” security requirements: a 10-year FTC retrospective. Computers & Security, 30(4):178–193, June 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810001124.

Bauer:2017:PBT

[BBC17] Stefan Bauer, Edward W. N. Bernroider, and Katharina Chudzikowski. Prevention is better than cure! Designing information security awareness programs to overcome users’ non-compliance with information security policies in banks. Computers & Security, 68(??):145–159, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300871.

Barradas:2019:FAC

[BBD^{+}19] Diogo Barradas, Tiago Brito, David Duarte, Nuno Santos, and Luís Rodrigues. Forensic analysis of communication records of messaging applications from physical memory. Computers & Security, 86(??):484–497, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311313.

Barenghi:2016:PPE

[BBDP16] Alessandro Barenghi, Michele Beretta, Alessandro Di Federico, and Gerardo Pelosi. A privacy-preserving encrypted OSN with stateless server interaction: the Snake design. Computers & Security, 63(??):67–84, November 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301031.

Bertino:2018:ESP

[BCCS18] Elisa Bertino, Valentina Casola, Aniello Castiglione, and Willy Susilo. Editorial: Security and privacy protection vs sustainable development. Computers & Security, 76(??):250– 251, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481830539X.

Blundo:2010:MKH

[BCD^{+}10] Carlo Blundo, Stelvio Cimato, Sabrina De Capitani di Vimercati, Alfredo De Santis, Sara Foresti, Stefano Paraboschi, and Pierangela Samarati. Managing key hierarchies for access control enforcement: Heuristic approaches.

Computers & Security, 29(5):533–547, July 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001448.

Bedford:2017:PSF

[BCD^{+}17] Andrew Bedford, Stephen Chong, Josée Desharnais, Elisavet Kozyri, and Nadia Tawbi. A progress-sensitive flow-sensitive inlined information-flow control monitor (extended version). Computers & Security, 71(??):114–131, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300676.

Burdon:2019:SSC

[BCK19] Mark Burdon and Lizzie Coles-Kemp. The significance of securing as a critical component of information security: an Australian narrative. Computers & Security, 87 (??):Article 101601, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301555.

Belguechi:2013:IFC

[BCRAA13] Rima Belguechi, Estelle Cherrier, Christophe Rosenberger, and Samy Ait-Aoudia. An integrated framework combining bio-hashed minutiae template and PKCS15 compliant card for a better secure management of fingerprint cancelable templates. Computers & Security, 39 (part B)(??):325–339, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001260.

Bajwa:2016:NTN

[BD16] Garima Bajwa and Ram Dantu. Neurokey: Towards a new paradigm of cancelable biometrics-based key generation using electroencephalograms. Computers & Security, 62(??):95–113, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300669.

Baruah:2018:TFA

[BD18] Barnana Baruah and Subhasish Dhal. A two-factor authentication scheme against FDM attack in IFTTT based smart home system. Computers & Security, 77(??):21–35, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818302402.

Bottarelli:2018:PCW

[BEB^{+}18] Mirko Bottarelli, Gregory Epiphaniou, Dhouha Kbaier Ben Ismail, Petros Karadimas, and Haider Al-Khateeb. Physical characteristics of wireless communication channels for secret key establishment: a survey of the research. Computers & Security, 78(??):454–476, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300841.

Biondi:2019:EER

[BEGW^{+}19] Fabrizio Biondi, Michael A. Enescu, Thomas Given-Wilson, Axel Legay, Lamine Noureddine, and Vivek Verma. Effective, efficient, and robust packing detection and classification. Computers & Security, 85(??):436–451, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311040.

Beltran:2018:IAA

[Bel18] Marta Beltrán. Identifying, authenticating and authorizing smart objects and end users to cloud services in Internet of Things. Computers & Security, 77(??):595–611, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818301366.

Bitton:2018:TMU

[BFS^{+}18] Ron Bitton, Andrey Finkelshtein, Lior Sidi, Rami Puzis, Lior Rokach, and Asaf Shabtai. Taxonomy of mobile users’ security awareness. Computers & Security, 73(??):266–293, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302316.

Burnap:2018:MCU

[BFTJ18] Pete Burnap, Richard French, Frederick Turner, and Kevin Jones. Malware classification using self organising feature maps and machine activity data. Computers & Security, 73(??):399–410, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302535.

Black:2018:SSB

[BGL18] Paul Black, Iqbal Gondal, and Robert Layton. A survey of similarities in banking malware behaviours. Computers & Security, 77(??):756–772, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730202X.

Bella:2017:TET

[BGLR17] Giampaolo Bella, Rosario Giustolisi, Gabriele Lenzini, and Peter Y. A. Ryan. Trustworthy exams without trusted parties. Computers & Security, 67(??):291–307, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301766.

Bella:2011:EPC

[BGR11] Giampaolo Bella, Rosario Giustolisi, and Salvatore Riccobene. Enforcing privacy in e-commerce by balancing anonymity and trust. Computers & Security, 30(8):705–718, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811001052.

Badenhop:2017:ZWR

[BGR^{+}17] Christopher W. Badenhop, Scott R. Graham, Benjamin W. Ramsey, Barry E. Mullins, and Logan O. Mailloux. The Z-Wave routing protocol and its security implications. Computers & Security, 68(??):112–129, July 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300792.

Barnes:2015:LEA

[BHC15] David J. Barnes and Julio Hernandez-Castro. On the limits of engine analysis for cheating detection in chess. Computers & Security, 48(??):58–73, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001485.

Baseri:2018:PPF

[BHC18] Yaser Baseri, Abdelhakim Hafid, and Soumaya Cherkaoui. Privacy preserving fine-grained location-based access control for mobile cloud. Computers & Security, 73(??):249–265, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302304.

Blasco:2012:BIL

[BHCTR12] Jorge Blasco, Julio Cesar Hernandez-Castro, Juan E. Tapiador, and Arturo Ribagorda. Bypassing information leakage protection with trusted applications. Computers & Security, 31(4):557–568, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000120.

Bou-Harb:2014:FPA

[BHDA14] Elias Bou-Harb, Mourad Debbabi, and Chadi Assi. On fingerprinting probing activities. Computers & Security, 43 (??):35–48, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404814000248.

Bauer:2010:RVC

[BJ10] Andreas Bauer and Jan Jürjens. Runtime verification of cryptographic protocols. Computers & Security, 29(3):315– 330, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404809001047.

Baracaldo:2013:ARM

[BJ13] Nathalie Baracaldo and James Joshi. An adaptive risk management and access control framework to mitigate insider threats. Computers & Security, 39 (part B)(??):237–254, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001119.

Bhandari:2017:AIA

[BJJ^{+}17] Shweta Bhandari, Wafa Ben Jaballah, Vineeta Jain, Vijay Laxmi, Akka Zemmari, Manoj Singh Gaur, Mohamed Mosbah, and Mauro Conti. Android inter-app communication threats and detection techniques. Computers & Security, 70(??):392–421, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301414.

Biondi:2017:ESC

[BJLS17] Fabrizio Biondi, Sébastien Josse, Axel Legay, and Thomas Sirvent. Effectiveness of synthesis in concolic deobfuscation. Computers & Security, 70(??):500–515, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301475.

BenOthmane:2019:SIS

[BJW19] Lotfi Ben Othmane, Doug Jacobson, and Edgar Weippl. Special issue on security and privacy in smart cyber-physical systems. Computers & Security, 87(??):Article 101611, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301609.

B:2019:FAB

[BK19] Selvakumar B. and Muneeswaran K. Firefly algorithm based feature selection for network intrusion detection. Computers & Security, 81(??):148–155, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303936.

Bu:2019:BSI

[BKK19] Lake Bu, Mark G. Karpovsky, and Michel A. Kinsy. Bulwark: Securing implantable medical devices communication channels. Computers & Security, 86(??):498–511, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818312525.

Basagiannis:2010:IMM

[BKP10] Stylianos Basagiannis, Panagiotis Katsaros, and Andrew Pombortsis. An intruder model with message inspection for model checking security protocols. Computers & Security, 29(1):16–34, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000856.

Bouffard:2015:UCF

[BL15] Guillaume Bouffard and Jean-Louis Lanet. The ultimate control flow transfer in a Java based smart card. Computers & Security, 50(??):33–46, May 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500005X.

Beheshti:2017:CMB

[BLR17] Seyed Mohammad Reza Saadat Beheshti, Panos Liatsis, and Muttukrishnan Rajarajan. A CAPTCHA model based on visual psychophysics: Using the brain to distinguish between human users and automated computer bots. Computers & Security, 70(??):596–617, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730161X.

Boshmaf:2016:ILV

[BLS^{+}16] Yazan Boshmaf, Dionysios Logothetis, Georgos Siganos, Jorge Lería, Jose Lorenzo, Matei Ripeanu, Konstantin Beznosov, and Hassan Halawa. Íntegro: Leveraging victim prediction for robust fake account detection in large scale OSNs. Computers & Security, 61(??):142–168, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300633.

Bartnes:2017:CIS

[BM17a] Maria Bartnes and Nils Brede Moe. Challenges in IT security preparedness exercises: a case study. Computers & Security, 67(??):280–290, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301705.

Bopche:2017:GSM

[BM17b] Ghanshyam S. Bopche and Babu M. Mehtre. Graph similarity metrics for assessing temporal changes in attack surface of dynamic networks. Computers & Security, 64(??):16–43, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301171.

Badhani:2019:CCE

[BM19] Shikha Badhani and Sunil K. Muttoo. CENDroid — a cluster-ensemble classifier for detecting malicious Android applications. Computers & Security, 85(??):25–40, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819300872.

Bartnes:2016:FIS

[BMH16] Maria Bartnes, Nils Brede Moe, and Poul E. Heegaard. The future of information security incident management training: a case study of electrical power companies. Computers & Security, 61(??):32–45, August 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300530.

Baddar:2018:SEA

[BMMP18] Sherenaz Al-Haj Baddar, Alessio Merlo, Mauro Migliardi, and Francesco Palmieri. Saving energy in aggressive intrusion detection through dynamic latency sensitivity recognition. Computers & Security, 76(??):311–326, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302638.

Bartoli:2018:ETW

[BMO18] Alberto Bartoli, Eric Medvet, and Filippo Onesti. Evil twins and WPA2 Enterprise: a coming security disaster? Computers & Security, 74(??):1–11, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302808.

Bongiovanni:2019:LSP

[Bon19] Ivano Bongiovanni. The least secure places in the universe? A systematic literature review on information security management in higher education. Computers & Security, 86(??):350–357, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301324.

Bailey:2014:UIA

[BOP14] Kyle O. Bailey, James S. Okolica, and Gilbert L. Peterson. User identification and authentication using multi-modal behavioral biometrics. Computers & Security, 43(??):77–89, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000340.

benOthmane:2015:IAC

[bORF^{+}15] Lotfi ben Othmane, Rohit Ranchal, Ruchith Fernando, Bharat Bhargava, and Eric Bodden. Incorporating attacker capabilities in risk estimation and mitigation. Computers & Security, 51(??):41–61, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000334.

BenSaied:2013:TMS

[BOZL13] Yosra Ben Saied, Alexis Olivereau, Djamal Zeghlache, and Maryline Laurent. Trust management system design for the Internet of Things: a context-aware and multi-service approach. Computers & Security, 39 (part B)(??):351–365, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001302.

Basagiannis:2011:QAC

[BPA^{+}11] S. Basagiannis, S. Petridou, N. Alexiou, G. Papadimitriou, and P. Katsaros. Quantitative analysis of a certified email protocol in mobile environments: a probabilistic model checking approach. Computers & Security, 30(4):257–272, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000435.

Bai:2019:MCN

[BPG^{+}19] Wei Bai, Zhisong Pan, Shize Guo, Zhe Chen, and Shiming Xia. MDC-Checker: a novel network risk assessment framework for multiple domain configurations. Computers & Security, 86(??):388–401, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307582.

Belyaev:2017:FDI

[BR17] Kirill Belyaev and Indrakshi Ray. On the formalization, design, and implementation of component-oriented access control in lightweight virtualized server environments. Computers & Security, 71(??):15–35, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301232.

Bushouse:2018:GCP

[BR18] Micah Bushouse and Douglas Reeves. Goalkeeper: Comprehensive process enforcement from the hypervisor. Computers & Security, 73(??):459–473, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302572.

Bouhaddi:2018:EID

[BRA18] Myria Bouhaddi, Mohammed Saïd Radjef, and Kamel Adi. An efficient intrusion detection in resource-constrained mobile ad-hoc networks. Computers & Security, 76(??):156–177, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818301652.

Bal:2015:SPR

[BRH15] Gökhan Bal, Kai Rannenberg, and Jason I. Hong. Styx: Privacy risk communication for the Android smartphone platform based on apps’ data-access behavior patterns. Computers & Security, 53(??):187–202, September 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000541.

Buchler:2018:SOA

[BRM^{+}18] Norbou Buchler, Prashanth Rajivan, Laura R. Marusich, Lewis Lightner, and Cleotilde Gonzalez. Sociometrics and observational assessment of teaming and leadership in a cyber security defense competition. Computers & Security, 73(??):114–136, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302298.

Bryant:2017:NKC

[BS17] Blake D. Bryant and Hossein Saiedian. A novel kill-chain framework for remote security log analysis with SIEM software. Computers & Security, 67(??):198–210, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300561.

Basuchoudhary:2019:SSC

[BS19] Atin Basuchoudhary and Nicola Searle. Snatched secrets: Cybercrime and trade secrets modelling a firm’s decision to report a theft of trade secrets. Computers & Security, 87(??):Article 101591, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300616.

Bendiab:2019:FNF

[BSBG19] Keltoum Bendiab, Stavros Shiaeles, Samia Boucherkha, and Bogdan Ghita. FCMDT: a novel fuzzy cognitive maps dynamic trust model for cloud federated identity management. Computers & Security, 86(??):270–290, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818312252.

Bui:2019:GEA

[BSL19] Thang Bui, Scott D. Stoller, and Jiajie Li. Greedy and evolutionary algorithms for mining relationship-based access control policies. Computers & Security, 80(??):317–333, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818304565.

Bhatia:2014:FGR

[BSMT14] Sajal Bhatia, Desmond Schmidt, George Mohay, and Alan Tickle. A framework for generating realistic traffic for distributed denial-of-service attacks and flash events. Computers & Security, 40(??):95–107, February 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001673.

Balachandran:2016:CFO

[BSTT16] Vivek Balachandran, Sufatrio, Darell J. J. Tan, and Vrizlynn L. L. Thing. Control flow obfuscation for Android applications. Computers & Security, 61(??):72–93, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300529.

Bouabana-Tebibel:2015:PSE

[BTK15] Thouraya Bouabana-Tebibel and Abdellah Kaci. Parallel search over encrypted data under attribute based encryption on the cloud computing. Computers & Security, 54(??):77–91, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000577.

Barton:2016:ISS

[BTLT16] Kevin A. Barton, Gurvirender Tejay, Michael Lane, and Steve Terrell. Information system security commitment: a study of external influences on senior management. Computers & Security, 59(??):9–25, June 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300104.

Bachlechner:2014:SCC

[BTM14] Daniel Bachlechner, Stefan Thalmann, and Ronald Maier. Security and compliance challenges in complex IT outsourcing arrangements: a multi-stakeholder perspective. Computers & Security, 40(??):38–59, February 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001533.

Beuran:2018:IFH

[BTP^{+}18] Razvan Beuran, Dat Tang, Cuong Pham, Ken ichi Chinen, Yasuo Tan, and Yoichi Shinoda. Integrated framework for hands-on cybersecurity training: CyTrONE. Computers & Security, 78(??):43–59, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306527.

Bock:2019:ADS

[BVWM19] Leon Böck, Emmanouil Vasilomanolakis, Jan Helge Wolf, and Max Mühlhäuser. Autonomously detecting sensors in fully distributed botnets. Computers & Security, 83(??):1– 13, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818312094.

Bashir:2017:PCC

[BWMG17] Masooda Bashir, Colin Wee, Nasir Memon, and Boyi Guo. Profiling cybersecurity competition participants: Self-efficacy, decision-making and interests predict effectiveness of competitions as a recruitment tool. Computers & Security, 65(??):153–165, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301389.

Barlow:2013:DME

[BWOD13] Jordan B. Barlow, Merrill Warkentin, Dustin Ormond, and Alan R. Dennis. Don’t make excuses! Discouraging neutralization to reduce IT policy violation. Computers & Security, 39 (part B)(??):145–159, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000898.

Bhatnagar:2012:NRA

[BWR12] Gaurav Bhatnagar, Q. M. Jonathan Wu, and Balasubramanian Raman. A new robust adjustable logo watermarking scheme. Computers & Security, 31(1):40–58, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001398.

Caelli:2012:CWT

[Cae12] William J. (Bill) Caelli. Cyber warfare — techniques, tactics and tools for security practitioners. Computers & Security, 31(4):637, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000430.

Cazorla:2015:TSA

[CAL15] Lorena Cazorla, Cristina Alcaraz, and Javier Lopez. A three-stage analysis of IDS for critical infrastructures. Computers & Security, 55(??):235–250, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001133.

Chiba:2018:NAC

[CAM^{+}18] Zouhair Chiba, Noureddine Abghour, Khalid Moussaid, Amina El Omri, and Mohamed Rida. A novel architecture combined with optimal parameters for back propagation neural networks applied to anomaly network intrusion detection. Computers & Security, 75(??):36–58, June 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300543.

Chakraborty:2019:TIP

[CAM19a] Nilesh Chakraborty, Vijay S. Anand, and Samrat Mondal. Towards identifying and preventing behavioral side channel attack on recording attack resilient unaided authentication services. Computers & Security, 84(??):193–205, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302062.

Chiba:2019:IAB

[CAM^{+}19b] Zouhair Chiba, Noreddine Abghour, Khalid Moussaid, Amina El omri, and Mohamed Rida. Intelligent approach to build a deep neural network based IDS for cloud environment using combination of machine learning algorithms. Computers & Security, 86(??):291–317, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301221.

Chiba:2018:DBA

[CAY^{+}18] Daiki Chiba, Mitsuaki Akiyama, Takeshi Yagi, Kunio Hato, Tatsuya Mori, and Shigeki Goto. DomainChroma: Building actionable threat intelligence from malicious domain names. Computers & Security, 77(??):138–161, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302967.

Cuppens-Boulahia:2012:PTO

[CB12] Nora Cuppens-Boulahia. Penetration tester’s open source toolkit. Computers & Security, 31(4):630–632, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000600.

Cherdantseva:2016:RCS

[CBB^{+}16] Yulia Cherdantseva, Pete Burnap, Andrew Blyth, Peter Eden, Kevin Jones, Hugh Soulsby, and Kristan Stoddart. A review of cyber security risk assessment methods for SCADA systems. Computers & Security, 56(??):1–27, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001388.

Choo:2018:ICT

[CBGN18] Kim-Kwang Raymond Choo, Matt Bishop, William Glisson, and Kara Nance. Internet- and cloud-of-things cybersecurity research challenges and advances. Computers & Security, 74(??):275–276, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481830097X.

Cheneau:2010:SIP

[CBL10] Tony Cheneau, Aymen Boudguiga, and Maryline Laurent. Significantly improved performances of the cryptographically generated addresses thanks to ECC and GPGPU. Computers & Security, 29(4):419–431, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001461.

Chamikara:2019:ESP

[CBL^{+}19] M. A. P. Chamikara, P. Bertok, D. Liu, S. Camtepe, and I. Khalil. An efficient and scalable privacy preserving algorithm for big data and data streams. Computers & Security, 87(??):Article 101570, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818313683.

Cohen:2014:TSI

[CC14] Fred Cohen and Don Cohen. Time and space interval record schedule consistency analysis for atomic items without interactions in open spaces with stationary locations. Computers & Security, 45(??):293–304, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000315.

Castiglione:2012:ESM

[CCC^{+}12] Aniello Castiglione, Giuseppe Cattaneo, Maurizio Cembalo, Alfredo De Santis, Pompeo Faruolo, Fabio Petagna, and Umberto Ferraro Petrillo. Engineering a secure mobile messaging framework. Computers & Security, 31(6):771–781, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481200096X.

Chang:2018:UFD

[CCC^{+}18] Bing Chang, Yao Cheng, Bo Chen, Fengwei Zhang, Wen-Tao Zhu, Yingjiu Li, and Zhan Wang. User-friendly deniable storage for mobile devices. Computers & Security, 72(??):163–174, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301943.

Carminati:2015:BDS

[CCM^{+}15] Michele Carminati, Roberto Caron, Federico Maggi, Ilenia Epifani, and Stefano Zanero. BankSealer: a decision support system for online banking fraud analysis and investigation. Computers & Security, 53(??):175–186, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000437.

Chen:2018:SRF

[CCM^{+}18] Chen Chen, Baojiang Cui, Jinxin Ma, Runpu Wu, Jianchao Guo, and Wenqian Liu. A systematic review of fuzzing techniques. Computers & Security, 75(??):118–137, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300658.

Chiew:2015:UWL

[CCST15] Kang Leng Chiew, Ee Hung Chang, San Nah Sze, and Wei King Tiong. Utilisation of website logo for phishing detection. Computers & Security, 54(??):16–26, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001145.

Charanek:2011:VEE

[CDE^{+}11] Elias Abou Charanek, Hoseb Dermanilian, Imad Elhajj, Ayman Kayssi, and Ali Chehab. E^{2} VoIP^{2}: Energy efficient voice over IP privacy. Computers & Security, 30(8):815–829, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000848.

Cabaj:2018:CEE

[CDKR18] Krzysztof Cabaj, Dulce Domingos, Zbigniew Kotulski, and Ana Respício. Cybersecurity education: Evolution of the discipline and analysis of master programs. Computers & Security, 75(??):24–35, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300373.

Colantonio:2010:TRM

[CDOV10] Alessandro Colantonio, Roberto Di Pietro, Alberto Ocello, and Nino Vincenzo Verde. Taming role mining complexity in RBAC. Computers & Security, 29(5):548–564, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000027.

Cheminod:2019:CAA

[CDS^{+}19] Manuel Cheminod, Luca Durante, Lucia Seno, Fulvio Valenza, and Adriano Valenzano. A comprehensive approach to the automatic refinement and verification of access control policies. Computers & Security, 80(??):186–199, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818303870.

Colantonio:2012:BDD

[CDV12] Alessandro Colantonio, Roberto Di Pietro, and Nino Vincenzo Verde. A business-driven decomposition methodology for role mining. Computers & Security, 31(7):844–855, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000090.

Carminati:2011:SWB

[CFH^{+}11] Barbara Carminati, Elena Ferrari, Raymond Heatherly, Murat Kantarcioglu, and Bhavani Thuraisingham. Semantic web-based social network access control. Computers & Security, 30(2–3):108–115, March 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000799.

Cid-Fuentes:2018:AFD

[CFSF18] Javier Álvarez Cid-Fuentes, Claudia Szabo, and Katrina Falkner. An adaptive framework for the detection of novel botnets. Computers & Security, 79(??):148–161, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818309805.

Camacho:2019:MBD

[CGGFGMF19] José Camacho, Jose Manuel García-Giménez, Noemí Marta Fuentes-García, and Gabriel Maciá-Fernández. Multivariate big data analysis for intrusion detection: 5 steps from the haystack to the needle. Computers & Security, 87 (??):Article 101603, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818307909.

Chong:2018:IPP

[CGLP18] Isis Chong, Huangyi Ge, Ninghui Li, and Robert W. Proctor. Influence of privacy priming and security framing on mobile app selection. Computers & Security, 78(??):143–154, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818305856.

Callegati:2018:CTM

[CGMP18] Franco Callegati, Saverio Giallorenzo, Andrea Melis, and Marco Prandini. Cloud-of-things meets mobility-as-a-service: an insider threat perspective. Computers & Security, 74(??):277–295, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302134.

Callegari:2017:ITM

[CGP17] Christian Callegari, Stefano Giordano, and Michele Pagano. An information-theoretic method for the detection of anomalies in network traffic. Computers & Security, 70(??):351–365, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301438.

Callegari:2011:CSW

[CGPP11] C. Callegari, S. Giordano, M. Pagano, and T. Pepe. Combining sketches and wavelet analysis for multi time-scale network anomaly detection. Computers & Security, 30(8):692–704, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811001064.

Callegari:2012:WCI

[CGPP12] C. Callegari, S. Giordano, M. Pagano, and T. Pepe. WAVE-CUSUM: Improving CUSUM performance in network anomaly detection by means of wavelet analysis. Computers & Security, 31(5):727–735, July 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000788.

Conti:2018:ESR

[CGR18] Mauro Conti, Ankit Gangwal, and Sushmita Ruj. On the economic significance of ransomware campaigns: a Bitcoin transactions perspective. Computers & Security, 79(??):162–189, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304334.

Cowley:2015:ENI

[CGW15] Jennifer A. Cowley, Frank L. Greitzer, and Bronwyn Woods. Effect of network infrastructure factors on information system risk judgments. Computers & Security, 52(??):142–158, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000619.

Chitrakar:2014:SCS

[CH14] Roshan Chitrakar and Chuanhe Huang. Selection of candidate support vectors in incremental SVM for network intrusion detection. Computers & Security, 45(??):231–241, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000996.

Cao:2019:TAA

[CHK^{+}19] Yan Cao, Zhiqiu Huang, Changbo Ke, Jian Xie, and Jinyong Wang. A topology-aware access control model for collaborative cyber-physical spaces: Specification and verification. Computers & Security, 87(??):Article 101478, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818310861.

Choo:2011:CTL

[Cho11] Kim-Kwang Raymond Choo. The cyber threat landscape: Challenges and future research directions. Computers & Security, 30(8):719–731, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001040.

Cho:2013:CFM

[Cho13a] Gyu-Sang Cho. A computer forensic method for detecting timestamp forgery in NTFS. Computers & Security, 34(??):36–46, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001721.

Choo:2013:NPM

[Cho13b] Kim-Kwang Raymond Choo. New payment methods: a review of 2010-2012 FATF mutual evaluation reports. Computers & Security, 36(??):12–26, July 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000278.

Cherdantseva:2016:MER

[CHRI16] Yulia Cherdantseva, Jeremy Hilton, Omer Rana, and Wendy Ivins. A multifaceted evaluation of the reference model of information assurance and security. Computers & Security, 63(??):45–66, November 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301146.

Chandrasekhar:2017:NAC

[CIS17] Santosh Chandrasekhar, Ahmed Ibrahim, and Mukesh Singhal. A novel access control protocol using proxy signatures for cloud-based health information exchange. Computers & Security, 67(??):73–88, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300342.

Chaudhary:2019:BBB

[CJA^{+}19] Rajat Chaudhary, Anish Jindal, Gagangeet Singh Aujla, Shubhani Aggarwal, Neeraj Kumar, and Kim-Kwang Raymond Choo. BEST: Blockchain-based secure energy trading in SDN-enabled intelligent transportation system. Computers & Security, 85(??):288–299, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831201X.

Crossler:2013:FDB

[CJL^{+}13] Robert E. Crossler, Allen C. Johnston, Paul Benjamin Lowry, Qing Hu, Merrill Warkentin, and Richard Baskerville. Future directions for behavioral information security research. Computers & Security, 32(??):90–101, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001460.

Cook:2017:ICS

[CJSM17] Allan Cook, Helge Janicke, Richard Smith, and Leandros Maglaras. The industrial control system cyber defence triage process. Computers & Security, 70(??):467–481, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301505.

Choi:2019:PTE

[CJW^{+}19] Hoyul Choi, Jongmin Jeong, Simon S. Woo, Kyungtae Kang, and Junbeom Hur. Password typographical error resilience in honey encryption. Computers & Security, 87(??):Article 101411, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311246.

Chen:2017:MLB

[CKA^{+}17] Yizheng Chen, Panagiotis Kintis, Manos Antonakakis, Yacin Nadji, David Dagon, and Michael Farrell. Measuring lower bounds of the financial abuse to online advertisers: a four year case study of the TDSS/TDL4 Botnet. Computers & Security, 67(??):164–180, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300366.

Choi:2017:OSL

[CKJ^{+}17] Changho Choi, Nohyun Kwak, Jinsoo Jang, Daehee Jang, Kuenwhee Oh, Kyungsoo Kwag, and Brent Byunghoon Kang. S-OpenSGX: a system-level platform for exploring SGX enclave-based computing. Computers & Security, 70(??):290–306, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301256.

Chabot:2011:EEP

[CKT11] Hugues Chabot, Raphaël Khoury, and Nadia Tawbi. Extending the enforcement power of truncation monitors using static analysis. Computers & Security, 30(4):194–207, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810001136.

Clarke:2017:NPP

[CLF17] N. Clarke, F. Li, and S. Furnell. A novel privacy preserving user identification approach for network traffic. Computers & Security, 70(??):335–350, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301384.

Cheng:2013:UVS

[CLL^{+}13] Lijiao Cheng, Ying Li, Wenli Li, Eric Holm, and Qingguo Zhai. Understanding the violation of IS security policy in organizations: an integrated model based on social control and deterrence theory. Computers & Security, 39 (part B)(??):447–459, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001387.

Choi:2010:WVE

[CLLK10] Y.-H. Choi, L. Li, P. Liu, and G. Kesidis. Worm virulence estimation for the containment of local worm outbreak. Computers & Security, 29(1):104–123, February 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000753.

Choi:2017:ART

[CLS17] Hyo Shin Choi, Won Sang Lee, and So Young Sohn. Analyzing research trends in personal information privacy using topic modeling. Computers & Security, 67(??):244–253, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300603.

Chakraborty:2017:DMU

[CM17] Nilesh Chakraborty and Samrat Mondal. On designing a modified-UI based honeyword generation approach for overcoming the existing limitations. Computers & Security, 66(??):155–168, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300226.

Chejerla:2019:IFA

[CM19] Brijesh Kashyap Chejerla and Sanjay. K. Madria. Information fusion architecture for secure cyber physical systems. Computers & Security, 85(??):122–137, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818314391.

Costa:2010:RMN

[CMM^{+}10] Gabriele Costa, Fabio Martinelli, Paolo Mori, Christian Schaefer, and Thomas Walter. Runtime monitoring for next generation Java ME platform. Computers & Security, 29(1):74–87, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000790.

Cerroni:2015:DDN

[CMPR15] Walter Cerroni, Gianluca Moro, Roberto Pasolini, and Marco Ramilli. Decentralized detection of network attacks through P2P data clustering of SNMP data. Computers & Security, 52(??):1–16, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000383.

Canfora:2016:HSE

[CMV16] Gerardo Canfora, Francesco Mercaldo, and Corrado Aaron Visaggio. An HMM and structural entropy based detector for Android malware: an empirical study. Computers & Security, 61(??):1–18, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300499.

Cohen:2010:MFA

[Coh10] Fred Cohen. A method for forensic analysis of control. Computers & Security, 29(8):891–902, November 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000374.

Carlin:2019:CAM

[COS19] Domhnall Carlin, Philip O’Kane, and Sakir Sezer. A cost analysis of machine learning using dynamic runtime opcodes for malware detection. Computers & Security, 85(??):138–155, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819300082.

Coulter:2018:IAD

[CP18] Rory Coulter and Lei Pan. Intelligent agents defending for an IoT world: a review. Computers & Security, 73(??):439–458, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302511.

Castillo-Perez:2013:ORC

[CPGA13] Sergio Castillo-Pérez and Joaquin Garcia-Alfaro. Onion routing circuit construction via latency graphs. Computers & Security, 37(??):197–214, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000485.

Camacho:2016:PBM

[CPVGTMF16] José Camacho, Alejandro Pérez-Villegas, Pedro García-Teodoro, and Gabriel Maciá-Fernández. PCA-based multivariate statistical network monitoring for anomaly detection. Computers & Security, 59(??):118–137, June 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300116.

Chun:2018:PPP

[CRJ18] Hu Chun, Kui Ren, and Wei Jiang. Privacy-preserving power usage and supply control in smart grid. Computers & Security, 77(??):709–719, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481830049X.

Crawford:2013:FCT

[CRS13] Heather Crawford, Karen Renaud, and Tim Storer. A framework for continuous, transparent mobile device authentication. Computers & Security, 39 (part B)(??):127–136, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000886.

Cheah:2018:BAS

[CSBW18] Madeline Cheah, Siraj A. Shaikh, Jeremy Bryans, and Paul Wooderson. Building an automotive security assurance case using systematic security evaluations. Computers & Security, 77(??):360–379, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303584.

Chen:2019:PGP

[CSS^{+}19] Jinyin Chen, Mengmeng Su, Shijing Shen, Hui Xiong, and Haibin Zheng. POBA-GA: Perturbation optimized black-box adversarial attacks via genetic algorithm. Computers & Security, 85(??):89–106, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818314378.

Chen:2019:UMP

[CSWX19] Kai Chen, Chen Song, Liming Wang, and Zhen Xu. Using memory propagation tree to improve performance of protocol fuzzer when testing ICS. Computers & Security, 87 (??):Article 101582, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301385.

Cheng:2015:DPM

[CSX^{+}15] Xiang Cheng, Sen Su, Shengzhi Xu, Peng Tang, and Zhengyi Li. Differentially private maximal frequent sequence mining. Computers & Security, 55(??):175–192, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001273.

Cheng:2015:DAD

[CSXL15] Xiang Cheng, Sen Su, Shengzhi Xu, and Zhengyi Li. DP-Apriori: a differentially private frequent itemset mining algorithm based on transaction splitting. Computers & Security, 50(??):74–90, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001850.

Chang:2010:EFB

[CTH10] Chin-Chen Chang, Hao-Chuan Tsai, and Yi-Pei Hsieh. An efficient and fair buyer-seller fingerprinting scheme for large scale networks. Computers & Security, 29(2):269–277, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000881.

Chakraborty:2018:SBD

[CVS18] Bodhi Chakraborty, Shekhar Verma, and Krishna Pratap Singh. Staircase based differential privacy with branching mechanism for location privacy preservation in wireless sensor networks. Computers & Security, 77(??):36–48, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481830227X.

Connolly:2019:RCR

[CW19] Lena Y. Connolly and David S. Wall. The rise of cryptoransomware in a changing cybercrime landscape: Taxonomising countermeasures. Computers & Security, 87 (??):Article 101568, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301336.

Chung:2012:HQB

[CWCC12] Yu-Chi Chung, Ming-Chuan Wu, Yih-Chang Chen, and Wen-Kui Chang. A hot query bank approach to improve detection performance against SQL injection attacks. Computers & Security, 31(2):233–248, March 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481100143X.

Cui:2015:PLT

[CWGD15] Baojiang Cui, Fuwei Wang, Tao Guo, and Guowei Dong. A practical off-line taint analysis framework and its application in reverse engineering of file format. Computers & Security, 51(??):1–15, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000218.

Coull:2012:UDR

[CWY^{+}12] Scott E. Coull, Andrew M. White, Ting-Fang Yen, Fabian Monrose, and Michael K. Reiter. Understanding domain registration abuses. Computers & Security, 31(7):806–815, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481200082X.

Chen:2018:APA

[CXF^{+}18] Sen Chen, Minhui Xue, Lingling Fan, Shuang Hao, Lihua Xu, Haojin Zhu, and Bo Li. Automated poisoning attacks and defenses in malware detection systems: an adversarial machine learning approach. Computers & Security, 73(??):326–344, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302444.

Chen:2016:CCI

[CY16] Xiao-Fan Chen and Shun-Zheng Yu. CIPA: a collaborative intrusion prevention architecture for programmable network and SDN. Computers & Security, 58(??):1–19, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001856.

Chen:2017:DNA

[CYLL17] Zhaomin Chen, Chai Kiat Yeo, Bu Sung Lee, and Chiew Tong Lau. Detection of network anomalies using Improved-MSPCA with sketches. Computers & Security, 65(??):314–328, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301419.

Cheng:2014:ECQ

[CZH^{+}14] Liang Cheng, Yang Zhang, Zhihui Han, Yi Deng, Xiaoshan Sun, and Dengguo Feng. Evaluating and comparing the quality of access control in different operating systems. Computers & Security, 47(??):26–40, November 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000716.

Chondros:2019:DEE

[CZZ^{+}19] Nikos Chondros, Bingsheng Zhang, Thomas Zacharias, Panos Diamantopoulos, Stathis Maneas, Christos Patsonakis, Alex Delis, Aggelos Kiayias, and Mema Roussopoulos. Distributed, end-to-end verifiable, and privacy-preserving Internet voting systems. Computers & Security, 83(??):268– 299, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818310290.

Darivandpour:2018:ESP

[DA18] Javad Darivandpour and Mikhail J. Atallah. Efficient and secure pattern matching with wildcards using lightweight cryptography. Computers & Security, 77(??):666–674, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481830021X.

deAlvarenga:2018:PMH

[dABM^{+}18] Sean Carlisto de Alvarenga, Sylvio Barbon, Rodrigo Sanches Miani, Michel Cukier, and Bruno Bogaz Zarpelão. Process mining and hierarchical clustering to help intrusion alert visualization. Computers & Security, 73(??):474–491, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817302584.

Diaz:2014:NXB

[DAR14] Jesus Diaz, David Arroyo, and Francisco B. Rodriguez. New x.509-based mechanisms for fair anonymity management. Computers & Security, 46(??):111–125, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001023.

Derawi:2013:GAR

[DB13] Mohammad Derawi and Patrick Bours. Gait and activity recognition using commercial phones. Computers & Security, 39 (part B)(??):137–144, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000953.

Do:2017:ADD

[DBH17] Quoc Huy Do, Richard Bubel, and Reiner Hähnle. Automatic detection and demonstrator generation for information flow leaks in object-oriented programs. Computers & Security, 67(??):335–349, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301730.

Davis:2011:DPA

[DC11] Jonathan J. Davis and Andrew J. Clark. Data preprocessing for anomaly based network intrusion detection: a review. Computers & Security, 30(6–7):353–375, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000691.

DOrazio:2016:AME

[DC16] Christian D’Orazio and Kim-Kwang Raymond Choo. An adversary model to evaluate DRM protection of video contents on iOS devices. Computers & Security, 56(??):94–110, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000954.

Das:2018:RRD

[DCC^{+}18] Sanjeev Das, Bihuan Chen, Mahintham Chandramohan, Yang Liu, and Wei Zhang. ROPSentry: Runtime defense against ROP attacks using hardware performance counters. Computers & Security, 73(??):374–388, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302481.

deClercq:2017:SSC

[dCGÜ^{+}17] Ruan de Clercq, Johannes Götzfried, David Übler, Pieter Maene, and Ingrid Verbauwhede. SOFIA: Software and control flow integrity architecture. Computers & Security, 68(??):16–35, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300664.

Dwivedi:2017:PPC

[DDSP17] Rudresh Dwivedi, Somnath Dey, Ramveer Singh, and Aditya Prasad. A privacy-preserving cancelable iris template generation scheme using decimal encoding and look-up table mapping. Computers & Security, 65(??):373–386, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301353.

Ding:2014:CFB

[DDYZ14] Yuxin Ding, Wei Dai, Shengli Yan, and Yumei Zhang. Control flow-based opcode behavior analysis for malware detection. Computers & Security, 44(??):65–74, July 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000558.

DaVeiga:2010:FAI

[DE10] A. Da Veiga and J. H. P. Eloff. A framework and assessment instrument for information security culture. Computers & Security, 29(2):196–207, March 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000923.

Dor:2016:MIS

[DE16] Daniel Dor and Yuval Elovici. A model of the information security investment decision-making process. Computers & Security, 63(??):1–13, November 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301043.

Denning:2014:FPA

[Den14] Dorothy E. Denning. Framework and principles for active cyber defense. Computers & Security, 40(??):108–113, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813001661.

DeMott:2013:SBF

[DEP13] Jared D. DeMott, Richard J. Enbody, and William F. Punch. Systematic bug finding and fault localization enhanced with input data tracking. Computers & Security, 32(??):130–157, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481200168X.

Davis:2016:AFE

[DF16] Jonathan J. Davis and Ernest Foo. Automated feature engineering for HTTP tunnel detection. Computers & Security, 59(??):166–185, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816000080.

Djamaludin:2013:EIT

[DFC13] C. I. Djamaludin, E. Foo, and P. Corke. Establishing initial trust in autonomous delay tolerant networks without centralised PKI. Computers & Security, 39 (part B)(??):299–314, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001247.

Djamaludin:2016:RUT

[DFCC16] C. I. Djamaludin, E. Foo, S. Camtepe, and P. Corke. Revocation and update of trust in autonomous delay tolerant networks. Computers & Security, 60(??):15–36, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630030X.

deFuentes:2017:PPP

[dFGMTPL17] José M. de Fuentes, Lorena González-Manzano, Juan Tapiador, and Pedro Peris-Lopez. PRACIS: Privacy-preserving and aggregatable cybersecurity information sharing. Computers & Security, 69(??):127–141, August 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301821.

DeCapitanidiVimercati:2013:EDW

[DFJ^{+}13] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Giovanni Livraga, Stefano Paraboschi, and Pierangela Samarati. Enforcing dynamic write privileges in data outsourcing. Computers & Security, 39 (Part A)(??):47–63, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000266.

Du:2019:SPA

[DHZ19] Shaoyong Du, Jingyu Hua, and Sheng Zhong. Securing peer-assisted indoor localization leveraging acoustic ranging. Computers & Security, 81(??):58–77, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818312392.

Dong:2017:FMA

[DJP17] Naipeng Dong, Hugo Jonker, and Jun Pang. Formal modelling and analysis of receipt-free auction protocols in applied pi calculus. Computers & Security, 65(??):405–432, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301006.

Damopoulos:2013:KTT

[DKG13] D. Damopoulos, G. Kambourakis, and S. Gritzalis. From keyloggers to touchloggers: Take the rough with the smooth. Computers & Security, 32(??):102–114, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001654.

Durkota:2019:HNA

[DLB^{+}19] Karel Durkota, Viliam Lisý, Branislav Bosanský, Christopher Kiekintveld, and Michal Pechoucek. Hardening networks against strategic attackers using attack graph games. Computers & Security, 87(??):Article 101578, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819300689.

Diaz-Lopez:2015:MXS

[DLDTGMMP15] Daniel Díaz-López, Ginés Dólera-Tormo, Félix Gómez-Mármol, and Gregorio Martínez-Pérez. Managing XACML systems in distributed environments through meta-policies. Computers & Security, 48(??):92–115, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001503.

deLeon:2018:ASD

[dLGHK18] Daniel Conte de Leon, Christopher E. Goes, Michael A. Haney, and Axel W. Krings. ADLES: Specifying, deploying, and sharing hands-on cyber-exercises. Computers & Security, 74(??):12–40, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302742.

Dai:2017:ORR

[DLH^{+}17] Ting Dai, Xiaolei Li, Behnaz Hassanshahi, Roland H. C. Yap, and Zhenkai Liang. R opp D roid: Robust permission redelegation prevention in Android inter-component communication. Computers & Security, 68(??):98–111, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300688.

Ding:2019:KBC

[DLLS19] Yan Ding, Nurbol Luktarhan, Keqin Li, and Wushour Slamu. A keyword-based combination approach for detecting phishing webpages. Computers & Security, 84(??):256–275, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819300707.

Dumas:2017:DPP

[DLOP17] Jean-Guillaume Dumas, Pascal Lafourcade, Jean-Baptiste Orfila, and Maxime Puys. Dual protocols for private multiparty matrix multiplication and trust computations. Computers & Security, 71(??):51–70, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300913.

Dong:2014:HFC

[DLS14] Yan Dong, Yongna Li, and Tingting Sun. Happy faces considered trustworthy irrespective of perceiver’s mood: Challenges to the mood congruency effect. Computers & Security, 47(??):85–93, November 2014. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000625.

Dimitrakakis:2017:NOB

[DM17] Christos Dimitrakakis and Aikaterini Mitrokotsa. Nearoptimal blacklisting. Computers & Security, 64(??):110–121, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000966.

Dodel:2019:IMA

[DM19] Matias Dodel and Gustavo Mesch. An integrated model for assessing cyber-safety behaviors: How cognitive, socioeconomic and digital determinants affect diverse safety practices. Computers & Security, 86(??):75–91, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303080.

Do:2015:EDA

[DMC15] Quang Do, Ben Martini, and Kim-Kwang Raymond Choo. Exfiltrating data from Android devices. Computers & Security, 48(??):74–91, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481400162X.

Do:2019:RAM

[DMC19] Quang Do, Ben Martini, and Kim-Kwang Raymond Choo. The role of the adversary model in applied security research. Computers & Security, 81(??):156–181, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306369.

DiFrancescoMaesa:2019:BBA

[DMR19] Damiano Di Francesco Maesa, Paolo Mori, and Laura Ricci. A blockchain based approach for the definition of auditable access control systems. Computers & Security, 84(??):93– 119, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818309398.

Dang-Pham:2015:CIA

[DPP15] Duy Dang-Pham and Siddhi Pittayachawan. Comparing intention to avoid malware across contexts in a BYOD-enabled Australian university: a protection motivation theory approach. Computers & Security, 48(??):281–297, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001643.

Dreier:2019:FPV

[DPP^{+}19] Jannik Dreier, Maxime Puys, Marie-Laure Potet, Pascal Lafourcade, and Jean-Louis Roch. Formally and practically verifying flow properties in industrial systems. Computers & Security, 86(??):453–470, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831441X.

Dang-Pham:2017:ASN

[DPPB17a] Duy Dang-Pham, Siddhi Pittayachawan, and Vince Bruno. Applications of social network analysis in behavioural information security research: Concepts and empirical analysis. Computers & Security, 68(??):1–15, July 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300639.

Dang-Pham:2017:IFI

[DPPB17b] Duy Dang-Pham, Siddhi Pittayachawan, and Vince Bruno. Investigation into the formation of information security influence: Network analysis of an emerging organisation. Computers & Security, 70(??):111–123, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301177.

delRio:2016:ABC

[dRMC^{+}16] Jose Sanchez del Rio, Daniela Moctezuma, Cristina Conde, Isaac Martin de Diego, and Enrique Cabello. Automated border control e-gates and facial recognition systems. Computers & Security, 62(??):49–72, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300736.

Dasgupta:2016:TDA

[DRN16] Dipankar Dasgupta, Arunava Roy, and Abhijit Nag. Toward the design of adaptive selection strategies for multifactor authentication. Computers & Security, 63(??):85–116, November 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481630102X.

Dube:2012:MTR

[DRP^{+}12] T. Dube, R. Raines, G. Peterson, K. Bauer, M. Grimaila, and S. Rogers. Malware target recognition via static heuristics. Computers & Security, 31(1):137–147, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001131.

Faria:2016:IPK

[dSFK16] Gerson de Souza Faria and Hae Yong Kim. Identification of pressed keys by time difference of arrivals of mechanical vibrations. Computers & Security, 57(??):93–105, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001662.

Dassouki:2017:PCB

[DSNH17] Khaled Dassouki, Haidar Safa, Mohamed Nassar, and Abbas Hijazi. Protecting from cloud-based SIP flooding attacks by leveraging temporal and structural fingerprints. Computers & Security, 70(??):618–633, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730158X.

Dhillon:2016:IIS

[DSP16] Gurpreet Dhillon, Romilla Syed, and Cristiane Pedron. Interpreting information security culture: an organizational transformation case study. Computers & Security, 56(??):63–69, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481500139X.

David:2019:EDF

[DT19] Jisa David and Ciza Thomas. Efficient DDoS flood attack detection using dynamic thresholding on flow-based network traffic. Computers & Security, 82(??):284–295, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307624.

Dalmazo:2018:TSM

[DVC18] Bruno L. Dalmazo, João P. Vilela, and Marilia Curado. Triple-similarity mechanism for alarm management in the cloud. Computers & Security, 78(??):33–42, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306515.

daVeiga:2015:IIS

[dVM15] Adéle da Veiga and Nico Martins. Improving the information security culture through monitoring and implementation actions illustrated through a case study. Computers & Security, 49(??):162–176, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001862.

daVeiga:2017:DID

[dVM17] Adéle da Veiga and Nico Martins. Defining and identifying dominant information security cultures and subcultures. Computers & Security, 70(??):72–94, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300937.

Dong:2019:SCA

[DWF^{+}19] Xuewen Dong, Feng Wu, Anter Faree, Deke Guo, Yulong Shen, and Jianfeng Ma. Selfholding: a combined attack model using selfish mining with block withholding attack. Computers & Security, 87(??):Article 101584, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818313038.

Ding:2019:OBK

[DWZ19] Yuxin Ding, Rui Wu, and Xiao Zhang. Ontology-based knowledge representation for malware individuals and families. Computers & Security, 87(??):Article 101574, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301373.

Ding:2018:MDM

[DXCL18] Yuxin Ding, Xiaoling Xia, Sheng Chen, and Ye Li. A malware detection method based on family behavior graph. Computers & Security, 73(??):73–86, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302146.

Dong:2014:AES

[DYL^{+}14] Xin Dong, Jiadi Yu, Yuan Luo, Yingying Chen, Guangtao Xue, and Minglu Li. Achieving an effective, scalable and privacy-preserving data sharing service in cloud computing. Computers & Security, 42(??):151–164, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001703.

Ding:2013:FMD

[DYT^{+}13] Yuxin Ding, Xuebing Yuan, Ke Tang, Xiao Xiao, and Yibin Zhang. A fast malware detection algorithm based on objective-oriented association mining. Computers & Security, 39 (part B)(??):315–324, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001259.

Dong:2015:SSS

[DYZ^{+}15] Xin Dong, Jiadi Yu, Yanmin Zhu, Yingying Chen, Yuan Luo, and Minglu Li. SECO: Secure and scalable data collaboration services in cloud computing. Computers & Security, 50(??):91–105, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000048.

Elyas:2015:DFR

[EAML15] Mohamed Elyas, Atif Ahmad, Sean B. Maynard, and Andrew Lonie. Digital forensic readiness: Expert perspectives on a theoretical framework. Computers & Security, 52(??):70–89, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000449.

El-Emam:2015:NDH

[EE15] Nameer N. El-Emam. New data-hiding algorithm based on adaptive neural networks with modified particle swarm optimization. Computers & Security, 55(??):21–45, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481500098X.

Ehlert:2010:SNS

[EGM10] Sven Ehlert, Dimitris Geneiatakis, and Thomas Magedanz. Survey of network security systems to counter SIPbased denial-of-service attacks. Computers & Security, 29(2):225–243, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001060.

El-Hajj:2016:SCW

[EHBH^{+}16] Wassim El-Hajj, Ghassen Ben Brahim, Hazem Hajj, Haidar Safa, and Ralph Adaimy. Security-by-construction in web applications development via database annotations. Computers & Security, 59(??):151–165, June 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001972.

Evans:2019:HNT

[EHMJ19] Mark Evans, Ying He, Leandros Maglaras, and Helge Janicke. HEART-IS: a novel technique for evaluating human error-related information security incidents. Computers & Security, 80(??):74–89, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301615.

Edwards:2017:PGA

[ELG^{+}17] Matthew Edwards, Robert Larson, Benjamin Green, Awais Rashid, and Alistair Baron. Panning for gold: Automatically analysing online social engineering attack surfaces. Computers & Security, 69(??):18–34, August 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301845.

Elkhadir:2019:CNA

[EM19] Zyad Elkhadir and Benattou Mohammed. A cyber network attack detection based on GM median nearest neighbors LDA. Computers & Security, 86(??):63–74, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301142.

Elhadi:2014:EDM

[EMBH14] Ammar Ahmed E. Elhadi, Mohd Aizaini Maarof, Bazara I. A. Barry, and Hentabli Hamza. Enhancing the detection of metamorphic malware using call graphs. Computers & Security, 46(??):62–78, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001060.

Eckhart:2019:STP

[EMWE19] Matthias Eckhart, Kristof Meixner, Dietmar Winkler, and Andreas Ekelhart. Securing the testing process for industrial automation software. Computers & Security, 85(??):156–180, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818314482.

Eterovic-Soric:2017:SSD

[ESCAM17] Brett Eterovic-Soric, Kim-Kwang Raymond Choo, Helen Ashman, and Sameera Mubarak. Stalking the stalkers — detecting and deterring stalking behaviours using technology: a review. Computers & Security, 70(??):278–289, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481730127X.

Elish:2015:PUT

[ESY^{+}15] Karim O. Elish, Xiaokui Shu, Danfeng (Daphne) Yao, Barbara G. Ryder, and Xuxian Jiang. Profiling usertrigger dependence for Android malware detection. Computers & Security, 49(??):255–273, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001631.

Flores:2014:ISK

[FAE14] Waldo Rocha Flores, Egil Antonsen, and Mathias Ekstedt. Information security knowledge sharing in organizations: Investigating the effect of behavioral information security governance and national culture. Computers & Security, 43(??):90–110, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000339.

Furnell:2018:ESB

[FanE^{+}18] Steven Furnell, Warut Khern am nuai, Rawan Esmael, Weining Yang, and Ninghui Li. Enhancing security behaviour by supporting the user. Computers & Security, 75(??):1– 9, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818300385.

Feizollah:2017:AAA

[FAS^{+}17] Ali Feizollah, Nor Badrul Anuar, Rosli Salleh, Guillermo Suarez-Tangil, and Steven Furnell. AndroDialysis: Analysis of Android intent effectiveness in malware detection. Computers & Security, 65(??):121–134, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301602.

Franke:2014:CSA

[FB14] Ulrik Franke and Joel Brynielsson. Cyber situational awareness — a systematic review of the literature. Computers & Security, 46(??):18–31, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001011.

Fadolalkarim:2019:PAP

[FB19] Daren Fadolalkarim and Elisa Bertino. A-PANDDE: Advanced provenance-based ANomaly detection of data exfiltration. Computers & Security, 84(??):276–287, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300823.

Furnell:2012:PPE

[FC12] Steven Furnell and Nathan Clarke. Power to the people? The evolving recognition of human aspects of security. Computers & Security, 31(8):983–988, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001228.

Farina:2016:MBP

[FCPA16] Paolo Farina, Enrico Cambiaso, Gianluca Papaleo, and Maurizio Aiello. Are mobile botnets a possible threat? The case of SlowBot Net. Computers & Security, 58(??):268–283, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300086.

Fernandez:2010:NAI

[FCSD10] Marcel Fernandez, Josep Cotrina, Miguel Soriano, and Neus Domingo. A note about the identifier parent property in Reed–Solomon codes. Computers & Security, 29(5):628–635, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809001503.

Flores:2016:SIR

[FE16] Waldo Rocha Flores and Mathias Ekstedt. Shaping intention to resist social engineering through transformational leadership, information security culture and awareness. Computers & Security, 59(??):26–44, June 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816000067.

Fu:2018:CES

[FFGL18] Qiang Fu, Bo Feng, Dong Guo, and Qiang Li. Combating the evolving spammers in online social networks. Computers & Security, 72(??):60–73, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301797.

Ferrer-Gomila:2018:SEC

[FGHHR18] Josep-Lluís Ferrer-Gomila, M. Francisca Hinarejos, and Llorenç Huguet-Rotger. A survey on electronic coupons. Computers & Security, 77(??):106–127, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302499.

Ferrer-Gomilla:2010:CEM

[FGOPL10] Josep Lluis Ferrer-Gomilla, Jose A. Onieva, Magdalena Payeras, and Javier Lopez. Certified electronic mail: Properties revisited. Computers & Security, 29(2):167–179, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000704.

Fuchs:2015:WMM

[FHF15] Karl-Peter Fuchs, Dominik Herrmann, and Hannes Federrath. Workload modelling for mix-based anonymity services. Computers & Security, 52(??):221–233, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500019X.

Fang:2014:PBA

[FHL14] Zheran Fang, Weili Han, and Yingjiu Li. Permission based Android security: Issues and countermeasures. Computers & Security, 43(??):205–218, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000261.

Forne:2010:PAA

[FHM^{+}10] Jordi Forné, Francisca Hinarejos, Andrés Marín, Florina Almenárez, Javier Lopez, Jose A. Montenegro, Marc Lacoste, and Daniel Díaz. Pervasive authentication and authorization infrastructures for mobile users. Computers & Security, 29(4):501–514, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000911.

Fu:2013:VSP

[FK13] Yulong Fu and Ousmane Koné. Validation of security protocol implementations from security objectives. Computers & Security, 36(??):27–39, July 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000229.

Firoozjaei:2019:TOO

[FKSK19] Mahdi Daghmehchi Firoozjaei, MinChang Kim, JaeSeung Song, and Hyoungshick Kim. O^{2} TR: Offline OTR messaging system under network disruption. Computers & Security, 82(??):227–240, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818309076.

Fattori:2015:HBM

[FLBK15] Aristide Fattori, Andrea Lanzi, Davide Balzarotti, and Engin Kirda. Hypervisor-based malware protection with AccessMiner. Computers & Security, 52(??):33–50, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000395.

Farhi:2019:MNU

[FNE19] Nitzan Farhi, Nir Nissim, and Yuval Elovici. Malboard: a novel user keystroke impersonation attack and trusted detection framework based on side-channel analysis. Computers & Security, 85(??):240–269, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818309957.

Fuchs:2011:RIS

[FPS11] L. Fuchs, G. Pernul, and R. Sandhu. Roles in information security — a survey and classification of the research area.

Computers & Security, 30(8):748–769, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481100099X.

Franke:2017:CIM

[Fra17] Ulrik Franke. The cyber insurance market in Sweden. Computers & Security, 68(??):130–144, July 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300883.

Fuller:2017:MBD

[FRRP17] Jonathan D. Fuller, Benjamin W. Ramsey, Mason J. Rice, and John M. Pecarina. Misuse-based detection of Z-Wave network attacks. Computers & Security, 64(??):44–58, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816301341.

Friedberg:2015:CAP

[FSSF15] Ivo Friedberg, Florian Skopik, Giuseppe Settanni, and Roman Fiedler. Combating advanced persistent threats: From network event correlation to incident detection. Computers & Security, 48(??):35–57, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001461.

Fan:2017:DPI

[FSZ^{+}17] Wenhao Fan, Yaohui Sang, Daishuai Zhang, Ran Sun, and Yuan’an Liu. DroidInjector: a process injection-based dynamic tracking system for runtime behaviors of Android applications. Computers & Security, 70(??):224–237, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301207.

Flowerday:2016:ISP

[FT16] Stephen V. Flowerday and Tite Tuyikeze. Information security policy development and implementation: the what, how and who. Computers & Security, 61(??):169–183, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300670.

Farley:2010:RBD

[FW10] Ryan Farley and Xinyuan Wang. Roving bugnet: Distributed surveillance threat and mitigation. Computers & Security, 29(5):592–602, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001400.

Fu:2018:FFN

[FXC18] Bo Fu, Yang Xiao, and Hui Chen. FNF: Flow-net based fingerprinting and its applications. Computers & Security, 75(??):167–181, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300877.

Fu:2017:SSB

[FYL^{+}17] Yulong Fu, Zheng Yan, Hui Li, Xiao Long Xin, and Jin Cao. A secure SDN based multi-RANs architecture for future 5G networks. Computers & Security, 70(??):648–662, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301785.

Fu:2012:EHA

[FZZ^{+}12] Anmin Fu, Yuqing Zhang, Zhenchao Zhu, Qi Jing, and Jingyu Feng. An efficient handover authentication scheme with privacy preservation for IEEE 802.16m network. Computers & Security, 31(6):741–749, September 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001009.

Geniola:2018:AAF

[GAA18] Alberto Geniola, Markku Antikainen, and Tuomas Aura. Automated analysis of freeware installers promoted by download portals. Computers & Security, 77(??):209–225, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818302797.

Garcia-Alfaro:2013:MSF

[GACCB^{+}13] Joaquin Garcia-Alfaro, Frédéric Cuppens, Nora Cuppens-Boulahia, Salvador Martinez, and Jordi Cabot. Management of stateful firewall misconfiguration. Computers & Security, 39 (Part A)(??):64–85, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000217.

Garfinkel:2013:DMT

[Gar13] Simson L. Garfinkel. Digital media triage with bulk data analysis and bulk extractor. Computers & Security, 32(??):56–72, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001472.

Garcia:2016:FVS

[Gar16] D. A. López García. A flexible e-voting scheme for debate tools. Computers & Security, 56(??):50–62, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001546.

Guri:2019:AJC

[GB19] Mordechai Guri and Dima Bykhovsky. aIR-Jumper: Covert air-gap exfiltration/infiltration via security cameras and infrared (IR). Computers & Security, 82(??):15–29, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307193.

Gratian:2018:CHT

[GBC^{+}18] Margaret Gratian, Sruthi Bandi, Michel Cukier, Josiah Dykstra, and Amy Ginther. Correlating human traits and cyber security behavior intentions. Computers & Security, 73(??):345–358, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302523.

Gupta:2019:DRB

[GBC19] Sandeep Gupta, Attaullah Buriro, and Bruno Crispo. Driver-Auth: a risk-based multi-modal biometric-based driver authentication scheme for ride-sharing platforms. Computers & Security, 83(??):122–139, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818310113.

Gratian:2019:IIU

[GBCD19] Margaret Gratian, Darshan Bhansali, Michel Cukier, and Josiah Dykstra. Identifying infected users via network traffic. Computers & Security, 80(??):306–316, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306680.

Guerrero:2018:SSS

[GCC18] Daniel Guerrero, Alin A. Carsteanu, and Julio B. Clempner. Solving Stackelberg security Markov games employing the bargaining Nash approach: Convergence analysis. Computers & Security, 74(??):240–257, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300221.

Gottwalt:2019:CFS

[GCD19] Florian Gottwalt, Elizabeth Chang, and Tharam Dillon. CorrCorr: a feature selection method for multivariate correlation network anomaly detection techniques. Computers & Security, 83(??):234–245, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306485.

Gritzalis:2010:Eb

[GD10] Dimitris Gritzalis and Sabrina De Capitani di Vimercati. Editorial. Computers & Security, 29(2):165–166, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001424.

Giot:2015:RPB

[GDR15] Romain Giot, Bernadette Dorizzi, and Christophe Rosenberger. A review on the public benchmark databases for static keystroke dynamics. Computers & Security, 55(??):46–61, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000942.

Gritzalis:2010:Ea

[GE10] Dimitris Gritzalis and Jan Eloff. Editorial. Computers & Security, 29(1):1–2, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001138.

Giot:2011:UKD

[GEAHR11] Romain Giot, Mohamad El-Abed, Baptiste Hemery, and Christophe Rosenberger. Unconstrained keystroke dynamics authentication with shared secret. Computers & Security, 30(6–7):427–445, September/October 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000502.

Geneiatakis:2015:PVA

[GFKS15] Dimitris Geneiatakis, Igor Nai Fovino, Ioannis Kounelis, and Paquale Stirparo. A permission verification approach for Android mobile applications. Computers & Security, 49(??):192–205, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001515.

Gu:2011:SSP

[GFN11] Qijun Gu, Christopher Ferguson, and Rizwan Noorani. A study of self-propagating mal-packets in sensor networks: Attacks and defenses. Computers & Security, 30(1):13–27, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000854.

GhasemiGol:2016:CAN

[GGBT16] Mohammad GhasemiGol, Abbas Ghaemi-Bafghi, and Hassan Takabi. A comprehensive approach for network attack forecasting. Computers & Security, 58(??):83–105, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001819.

Garcia:2014:ECB

[GGSZ14] S. García, M. Grill, J. Stiborek, and A. Zunino. An empirical comparison of botnet detection methods. Computers & Security, 45(??):100–123, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000923.

Gerber:2018:EPP

[GGV18] Nina Gerber, Paul Gerber, and Melanie Volkamer. Explaining the privacy paradox: a systematic review of literature investigating privacy attitude and behavior. Computers & Security, 77(??):226–261, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303031.

Gope:2015:RLA

[GH15] Prosanta Gope and Tzonelih Hwang. A realistic lightweight authentication protocol preserving strong anonymity for securing RFID system. Computers & Security, 55(??):271–280, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000681.

Gomez-Hernandez:2018:RLT

[GHÁGGT18] J. A. Gómez-Hernández, L. Álvarez-González, and P. García-Teodoro. R-Locker: Thwarting ransomware action through a honeyfile-based approach. Computers & Security, 73(??): 389–398, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404817302560.

Guerrero-Higueras:2017:EAC

[GHDGRLM17] Ángel Manuel Guerrero-Higueras, Noemí DeCastro-García, Francisco Javier Rodríguez-Lera, and Vicente Matellán. Empirical analysis of cyber-attacks to an indoor real time localization system for autonomous robots. Computers & Security, 70(??):422–435, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301396.

Guo:2017:SPA

[GHL^{+}17] Zeqing Guo, Weili Han, Liangxing Liu, Wenyuan Xu, Minyue Ni, Yunlei Zhao, and Xiaoyang Sean Wang. Socialized policy administration. Computers & Security, 67(??):232–243, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300585.

Gritzaliz:2011:E

[GJ11] Dimitris Gritzaliz and James Joshi. Editorial. Computers & Security, 30(2–3):89–90, March 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810001094.

Goel:2018:MPA

[GJ18] Diksha Goel and Ankit Kumar Jain. Mobile phishing attacks and defence mechanisms: State of art and open research challenges. Computers & Security, 73(??):519–544, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302717.

Geetha:2018:OVC

[GJJ18] P. Geetha, V. S. Jayanthi, and A. N. Jayanthi. Optimal visual cryptographic scheme with multiple share creation for multimedia applications. Computers & Security, 78(??):301–320, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308241.

Gowtham:2014:CEA

[GK14] R. Gowtham and Ilango Krishnamurthi. A comprehensive and efficacious architecture for detecting phishing webpages. Computers & Security, 40(??):23–37, February 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001442.

Gupta:2014:FPM

[GKB14] Aditi Gupta, Michael S. Kirkpatrick, and Elisa Bertino. A formal proximity model for RBAC systems. Computers & Security, 41(??):52–67, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001296.

Gritzalis:2010:Ed

[GL10] Dimitris Gritzalis and Javier Lopez. Editorial. Computers & Security, 29(5):521–522, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000337.

Gao:2018:SUP

[GL18] Tianchong Gao and Feng Li. Studying the utility preservation in social network anonymization via persistent homology. Computers & Security, 77(??):49–64, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303122.

Gouglidis:2012:DAC

[GM12] Antonios Gouglidis and Ioannis Mavridis. domRBAC: an access control model for modern collaborative systems. Computers & Security, 31(4):540–556, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000144.

Gonzalez-Manzano:2019:DRO

[GMdF19] Lorena González-Manzano and Jose M. de Fuentes. Design recommendations for online cybersecurity courses. Computers & Security, 80(??):238–256, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302050.

Gonzalez-Manzano:2014:CCO

[GMGTdFR14a] Lorena González-Manzano, Ana I. González-Tablas, José M. de Fuentes, and Arturo Ribagorda. CooPeD: Co-owned personal data management. Computers & Security, 47(??): 41–65, November 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404814000960.

Gonzalez-Manzano:2014:SEU

[GMGTdFR14b] Lorena González-Manzano, Ana I. González-Tablas, José M. de Fuentes, and Arturo Ribagorda. SoNeUCON_{ABC}, an expressive usage control model for Web-based social networks. Computers & Security, 43(??):159–187, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000480.

Guo:2011:ENI

[GMLZ11] Hua Guo, Yi Mu, Zhoujun Li, and Xiyong Zhang. An efficient and non-interactive hierarchical key agreement protocol. Computers & Security, 30(1):28–34, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000866.

Gunson:2011:UPS

[GMMJ11] Nancie Gunson, Diarmid Marshall, Hazel Morton, and Mervyn Jack. User perceptions of security and usability of single-factor and two-factor authentication in automated telephone banking. Computers & Security, 30(4):208–220, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810001148.

Guerar:2018:ICU

[GMMP18] Meriem Guerar, Alessio Merlo, Mauro Migliardi, and Francesco Palmieri. Invisible CAPPCHA: a usable mechanism to distinguish between malware and humans on the mobile IoT. Computers & Security, 78(??):255–266, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818307557.

Gascon:2011:AUD

[GOB11] Hugo Gascon, Agustin Orfila, and Jorge Blasco. Analysis of update delays in signature-based network intrusion detection systems. Computers & Security, 30(8):613–624, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811001106.

Gal-Oz:2019:MMR

[GOGG19] Nurit Gal-Oz, Yaron Gonen, and Ehud Gudes. Mining meaningful and rare roles from web application usage patterns. Computers & Security, 82(??):296–313, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301202.

Gope:2019:LLA

[Gop19] Prosanta Gope. LAAP: Lightweight anonymous authentication protocol for D2D-aided fog computing paradigm. Computers & Security, 86(??):223–237, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831160X.

Goseva-Popstojanova:2014:CCM

[GPAD^{+}14] Katerina Goseva-Popstojanova, Goce Anastasovski, Ana Dimitrijevikj, Risto Pantev, and Brandon Miller. Characterization and classification of malicious Web traffic. Computers & Security, 42(??):92–115, May 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000170.

Giboney:2016:SEA

[GPGV16] Justin Scott Giboney, Jeffrey Gainer Proudfoot, Sanjay Goel, and Joseph S. Valacich. The security expertise assessment measure (SEAM): Developing a scale for hacker expertise. Computers & Security, 60(??):37–51, July 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300323.

Gritzalis:2010:Ec

[GS10] Dimitris Gritzalis and Pierangela Samarati. Editorial. Computers & Security, 29(4):381–382, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000076.

Garg:2018:EAA

[GSA18] Urvashi Garg, Geeta Sikka, and Lalit K. Awasthi. Empirical analysis of attack graphs for mitigating critical paths and vulnerabilities. Computers & Security, 77(??):349–359, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818303493.

Gutierrez:2018:RRD

[GSBY18] Christopher N. Gutierrez, Eugene H. Spafford, Saurabh Bagchi, and Thomas Yurek. Reactive redundancy for data destruction protection (R2D2). Computers & Security, 74(??):184–201, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730281X.

Guan:2019:DAI

[GSL^{+}19] Chong Guan, Kun Sun, Lingguang Lei, Pingjian Wang, Yuewu Wang, and Wei Chen. DangerNeighbor attack: Information leakage via postMessage mechanism in HTML5. Computers & Security, 80(??):291–305, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308484.

Gu:2018:GTH

[GSZX18] Zhongshu Gu, Brendan Saltaformaggio, Xiangyu Zhang, and Dongyan Xu. G emini: Guest-transparent honey files via hypervisor-level access redirection. Computers & Security, 77(??):737–744, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818301433.

Gritzalis:2013:CDE

[GT13] Dimitris Gritzalis and Gurvirender Tejay. Cybercrime in the digital economy — editorial. Computers & Security, 38(??):1–2, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001120.

Garcia-Teodoro:2015:AGH

[GTDVTSH15] P. Garcia-Teodoro, J. E. Diaz-Verdejo, J. E. Tapiador, and R. Salazar-Hernandez. Automatic generation of HTTP intrusion signatures by selective identification of anomalies. Computers & Security, 55(??):159–174, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001297.

GhasemiGol:2016:FMI

[GTGB16] Mohammad GhasemiGol, Hassan Takabi, and Abbas Ghaemi-Bafghi. A foresight model for intrusion response management. Computers & Security, 62(??):73–94, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300700.

Guo:2013:SRB

[Guo13] Ken H. Guo. Security-related behavior in using information systems in the workplace: a review and synthesis. Computers & Security, 32(??):242–251, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001666.

Grothoff:2018:TSN

[GWEA18] Christian Grothoff, Matthias Wachs, Monika Ermert, and Jacob Appelbaum. Toward secure name resolution on the Internet. Computers & Security, 77(??):694–708, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300403.

Gheisari:2019:CAP

[GWKFC19] Mehdi Gheisari, Guojun Wang, Wazir Zada Khan, and Christian Fernández-Campusano. A context-aware privacypreserving method for IoT-based smart city using Software Defined Networking. Computers & Security, 87 (??):Article 101470, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313336.

Gezer:2019:FBA

[GWWS19] Ali Gezer, Gary Warner, Clifford Wilson, and Prakash Shrestha. A flow-based approach for Trickbot banking Trojan detection. Computers & Security, 84(??):179–192, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818309568.

Gu:2019:NAI

[GWWW19] Jie Gu, Lihong Wang, Huiwen Wang, and Shanshan Wang. A novel approach to intrusion detection using SVM ensemble with feature augmentation. Computers & Security, 86(??):53–62, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301154.

Guinde:2010:EHS

[GZ10] Nitesh B. Guinde and Sotirios G. Ziavras. Efficient hardware support for pattern matching in network intrusion detection. Computers & Security, 29(7):756–769, October 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000350.

Guo:2018:LLP

[GZ18] Yimin Guo and Zhenfeng Zhang. LPSE: Lightweight password-strength estimation for password meters. Computers & Security, 73(??):507–518, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic).  URL https://www.sciencedirect.com/
science/article/pii/S0167404817301530. See corrigen-

dum [GZ20].

Guo:2020:CLL

[GZ20] Yimin Guo and Zhenfeng Zhang. Corrigendum to “LPSE: Lightweight password-strength estimation for password meters” [Computers and Security, Volume 73, 2018, Pages 507– 518]. Computers & Security, 94(??):Article 101879, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301528. See [GZ18].

Guo:2019:ONP

[GZG19] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. Optiwords: a new password policy for creating memorable and strong passwords. Computers & Security, 85(??):423–435, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301105. See [GZG20].

Guo:2020:CON

[GZG20] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. Corrigendum to “Optiwords: a new password policy for creating memorable and strong password” [Computers and Security, Volume 85, 2019, Pages 423–435]. Computers & Security, 94(??):Article 101880, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482030153X. See [GZG19].

Guo:2013:EID

[GZP^{+}13] Chun Guo, Ya-Jian Zhou, Yuan Ping, Shou-Shan Luo, Yu-Ping Lai, and Zhong-Kun Zhang. Efficient intrusion detection using representative instances. Computers & Security, 39 (part B)(??):255–267, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481300120X.

Genga:2019:DRE

[GZS19] Laura Genga, Nicola Zannone, and Anna Squicciarini. Discovering reliable evidence of data misuse by exploiting rule redundancy. Computers & Security, 87(??):Article 101577, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818302025.

Hamid:2014:APP

[HA14] Isredza Rahmi A. Hamid and Jemal H. Abawajy. An approach for profiling phishing activities. Computers & Security, 45(??):27–41, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000546.

Holm:2015:EBI

[HA15] Hannes Holm and Khalid Khan Afridi. An expert-based investigation of the common vulnerability scoring system. Computers & Security, 53(??):18–30, September 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000620.

Hernandez-Ardieta:2013:TSA

[HAGTdFR13] Jorge L. Hernandez-Ardieta, Ana I. Gonzalez-Tablas, Jose M. de Fuentes, and Benjamin Ramos. A taxonomy and survey of attacks on digital signatures. Computers & Security, 34(??):67–112, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001794.

Hussain:2018:SFM

[HAHZ^{+}18] Muzammil Hussain, Ahmed Al-Haiqi, A. A. Zaidan, B. B. Zaidan, M. Kiah, Salman Iqbal, S. Iqbal, and Mohamed Abdulnabi. A security framework for mHealth apps on Android platform. Computers & Security, 75(??):191–217, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300798.

Hatfield:2018:SEC

[Hat18] Joseph M. Hatfield. Social engineering in cybersecurity: the evolution of a concept. Computers & Security, 73(??):102–113, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302249.

Hatfield:2019:VHH

[Hat19] Joseph M. Hatfield. Virtuous human hacking: the ethics of social engineering in penetration-testing. Computers & Security, 83(??):354–366, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481831174X.

Hellal:2016:MCF

[HB16] Aya Hellal and Lotfi Ben Romdhane. Minimal contrast frequent pattern mining for malware detection. Computers & Security, 62(??):19–32, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300694.

Halabi:2018:BBF

[HB18] Talal Halabi and Martine Bellaiche. A broker-based framework for standardization and management of cloud Security-SLAs. Computers & Security, 75(??):59–71, June 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300476.

Hasan:2012:PPF

[HBB12] Omar Hasan, Lionel Brunie, and Elisa Bertino. Preserving privacy of feedback providers in decentralized reputation systems. Computers & Security, 31(7):816–826, October 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001465.

Harichandran:2016:CFN

[HBBM16] Vikram S. Harichandran, Frank Breitinger, Ibrahim Baggili, and Andrew Marrington. A cyber forensics needs analysis survey: Revisiting the domain’s needs a decade later. Computers & Security, 57(??):1–13, March 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001595.

Bang:2017:ADN

[hBCK17] June ho Bang, Young-Jong Cho, and Kyungran Kang. Anomaly detection of network-initiated LTE signaling traffic in wireless sensor and actuator networks based on a hidden semi-Markov model. Computers & Security, 65(??):108–120, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301614.

Herrmann:2013:BBT

[HBF13] Dominik Herrmann, Christian Banse, and Hannes Federrath. Behavior-based tracking: Exploiting characteristic patterns in DNS traffic. Computers & Security, 39 (Part A)(??):17–33, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000576.

Henriksen-Bulmer:2019:PRA

[HBFJ19] Jane Henriksen-Bulmer, Shamal Faily, and Sheridan Jeary. Privacy risk assessment in context: a meta-model based on contextual integrity. Computers & Security, 82(??):270–283, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818301998.

Hassan:2018:STB

[HBHA18] Shaikh Shahriar Hassan, Soumik Das Bibon, Md Shohrab Hossain, and Mohammed Atiquzzaman. Security threats in Bluetooth technology. Computers & Security, 74(??):308– 322, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817300615.

Horak:2019:OHS

[HBT^{+}19] Karel Horák, Branislav Bosanský, Petr Tomásek, Christopher Kiekintveld, and Charles Kamhoua. Optimizing honeypot strategies against dynamic lateral movement using partially observable stochastic games. Computers & Security, 87(??):Article 101579, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300665.

Hernandez-Castro:2010:PCD

[HCR10] Carlos Javier Hernandez-Castro and Arturo Ribagorda. Pitfalls in CAPTCHA design and implementation: the math CAPTCHA, a case study. Computers & Security, 29(1):141–157, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000728.

Hernandez-Castro:2017:UML

[HCRMBG17] Carlos Javier Hernández-Castro, María D. R-Moreno, David F. Barrero, and Stuart Gibson. Using machine learning to identify common flaws in CAPTCHA design: FunCAPTCHA case analysis. Computers & Security, 70(??):744–756, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301128.

Hamed:2018:NID

[HDK18] Tarfa Hamed, Rozita Dara, and Stefan C. Kremer. Network intrusion detection system based on recursive feature addition and bigram technique. Computers & Security, 73(??):137–155, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302274.

Hilgenstieler:2010:ESP

[HDMKS10] Egon Hilgenstieler, Elias P. Duarte, Glenn Mansfield-Keeni, and Norio Shiratori. Extensions to the source path isolation engine for precise and efficient log-based IP traceback. Computers & Security, 29(4):383–392, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001497.

Heigl:2019:RPS

[HDT^{+}19] Michael Heigl, Laurin Doerr, Nicolas Tiefnig, Dalibor Fiala, and Martin Schramm. A resource-preserving self-regulating uncoupled MAC algorithm to be applied in incident detection. Computers & Security, 85(??):270–287, August 2019.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301099.

Hong:2018:DSM

[HEK^{+}18] Jin B. Hong, Simon Yusuf Enoch, Dong Seong Kim, Armstrong Nhlabatsi, Noora Fetais, and Khaled M. Khan. Dynamic security metrics for measuring the effectiveness of moving target defense techniques. Computers & Security, 79(??):33–52, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304449.

Herrmann:2017:EII

[HF17] Dominik Herrmann and Hannes Federrath. Editorial: 30th IFIP International Information Security Conference (IFIP SEC 2015). Computers & Security, 67(??):266, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730069X.

Hocking:2013:COU

[HFCR13] C. G. Hocking, S. M. Furnell, N. L. Clarke, and P. L. Reynolds. Co-operative user identity verification using an authentication aura. Computers & Security, 39 (part B)(??):486–502, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001417.

Huber:2019:FCL

[HHAW19] Manuel Huber, Julian Horsch, Junaid Ali, and Sascha Wessel. Freeze and crypt: Linux kernel support for main memory encryption. Computers & Security, 86(??):420–436, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818310435.

Hammi:2018:BTD

[HHBS18] Mohamed Tahar Hammi, Badis Hammi, Patrick Bellot, and Ahmed Serhrouchni. Bubbles of trust: a decentralized blockchain-based authentication system for IoT. Computers & Security, 78(??):126–142, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300890.

Hu:2019:TRF

[HHK^{+}19] Hongxin Hu, Wonkyu Han, Sukwha Kyung, Juan Wang, Gail-Joon Ahn, Ziming Zhao, and Hongda Li. Towards a reliable firewall for software-defined networks. Computers & Security, 87(??):Article 101597, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481930152X.

Haenni:2013:GAP

[HK13] Rolf Haenni and Reto E. Koenig. A generic approach to prevent board flooding attacks in coercion-resistant electronic voting schemes. Computers & Security, 33(??):59–69, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001381.

Hasan:2017:UAF

[HK17] Ragib Hasan and Rasib Khan. Unified authentication factors and fuzzy service access using interaction provenance. Computers & Security, 67(??):211–231, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300408.

Han:2017:IMI

[HKK17] JinYoung Han, Yoo Jung Kim, and Hyungjin Kim. An integrative model of information security policy compliance with psychological contract: Examining a bilateral perspective. Computers & Security, 66(??):52–65, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300032.

Hsieh:2012:EHF

[HL12] Wen-Bin Hsieh and Jenq-Shiou Leu. Exploiting hash functions to intensify the remote user authentication scheme.

Computers & Security, 31(6):791–798, September 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000910.

Heartfield:2018:DSS

[HL18] Ryan Heartfield and George Loukas. Detecting semantic social engineering attacks with the weakest link: Implementation and empirical evaluation of a human-as-a-security-sensor framework. Computers & Security, 76(??):101–127, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818301780.

Heartfield:2018:TCP

[HLB^{+}18] Ryan Heartfield, George Loukas, Sanja Budimir, Anatolij Bezemskij, Johnny R. J. Fontaine, Avgoustinos Filippoupolitis, and Etienne Roesch. A taxonomy of cyberphysical threats and impact in the smart home. Computers & Security, 78(??):398–428, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304875.

Heydemann:2019:FVS

[HLB19] Karine Heydemann, Jean-François Lalande, and Pascal Berthomé. Formally verified software countermeasures for control-flow integrity of smart card C code. Computers & Security, 85(??):202–224, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302037.

Huang:2016:SWS

[HLFZ16] Cheng Huang, JiaYong Liu, Yong Fang, and Zheng Zuo. A study on Web security incidents in China by analyzing vulnerability disclosure platforms. Computers & Security, 58(??):47–62, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001820.

Huang:2017:ESI

[HLL17] Xinyi Huang, Joseph Liu, and Javier Lopez. Editorial: Special issue on security data science and cyber threat management. Computers & Security, 69(??):1–2, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300949.

Hashemi:2019:SDN

[HM19] Atiye Sadat Hashemi and Saeed Mozaffari. Secure deep neural networks using adversarial image generation and training with Noise-GAN. Computers & Security, 86(??):372–387, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481930121X.

Hinarejos:2010:PEC

[HMFE10] M. Francisca Hinarejos, Jose L. Muñoz, Jordi Forné, and Oscar Esparza. PREON: an efficient cascade revocation mechanism for delegation paths. Computers & Security, 29(6):697–711, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000222.

Herranz:2010:CDP

[HMNT10] Javier Herranz, Stan Matwin, Jordi Nin, and Vicenç Torra. Classifying data from protected statistical datasets. Computers & Security, 29(8):875–890, November 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000507.

Hong:2018:PBI

[HNSD18] Sanghyun Hong, Alina Nicolae, Abhinav Srivastava, and Tudor Dumitras. Peek-a-boo: Inferring program behaviors in a virtualized infrastructure without introspection. Computers & Security, 79(??):190–207, November 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481830244X.

Holm:2012:PAN

[Hol12] Hannes Holm. Performance of automated network vulnerability scanning at remediating security issues. Computers & Security, 31(2):164–175, March 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001696.

Horsman:2016:CSR

[Hor16] Graeme Horsman. The challenges surrounding the regulation of anonymous communication provision in the United Kingdom. Computers & Security, 56(??):151–162, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000917.

Horsman:2018:FRE

[Hor18] Graeme Horsman. Framework for reliable experimental design (FRED): a research framework to ensure the dependable interpretation of digital data for digital forensics. Computers & Security, 73(??):294–306, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302468.

Ho:2017:TCC

[HOVB17] Shuyuan Mary Ho, Mónica Ocasio-Velázquez, and Cheryl Booth. Trust or consequences? Causal effects of perceived risk and subjective norms on cloud technology adoption. Computers & Security, 70(??):581–595, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301591.

Horcas:2016:ADM

[HPF^{+}16] Jose-Miguel Horcas, Mónica Pinto, Lidia Fuentes, Wissam Mallouli, and Edgardo Montes de Oca. An approach for deploying and monitoring dynamic security policies. Computers & Security, 58(??):20–38, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001832.

Hadlington:2019:ERW

[HPJ^{+}19] Lee Hadlington, Masa Popovac, Helge Janicke, Iryna Yevseyeva, and Kevin Jones. Exploring the role of work identity and work locus of control in information security awareness. Computers & Security, 81(??):41–48, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308897.

Hamedheidari:2013:NAB

[HR13] Sina Hamedheidari and Reza Rafeh. A novel agent-based approach to detect sinkhole attacks in wireless sensor networks. Computers & Security, 37(??):1–14, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000783.

Roh:2013:WBC

[hRKRR13] Byeong hee Roh, Ju Wan Kim, Ki-Yeol Ryu, and Jea-Tek Ryu. A whitelist-based countermeasure scheme using a Bloom filter against SIP flooding attacks. Computers & Security, 37(??):46–61, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000771.

Huang:2011:MDU

[HS11] Lin Huang and Mark Stamp. Masquerade detection using profile hidden Markov models. Computers & Security, 30(8):732–747, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001003.

Hassanzadeh:2013:OCI

[HS13] Amin Hassanzadeh and Radu Stoleru. On the optimality of cooperative intrusion detection for resource constrained wireless networks. Computers & Security, 34(??):16–35, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000060.

Habler:2018:ULE

[HS18] Edan Habler and Asaf Shabtai. Using LSTM encoderdecoder algorithm for detecting anomalous ADS-B messages. Computers & Security, 78(??):155–173, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303729.

Hina:2019:IGP

[HSL19] Sadaf Hina, Dhanapal Durai Dominic Panneer Selvam, and Paul Benjamin Lowry. Institutional governance and protection motivation: Theoretical insights into shaping employees’ security compliance behavior in higher education institutions in the developing world. Computers & Security, 87(??):Article 101594, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308320.

Hassanzadeh:2014:RTA

[HSPX14] Amin Hassanzadeh, Radu Stoleru, Michalis Polychronakis, and Geoffrey Xie. RAPID: Traffic-agnostic intrusion detection for resource-constrained wireless mesh networks. Computers & Security, 46(??):1–17, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001047.

Hatamian:2019:RUM

[HSR19] Majid Hatamian, Jetzabel Serna, and Kai Rannenberg. Revealing the unrevealed: Mining smartphone users privacy perception on app markets. Computers & Security, 83(??):332–353, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313051.

Hudic:2017:SAA

[HSW17] Aleksandar Hudic, Paul Smith, and Edgar R. Weippl. Security assurance assessment methodology for hybrid clouds. Computers & Security, 70(??):723–743, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300627.

Hjorth:2012:TDS

[HT12a] Theis Solberg Hjorth and Rune Torbensen. Trusted domain: a security platform for home automation. Computers & Security, 31(8):940–955, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001046.

Huang:2012:FQE

[HT12b] Yu-Lun Huang and Hsin-Yi Tsai. A framework for quantitative evaluation of parallel control-flow obfuscation. Computers & Security, 31(8):886–896, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001198.

Hubballi:2017:CLD

[HT17] Neminath Hubballi and Nikhil Tripathi. A closer look into DHCP starvation attack in wireless networks. Computers & Security, 65(??):387–404, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301262.

Hudic:2012:PCC

[HW12] Aleksandar Hudic and Edgar Weippl. Private cloud computing: Consolidation, virtualization, and service-oriented infrastructure. Computers & Security, 31(4):629, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000612.

Hamid:2018:ESS

[HW18] Brahim Hamid and Donatus Weber. Engineering secure systems: Models, patterns and empirical validation. Computers & Security, 77(??):315–348, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303043.

Hammad:2019:PSF

[HW19] Mohamed Hammad and Kuanquan Wang. Parallel score fusion of ECG and fingerprint for human authentication based on convolution neural network. Computers & Security, 81(??):107–122, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308411.

Holbl:2010:TPI

[HWB10] Marko Hölbl, Tatjana Welzer, and Bostjan Brumen. Two proposed identity-based three-party authenticated key agreement protocols from pairings. Computers & Security, 29(2):244–252, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480900090X.

Heckman:2013:ACD

[HWS^{+}13] Kristin E. Heckman, Michael J. Walsh, Frank J. Stech, Todd A. O’Boyle, Stephen R. DiCato, and Audra F. Herber. Active cyber defense with denial and deception: a cyberwargame experiment. Computers & Security, 37(??):72–77, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481300076X.

Han:2019:RRO

[HWX^{+}19] Xiaohui Han, Lianhai Wang, Shujiang Xu, Dawei Zhao, and Guangqi Liu. Recognizing roles of online illegal gambling participants: an ensemble learning approach. Computers & Security, 87(??):Article 101588, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301415.

Han:2017:NNT

[HX17] Wenlin Han and Yang Xiao. NFD: Non-technical loss fraud detection in smart grid. Computers & Security, 65(??):187–201, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301626.

Hassanzadeh:2016:PPI

[HXS^{+}16] Amin Hassanzadeh, Zhaoyan Xu, Radu Stoleru, Guofei Gu, and Michalis Polychronakis. PRIDE: a practical intrusion detection system for resource constrained wireless mesh networks. Computers & Security, 62(??):114–132, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300724.

Han:2019:MDE

[HXW^{+}19] Weijie Han, Jingfeng Xue, Yong Wang, Lu Huang, Zixiao Kong, and Limin Mao. MalDAE: Detecting and explaining malware based on correlation and fusion of static and dynamic characteristics. Computers & Security, 83(??):208– 233, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481831246X.

Hu:2019:AMP

[HYLY19] Jen-Wei Hu, Lo-Yao Yeh, Shih-Wei Liao, and Chu-Sing Yang. Autonomous and malware-proof blockchain-based firmware update platform with efficient batch verification for Internet of Things devices. Computers & Security, 86(??):238–252, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481831438X.

Hu:2018:TPP

[HYZ18] Zhaowei Hu, Jing Yang, and Jianpei Zhang. Trajectory privacy protection method based on the time interval divided. Computers & Security, 77(??):488–499, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304899.

Hu:2019:BBA

[HZG19] Kexin Hu, Zhenfeng Zhang, and Kaiwen Guo. Breaking the binding: Attacks on the Merkle approach to prove liabilities and its applications. Computers & Security, 87 (??):Article 101585, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/

S0167404818314093. See corrigendum [HZG20].

Hu:2020:CBB

[HZG20] Kexin Hu, Zhenfeng Zhang, and Kaiwen Guo. Corrigendum to “Breaking the binding: Attacks on the Merkle approach to prove liabilities and its applications” [Computers and Security, Volume 87, 2019, 101585]. Computers & Security, 94(??):Article 101878, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301516. See [HZG19].

Huang:2018:MCA

[HZH^{+}18] Tianqiang Huang, Xueli Zhang, Wei Huang, Lingpeng Lin, and Weifeng Su. A multi-channel approach through fusion of audio for detecting video inter-frame forgery. Computers & Security, 77(??):412–426, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304243.

Hu:2018:TSS

[HZMH18] Qiao Hu, Jingyi Zhang, Aikaterini Mitrokotsa, and Gerhard Hancke. Tangible security: Survey of methods supporting secure ad-hoc connects of edge devices with physical context. Computers & Security, 78(??):281–300, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307739.

Heurix:2015:TPE

[HZNF15] Johannes Heurix, Peter Zimmermann, Thomas Neubauer, and Stefan Fenz. A taxonomy for privacy enhancing technologies. Computers & Security, 53(??):1–17, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000668.

Ibrahim:2019:RAM

[IAA^{+}19] Tahir Musa Ibrahim, Shafi’i Muhammad Abdulhamid, Ala Abdusalam Alarood, Haruna Chiroma, Mohammed Ali Al-garadi, Nadim Rana, Amina Nuhu Muhammad, Adamu Abubakar, Khalid Haruna, and Lubna A. Gabralla. Recent advances in mobile touch screen security authentication methods: a systematic literature review. Computers & Security, 85(??):1–24, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818305030.

Idrus:2014:SBK

[ICRB14] Syed Zulkarnain Syed Idrus, Estelle Cherrier, Christophe Rosenberger, and Patrick Bours. Soft biometrics for keystroke dynamics: Profiling individuals while typing passwords. Computers & Security, 45(??):147–155, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000893.

Ifinedo:2012:UIS

[Ifi12] Princely Ifinedo. Understanding information systems security policy compliance: an integration of the theory of planned behavior and the protection motivation theory. Computers & Security, 31(1):83–95, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001337.

Ismail:2017:PRT

[IGS17] Hatem Ismail, Daniel Germanus, and Neeraj Suri. P2P routing table poisoning: a quorum-based sanitizing approach. Computers & Security, 65(??):283–299, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630178X.

Ibrahim:2019:TCC

[IH19] Fady A. M. Ibrahim and Elsayed E. Hemayed. Trusted cloud computing architectures for infrastructure as a service: Survey and systematic literature review. Computers & Security, 82(??):196–226, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302712.

Ibrahim:2015:RJS

[IHED15] Doaa H. Ibrahim, Emad S. Hassan, and Sami A. El-Dolil. Relay and jammer selection schemes for improving physical layer security in two-way cooperative networks. Computers & Security, 50(??):47–59, May 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000140.

Idrees:2017:PNA

[IRC^{+}17] Fauzia Idrees, Muttukrishnan Rajarajan, Mauro Conti, Thomas M. Chen, and Yogachandran Rahulamathavan. PIndroid: a novel Android malware detection system using ensemble learning methods. Computers & Security, 68(??):36– 46, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817300640.

Junior:2019:BSP

[JAAS19] Edson Floriano S. Junior, Eduardo Alchieri, Diego F. Aranha, and Priscila Solis. Building secure protocols for extensible distributed coordination through secure extensions. Computers & Security, 87(??):Article 101583, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301397.

Jin:2011:PCA

[JAH^{+}11] Jing Jin, Gail-Joon Ahn, Hongxin Hu, Michael J. Covington, and Xinwen Zhang. Patient-centric authorization framework for electronic healthcare services. Computers & Security, 30(2–3):116–127, March 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000805.

Jandel:2014:DSR

[Jän14] Magnus Jändel. Decision support for releasing anonymised data. Computers & Security, 46(??):48–61, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001035.

Jones:2019:SDR

[JAN19] A. Jones, O. Angelopoulou, and L. Noriega. Survey of data remaining on second hand memory cards in the UK. Computers & Security, 84(??):239–243, July 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307077.

Jang:2014:DSI

[JC14] Young-Su Jang and Jin-Young Choi. Detecting SQL injection attacks using query result size. Computers & Security, 44(??):104–118, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000595.

Jurcut:2014:DGS

[JCD14] Anca D. Jurcut, Tom Coffey, and Reiner Dojen. Design guidelines for security protocols to prevent replay and parallel session attacks. Computers & Security, 45(??):255–273, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000911.

Jia:2015:MBC

[JCD^{+}15] Yaoqi Jia, Yue Chen, Xinshu Dong, Prateek Saxena, Jian Mao, and Zhenkai Liang. Man-in-the-browser-cache: Persisting HTTPS attacks via browser cache poisoning. Computers & Security, 55(??):62–80, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001121.

Jardi-Cedo:2012:SPS

[JCPACRV12] Roger Jardí-Cedó, Jordi Pujol-Ahulló, Jordi Castellà-Roca, and Alexandre Viejo. Study on poll-site voting and verification systems. Computers & Security, 31(8):989–1010, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001162.

Joaquim:2013:EEE

[JFR13] Rui Joaquim, Paulo Ferreira, and Carlos Ribeiro. EVIV: an end-to-end verifiable Internet voting system. Computers & Security, 32(??):170–191, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001642.

Johnson:2016:TBV

[JGLE16] Pontus Johnson, Dan Gorton, Robert Lagerström, and Mathias Ekstedt. Time between vulnerability disclosures: a measure of software product vulnerability. Computers & Security, 62(??):278–295, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300955.

Jammalamadaka:2013:MAO

[JGM^{+}13] Ravi Chandra Jammalamadaka, Roberto Gamboni, Sharad Mehrotra, Kent Seamons, and Nalini Venkatasubramanian. A middleware approach for outsourcing data securely. Computers & Security, 32(??):252–266, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481200106X.

Ji:2016:CEM

[JHJ^{+}16] Yuede Ji, Yukun He, Xinyang Jiang, Jian Cao, and Qiang Li. Combating the evasion mechanisms of social bots. Computers & Security, 58(??):230–249, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300025.

Jung:2012:CCC

[JJ12] Youna Jung and James B. D. Joshi. CRiBAC: Communitycentric role interaction based access control model. Computers & Security, 31(4):497–523, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000351.

Jung:2014:CPB

[JJ14] Youna Jung and James B. D. Joshi. CPBAC: Property-based access control model for secure cooperation in online social networks. Computers & Security, 41(??):19–39, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001211.

Jin:2013:MFB

[JJA13] Lei Jin, James B. D. Joshi, and Mohd Anwar. Mutualfriend based attacks in social network systems. Computers & Security, 37(??):15–30, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000795.

Jang:2019:RAE

[JJL^{+}19] Daehee Jang, Yunjong Jeong, Sungman Lee, Minjoon Park, Kuenhwan Kwak, Donguk Kim, and Brent Byunghoon Kang. Rethinking anti-emulation techniques for large-scale software deployment. Computers & Security, 83(??):182–200, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818310216.

Jang:2019:SCC

[JK19a] Jinsoo Jang and Brent Byunghoon Kang. Securing a communication channel for the trusted execution environment. Computers & Security, 83(??):79–92, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306060.

Jolfaei:2019:LIP

[JK19b] Alireza Jolfaei and Krishna Kant. A lightweight integrity protection scheme for low latency smart grid applications. Computers & Security, 86(??):471–483, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831126X.

Jamil:2018:SPU

[JKA^{+}18] Fuzel Jamil, Abid Khan, Adeel Anjum, Mansoor Ahmed, Farhana Jabeen, and Nadeem Javaid. Secure provenance using an authenticated data structure approach. Computers & Security, 73(??):34–56, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302122.

James:2013:UNB

[JKC^{+}13] Tabitha L. James, Lara Khansa, Deborah F. Cook, Olga Bruyaka, and Kellie B. Keeling. Using network-based text analysis to analyze trends in Microsoft’s security innovations. Computers & Security, 36(??):49–67, July 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000333.

Jarraya:2010:SPP

[JL10] Houssem Jarraya and Maryline Laurent. A secure peerto-peer backup service keeping great autonomy while under the supervision of a provider. Computers & Security, 29(2):180–195, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001102.

Jain:2019:EDD

[JLGM19] Vineeta Jain, Vijay Laxmi, Manoj Singh Gaur, and Mohamed Mosbah. ETGuard: Detecting D2D attacks using wireless Evil Twins. Computers & Security, 83(??):389–405, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818310708.

Junaid:2016:DDM

[JLK16] Mohsin Junaid, Donggang Liu, and David Kung. Dexteroid: Detecting malicious behaviors in Android apps using reverse-engineered life cycle models. Computers & Security, 59(??):92–117, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300037.

Jin:2018:TLI

[JLPJ18] Lei Jin, Chao Li, Balaji Palanisamy, and James Joshi. k-Trustee: Location injection attack-resilient anonymization for location privacy. Computers & Security, 78(??):212–230, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818305200.

Jerschow:2013:MSR

[JM13] Yves Igor Jerschow and Martin Mauve. Modular square root puzzles: Design of non-parallelizable and non-interactive client puzzles. Computers & Security, 35(??):25–36, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001782.

Jiang:2017:SCF

[JMGW17] Peng Jiang, Yi Mu, Fuchun Guo, and Qiaoyan Wen. Securechannel free keyword search with authorization in managercentric databases. Computers & Security, 69(??):50–64, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816301687.

Jha:2014:SAT

[JSVA14] Sadhana Jha, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Security analysis of temporal RBAC under an administrative model. Computers & Security, 46(??):154–172, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001217.

Julisch:2011:CDB

[JSWZ11] Klaus Julisch, Christophe Suter, Thomas Woitalla, and Olaf Zimmermann. Compliance by design — bridging the chasm between auditors and IT architects. Computers & Security, 30(6–7):410–426, September/October 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000514.

Jaeger:2011:CUC

[JvOW11] Trent Jaeger, Paul C. van Oorschot, and Glenn Wurster. Countering unauthorized code execution on commodity kernels: a survey of common interfaces allowing kernel code modification. Computers & Security, 30(8):571–579, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811001143.

Jeske:2017:FIT

[JvS17] Debora Jeske and Paul van Schaik. Familiarity with Internet threats: Beyond awareness. Computers & Security, 66(??):129–141, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300214.

Jiang:2016:SMS

[JWB16] Rui Jiang, Xianglong Wu, and Bharat Bhargava. SDSS-MAC: Secure data sharing scheme in multi-authority cloud storage systems. Computers & Security, 62(??):193–212, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300864.

Jiang:2019:PPW

[JWZX19] Han Jiang, Hao Wang, Zhihua Zheng, and Qiuliang Xu. Privacy preserved wireless sensor location protocols based on mobile edge computing. Computers & Security, 84(??):393– 401, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818311295.

Jiang:2018:SCU

[JXK18] Jun Jiang, Bowei Xi, and Murat Kantarcioglu. Spatial counts under differential privacy mechanism on changing spatial scales. Computers & Security, 77(??):642–652, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302559.

Ki-Aries:2017:PCI

[KAF17] Duncan Ki-Aries and Shamal Faily. Persona-centred information security awareness. Computers & Security, 70(??):663–674, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301566.

Kalech:2019:CAD

[Kal19] Meir Kalech. Cyber-attack detection in SCADA systems using temporal pattern recognition techniques. Computers & Security, 84(??):225–238, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818307594.

Kammuller:2019:ATI

[Kam19] Florian Kammüller. Attack trees in Isabelle extended with probabilities for quantum cryptography. Computers & Security, 87(??):Article 101572, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481930135X.

Karadsheh:2012:ASP

[Kar12] Louay Karadsheh. Applying security policies and service level agreement to IaaS service model to enhance security and transition. Computers & Security, 31(3):315–326, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000077.

Khambhammettu:2013:FRA

[KBAL13] Hemanth Khambhammettu, Sofiene Boulares, Kamel Adi, and Luigi Logrippo. A framework for risk assessment in access control systems. Computers & Security, 39 (Part A)(??):86–103, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000552.

Knowles:2016:SSA

[KBM16] William Knowles, Alistair Baron, and Tim McGarr. The simulated security assessment ecosystem: Does penetration testing need standardisation? Computers & Security, 62(??):296–316, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300906.

Kim:2013:ALG

[KCB13] Su Yong Kim, Sungdeok Cha, and Doo-Hwan Bae. Automatic and lightweight grammar generation for fuzz testing. Computers & Security, 36(??):1–11, July 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000254.

Khansa:2012:IHP

[KCJB12] Lara Khansa, Deborah F. Cook, Tabitha James, and Olga Bruyaka. Impact of HIPAA provisions on the stock market value of healthcare institutions, and information security and other information technology firms. Computers & Security, 31(6):750–770, September 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000995.

Kolkowska:2013:OPI

[KD13] Ella Kolkowska and Gurpreet Dhillon. Organizational power and information security rule compliance. Computers & Security, 33(??):3–11, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001010.

Kajzer:2014:EIM

[KDC^{+}14] Mitchell Kajzer, John D’Arcy, Charles R. Crowell, Aaron Striegel, and Dirk Van Bruggen. An exploratory investigation of message-person congruence in information security awareness campaigns. Computers & Security, 43(??):64–76, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000327.

Kohn:2013:IDF

[KEE13] M. D. Kohn, M. M. Eloff, and J. H. P. Eloff. Integrated digital forensic process model. Computers & Security, 38(??):103–115, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000849.

Kolar:2019:MSD

[KFGL19] Martin Kolar, Carmen Fernandez-Gago, and Javier Lopez. A model specification for the design of trust negotiations. Computers & Security, 84(??):288–300, July 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818310484.

Kandias:2013:MPT

[KG13] Miltiadis Kandias and Dimitris Gritzalis. Metasploit: the penetration tester’s guide. Computers & Security, 32(??):268–269, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001459.

Kandias:2017:SLD

[KGSN17] Miltiadis Kandias, Dimitris Gritzalis, Vasilis Stavrou, and Kostas Nikoloulis. Stress level detection via OSN usage pattern and chronicity analysis: an OSINT threat intelligence module. Computers & Security, 69(??):3–17, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301742.

Kim:2012:PSP

[KH12] Hyoungshick Kim and Jun Ho Huh. PIN selection policies: Are they really effective? Computers & Security, 31(4):484– 496, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000363.

Khan:2012:PAB

[Kha12] Muhammad Naeem Ahmed Khan. Performance analysis of Bayesian networks and neural networks in classification of file system activities. Computers & Security, 31(4):391–401, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000533.

Karlsson:2017:PBD

[KHG17] Fredrik Karlsson, Karin Hedström, and Göran Goldkuhl. Practice-based discourse analysis of information security policies. Computers & Security, 67(??):267–279, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301833.

Khan:2011:DCB

[KJKM11] Hassan Khan, Mobin Javed, Syed Ali Khayam, and Fauzan Mirza. Designing a cluster-based covert channel to evade disk investigation and forensics. Computers & Security, 30(1):35–49, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481000088X.

Kim:2019:SLF

[KJP^{+}19] Deokjin Kim, Daehee Jang, Minjoon Park, Yunjong Jeong, Jonghwan Kim, Seokjin Choi, and Brent Byunghoon Kang. SGX-LEGO: Fine-grained SGX controlled-channel attack and its countermeasure. Computers & Security, 82(??):118– 139, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818302347.

Kantzavelou:2010:GBI

[KK10a] Ioanna Kantzavelou and Sokratis Katsikas. A game-based intrusion detection mechanism to confront internal attackers. Computers & Security, 29(8):859–874, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000611.

Kavallaris:2010:DPS

[KK10b] Theodoros Kavallaris and Vasilios Katos. On the detection of pod slurping attacks. Computers & Security, 29(6):680–685, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000052.

Kearney:2016:CPD

[KK16] W. D. Kearney and H. A. Kruger. Can perceptual differences account for enigmatic information security behaviour in an organisation? Computers & Security, 61(??):46–58, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300645.

Khammassi:2017:GLW

[KK17] Chaouki Khammassi and Saoussen Krichen. A GA-LR wrapper approach for feature selection in network intrusion detection. Computers & Security, 70(??):255–277, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301244.

Kim:2016:DBM

[KKJ^{+}16] Yonggon Kim, Ohmin Kwon, Jinsoo Jang, Seongwook Jin, Hyeongboo Baek, Brent Byunghoon Kang, and Hyunsoo Yoon. On-demand bootstrapping mechanism for isolated cryptographic operations on commodity accelerators. Computers & Security, 62(??):33–48, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300712.

Kim:2017:QPI

[KKK17] Seung-Hyun Kim, In-Young Ko, and Soo-Hyung Kim. Quality of private information (QoPI) model for effective representation and prediction of privacy controls in mobile computing. Computers & Security, 66(??):1–19, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300056.

Kim:2018:MUP

[KKK18] Sungjin Kim, Jinkook Kim, and Brent ByungHoon Kang. Malicious URL protection based on attackers’ habitual behavioral analysis. Computers & Security, 77(??):790–806, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818300348.

Krzywiecki:2018:POD

[KKKP18] Lukasz Krzywiecki, Kamil Kluczniak, Patryk Kozieł, and Nisha Panwar. Privacy-oriented dependency via deniable SIGMA protocol. Computers & Security, 79(??):53–67, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818301962.

Kolias:2011:SII

[KKM11] C. Kolias, G. Kambourakis, and M. Maragoudakis. Swarm intelligence in intrusion detection: a survey. Computers & Security, 30(8):625–642, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481100109X.

Kubilay:2019:CNP

[KKM19] Murat Yasin Kubilay, Mehmet Sabir Kiraz, and Haci Ali Mantar. CertLedger: a new PKI model with certificate transparency based on blockchain. Computers & Security, 85(??):333–352, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313014.

Kolter:2010:CPM

[KKP10] Jan Kolter, Thomas Kernchen, and Günther Pernul. Collaborative privacy management. Computers & Security, 29(5):580–591, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480900145X.

Kwon:2018:FSE

[KKP^{+}18] Hyun Kwon, Yongchul Kim, Ki-Woong Park, Hyunsoo Yoon, and Daeseon Choi. Friend-safe evasion attack: an adversarial example that is correctly recognized by a friendly classifier. Computers & Security, 78(??):380–397, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308770.

Kussul:2013:AST

[KKS13] Olga Kussul, Nataliia Kussul, and Sergii Skakun. Assessing security threat scenarios for utility-based reputation model in grids. Computers & Security, 34(??):1–15, May 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000230.

Kim:2018:SSR

[KL18] Seulbae Kim and Heejo Lee. Software systems at risk: an empirical study of cloned vulnerabilities in practice. Computers & Security, 77(??):720–736, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300944.

Kent:2015:AGA

[KLN15] Alexander D. Kent, Lorie M. Liebrock, and Joshua C. Neil. Authentication graphs: Analyzing user behavior within an enterprise network. Computers & Security, 48(??):150–166, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001321.

Kandi:2017:ELC

[KMG17] Haribabu Kandi, Deepak Mishra, and Subrahmanyam R. K. Sai Gorthi. Exploring the learning capabilities of convolutional neural networks for robust image watermarking. Computers & Security, 65(??):247–268, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301699.

Karopoulos:2013:UCS

[KMM13] Georgios Karopoulos, Paolo Mori, and Fabio Martinelli. Usage control in SIP-based multimedia delivery. Computers & Security, 39 (part B)(??):406–418, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481300134X.

Klobas:2019:HPS

[KMW19] Jane E. Klobas, Tanya McGill, and Xuequn Wang. How perceived security risk affects intention to use smart home devices: a reasoned action explanation. Computers & Security, 87(??):Article 101571, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301348.

Kwon:2014:TAD

[KN14] Taekyoung Kwon and Sarang Na. TinyLock: Affordable defense against smudge attacks on smartphone pattern lock systems. Computers & Security, 42(??):137–150, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001697.

Kulyk:2017:CRP

[KNM^{+}17] Oksana Kulyk, Stephan Neumann, Karola Marky, Jurlind Budurushi, and Melanie Volkamer. Coercion-resistant proxy voting. Computers & Security, 71(??):88–99, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301268.

Karopoulos:2018:MMP

[KNX18] Georgios Karopoulos, Christoforos Ntantogian, and Christos Xenakis. MASKER: Masking for privacy-preserving aggregation in the smart grid ecosystem. Computers & Security, 73(??):307–325, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302456.

Kokolakis:2017:PAP

[Kok17] Spyros Kokolakis. Privacy attitudes and privacy behaviour: a review of current research on the privacy paradox phenomenon. Computers & Security, 64(??):122–134, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815001017.

Klonowski:2018:LWS

[KP18] Marek Klonowski and Ania M. Piotrowska. Light-weight and secure aggregation protocols based on Bloom filters. Computers & Security, 72(??):107–121, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301815.

Komurcu:2015:EGM

[KPD15] Giray Kömürcü, Ali Emre Pusane, and Günhan Dündar. An efficient grouping method and error probability analysis for RO-PUFs. Computers & Security, 49(??):123–131, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001801.

Kim:2015:DAE

[KPKS15] Eunhyun Kim, Kyungwon Park, Hyoungshick Kim, and Jaeseung Song. Design and analysis of enumeration attacks on finding friends with phone numbers: a case study with KakaoTalk. Computers & Security, 52(??):267–275, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000589.

Kar:2016:SDS

[KPS16] Debabrata Kar, Suvasini Panigrahi, and Srikanth Sundararajan. SQLiGoT: Detecting SQL injection attacks using graph of tokens and SVM. Computers & Security, 60(??):206–225, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300451.

Kortesniemi:2014:SCU

[KS14] Yki Kortesniemi and Mikko Särelä. Survey of certificate usage in distributed access control. Computers & Security, 44(??):16–32, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000534.

Korac:2019:FMU

[KS19] Dragan Korać and Dejan Simić. Fishbone model and universal authentication framework for evaluation of multifactor authentication in mobile environment. Computers & Security, 85(??):313–332, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818307375.

Khurat:2017:PPV

[KSG17] Assadarat Khurat, Boontawee Suntisrivaraporn, and Dieter Gollmann. Privacy policies verification in composite services using OWL. Computers & Security, 67(??):122–141, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300512.

Ksiezopolski:2012:QMQ

[Ksi12] Bogdan Ksiezopolski. QoP-ML: Quality of protection modelling language for cryptographic protocols. Computers & Security, 31(4):569–596, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000107.

Konradt:2016:PEA

[KSW16] Christian Konradt, Andreas Schilling, and Brigitte Werners. Phishing: an economic analysis of cybercrime perpetrators. Computers & Security, 58(??):39–46, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001844.

Kaghazgaran:2019:CEP

[KTCS19] Parisa Kaghazgaran, Hassan Takabi, Flannery Hope Currin, and Armando Soriano. Communication-efficient private distance calculation based on oblivious transfer extensions. Computers & Security, 83(??):93–105, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307053.

Kuang:2018:EVM

[KTG^{+}18] Kaiyuan Kuang, Zhanyong Tang, Xiaoqing Gong, Dingyi Fang, Xiaojiang Chen, and Zheng Wang. Enhance virtualmachine-based code obfuscation security through dynamic bytecode scheduling. Computers & Security, 74(??):202–220, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818300270.

Khalid:2018:TMN

[KUA^{+}18] Waqar Khalid, Zahid Ullah, Naveed Ahmed, Yue Cao, Muhammad Khalid, Muhammad Arshad, Farhan Ahmad, and Haitham Cruickshank. A taxonomy on misbehaving nodes in delay tolerant networks. Computers & Security, 77(??):442–471, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304681.

Kulekci:2012:SBW

[Kül12] M. Oguzhan Külekci. On scrambling the Burrows–Wheeler transform to provide privacy in lossless compression. Computers & Security, 31(1):26–32, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001416.

Krishnan:2015:EPR

[KV15] Padmanabhan Krishnan and Kostyantyn Vorobyov. Enforcement of privacy requirements. Computers & Security, 52(??):164–177, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481500036X.

Kritzinger:2010:CSH

[KvS10] E. Kritzinger and S. H. von Solms. Cyber security for home users: a new way of protection through awareness enforcement. Computers & Security, 29(8):840–847, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000775.

Kirchner:2014:SCM

[KW14] Martin Kirchner and Edgar Weippl. Securing cloud and mobility. Computers & Security, 44(??):64, July 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000224.

Kim:2015:DFA

[KYZ15] Dae Wook Kim, Peiying Yan, and Junjie Zhang. Detecting fake anti-virus software distribution webpages. Computers & Security, 49(??):95–106, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001709.

Lueks:2017:FRA

[LAHV17] Wouter Lueks, Gergely Alpár, Jaap-Henk Hoepman, and Pim Vullers. Fast revocation of attribute-based credentials for both users and verifiers. Computers & Security, 67(??):308–323, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301729.

Lopez:2013:SCO

[LAR13] Javier Lopez, Cristina Alcaraz, and Rodrigo Roman. Smart control of operational threats in control substations. Computers & Security, 38(??):14–27, October 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000588.

Laszka:2019:DMA

[LAVK19] Aron Laszka, Waseem Abbas, Yevgeniy Vorobeychik, and Xenofon Koutsoukos. Detection and mitigation of attacks on transportation networks as a multi-stage security game. Computers & Security, 87(??):Article 101576, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819300628.

Liu:2015:RBF

[LC15] Bing Liu and Chao-Hsien Chu. Relationship-based federated access control model for EPC discovery service. Computers & Security, 55(??):251–270, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000899.

Lemay:2018:SPA

[LCMF18] Antoine Lemay, Joan Calvet, François Menet, and José M. Fernandez. Survey of publicly available reports on advanced persistent threat actors. Computers & Security, 72(??):26–59, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301608.

Liu:2017:POP

[LCZ17] Zhe Liu, Kim-Kwang Raymond Choo, and Minghao Zhao. Practical-oriented protocols for privacy-preserving outsourced big data analysis: Challenges and future research directions. Computers & Security, 69(??):97–113, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816301778.

Lallie:2018:EPC

[LDB18] Harjinder Singh Lallie, Kurt Debattista, and Jay Bal. Evaluating practitioner cyber-security attack graph configuration preferences. Computers & Security, 79(??):117–131, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818306163.

Lasc:2013:DDA

[LDC13] Ioana Lasc, Reiner Dojen, and Tom Coffey. On the detection of desynchronisation attacks against security protocols that use dynamic shared secrets. Computers & Security, 32(??):115–129, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001678.

Lent:2012:EMB

[Len12] Ricardo Lent. Evaluating a migration-based response to DoS attacks in a system of distributed auctions. Computers & Security, 31(3):327–343, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000053.

Leszczyna:2018:RSC

[Les18] Rafał Leszczyna. A review of standards with cybersecurity requirements for smart grid. Computers & Security, 77(??):262–276, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302803.

Luckner:2014:SWS

[LGS14] Marcin Luckner, Michał Gad, and Paweł Sobkowiak. Stable web spam detection using features based on lexical items. Computers & Security, 46(??):79–93, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001151.

Li:2019:SCE

[LGZ^{+}19] Dong Li, Huaqun Guo, Jianying Zhou, Luying Zhou, and Jun Wen Wong. SCADAWall: a CPI-enabled firewall model for SCADA security. Computers & Security, 80(??):134–154, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818308071.

Lee:2011:IMS

[LH11] Jung-San Lee and Ming-Huang Hsieh. An interactive mobile SMS confirmation method using secret sharing technique. Computers & Security, 30(8):830–839, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000836.

Li:2019:SAD

[LHI^{+}19] Yu Li, Jin Huang, Ademola Ikusan, Milliken Mitchell, Junjie Zhang, and Rui Dai. ShellBreaker: Automatically detecting PHP-based malicious web shells. Computers & Security, 87(??):Article 101595, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301506.

Lerch-Hostalot:2013:LMS

[LHM13] Daniel Lerch-Hostalot and David Megías. LSB matching steganalysis based on patterns of pixel differences and random embedding. Computers & Security, 32(??):192–206, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001745.

Lin:2018:UYU

[LHS^{+}18] Dan Lin, Nicholas Hilbert, Christian Storer, Wei Jiang, and Jianping Fan. UFace: Your universal password that no one can see. Computers & Security, 77(??):627–641, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302067.

Lu:2015:TCA

[LJJL15] Jianfeng Lu, James B. D. Joshi, Lei Jin, and Yiding Liu. Towards complexity analysis of user authorization query problem in RBAC. Computers & Security, 48(??):116–130, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001497.

Luh:2019:ADC

[LJS19] Robert Luh, Helge Janicke, and Sebastian Schrittwieser. AIDIS: Detecting and classifying anomalous behavior in ubiquitous kernel processes. Computers & Security, 84(??):120–147, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818314457.

Liu:2010:SVE

[LK10] Fuwen Liu and Hartmut Koenig. A survey of video encryption algorithms. Computers & Security, 29(1):3–15, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000698.

Lencse:2018:MIP

[LK18] Gábor Lencse and Youki Kadobayashi. Methodology for the identification of potential security issues of different IPv6 transition technologies: Threat analysis of DNS64 and stateful NAT64. Computers & Security, 77(??):397–411, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818303663.

Le:2014:CEA

[LKJ14] Meixing Le, Krishna Kant, and Sushil Jajodia. Consistency and enforcement of access rules in cooperative data sharing environment. Computers & Security, 41(??):3–18, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813001284.

Lee:2018:DPC

[LKPK18] Hojoon Lee, Minsu Kim, Yunheung Paek, and Brent Byunghoon Kang. A dynamic per-context verification of kernel address integrity from external monitors. Computers & Security, 77(??):824–837, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818301421.

Lin:2013:IAM

[LLCT13] Ying-Dar Lin, Yuan-Cheng Lai, Chien-Hung Chen, and Hao-Chuan Tsai. Identifying Android malicious repackaged applications by thread-grained system call sequences. Computers & Security, 39 (part B)(??):340–350, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001272.

Liu:2019:WHC

[LLD^{+}19] Ximing Liu, Yingjiu Li, Robert H. Deng, Bing Chang, and Shujun Li. When human cognitive modeling meets PINs: User-independent inter-keystroke timing attacks. Computers & Security, 80(??):90–107, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302736.

Lyvas:2018:DDP

[LLG18] Christos Lyvas, Costas Lambrinoudakis, and Dimitris Geneiatakis. Dypermin: Dynamic permission mining framework for Android platform. Computers & Security, 77(??):472–487, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304954.

Lin:2013:CBW

[LLHT13] Ying-Dar Lin, Yuan-Cheng Lai, Cheng-Yuan Ho, and Wei-Hsuan Tai. Creditability-based weighted voting for reducing false positives and negatives in intrusion detection. Computers & Security, 39 (part B)(??):460–474, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001399.

Lee:2016:UIS

[LLK16] Chunghun Lee, Choong C. Lee, and Suhyun Kim. Understanding information security stress: Focusing on the type of information security compliance activity. Computers & Security, 59(??):60–70, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300074.

Lee:2015:SSA

[LLL15] Jehyun Lee, Suyeon Lee, and Heejo Lee. Screening smartphone applications using malware family signatures. Computers & Security, 52(??):234–249, July 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000188.

Liu:2018:KSP

[LLL^{+}18] Wangtong Liu, Senlin Luo, Yu Liu, Limin Pan, and Qamas Gul Khan Safi. A kernel stack protection model against attacks from kernel execution units. Computers & Security, 72(??):96–106, January 2018. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301979.

Liang:2017:VVB

[LLX^{+}17] Hongliang Liang, Mingyu Li, Jian Xu, Wenying Hu, Xiaoxiao Pei, Xiaodong Jia, and Yan Song. vmOS: a virtualization-based, secure desktop system. Computers & Security, 65(??):329–343, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301390.

Le:2017:FAN

[LLY17] Junqing Le, Xiaofeng Liao, and Bo Yang. Full autonomy: a novel individualized anonymity model for privacy preserving. Computers & Security, 66(??):204–217, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630181X.

Li:2015:PLA

[LLYD15] Yan Li, Yingjiu Li, Qiang Yan, and Robert H. Deng. Privacy leakage analysis in online social networks. Computers & Security, 49(??):239–254, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001588.

Lee:2010:SIS

[LM10] Seok-Won Lee and Mattia Monga. Special issue on software engineering for secure systems. Computers & Security, 29(3):299–301, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000064.

Lykou:2018:NMT

[LMG18] Georgia Lykou, Despina Mentzelioti, and Dimitris Gritzalis. A new methodology toward effectively assessing data center sustainability. Computers & Security, 76(??):327–340, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817302754.

Li:2016:SER

[LML^{+}16] Qi Li, Jianfeng Ma, Rui Li, Ximeng Liu, Jinbo Xiong, and Danwei Chen. Secure, efficient and revocable multiauthority access control system in cloud storage. Computers & Security, 59(??):45–59, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300050.

Lee:2017:DPP

[LML^{+}17] Min Cherng Lee, Robin Mitra, Emmanuel Lazaridis, An-Chow Lai, Yong Kheng Goh, and Wun-She Yap. Data privacy preserving scheme using generalised linear models. Computers & Security, 69(??):142–154, August 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301808.

Li:2016:MTP

[LMLK16] Wenjuan Li, Weizhi Meng, Xiapu Luo, and Lam For Kwok. MVPSys: Toward practical multi-view based false alarm reduction system in network intrusion detection. Computers & Security, 60(??):177–192, July 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300475.

Li:2019:PPF

[LMLL19] Shaoxin Li, Nankun Mu, Junqing Le, and Xiaofeng Liao. Privacy preserving frequent itemset mining: Maximizing data utility based on database reconstruction. Computers & Security, 84(??):17–34, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302748.

Lacey:2012:RUR

[LMM^{+}12] T. H. Lacey, R. F. Mills, B. E. Mullins, R. A. Raines, M. E. Oxley, and S. K. Rogers. RIPsec — using reputationbased multilayer security to protect MANETs. Computers & Security, 31(1):122–136, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001167.

Lopez:2019:ACU

[LMNM^{+}19] Alberto Urueña López, Fernando Mateo, Julio Navío-Marco, José María Martínez-Martínez, Juan Gómez-Sanchís, Joan Vila-Francés, and Antonio José Serrano-López. Analysis of computer user behavior, security incidents and fraud using self-organizing maps. Computers & Security, 83(??):38–51, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818300828.

LaBarge:2014:ASR

[LMS14] Ralph LaBarge, Thomas A. Mazzuchi, and Shahram Sarkani. An automated system for rapid and secure device sanitization. Computers & Security, 42(??):77–91, May 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000194.

Lee:2016:SBP

[LNK16] Mun-Kyu Lee, Hyeonjin Nam, and Dong Kyue Kim. Secure bimodal PIN-entry method using audio signals. Computers & Security, 56(??):140–150, February 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000929.

Li:2013:EAF

[LNNH13] Celia Li, Uyen Trang Nguyen, Hoang Lan Nguyen, and Nurul Huda. Efficient authentication for fast handover in wireless mesh networks. Computers & Security, 37(??):124–142, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000916.

Lopriore:2016:ACL

[Lop16] Lanfranco Lopriore. Access control lists in password capability environments. Computers & Security, 62(??):317–327, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300967.

Lin:2018:EDM

[LPL18] Chih-Hung Lin, Hsing-Kuo Pao, and Jian-Wei Liao. Efficient dynamic malware analysis using virtual time control mechanics. Computers & Security, 73(??):359–373, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481730247X.

Leitold:2010:RES

[LPR10] Herbert Leitold, Reinhard Posch, and Thomas Rössler. Reconstruction of electronic signatures from eDocument printouts. Computers & Security, 29(5):523–532, July 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001254.

Liew:2019:ESA

[LSA^{+}19] Seow Wooi Liew, Nor Fazlida Mohd Sani, Mohd. Taufik Abdullah, Razali Yaakob, and Mohd Yunus Sharum. An effective security alert mechanism for real-time phishing tweet detection on Twitter. Computers & Security, 83(??):201– 207, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818309040.

Lagazio:2014:MLA

[LSC14] Monica Lagazio, Nazneen Sherif, and Mike Cushman. A multi-level approach to understanding the impact of cyber crime on the financial sector. Computers & Security, 45(??):58–74, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481400087X.

Liginlal:2012:HPR

[LSKF12] Divakaran Liginlal, Inkook Sim, Lara Khansa, and Paul Fearn. HIPAA privacy rule compliance: an interpretive study using Norman’s action theory. Computers & Security, 31(2):206–220, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001453.

Lin:2015:IAA

[LSQ15] Xi-Jun Lin, Lin Sun, and Haipeng Qu. Insecurity of an anonymous authentication for privacy-preserving IoT targetdriven applications. Computers & Security, 48(??):142–149, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001229.

Langer:2016:ONA

[LSSK16] Lucie Langer, Florian Skopik, Paul Smith, and Markus Kammerstetter. From old to new: Assessing cybersecurity risks for an evolving smart grid. Computers & Security, 62(??):165–176, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300876.

Tan:2019:OSS

[lTLqZqC19] Jing lei Tan, Cheng Lei, Hong qi Zhang, and Yu qiao Cheng. Optimal strategy selection approach to moving target defense based on Markov robust game. Computers & Security, 85(??):63–76, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313919.

Leithner:2012:AF

[LW12a] Manuel Leithner and Edgar Weippl. Android forensics. Computers & Security, 31(1):3, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001301.

Leithner:2012:CPT

[LW12b] Manuel Leithner and Edgar Weippl. Coding for penetration testers. Computers & Security, 31(2):252, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001490.

Leithner:2012:LTH

[LW12c] Manuel Leithner and Edgar Weippl. Low tech hacking. Computers & Security, 31(5):738, July 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000806.

Leithner:2012:XFD

[LW12d] Manuel Leithner and Edgar Weippl. XBOX 360 forensics: a digital forensics guide to examining artifacts. Computers & Security, 31(4):639, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000375.

Leithner:2012:TMS

[LW12e] Manuel Leithner and Edgar R. Weippl. Thor’s Microsoft security Bible. Computers & Security, 31(4):638, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000405.

Lu:2019:DFB

[LW19] Tingting Lu and Junfeng Wang. Data-flow bending: On the effectiveness of data-flow integrity. Computers & Security, 84(??):365–375, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818312604.

Landauer:2018:DLF

[LWS^{+}18] Max Landauer, Markus Wurzenberger, Florian Skopik, Giuseppe Settanni, and Peter Filzmoser. Dynamic log file analysis: an unsupervised cluster evolution approach for anomaly detection. Computers & Security, 79(??):94–116, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818306333.

Li:2019:OSA

[LWS^{+}19] Tengyao Li, Buhong Wang, Fute Shang, Jiwei Tian, and Kunrui Cao. Online sequential attack detection for ADS-B data based on hierarchical temporal memory. Computers & Security, 87(??):Article 101599, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301531.

Liu:2017:ACI

[LWZ^{+}17] Shigang Liu, Yu Wang, Jun Zhang, Chao Chen, and Yang Xiang. Addressing the class imbalance problem in Twitter spam detection using ensemble learning. Computers & Security, 69(??):35–49, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301754.

Li:2016:FSL

[LXDW16] Weize Li, Lun Xie, Zulan Deng, and Zhiliang Wang. False sequential logic attack on SCADA system and its physical impact analysis. Computers & Security, 58(??):149–159, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816000031.

Luo:2018:EAS

[LXH^{+}18] Yuchuan Luo, Ming Xu, Kai Huang, Dongsheng Wang, and Shaojing Fu. Efficient auditing for shared data in the cloud with secure user revocation and computations outsourcing. Computers & Security, 73(??):492–506, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302663.

Lv:2019:ACD

[LZ19] Denglong Lv and Shibing Zhu. Achieving correlated differential privacy of big data publication. Computers & Security, 82(??):184–195, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481830289X.

Luo:2013:IPV

[LZBS13] Xin (Robert) Luo, Wei Zhang, Stephen Burd, and Alessandro Seazzu. Investigating phishing victimization with the heuristic-systematic model: a theoretical framework and an exploration. Computers & Security, 38(??):28–38, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001927.

Li:2018:SAB

[LZCX18] Jin Li, Yinghui Zhang, Xiaofeng Chen, and Yang Xiang. Secure attribute-based data sharing for resource-limited users in cloud computing. Computers & Security, 72(??):1–12, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301621.

Liu:2016:TRR

[LZS^{+}16] Gang Liu, Runnan Zhang, Huimin Song, Can Wang, Jinhui Liu, and Aijun Liu. Ts-RBAC: a RBAC model with transformation. Computers & Security, 60(??):52–61, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300281.

Mavani:2017:MAI

[MA17] Monali Mavani and Krishna Asawa. Modeling and analyses of IP spoofing attack in 6LoWPAN network. Computers & Security, 70(??):95–110, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301116.

Maiorca:2015:SAE

[MAC^{+}15] Davide Maiorca, Davide Ariu, Igino Corona, Marco Aresu, and Giorgio Giacinto. Stealth attacks: an extended insight into the obfuscation effects on Android malware. Computers & Security, 51(??):16–31, June 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500022X.

Moriano:2017:ICS

[MAC17] Pablo Moriano, Soumya Achar, and L. Jean Camp. Incompetents, criminals, or spies: Macroeconomic analysis of routing anomalies. Computers & Security, 70(??):319–334, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301372.

Mohaisen:2015:AHF

[MAM15] Aziz Mohaisen, Omar Alrawi, and Manar Mohaisen. AMAL: High-fidelity, behavior-based automated malware analysis and classification. Computers & Security, 52(??):251–266, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000425.

Mohsin:2017:IDD

[MAZAS17] Mujahid Mohsin, Zahid Anwar, Farhat Zaman, and Ehab Al-Shaer. IoTChecker: a data-driven framework for security analytics of Internet of Things configurations. Computers & Security, 70(??):199–223, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301190.

Meszaros:2017:IOF

[MB17] Jan Meszaros and Alena Buchalcevova. Introducing OSSF: a framework for online service cybersecurity risk management. Computers & Security, 65(??):300–313, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301791.

Molinaro:2018:EAD

[MB18] Kylie A. Molinaro and Matthew L. Bolton. Evaluating the applicability of the double system lens model to the analysis of phishing email judgments. Computers & Security, 77(??):128–137, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302955.

Moller:2011:MBU

[MBAE^{+}11] Sebastian Möller, Noam Ben-Asher, Klaus-Peter Engelbrecht, Roman Englert, and Joachim Meyer. Modeling the behavior of users who are confronted with security mechanisms. Computers & Security, 30(4):242–256, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000423.

Macfarlane:2012:FSP

[MBE^{+}12] Richard Macfarlane, William Buchanan, Elias Ekonomou, Omair Uthmani, Lu Fan, and Owen Lo. Formal security policy implementations in network firewalls. Computers & Security, 31(2):253–270, March 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001192.

Mainardi:2019:PRA

[MBP19] Nicholas Mainardi, Alessandro Barenghi, and Gerardo Pelosi. Plaintext recovery attacks against linearly decryptable fully homomorphic encryption schemes. Computers & Security, 87(??):Article 101587, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301403.

Murnion:2018:MLS

[MBSR18] Shane Murnion, William J. Buchanan, Adrian Smales, and Gordon Russell. Machine learning and semantic analysis of in-game chat for cyberbullying. Computers & Security, 76(??):197–213, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818301597.

Mancuhan:2018:SVC

[MC18]  Koray Mancuhan and Chris Clifton.  Support vec-

tor classification with l-diversity. Computers & Security, 77(??):653–665, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302791.

Movahedi:2019:VPC

[MCG19] Yazdan Movahedi, Michel Cukier, and Ilir Gashi. Vulnerability prediction capability: a comparison between vulnerability discovery models and neural network models. Computers & Security, 87(??):Article 101596, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301518.

Montanari:2013:DSP

[MCL^{+}13] Mirko Montanari, Ellick Chan, Kevin Larson, Wucherl Yoo, and Roy H. Campbell. Distributed security policy conformance. Computers & Security, 33(??):28–40, March 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001770.

Mander:2010:PSD

[MCN10] Todd Mander, Richard Cheung, and Farhad Nabhani. Power system DNP3 data object security using data sets. Computers & Security, 29(4):487–500, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001084.

Mhenni:2019:DSA

[MCRB19] Abir Mhenni, Estelle Cherrier, Christophe Rosenberger, and Najoua Essoukri Ben Amara. Double serial adaptation mechanism for keystroke dynamics authentication based on a single password. Computers & Security, 83(??):151–166, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818306059.

Mao:2017:SIA

[MCT^{+}17] Weixuan Mao, Zhongmin Cai, Don Towsley, Qian Feng, and Xiaohong Guan. Security importance assessment for system objects and malware detection. Computers & Security, 68(??):47–68, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300354.

Mao:2018:BDK

[MCY^{+}18] Weixuan Mao, Zhongmin Cai, Yuan Yang, Xiaohong Shi, and Xiaohong Guan. From big data to knowledge: a spatiotemporal approach to malware detection. Computers & Security, 74(??):167–183, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302705.

Malina:2018:SET

[MDHM18] Lukas Malina, Petr Dzurenda, Jan Hajny, and Zdenek Martinasek. Secure and efficient two-factor zero-knowledge authentication solution for access control systems. Computers & Security, 77(??):500–513, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304942.

Medkova:2018:CAA

[Med18] Jana Medková. Composition attack against social network data. Computers & Security, 74(??):115–129, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300051.

Ma:2011:UCC

[MFBC11] Xiongfeng Ma, Chi-Hang Fred Fung, Jean-Christian Boileau, and H. F. Chau. Universally composable and customizable post-processing for practical quantum key distribution. Computers & Security, 30(4):172–177, June 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810001021.

Macia-Fernandez:2018:UND

[MFCMC^{+}18] Gabriel Maciá-Fernández, José Camacho, Roberto Magán-Carrión, Pedro García-Teodoro, and Roberto Therón. UGR‘16: a new dataset for the evaluation of cyclostationaritybased network IDSs. Computers & Security, 73(??):411–424, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302353.

Mylonas:2012:PMA

[MG12] Alexis Mylonas and Dimitris Gritzalis. Practical malware analysis: the hands-on guide to dissecting malicious software. Computers & Security, 31(6):802–803, September 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000818.

Miller:2018:IPR

[MGYC18] Daniel Bradford Miller, William Bradley Glisson, Mark Yampolskiy, and Kim-Kwang Raymond Choo. Identifying 3D printer residual data via open-source documentation. Computers & Security, 75(??):10–23, June 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300324.

Muller:2017:ECL

[MHL^{+}17] Steve Muller, Carlo Harpes, Yves Le Traon, Sylvain Gombault, and Jean-Marie Bonnin. Efficiently computing the likelihoods of cyclically interdependent risk scenarios. Computers & Security, 64(??):59–68, January 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301158.

Miyamoto:2017:WCR

[MHS17] Inez Miyamoto, Thomas H. Holzer, and Shahryar Sarkani. Why a counterfeit risk avoidance strategy fails. Computers & Security, 66(??):81–96, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300020.

Michael:2012:BIS

[Mic12a] Katina Michael. The basics of information security: Understanding the fundamentals of InfoSec in theory and practice. Computers & Security, 31(4):634–635, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000557.

Michael:2012:HNG

[Mic12b] Katina Michael. Hacking: the next generation. Computers & Security, 31(6):799, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000971.

Michael:2012:HSC

[Mic12c] Katina Michael. Handbook on securing cyber-physical critical infrastructure: Foundations and challenges. Computers & Security, 31(8):1013, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001150.

Michael:2012:SCC

[Mic12d] Katina Michael. Securing the cloud: Cloud computer security techniques and tactics. Computers & Security, 31(4): 633, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000569.

Michael:2012:SRM

[Mic12e] Katina Michael. Security risk management: Building an information security risk management program from the ground up. Computers & Security, 31(2):249–250, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811001660.

Masdari:2015:TEC

[MJBKZ15] Mohammad Masdari, Sam Jabbehdari, Jamshid Bagherzadeh, and Ahmad Khadem-Zadeh. Towards efficient certificate status validations with E-ADOPT in mobile ad hoc networks. Computers & Security, 49(??):17–27, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001667.

Manes:2018:DIK

[MJRK18] Valentin J. M. Manès, Daehee Jang, Chanho Ryu, and Brent Byunghoon Kang. Domain isolated kernel: a lightweight sandbox for untrusted kernel extensions. Computers & Security, 74(??):130–143, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300282.

Meng:2018:JDJ

[MJW^{+}18] Weizhi Meng, Lijun Jiang, Yu Wang, Jin Li, Jun Zhang, and Yang Xiang. JFCGuard: Detecting juice filming charging attack via processor usage analysis on smartphones. Computers & Security, 76(??):252–264, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302493.

Makri:2011:CRG

[MK11] Eleftheria Makri and Elisavet Konstantinou. Constant round group key agreement protocols: a comparative study. Computers & Security, 30(8):643–678, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001088.

Manesh:2019:CAU

[MK19] Mohsen Riahi Manesh and Naima Kaabouch. Cyberattacks on unmanned aerial system networks: Detection, countermeasure, and future research directions. Computers & Security, 85(??):386–401, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300963.

Mylonas:2013:DSU

[MKG13] Alexios Mylonas, Anastasia Kastania, and Dimitris Gritzalis. Delegate the smartphone user? Security awareness in smartphone platforms. Computers & Security, 34(??):47– 66, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812001733.

Morovati:2016:NBD

[MKG16] Kamran Morovati, Sanjay Kadam, and Ali Ghorbani. A network based document management model to prevent data extrusion. Computers & Security, 59(??):71–91, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300062.

Mavrogiannopoulos:2011:TSM

[MKP11] Nikos Mavrogiannopoulos, Nessim Kisserli, and Bart Preneel. A taxonomy of self-modifying code for obfuscation. Computers & Security, 30(8):679–691, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001076.

Mana:2012:TNB

[MKP12] Antonio Maña, Hristo Koshutanski, and Ernesto J. Pérez. A trust negotiation based security framework for service provisioning in load-balancing clusters. Computers & Security, 31(1):4–25, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001428.

Montenegro:2014:PSS

[ML14] Jose A. Montenegro and Javier Lopez. A practical solution for sealed bid and multi-currency auctions. Computers & Security, 45(??):186–198, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000972.

Muller:2018:TRA

[MLH^{+}18] Steve Muller, Jean Lancrenon, Carlo Harpes, Yves Le Traon, Sylvain Gombault, and Jean-Marie Bonnin. A training-resistant anomaly detection system. Computers & Security, 76(??):1–11, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481830155X.

Meng:2014:EEP

[MLK14] Weizhi Meng, Wenjuan Li, and Lam-For Kwok. EFM: Enhancing the performance of signature-based network intrusion detection systems using enhanced filter mechanism. Computers & Security, 43(??):189–204, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481400025X.

Meng:2017:TEC

[MLKC17] Weizhi Meng, Wenjuan Li, Lam-For Kwok, and Kim-Kwang Raymond Choo. Towards enhancing click-draw based graphical passwords using multi-touch behaviours on smartphones. Computers & Security, 65(??):213–229, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301638.

Mesbah:2017:REJ

[MLM17] Abdelhak Mesbah, Jean-Louis Lanet, and Mohamed Mezghiche. Reverse engineering a Java Card memory management algorithm. Computers & Security, 66(??):97–114, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300093.

Mouton:2016:SEA

[MLV16] Francois Mouton, Louise Leenen, and H. S. Venter. Social engineering attack examples, templates and scenarios. Computers & Security, 59(??):186–209, June 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300268.

Mesquida:2015:IIS

[MM15a] Antoni Lluís Mesquida and Antonia Mas. Implementing information security best practices on software lifecycle processes: the ISO/IEC 15504 Security Extension. Computers & Security, 48(??):19–34, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001345.

Montesdioca:2015:MUS

[MM15b] Gustavo Percio Zimmermann Montesdioca and Antônio Carlos Gastaud Maçada. Measuring user satisfaction with information security practices. Computers & Security, 48(??):267–280, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001618.

Motzek:2017:CBF

[MM17] Alexander Motzek and Ralf Möller. Context- and bias-free probabilistic mission impact assessment. Computers & Security, 65(??):166–186, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301584.

Martinez-Moyano:2011:MBC

[MMCA11] Ignacio J. Martinez-Moyano, Stephen H. Conrad, and David F. Andersen. Modeling behavioral considerations related to information security. Computers & Security, 30(6– 7):397–409, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000472.

Milosevic:2019:TAP

[MMF19] Jelena Milosevic, Miroslaw Malek, and Alberto Ferrante. Time, accuracy and power consumption tradeoff in mobile malware detection systems. Computers & Security, 82(??):314–328, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818307880.

Michalopoulos:2014:GRT

[MMJ14] Dimitrios Michalopoulos, Ioannis Mavridis, and Marija Jankovic. GARS: Real-time system for identification, assessment and control of cyber grooming attacks. Computers & Security, 42(??):177–190, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001806.

Mouton:2015:NES

[MMKV15] Francois Mouton, Mercia M. Malan, Kai K. Kimppa, and H. S. Venter. Necessity for ethics in social engineering research. Computers & Security, 55(??):114–127, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815001224.

Maeng:2015:TAU

[MMLN15] YoungJae Maeng, Aziz Mohaisen, Mun-Kyu Lee, and Dae-Hun Nyang. Transaction authentication using complementary colors. Computers & Security, 48(??):167–181, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001473.

Mylonas:2013:SSD

[MMMG13] Alexios Mylonas, Vasilis Meletiadis, Lilian Mitrou, and Dimitris Gritzalis. Smartphone sensor data as digital evidence. Computers & Security, 38(??):51–75, October 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000527.

Molin:2018:SUC

[MMPC18] Eric Molin, Kirsten Meeuwisse, Wolter Pieters, and Caspar Chorus. Secure or usable computers? Revealing employees’ perceptions and trade-offs by means of a discrete choice experiment. Computers & Security, 77(??):65–78, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302396.

Mitrokotsa:2014:LLD

[MOV14] Aikaterini Mitrokotsa, Cristina Onete, and Serge Vaudenay. Location leakage in distance bounding: Why location privacy does not work. Computers & Security, 45(??):199–209, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000947.

Menges:2018:CAI

[MP18] Florian Menges and Günther Pernul. A comparative analysis of incident reporting formats. Computers & Security, 73(??):87–101, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302250.

Mamolar:2018:TTD

[MPCK18] Ana Serrano Mamolar, Zeeshan Pervez, Jose M. Alcaraz Calero, and Asad Masood Khattak. Towards the transversal detection of DDoS network attacks in 5G multi-tenant overlay networks. Computers & Security, 79(??):132–147, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818309313.

Maier:2015:GDM

[MPM15] Dominik Maier, Mykola Protsenko, and Tilo Müller. A game of droid and mouse: the threat of split-personality malware on Android. Computers & Security, 54(??):2–15, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000656.

Molina:2012:OEA

[MPORBR12] Maurizio Molina, Ignasi Paredes-Oliva, Wayne Routly, and Pere Barlet-Ros. Operational experiences with anomaly detection in backbone networks. Computers & Security, 31(3):273–285, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000132.

Mut-Puigserver:2012:SET

[MPPCFG^{+}12] Macià Mut-Puigserver, M. Magdalena Payeras-Capellà, Josep-Lluís Ferrer-Gomila, Arnau Vives-Guasch, and Jordi Castellà-Roca. A survey of electronic ticketing applied to transport. Computers & Security, 31(8):925–939, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001058.

Millan:2010:PBT

[MPPS10] Gabriel López Millán, Manuel Gil Pérez, Gregorio Martínez Pérez, and Antonio F. Gómez Skarmeta. PKI-based trust management in inter-domain scenarios. Computers & Security, 29(2):278–290, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000868.

Martin:2011:CUA

[MR11] Nigel Martin and John Rice. Cybercrime: Understanding and addressing the concerns of stakeholders. Computers & Security, 30(8):803–814, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481100085X.

Magazinius:2012:FID

[MRS12] Jonas Magazinius, Alejandro Russo, and Andrei Sabelfeld. On-the-fly inlining of dynamic security monitors. Computers & Security, 31(7):827–843, October 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001180.

Mondal:2011:SAG

[MSA11] Samrat Mondal, Shamik Sural, and Vijayalakshmi Atluri. Security analysis of GTRBAC and its variants using model checking. Computers & Security, 30(2–3):128–147, March 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404810000817.

Moreno-Sanchez:2018:CEI

[MSMK18] Pedro Moreno-Sanchez, Uzair Mahmood, and Aniket Kate. ClearChart: Ensuring integrity of consumer ratings in online marketplaces. Computers & Security, 78(??):90–102, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818304401.

Masood:2019:DMM

[MSR19] Zaheer Masood, Raza Samar, and Muhammad Asif Zahoor Raja. Design of a mathematical model for the Stuxnet virus in a network of critical control infrastructure. Computers & Security, 87(??):Article 101565, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831229X.

Myers:2018:ADI

[MSRF18] David Myers, Suriadi Suriadi, Kenneth Radke, and Ernest Foo. Anomaly detection for industrial control systems using process mining. Computers & Security, 78(??):103–125, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818306795.

Martinez:2012:SAM

[MSV12] Sergio Martínez, David Sánchez, and Aida Valls. Semantic adaptive microaggregation of categorical microdata. Computers & Security, 31(5):653–672, July 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000648.

Mitra:2016:MTR

[MSVA16] Barsha Mitra, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Mining temporal roles using many-valued concepts. Computers & Security, 60(??):79–94, July 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300335.

Mohammed:2017:EPC

[MT17] Zareef A. Mohammed and Gurvirender P. Tejay. Examining privacy concerns and ecommerce adoption in developing countries: the impact of culture in shaping individuals’ perceptions toward technology. Computers & Security, 67(??):254–265, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300548.

Mousavi:2019:CCB

[MT19] Nima Mousavi and Mahesh Tripunitara. Constructing cascade Bloom filters for efficient access enforcement. Computers & Security, 81(??):1–14, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818311271.

Meng:2018:SAE

[MTC^{+}18] Huasong Meng, Vrizlynn L. L. Thing, Yao Cheng, Zhongmin Dai, and Li Zhang. A survey of Android exploits in the wild. Computers & Security, 76(??):71–91, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301664.

Modic:2016:NET

[MTT^{+}16] Jolanda Modic, Ruben Trapero, Ahmed Taha, Jesus Luna, Miha Stopar, and Neeraj Suri. Novel efficient techniques for real-time cloud security assessment. Computers & Security, 62(??):1–18, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300682.

Marett:2019:QTA

[MVD19] Kent Marett, Ali Vedadi, and Alexandra Durcikova. A quantitative textual analysis of three types of threat communication and subsequent maladaptive responses. Computers & Security, 80(??):25–35, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308629.

Menesidou:2016:AKE

[MVK16] Sofia Anna Menesidou, Dimitrios Vardalis, and Vasilios Katos. Automated key exchange protocol evaluation in delay tolerant networks. Computers & Security, 59(??):1–8, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300098.

Mallesh:2011:ASD

[MW11] Nayantara Mallesh and Matthew Wright. An analysis of the statistical disclosure attack and receiver-bound cover. Computers & Security, 30(8):597–612, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001118.

Miller:2016:MCS

[MWAG16] Simon Miller, Christian Wagner, Uwe Aickelin, and Jonathan M. Garibaldi. Modelling cyber-security experts’ decision making processes using aggregation operators. Computers & Security, 62(??):229–245, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630089X.

Menard:2018:ICP

[MWL18] Philip Menard, Merrill Warkentin, and Paul Benjamin Lowry. The impact of collectivism and psychological ownership on protection motivation: a cross-cultural examination. Computers & Security, 75(??):147–166, June 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300488.

Mou:2012:FEC

[MZJ^{+}12] Sheng Mou, Zhiwen Zhao, Sisi Jiang, Zushun Wu, and Jiaojiao Zhu. Feature extraction and classification algorithm for detecting complex covert timing channel. Computers & Security, 31(1):70–82, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001349.

Mao:2018:LPP

[MZZ^{+}18] Yunlong Mao, Yuan Zhang, Xiaoyan Zhang, Fengyuan Xu, and Sheng Zhong. Location privacy in public access points positioning: an optimization and geometry approach. Computers & Security, 73(??):425–438, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302365.

Nakhila:2018:GIU

[NADZ18] Omar Nakhila, Muhammad Faisal Amjad, Erich Dondyk, and Cliff Zou. Gateway independent user-side wi-fi Evil Twin attack detection using virtual wireless clients. Computers & Security, 74(??):41–54, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302766.

Ntantogian:2014:EPA

[NAMX14] Christoforos Ntantogian, Dimitris Apostolopoulos, Giannis Marinakis, and Christos Xenakis. Evaluating the privacy of Android mobile applications under forensic analysis. Computers & Security, 42(??):66–76, May 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000157.

Nadler:2019:DML

[NAS19] Asaf Nadler, Avi Aminov, and Asaf Shabtai. Detection of malicious and low throughput data exfiltration over the DNS protocol. Computers & Security, 80(??):36–53, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304000.

Navas:2019:UMO

[NB19] Jorge Navas and Marta Beltrán. Understanding and mitigating OpenID connect threats. Computers & Security, 84(??): 1–16, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818312781.

Nissim:2015:DMP

[NCGE15] Nir Nissim, Aviad Cohen, Chanan Glezer, and Yuval Elovici. Detection of malicious PDF files and directions for enhancements: a state-of-the art survey. Computers & Security, 48(??):246–266, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001606.

Nguyen:2014:STB

[NCS14] Vu Duc Nguyen, Yang-Wai Chow, and Willy Susilo. On the security of text-based 3D CAPTCHAs. Computers & Security, 45(??):84–99, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000856.

Ngo:2015:DDX

[NDdL15] Canh Ngo, Yuri Demchenko, and Cees de Laat. Decision diagrams for XACML policy evaluation and management. Computers & Security, 49(??):1–16, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001655.

Nakaguro:2013:DLN

[NDMM13] Yoichi Nakaguro, Matthew N. Dailey, Sanparith Marukatat, and Stanislav S. Makhanov. Defeating line-noise CAPTCHAs with multiple quadratic snakes. Computers & Security, 37(??):91–110, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000862.

Navarro:2018:SSM

[NDP18] Julio Navarro, Aline Deruyver, and Pierre Parrend. A systematic survey on multi-step attack detection. Computers & Security, 76(??):214–249, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302141.

Nacer:2017:DAM

[NDSA17] Hassina Nacer, Nabil Djebari, Hachem Slimani, and Djamil Aissani. A distributed authentication model for composite Web services. Computers & Security, 70(??):144–178, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301153.

Ngoqo:2015:ISB

[NF15] Bukelwa Ngoqo and Stephen V. Flowerday. Information security behaviour profiling framework (ISBPF) for student mobile phone users. Computers & Security, 53(??):132–142, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000851.

Nunez:2016:EMA

[NFGL16] David Nuñez, Carmen Fernández-Gago, and Jesús Luna. Eliciting metrics for accountability of cloud systems. Computers & Security, 62(??):149–164, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630075X.

Nergiz:2014:HA

[NG14] Mehmet Ercan Nergiz and Muhammed Zahit Gök. Hybrid kanonymity. Computers & Security, 44(??):51–63, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000455.

Nguyen:2016:EIS

[NGSV16] Hai Nguyen, Vinod Ganapathy, Abhinav Srivastava, and Shivaramakrishnan Vaidyanathan. Exploring infrastructure support for app-based services on cloud platforms. Computers & Security, 62(??):177–192, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300888.

Nguyen:2018:PPM

[Ngu18] Huu Hiep Nguyen. Privacy-preserving mechanisms for kmodes clustering. Computers & Security, 78(??):60–75, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818304218.

Niemimaa:2017:ISC

[Nie17] Marko Niemimaa. Information systems continuity process: Conceptual foundations for the study of the “social”. Computers & Security, 65(??):1–13, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301547.

Nyang:2018:TTP

[NKL^{+}18] DaeHun Nyang, Hyoungshick Kim, Woojoo Lee, Sung bae Kang, Geumhwan Cho, Mun-Kyu Lee, and Aziz Mohaisen. Two-Thumbs-Up: Physical protection for PIN entry secure against recording attacks. Computers & Security, 78(??):1–15, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818305789.

Nissim:2019:VMA

[NLC^{+}19] Nir Nissim, Omri Lahav, Aviad Cohen, Yuval Elovici, and Lior Rokach. Volatile memory analysis using the MinHash method for efficient and secured detection of malware in private cloud. Computers & Security, 87(??):Article 101590, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481831188X.

Nguyen:2018:TBU

[NM18] Toan Nguyen and Nasir Memon. Tap-based user authentication for smartwatches. Computers & Security, 78(??):174–186, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818303778.

Naicker:2019:ECM

[NM19] V. Naicker and M. Mafaiti. The establishment of collaboration in managing information security through multisourcing. Computers & Security, 80(??):224–237, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303067.

Norta:2019:SFB

[NML19] Alex Norta, Raimundas Matulevicius, and Benjamin Leiding. Safeguarding a formalized blockchain-enabled identityauthentication protocol by applying security risk-oriented patterns. Computers & Security, 86(??):253–269, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818302670.

Ntantogian:2015:GTF

[NMX15] Christoforos Ntantogian, Stefanos Malliaros, and Christos Xenakis. Gaithashing: a two-factor authentication scheme based on gait features. Computers & Security, 52(??):17– 32, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000413.

Ntantogian:2019:EPH

[NMX19] Christoforos Ntantogian, Stefanos Malliaros, and Christos Xenakis. Evaluation of password hashing schemes in open source web platforms. Computers & Security, 84(??):206– 224, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818308332.

Navarro:2018:LOM

[NNG^{+}18] Luiz C. Navarro, Alexandre K. W. Navarro, André Grégio, Anderson Rocha, and Ricardo Dahab. Leveraging ontologies and machine-learning techniques for malware analysis into Android permissions ecosystems. Computers & Security, 78(??):429–453, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302311.

Nguyen:2018:ADS

[NNNQ18] Minh Hai Nguyen, Dung Le Nguyen, Xuan Mao Nguyen, and Tho Thanh Quan. Auto-detection of sophisticated malware using lazy-binding control flow graph and deep learning. Computers & Security, 76(??):128–155, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300889.

Nazir:2017:AAS

[NPP17] Sajid Nazir, Shushma Patel, and Dilip Patel. Assessing and augmenting SCADA cyber security: a survey of techniques. Computers & Security, 70(??):436–454, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301293.

Ni:2018:MIU

[NQZ18] Sang Ni, Quan Qian, and Rui Zhang. Malware identification using visualization images and deep learning. Computers & Security, 77(??):871–885, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303481.

Nguyen:2019:KPT

[NRM19] Toan Nguyen, Aditi Roy, and Nasir Memon. Kid on the phone! Toward automatic detection of children on mobile devices. Computers & Security, 84(??):334–348, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303407.

Nguyen:2017:DPA

[NSBM17] Toan Van Nguyen, Napa Sae-Bae, and Nasir Memon. DRAW-a-PIN: Authentication using finger-drawn PIN on touch devices. Computers & Security, 66(??):115–128, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300123.

Neisse:2015:SMB

[NSFB15] Ricardo Neisse, Gary Steri, Igor Nai Fovino, and Gianmarco Baldini. SecKit: a model-based security toolkit for the Internet of Things. Computers & Security, 54(??):60–76, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000887.

Neisse:2016:PEF

[NSGF16] Ricardo Neisse, Gary Steri, Dimitris Geneiatakis, and Igor Nai Fovino. A privacy enforcing framework for Android applications. Computers & Security, 62(??):257–277, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300840.

Nur:2018:RRI

[NT18] Abdullah Yasin Nur and Mehmet Engin Tozal. Record route IP traceback: Combating DoS attacks and the variants. Computers & Security, 72(??):13–25, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301773.

Nurain:2019:PAI

[NTK^{+}19] Novia Nurain, Suraiya Tairin, Taslim Arefin Khan, Shahad Ishraq, and A. B. M. Alim Al Islam. Power attack: an imminent security threat in real-time system for detecting missing rail blocks in developing countries. Computers & Security, 84(??):35–52, July 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308794.

Nguyen-Vu:2019:AFM

[NVAJ19] Long Nguyen-Vu, Jinung Ahn, and Souhwan Jung. Android fragmentation in malware detection. Computers & Security, 87(??):Article 101573, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301361.

Nicholson:2012:SSL

[NWD^{+}12] A. Nicholson, S. Webber, S. Dyer, T. Patel, and H. Janicke. SCADA security in the light of cyber-warfare. Computers & Security, 31(4):418–436, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000429.

Ntantogian:2010:GME

[NXS10] Christoforos Ntantogian, Christos Xenakis, and Ioannis Stavrakakis. A generic mechanism for efficient authentication in b3g networks. Computers & Security, 29(4):460–475, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809001242.

Nissim:2017:UBA

[NYE17] Nir Nissim, Ran Yahalom, and Yuval Elovici. USB-based attacks. Computers & Security, 70(??):675–688, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301578.

Ni:2016:FMR

[NZGH16] Siru Ni, Yi Zhuang, Jingjing Gu, and Ying Huo. A formal model and risk assessment method for security-critical realtime embedded systems. Computers & Security, 58(??):199– 215, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816000079.

Oracevic:2017:SRO

[OAO17] Alma Oracevic, Serkan Akbas, and Suat Ozdemir. Secure and reliable object tracking in wireless sensor networks. Computers & Security, 70(??):307–318, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301281.

Ozcelik:2015:DEB

[ÖB15] Ilker Özçelik and Richard R. Brooks. Deceiving entropy based DoS detection. Computers & Security, 48(??):234–245, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481400159X.

Onarlioglu:2015:SSL

[OBRK15] Kaan Onarlioglu, Ahmet Salih Buyukkayhan, William Robertson, and Engin Kirda. SENTINEL: Securing legacy Firefox extensions. Computers & Security, 49(??):147–161, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001825.

Oluwatimi:2018:CAS

[ODB18] Oyindamola Oluwatimi, Maria Luisa Damiani, and Elisa Bertino. A context-aware system to secure enterprise content: Incorporating reliability specifiers. Computers & Security, 77(??):162–178, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818303018.

Ouedraogo:2013:SCS

[OM13] Moussa Ouedraogo and Haralambos Mouratidis. Selecting a cloud service provider in the age of cybercrime.

Computers & Security, 38(??):3–13, October 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000242.

Opasiak:2019:SAD

[OM19] Krzysztof Opasiak and Wojciech Mazurczyk. (in)secure Android debugging: Security analysis and lessons learned. Computers & Security, 82(??):80–98, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481831023X.

Okolica:2011:WDM

[OP11] James S. Okolica and Gilbert L. Peterson. Windows driver memory analysis: a reverse engineering methodology. Computers & Security, 30(8):770–779, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000988.

Obert:2016:PAE

[OPHC16] James Obert, Inna Pivkina, Hong Huang, and Huiping Cao. Proactively applied encryption in multipath networks. Computers & Security, 58(??):106–124, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001960.

Oza:2014:HAD

[ORLS14] Aditya Oza, Kevin Ross, Richard M. Low, and Mark Stamp. HTTP attack detection using n-gram analysis. Computers & Security, 45(??):242–254, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000959.

Osborn:2017:SSI

[OS17] Emma Osborn and Andrew Simpson. On small-scale IT users’ system architectures and cyber security: a UK case study. Computers & Security, 70(??):27–50, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300925.

Ohtake:2019:OSA

[OSNZ19] Go Ohtake, Reihaneh Safavi-Naini, and Liang Feng Zhang. Outsourcing scheme of ABE encryption secure against malicious adversary. Computers & Security, 86(??):437–452, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819300525.

Odelu:2019:EPP

[OSP^{+}19] Vanga Odelu, Sourav Saha, Rajendra Prasath, Lakshminarayana Sadineni, Mauro Conti, and Minho Jo. Efficient privacy preserving device authentication in WBANs for industrial e-health applications. Computers & Security, 83(??):300–312, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308204.

Ogutcu:2016:API

[ÖTC16] Gizem Ögütçü, Özlem Müge Testik, and Oumout Chouseinoglou. Analysis of personal information security behavior and awareness. Computers & Security, 56(??):83–93, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815001406.

Padayachee:2012:TCI

[Pad12] Keshnee Padayachee. Taxonomy of compliant information security behavior. Computers & Security, 31(5):673–680, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481200065X.

Patel:2015:NMT

[PAJAP15] Ahmed Patel, Samaher Al-Janabi, Ibrahim AlShourbaji, and Jens Pedersen. A novel methodology towards a trusted environment in mashup web applications. Computers & Security, 49(??):107–122, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001552.

Palmer:2010:ASM

[Pal10] Anthony J. Palmer. Approach for selecting the most suitable automated personal identification mechanism (ASMSA). Computers & Security, 29(7):785–806, October 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000325.

Patel:2017:NCI

[PAP^{+}17] Ahmed Patel, Hitham Alhussian, Jens Myrup Pedersen, Bouchaib Bounabat, Joaquim Celestino Júnior, and Sokratis Katsikas. A nifty collaborative intrusion detection and prevention architecture for smart grid ecosystems. Computers & Security, 64(??):92–109, January 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300748.

Park:2018:OTP

[Par18] Chang-Seop Park. One-time password based on hash chain without shared secret and re-registration. Computers & Security, 75(??):138–146, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818301391.

Polverini:2018:REP

[PAS^{+}18] D. Polverini, F. Ardente, I. Sanchez, F. Mathieux, P. Tecchio, and L. Beslay. Resource efficiency, privacy and security by design: a first experience on enterprise servers and data storage products triggered by a policy process. Computers & Security, 76(??):295–310, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302614.

Posey:2011:UMA

[PBR11] Clay Posey, Rebecca J. Bennett, and Tom L. Roberts. Understanding the mindset of the abusive insider: an examination of insiders’ causal reasoning following internal security changes. Computers & Security, 30(6–7): 486–497, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000630.

Pfleeger:2012:LBS

[PC12] Shari Lawrence Pfleeger and Deanna D. Caputo. Leveraging behavioral science to mitigate cyber security risk. Computers & Security, 31(4):597–611, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001659.

Pizzolante:2018:PCG

[PCC^{+}18] Raffaele Pizzolante, Arcangelo Castiglione, Bruno Carpentieri, Alfredo De Santis, Francesco Palmieri, and Aniello Castiglione. On the protection of consumer genomic data in the Internet of living things. Computers & Security, 74(??):384– 400, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301220.

Pierazzi:2016:ESA

[PCCM16] Fabio Pierazzi, Sara Casolari, Michele Colajanni, and Mirco Marchetti. Exploratory security analytics for anomaly detection. Computers & Security, 56(??):28–49, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001480.

Parsons:2017:HAI

[PCP^{+}17] Kathryn Parsons, Dragana Calic, Malcolm Pattinson, Marcus Butavicius, Agata McCormac, and Tara Zwaans. The human aspects of information security questionnaire (HAIS-Q): Two further validation studies. Computers & Security, 66(??):40–51, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300081.

Patsakis:2018:MRT

[PCP^{+}18] Constantinos Patsakis, Athanasios Charemis, Achilleas Papageorgiou, Dimitrios Mermigas, and Sotirios Pirounias. The market’s response toward privacy and mass surveillance: the Snowden aftermath. Computers & Security, 73(??):194–206, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481730233X.

Patsakis:2014:TDS

[PDB14] Constantinos Patsakis, Kleanthis Dellios, and Mélanie Bouroche. Towards a distributed secure in-vehicle communication architecture for modern vehicles. Computers & Security, 40(??):60–74, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481300165X.

Patil:2019:DES

[PDM19] Rajendra Patil, Harsha Dudeja, and Chirag Modi. Designing an efficient security framework for detecting intrusions in virtual network of cloud computing. Computers & Security, 85(??):402–422, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818310629.

Palmieri:2010:NAD

[PF10] Francesco Palmieri and Ugo Fiore. Network anomaly detection through nonlinear analysis. Computers & Security, 29(7):737–755, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000362.

Paladi:2019:SAC

[PG19] Nicolae Paladi and Christian Gehrmann. SDN access control for the masses. Computers & Security, 80(??):155–172, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818306047.

Pisani:2016:ETU

[PGdCL16] Paulo Henrique Pisani, Romain Giot, André C. P. L. F. de Carvalho, and Ana Carolina Lorena. Enhanced template update: Application to keystroke dynamics. Computers & Security, 60(??):134–153, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481630044X.

Pasquale:2016:AEC

[PHMN16] Liliana Pasquale, Sorren Hanvey, Mark Mcgloin, and Bashar Nuseibeh. Adaptive evidence collection in the cloud using attack scenarios. Computers & Security, 59(??):236–254, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300232.

Pan:2019:CAI

[PHP19] Zhiwen Pan, Salim Hariri, and Jesus Pacheco. Context aware intrusion detection for building automation systems. Computers & Security, 85(??):181–201, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300926.

Pham:2017:TCB

[PK17] Dang Vinh Pham and Dogan Kesdogan. Towards a causality based analysis of anonymity protection in indeterministic mix systems. Computers & Security, 67(??):350–368, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300378.

Park:2014:TOS

[PKK^{+}14] Joon S. Park, Kevin A. Kwiat, Charles A. Kamhoua, Jonathan White, and Sookyung Kim. Trusted online social network (OSN) services with optimal data management. Computers & Security, 42(??):116–136, May 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000236.

Park:2017:RIS

[PKP17] Eun Hee Park, Jongwoo Kim, and Young Soon Park. The role of information security learning and individual factors in disclosing patients’ health information. Computers & Security, 65(??):64–76, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301444.

Park:2019:FAI

[PKWP19] Eun Hee Park, Jongwoo Kim, Lynn L. Wiles, and Young Soon Park. Factors affecting intention to disclose patients’ health information. Computers & Security, 87 (??):Article 101340, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304917.

Parsons:2014:DEA

[PMB^{+}14] Kathryn Parsons, Agata McCormac, Marcus Butavicius, Malcolm Pattinson, and Cate Jerram. Determining employee awareness using the human aspects of information security questionnaire (HAIS-Q). Computers & Security, 42(??):165– 176, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481300179X.

Pavlich-Mariscal:2010:FSA

[PMDM10a] Jaime A. Pavlich-Mariscal, Steven A. Demurjian, and Laurent D. Michel. A framework for security assurance of access control enforcement code. Computers & Security, 29(7):770–784, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000349.

Pavlich-Mariscal:2010:FCA

[PMDM10b] Jaime A. Pavlich-Mariscal, Steven A. Demurjian, and Laurent D. Michel. A framework of composable access control features: Preserving separation of access control concerns from models to code. Computers & Security, 29(3):350–379, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809001382.

Pereniguez:2011:PUP

[PMLK^{+}11] F. Pereniguez, R. Marin-Lopez, G. Kambourakis, S. Gritzalis, and A. F. Gomez. PrivaKERB: a user privacy framework for Kerberos. Computers & Security, 30(6– 7):446–463, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000617.

Parsons:2015:DPS

[PMP^{+}15] Kathryn Parsons, Agata McCormac, Malcolm Pattinson, Marcus Butavicius, and Cate Jerram. The design of phishing studies: Challenges for researchers. Computers & Security, 52(??):194–206, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000231.

Putz:2019:SAL

[PMP19] Benedikt Putz, Florian Menges, and Günther Pernul. A secure and auditable logging infrastructure based on a permissioned blockchain. Computers & Security, 87 (??):Article 101602, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313907.

Park:2017:ISR

[PNKK17] Jaehyun Park, Jiseong Noh, Myungchul Kim, and Brent Byunghoon Kang. Invi-server: Reducing the attack surfaces by making protected server invisible on networks. Computers & Security, 67(??):89–106, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730038X.

Pieterse:2014:BCC

[PO14] Heloise Pieterse and Martin S. Olivier. Bluetooth command and control channel. Computers & Security, 45(??):75–83, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000881.

Pisani:2017:SNA

[PPdCL17] Paulo Henrique Pisani, Norman Poh, André C. P. L. F. de Carvalho, and Ana Carolina Lorena. Score normalization applied to adaptive biometric systems. Computers & Security, 70(??):565–580, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301554.

Park:2019:ESE

[PPKK19] Jaemin Park, Sungjin Park, Brent Byunghoon Kang, and Kwangjo Kim. eMotion: an SGX extension for migrating enclaves. Computers & Security, 80(??):173–185, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818304279.

Pham:2010:TD

[PRMD10] Quan Pham, Jason Reid, Adrian McCullagh, and Edward Dawson. On a taxonomy of delegation. Computers & Security, 29(5):565–579, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001473.

Park:2013:DCM

[PRS13] Younghee Park, Douglas S. Reeves, and Mark Stamp. Deriving common malware behavior through graph clustering. Computers & Security, 39 (part B)(??):419–430, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813001351.

Pironti:2010:PCJ

[PS10] Alfredo Pironti and Riccardo Sisto. Provably correct Java implementations of Spi Calculus security protocols specifications. Computers & Security, 29(3):302–314, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000832.

Palumbo:2017:PAM

[PSK^{+}17] Paolo Palumbo, Luiza Sayfullina, Dmitriy Komashinskiy, Emil Eirola, and Juha Karhunen. A pragmatic Android malware detection procedure. Computers & Security, 70(??):689–701, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301542.

Paul:2018:ADC

[PSMG18] Padma Polash Paul, Madeena Sultana, Sorin Adam Matei, and Marina Gavrilova. Authorship disambiguation in a collaborative editing environment. Computers & Security, 77(??):675–693, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300294.

Pellet:2019:LSN

[PSS19] Hector Pellet, Stavros Shiaeles, and Stavros Stavrou. Localising social network users and profiling their movement. Computers & Security, 81(??):49–57, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301524.

Pipyros:2018:NSI

[PTM^{+}18] Kosmas Pipyros, Christos Thraskias, Lilian Mitrou, Dimitris Gritzalis, and Theodoros Apostolopoulos. A new strategy for improving cyber-attacks evaluation in the context of Tallinn Manual. Computers & Security, 74(??):371–383, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300822.

Paulin:2013:USF

[PW13] Alois Paulin and Tatjana Welzer. A universal system for fair non-repudiable certified e-mail without a trusted third party. Computers & Security, 32(??):207–218, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001757.

Panda:2019:PIA

[PWL^{+}19] Sakshyam Panda, Daniel W. Woods, Aron Laszka, Andrew Fielder, and Emmanouil Panaousis. Post-incident audits on cyber insurance discounts. Computers & Security, 87 (??):Article 101593, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819300641.

Peng:2014:PMS

[PWWY14] Sancheng Peng, Min Wu, Guojun Wang, and Shui Yu. Propagation model of smartphone worms based on semi-Markov process and social relationship graph. Computers & Security, 44(??):92–103, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000583.

Peng:2019:SIP

[PYLL19] Fei Peng, Jing Yang, Zi-Xing Lin, and Min Long. Source identification of 3D printed objects based on inherent equipment distortion. Computers & Security, 82(??):173–183, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818309404.

Pang:2015:NAC

[PZ15] Jun Pang and Yang Zhang. A new access control scheme for Facebook-style social networks. Computers & Security, 54(??):44–59, October 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000632.

Qamar:2017:DDA

[QAR^{+}17] Sara Qamar, Zahid Anwar, Mohammad Ashiqur Rahman, Ehab Al-Shaer, and Bei-Tseng Chu. Data-driven analytics for cyber-threat intelligence and information sharing. Computers & Security, 67(??):35–58, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300287.

Qian:2012:MIS

[QFG12] Ying Qian, Yulin Fang, and Jose J. Gonzalez. Managing information security risks during new technology adoption. Computers & Security, 31(8):859–869, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001368.

Qin:2019:TPB

[QHW^{+}19] Maoyuan Qin, Wei Hu, Xinmu Wang, Dejun Mu, and Baolei Mao. Theorem proof based gate level information flow tracking for hardware security verification. Computers & Security, 85(??):225–239, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819300975.

Qiu:2019:CPT

[QLZ19] Jian Qiu, Hengjian Li, and Chuan Zhao. Cancelable palmprint templates based on random measurement and noise data for security and privacy-preserving authentication. Computers & Security, 82(??):1–14, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306618.

Zhang:2017:NMT

[qZLxCjY17] Hong qi Zhang, Cheng Lei, De xian Chang, and Ying jie Yang. Network moving target defense technique based on collaborative mutation. Computers & Security, 70(??):51–71, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301141.

Razzaq:2014:OAD

[RAA^{+}14] Abdul Razzaq, Zahid Anwar, H. Farooq Ahmad, Khalid Latif, and Faisal Munir. Ontology for attack detection: an intelligent approach to web application security. Computers & Security, 45(??):124–146, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000868.

Ramaki:2015:RRT

[RAA15] Ali Ahmadian Ramaki, Morteza Amini, and Reza Ebrahimi Atani. RTECA: Real time episode correlation algorithm for multi-step attack scenarios detection. Computers & Security, 49(??):206–219, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001527.

Ruiz:2017:SKR

[RAMR17] Jose Fran. Ruiz, Marcos Arjona, Antonio Maña, and Carsten Rudolph. Security knowledge representation artifacts for creating secure IT systems. Computers & Security, 64(??):69–91, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300992.

Rubio:2019:CCD

[RARL19] Juan Enrique Rubio, Cristina Alcaraz, Rodrigo Roman, and Javier Lopez. Current cyber-defense trends in industrial control systems. Computers & Security, 87(??):Article 101561, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301245.

Rekhis:2011:LBA

[RB11] Slim Rekhis and Noureddine Boudriga. Logic-based approach for digital forensic investigation in communication networks. Computers & Security, 30(6–7):376–396, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000447.

Rekhis:2012:HVT

[RB12] Slim Rekhis and Noureddine Boudriga. A hierarchical visibility theory for formal digital investigation of anti-forensic attacks. Computers & Security, 31(8):967–982, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001022.

Rathgeb:2014:CMB

[RB14] C. Rathgeb and C. Busch. Cancelable multi-biometrics: Mixing iris-codes based on adaptive Bloom filters. Computers & Security, 42(??):1–12, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000029.

Rhode:2018:ESM

[RBJ18] Matilda Rhode, Pete Burnap, and Kevin Jones. Earlystage malware prediction using recurrent neural networks. Computers & Security, 77(??):578–594, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818305546.

Rahman:2015:SIS

[RC15] Nurul Hidayah Ab Rahman and Kim-Kwang Raymond Choo. A survey of information security incident handling in the cloud. Computers & Security, 49(??):45–69, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001680.

Ramaswamy:2014:E

[RCB^{+}14] Lakshmish Ramaswamy, Barbara Carminati, Lujo Bauer, Dongwan Shin, James Joshi, Calton Pu, and Dimitris Gritzalis. Editorial. Computers & Security, 41(??):1–2, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000078.

Rashidibajgan:2019:PPH

[RD19] Samaneh Rashidibajgan and Robin Doss. Privacy-preserving history-based routing in opportunistic networks. Computers & Security, 84(??):244–255, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818312148.

Rahman:2019:SDA

[RDAS19] Mohammad Ashiqur Rahman, Amarjit Datta, and Ehab Al-Shaer. Security design against stealthy attacks on power system state estimation: a formal approach. Computers & Security, 84(??):301–317, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300725.

Rajab:2019:EEP

[RE19] Majed Rajab and Ali Eydgahi. Evaluating the explanatory power of theoretical frameworks on intention to comply with information security policies in higher education. Computers & Security, 80(??):211–223, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311325.

Rashidi:2017:ARU

[RFB17] Bahman Rashidi, Carol Fung, and Elisa Bertino. Android resource usage risk assessment using hidden Markov model and online learning. Computers & Security, 65(??):90–107, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816301596.

Rios:2018:MPA

[RFGL18] Ruben Rios, Carmen Fernandez-Gago, and Javier Lopez. Modelling privacy-aware trust negotiations. Computers & Security, 77(??):773–789, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302043.

Renaud:2018:RCS

[RFW^{+}18] Karen Renaud, Stephen Flowerday, Merrill Warkentin, Paul Cockshott, and Craig Orgeron. Is the responsibilization of the cyber security risk reasonable and judicious? Computers & Security, 78(??):198–211, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303262.

Raju:2019:STB

[RG19] BKSP Kumar Raju and G. Geethakumari. SNAPS: Towards building snapshot based provenance system for virtual machines in the cloud environment. Computers & Security, 86(??):92–111, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301895.

Ruohonen:2015:SGO

[RHL15] Jukka Ruohonen, Sami Hyrynsalmi, and Ville Leppänen. The sigmoidal growth of operating system security vulnerabilities: an empirical revisit. Computers & Security, 55(??):1–20, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001005.

Riley:2013:FPT

[Ril13] Ryan Riley. A framework for prototyping and testing dataonly rootkit attacks. Computers & Security, 37(??):62–71, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000825.

Robinson:2015:CWI

[RJJ15] Michael Robinson, Kevin Jones, and Helge Janicke. Cyber warfare: Issues and challenges. Computers & Security, 49(??):70–94, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001692.

Rao:2011:FGI

[RLB^{+}11] Prathima Rao, Dan Lin, Elisa Bertino, Ninghui Li, and Jorge Lobo. Fine-grained integration of access control policies. Computers & Security, 30(2–3):91–107, March 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000891.

Ruzicic:2017:CSN

[RM17] Vesna S. Ruzicić and Zivadin M. Micić. Creating a strategic national knowledge architecture: a comparative analysis of knowledge source innovation in the ICS subfields of multimedia and IT security. Computers & Security, 70(??):455–466, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817301487.

Rrushi:2011:EBV

[RMG11] Julian Rrushi, Ehsan Mokhtari, and Ali A. Ghorbani. Estimating botnet virulence within mathematical models of botnet propagation dynamics. Computers & Security, 30(8):791–802, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000964.

Rodero-Merino:2012:BSP

[RMVC^{+}12] Luis Rodero-Merino, Luis M. Vaquero, Eddy Caron, Adrian Muresan, and Frédéric Desprez. Building safe PaaS clouds: a survey on security in multitenant software platforms. Computers & Security, 31(1):96–108, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001313.

Ruighaver:2010:EDM

[RMW10] A. B. Ruighaver, S. B. Maynard, and M. Warren. Ethical decision making: Improving the quality of acceptable use policies. Computers & Security, 29(7):731–736, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000386.

Rodriguez:2017:TAS

[RNdP17] David Martínez Rodríguez, Jordi Nin, and Miguel Nuñez del Prado. Towards the adaptation of SDC methods to stream mining. Computers & Security, 70(??):702–722, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301761.

Roumani:2015:TSM

[RNR15] Yaman Roumani, Joseph K. Nwankpa, and Yazan F. Roumani. Time series modeling of vulnerabilities. Computers & Security, 51(??):32–40, June 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000358.

Rios:2013:CCT

[ROL13] Ruben Rios, Jose A. Onieva, and Javier Lopez. Covert communications through network configuration messages. Computers & Security, 39 (Part A)(??):34–46, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000497.

Renaud:2019:WCM

[ROW19] Karen Renaud, Robert Otondo, and Merrill Warkentin. “This is the way ‘I’ create my passwords” … does the endowment effect deter people from changing the way they create their passwords? Computers & Security, 82(??):241–260, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481830909X.

Rao:2019:JPI

[RP19] Routhu Srinivasa Rao and Alwyn Roshan Pais. Jail-Phish: an improved search engine based phishing detection system. Computers & Security, 83(??):246–267, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304280.

Rughinis:2014:NVN

[RR14] Cosima Rughinis and Razvan Rughinis. Nothing ventured, nothing gained. Profiles of online activity, cyber-crime exposure, and security measures of end-users in European Union. Computers & Security, 43(??):111–125, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000479.

Razaque:2016:TDP

[RR16] Abdul Razaque and Syed S. Rizvi. Triangular data privacypreserving model for authenticating all key stakeholders in a cloud environment. Computers & Security, 62(??):328–347, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300979.

Razaque:2017:SDA

[RR17] Abdul Razaque and Syed S. Rizvi. Secure data aggregation using access control and authentication for wireless sensor networks. Computers & Security, 70(??):532–545, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301402.

Rhee:2012:UOI

[RRK12] Hyeun-Suk Rhee, Young U. Ryu, and Cheong-Tag Kim. Unrealistic optimism on information security management. Computers & Security, 31(2):221–232, March 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001441.

Razaque:2017:SQS

[RRK^{+}17] Abdul Razaque, Syed S. Rizvi, Meer J. Khan, Qassim B. Hani, Julius P. Dichter, and Reza M. Parizi. Secure and quality-of-service-supported service-oriented architecture for mobile cloud handoff process. Computers & Security, 66(??):169–184, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730010X.

Rrushi:2016:NDT

[Rru16] Julian L. Rrushi. NIC displays to thwart malware attacks mounted from within the OS. Computers & Security, 61(??):59–71, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300517.

Reece:2015:PIS

[RS15] R. P. Reece and B. C. Stahl. The professionalisation of information security: Perspectives of UK practitioners. Computers & Security, 48(??):182–195, February 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001539.

Ring:2019:FBN

[RSLH19] Markus Ring, Daniel Schlör, Dieter Landes, and Andreas Hotho. Flow-based network traffic generation using generative adversarial networks. Computers & Security, 82(??):156– 172, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818308393.

Ruan:2017:ICU

[Rua17] Keyun Ruan. Introducing cybernomics: a unifying economic framework for measuring cyber risk. Computers & Security, 65(??):77–89, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301407.

Reddy:2013:ADF

[RV13] K. Reddy and H. S. Venter. The architecture of a digital forensic readiness management system. Computers & Security, 32(??):73–89, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001447.

Rossudowski:2010:SPA

[RVEK10] A. M. Rossudowski, H. S. Venter, J. H. P. Eloff, and D. G. Kourie. A security privacy aware architecture and protocol for a single smart card used for multiple services. Computers & Security, 29(4):393–409, June 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001394.

Ramanathan:2013:PDI

[RW13] Venkatesh Ramanathan and Harry Wechsler. Phishing detection and impersonated entity discovery using Conditional Random Field and Latent Dirichlet Allocation. Computers & Security, 34(??):123–139, May 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001812.

Ring:2019:SNB

[RWS^{+}19] Markus Ring, Sarah Wunderlich, Deniz Scheuring, Dieter Landes, and Andreas Hotho. A survey of networkbased intrusion detection data sets. Computers & Security, 86(??):147–167, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481930118X.

Spanos:2016:IIS

[SA16] Georgios Spanos and Lefteris Angelis. The impact of information security events to the stock market: a systematic literature review. Computers & Security, 58(??):216–229, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300013.

Senarath:2019:DMM

[SA19] Awanthika Senarath and Nalin Asanka Gamagedara Arachchilage. A data minimization model for embedding privacy into software systems. Computers & Security, 87 (??):Article 101605, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818309106.

Shaikh:2012:DRB

[SAL12] Riaz Ahmed Shaikh, Kamel Adi, and Luigi Logrippo. Dynamic risk-based decision methods for access control systems. Computers & Security, 31(4):447–464, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000399.

Savola:2013:QSM

[Sav13] Reijo M. Savola. Quality of security metrics and measurements. Computers & Security, 37(??):78–90, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000850.

Silic:2014:SIV

[SB14] Mario Silic and Andrea Back. Shadow IT — a view from behind the curtain. Computers & Security, 45(??):274–283, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481400100X.

Spyra:2017:SPA

[SBE17] Grzegorz Spyra, William J. Buchanan, and Elias Ekonomou. Sticky policies approach within cloud computing. Computers & Security, 70(??):366–375, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730144X.

Scholte:2012:TCN

[SBK12] Theodoor Scholte, Davide Balzarotti, and Engin Kirda. Have things changed now? An empirical study on input validation vulnerabilities in web applications. Computers & Security, 31(3):344–356, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001684.

Sai:2019:ASI

[SBL19] Ashish Rajendra Sai, Jim Buckley, and Andrew Le Gear. Assessing the security implication of Bitcoin exchange rates. Computers & Security, 86(??):206–222, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818312112.

Symeonidis:2018:CDF

[SBS^{+}18] Iraklis Symeonidis, Gergely Biczók, Fatemeh Shirazi, Cristina Pérez-Solà, Jessica Schroers, and Bart Preneel. Collateral damage of Facebook third-party applications: a comprehensive study. Computers & Security, 77(??):179–208, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481830302X.

Stakhanova:2010:SSB

[SBW10] Natalia Stakhanova, Samik Basu, and Johnny Wong. On the symbiosis of specification-based and anomaly-based detection. Computers & Security, 29(2):253–268, March 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000893.

Sahay:2017:ASB

[SBZD17] Rishikesh Sahay, Gregory Blanc, Zonghua Zhang, and Hervé Debar. ArOMA: an SDN based autonomic DDoS mitigation framework. Computers & Security, 70(??):482–499, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301499.

Sindiren:2019:AMP

[SC19] Erhan Sindiren and Bünyamin Ciylan. Application model for privileged account access control system in enterprise networks. Computers & Security, 83(??):52–67, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304395.

Salah:2013:ASW

[SCB^{+}13] Khaled Salah, Jose M. Alcaraz Calero, Jorge Bernal Bernabé, Juan M. Marín Perez, and Sherali Zeadally. Analyzing the security of Windows 7 and Linux for cloud computing. Computers & Security, 34(??):113–122, May 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001800.

Saevanee:2015:CUA

[SCFB15] Hataichanok Saevanee, Nathan Clarke, Steven Furnell, and Valerio Biscione. Continuous user authentication using multi-modal biometrics. Computers & Security, 53(??):234–246, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000875.

Shen:2014:PEA

[SCGM14] Chao Shen, Zhongmin Cai, Xiaohong Guan, and Roy Maxion. Performance evaluation of anomaly-detection algorithms for mouse dynamics. Computers & Security, 45(??):156–171, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000807.

Sabottke:2019:HTB

[SCLD19] Carl Sabottke, Daniel Chen, Lucas Layman, and Tudor Dumitras. How to trick the Borg: threat models against manual and automated techniques for detecting network attacks. Computers & Security, 81(??):25–40, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311283.

Smith-Creasey:2019:NWI

[SCR19] Max Smith-Creasey and Muttukrishnan Rajarajan. A novel word-independent gesture-typing continuous authentication scheme for mobile devices. Computers & Security, 83(??):140–150, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818306552.

Sciarretta:2017:AFS

[SCRA17] Giada Sciarretta, Roberto Carbone, Silvio Ranise, and Alessandro Armando. Anatomy of the Facebook solution for mobile single sign-on: Security assessment and improvements. Computers & Security, 71(??):71–86, November 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300895.

S:2017:EOB

[SD17] Shitharth S. and Prince Winston D. An enhanced optimization based algorithm for intrusion detection in SCADA network. Computers & Security, 70(??):16–26, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300901.

Stavova:2018:ELS

[SDM^{+}18] Vlasta Stavova, Lenka Dedkova, Vashek Matyas, Mike Just, David Smahel, and Martin Ukrop. Experimental large-scale review of attractors for detection of potentially unwanted applications. Computers & Security, 76(??):92–100, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301640.

Sachan:2011:RVD

[SE11] Amit Sachan and Sabu Emmanuel. Rights violation detection in multi-level digital rights management system. Computers & Security, 30(6–7):498–513, September/October 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000642.

Sommestad:2010:PRM

[SEJ10] Teodor Sommestad, Mathias Ekstedt, and Pontus Johnson. A probabilistic relational model for security risk analysis. Computers & Security, 29(6):659–679, September 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000209.

Sallam:2012:EBM

[SERF12] Ahmed I. Sallam, El-Sayed El-Rabaie, and Osama S. Faragallah. Encryption-based multilevel model for DBMS. Computers & Security, 31(4):437–446, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000417.

Scanlon:2015:NIM

[SFK15] Mark Scanlon, Jason Farina, and M-Tahar Kechadi. Network investigation methodology for BitTorrent sync: a peer-to-peer based file synchronisation service. Computers & Security, 54(??):27–43, October 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500067X.

Seewald:2010:DIB

[SG10a] Alexander K. Seewald and Wilfried N. Gansterer. On the detection and identification of botnets. Computers & Security, 29(1):45–58, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000820.

Soupionis:2010:ACE

[SG10b] Yannis Soupionis and Dimitris Gritzalis. Audio CAPTCHA: Existing solutions assessment and a new implementation for VoIP telephony. Computers & Security, 29(5):603–618, July 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809001412.

Soupionis:2013:HV

[SG13] Yannis Soupionis and Dimitris Gritzalis. Hacking VoIP. Computers & Security, 32(??):267, February 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812001423.

Stavrou:2014:ICN

[SG14a] Vasilis Stavrou and Dimitris Gritzalis. Introduction to computer and network security: Navigating shades of Gray. Computers & Security, 46(??):173, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000303.

Stergiopoulos:2014:BHD

[SG14b] George Stergiopoulos and Dimitris Gritzalis. A bug Hunter’s diary. Computers & Security, 43(??):188, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000200.

Stavrou:2015:BRI

[SG15a] Vasilis Stavrou and Dimitris Gritzalis. Book review: Introduction to Social Media Investigation — a hands-on Approach, Jennifer Golbeck, Elsevier Publications, USA (2015).

Computers & Security, 55(??):128–129, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001194.

Stergiopoulos:2015:HPT

[SG15b] George Stergiopoulos and Dimitris Gritzalis. Hacking and penetration testing with low power devices. Computers & Security, 49(??):274–275, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000024.

Such:2016:IAT

[SGK^{+}16] Jose M. Such, Antonios Gouglidis, William Knowles, Gaurav Misra, and Awais Rashid. Information assurance techniques: Perceived cost effectiveness. Computers & Security, 60(??):117–133, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300311.

Saini:2016:CBE

[SGLC16] Anil Saini, Manoj Singh Gaur, Vijay Laxmi, and Mauro Conti. Colluding browser extension attack on user privacy and its implication for web browsers. Computers & Security, 63(??):14–28, November 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301018.

Shen:2015:DHS

[SH15] Shu-Yuan Shen and Li-Hong Huang. A data hiding scheme using pixel value differencing and improving exploiting modification directions. Computers & Security, 48(??):131–141, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001199.

Sun:2012:SBF

[SHB12] San-Tsai Sun, Kirstie Hawkey, and Konstantin Beznosov. Systematically breaking and fixing OpenID security: Formal analysis, semi-automated empirical evaluation, and practical countermeasures. Computers & Security, 31(4):465–483, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000387.

Shittu:2015:IAP

[SHGH^{+}15] Riyanat Shittu, Alex Healing, Robert Ghanea-Hercock, Robin Bloomfield, and Muttukrishnan Rajarajan. Intrusion alert prioritisation and attack detection using postcorrelation analysis. Computers & Security, 50(??):1–15, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001837.

Silic:2013:DUO

[Sil13] Mario Silic. Dual-use open source security software in organizations — dilemma: Help or hinder? Computers & Security, 39 (part B)(??):386–395, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001326.

Sillic:2019:CIO

[Sil19] Mario Sillic. Critical impact of organizational and individual inertia in explaining non-compliant security behavior in the shadow IT context. Computers & Security, 80(??):108–119, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818306114.

Sapegin:2017:TSC

[SJCM17] Andrey Sapegin, David Jaeger, Feng Cheng, and Christoph Meinel. Towards a system for complex analysis of security events in large-scale networks. Computers & Security, 67(??):16–34, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730024X.

Spathoulas:2010:RFP

[SK10] Georgios P. Spathoulas and Sokratis K. Katsikas. Reducing false positives in intrusion detection systems. Computers & Security, 29(1):35–44, February 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809000844.

Spathoulas:2013:EIP

[SK13] Georgios P. Spathoulas and Sokratis K. Katsikas. Enhancing IDS performance through comprehensive alert postprocessing. Computers & Security, 37(??):176–196, September 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813000503.

Sepczuk:2018:NRB

[SK18] Mariusz Sepczuk and Zbigniew Kotulski. A new risk-based authentication management model oriented on user’s experience. Computers & Security, 73(??):17–33, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302079.

Sajjad:2019:EPP

[SKA^{+}19] Haider Sajjad, Tehsin Kanwal, Adeel Anjum, Saif ur Rehman Malik, Ahmed Khan, Abid Khan, and Umar Manzoor. An efficient privacy preserving protocol for dynamic continuous data collection. Computers & Security, 86(??):358–371, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301312.

Stergiopoulos:2017:PAR

[SKG17] George Stergiopoulos, Panayiotis Katsaros, and Dimitris Gritzalis. Program analysis with risk-based classification of dynamic invariants for logical error detection. Computers & Security, 71(??):36–50, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300330.

Schrittwieser:2014:CCH

[SKK^{+}14] Sebastian Schrittwieser, Stefan Katzenbeisser, Peter Kieseberg, Markus Huber, Manuel Leithner, Martin Mulazzani, and Edgar Weippl. Covert computation — hiding code in code through compile-time obfuscation. Computers & Security, 42(??):13–26, May 2014. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000030.

Shiaeles:2012:RTD

[SKKP12] Stavros N. Shiaeles, Vasilios Katos, Alexandros S. Karakos, and Basil K. Papadopoulos. Real time DDoS detection using fuzzy estimators. Computers & Security, 31(6):782–790, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000922.

Squicciarini:2014:IHS

[SKLD14] Anna Squicciarini, Sushama Karumanchi, Dan Lin, and Nicole DeSisto. Identifying hidden social circles for advanced privacy configuration. Computers & Security, 41(??):40–51, March 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001107.

Spyridopoulos:2013:GTD

[SKTO13] T. Spyridopoulos, G. Karanikas, T. Tryfonas, and G. Oikonomou. A game theoretic defence framework against DoS/DDoS cyber attacks. Computers & Security, 38(??):39–50, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481300059X.

Schmidt:2013:TFV

[SLB^{+}13] Andreas U. Schmidt, Andreas Leicher, Andreas Brett, Yogendra Shah, and Inhyok Cha. Tree-formed verification data for trusted platforms. Computers & Security, 32(??):19–35, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481200140X.

Shi:2012:SED

[SLD12] Jie Shi, Yingjiu Li, and Robert H. Deng. A secure and efficient discovery service system in EPCglobal network. Computers & Security, 31(8):870–885, November 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481200123X.

Sharafaldin:2019:EFN

[SLG19] Iman Sharafaldin, Arash Habibi Lashkari, and Ali A. Ghorbani. An evaluation framework for network security visualizations. Computers & Security, 84(??):70–92, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818308952.

Seo:2014:CCS

[SLH14] Ilju Seo, Heejo Lee, and Seung Chul Han. Cylindrical coordinates security visualization for multiple domain command and control botnet detection. Computers & Security, 46(??):141–153, October 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001187.

Seo:2015:MMS

[SLN^{+}15] Hwajeong Seo, Zhe Liu, Yasuyuki Nogami, Jongseok Choi, and Howon Kim. Montgomery multiplication and squaring for optimal prime fields. Computers & Security, 52(??):276– 291, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000371.

Seo:2013:AAP

[SLP13] Dongwon Seo, Heejo Lee, and Adrian Perrig. APFS: Adaptive probabilistic filter scheduling against distributed denial-of-service attacks. Computers & Security, 39 (part B)(??):366–385, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001314.

Shen:2018:TES

[SLS^{+}18] Dong Shen, Zhoujun Li, Xiaojing Su, Jinxin Ma, and Robert Deng. TinyVisor: an extensible secure framework on Android platforms. Computers & Security, 72(??):145–162, January 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301955.

Sorniotti:2010:PSS

[SM10] Alessandro Sorniotti and Refik Molva. A provably secure secret handshake with dynamic controlled matching. Computers & Security, 29(5):619–627, July 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001370.

Samarasinghe:2019:ALT

[SM19a] Nayanamana Samarasinghe and Mohammad Mannan. Another look at TLS ecosystems in networked devices vs. Web servers. Computers & Security, 80(??):1–13, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818306916.

Samarasinghe:2019:TGP

[SM19b] Nayanamana Samarasinghe and Mohammad Mannan. Towards a global perspective on web tracking. Computers & Security, 87(??):Article 101569, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818314007.

Sucasas:2018:PEO

[SMA^{+}18] Victor Sucasas, Georgios Mantas, Saud Althunibat, Leonardo Oliveira, Angelos Antonopoulos, Ifiok Otung, and Jonathan Rodriguez. A privacy-enhanced OAuth 2.0 based protocol for smart city mobile applications. Computers & Security, 74(??):258–274, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300361.

Sieklik:2016:ETD

[SMB16] Boris Sieklik, Richard Macfarlane, and William J. Buchanan. Evaluation of TFTP DDoS amplification attack. Computers & Security, 57(??):67–92, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001285.

Shen:2018:SEB

[SMC^{+}18] Lixiang Shen, Dejun Mu, Guo Cao, Maoyuan Qin, Jeremy Blackstone, and Ryan Kastner. Symbolic execution based test-patterns generation algorithm for hardware Trojan detection. Computers & Security, 78(??):267–280, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818305352.

Schmidt:2018:EIT

[SMK^{+}18a] Sabine Schmidt, Wojciech Mazurczyk, Radoslaw Kulesza, Jörg Keller, and Luca Caviglione. Exploiting IP telephony with silence suppression for hidden data transfers. Computers & Security, 79(??):17–32, November 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818305777.

Sundararaj:2018:OCF

[SMK18b] Vinu Sundararaj, Selvi Muthukumar, and R. S. Kumar. An optimal cluster formation based energy efficient dynamic scheduling hybrid MAC protocol for heavy traffic load in wireless sensor networks. Computers & Security, 77(??):277–288, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818303754.

Scalas:2019:ESA

[SMM^{+}19] Michele Scalas, Davide Maiorca, Francesco Mercaldo, Corrado Aaron Visaggio, Fabio Martinelli, and Giorgio Giacinto. On the effectiveness of system API-related information for Android ransomware detection. Computers & Security, 86(??):168–182, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301178.

Schreuders:2013:SAA

[SMP13] Z. Cliffe Schreuders, Tanya McGill, and Christian Payne. The state of the art of application restrictions and sandboxes: a survey of application-oriented access controls and their shortfalls. Computers & Security, 32(??):219–241, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001435.

Sucasas:2016:APP

[SMS^{+}16] Victor Sucasas, Georgios Mantas, Firooz B. Saghezchi, Ayman Radwan, and Jonathan Rodriguez. An autonomous privacy-preserving authentication scheme for intelligent transportation systems. Computers & Security, 60(??):193– 205, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300463.

Siadati:2017:MYS

[SNG^{+}17] Hossein Siadati, Toan Nguyen, Payas Gupta, Markus Jakobsson, and Nasir Memon. Mind your SMSes: Mitigating social engineering in second factor authentication. Computers & Security, 65(??):14–28, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481630116X.

Solic:2015:ISS

[SOG15] Kresimir Solic, Hrvoje Ocevcic, and Marin Golub. The information systems’ security level assessment model based on an ontology and evidential reasoning approach. Computers & Security, 55(??):100–112, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001212.

Serwadda:2015:WMD

[SP15] Abdul Serwadda and Vir V. Phoha. When mice devour the elephants: a DDoS attack against size-based scheduling schemes in the Internet. Computers & Security, 53(??):31–43, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000644.

Spafford:2010:Ea

[Spa10a] Eugene H. Spafford. Editorial. Computers & Security, 29(6):641–642, September 2010. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000519.

Spafford:2010:Eb

[Spa10b] Eugene H. Spafford. Editorial. Computers & Security, 29(7):729–730, October 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000714.

Spafford:2010:STP

[Spa10c] Eugene H. Spafford. Security, technology, publishing, and ethics (part i). Computers & Security, 29(8):813–814, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000878.

Spafford:2011:Ea

[Spa11a] Eugene H. Spafford. Editorial. Computers & Security, 30(1): 1, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481000115X.

Spafford:2011:Eb

[Spa11b] Eugene H. Spafford. Editorial. Computers & Security, 30(4): 171, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000496.

Spafford:2011:Ec

[Spa11c] Eugene H. Spafford. Editorial. Computers & Security, 30(6–7):351–352, September/October 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000733.

Spafford:2011:Ef

[Spa11d] Eugene H. Spafford. Editorial for 30/8. Computers & Security, 30(8):553–554, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001209.

Spafford:2011:STP

[Spa11e] Eugene H. Spafford. Security, technology, publishing, and ethics (part II). Computers & Security, 30(1):2–3, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404810001161.

Spafford:2012:Ea

[Spa12a] Eugene H. Spafford. Editorial. Computers & Security, 31(1):1–2, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001520.

Spafford:2012:Eb

[Spa12b] Eugene H. Spafford. Editorial. Computers & Security, 31(2):149–150, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000168.

Spafford:2012:Ec

[Spa12c] Eugene H. Spafford. Editorial. Computers & Security, 31(5):641–642, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000934.

Spafford:2012:Ed

[Spa12d] Eugene H. Spafford. Editorial. Computers & Security, 31(6):739–740, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001149.

Spafford:2012:Ee

[Spa12e] Gene Spafford. Editorial. Computers & Security, 31(8): 857–858, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404812001691.

Spafford:2013:Ea

[Spa13a] Eugene H. Spafford. Editorial. Computers & Security, 32(??):v, February 2013. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000035.

Spafford:2013:Eb

[Spa13b] Eugene H. Spafford. Editorial. Computers & Security, 34(??): iv, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813000473.

Spafford:2014:E

[Spa14] Eugene H. Spafford. Editorial. Computers & Security, 43(??): iv, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000704.

Spafford:2019:E

[Spa19] Eugene H. Spafford. Editorial. Computers & Security, 81(??):182–183, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818314548.

Signes-Pont:2018:MMP

[SPCCMMS18] María Teresa Signes-Pont, Antonio Cortés-Castillo, Higinio Mora-Mora, and Julian Szymanski. Modelling the malware propagation in mobile computer devices. Computers & Security, 79(??):80–93, November 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302773.

Stevanovic:2015:GTP

[SPD^{+}15] Matija Stevanovic, Jens Myrup Pedersen, Alessandro D’Alconzo, Stefan Ruehrup, and Andreas Berger. On the ground truth problem of malicious DNS traffic analysis. Computers & Security, 55(??):142–158, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500125X.

Sauerwein:2019:ACP

[SPFB19] Clemens Sauerwein, Irdin Pekaric, Michael Felderer, and Ruth Breu. An analysis and classification of public information security data sources used in research and practice. Computers & Security, 82(??):140–155, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304978.

Skracic:2017:AAU

[SPK17] Kristian Skracić, Predrag Pale, and Zvonko Kostanjcar. Authentication approach using one-time challenge generation based on user behavior patterns captured in transactional data sets. Computers & Security, 67(??):107–121, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730055X.

Stiborek:2018:PAD

[SPR18] Jan Stiborek, Tomás Pevný, and Martin Rehák. Probabilistic analysis of dynamic malware traces. Computers & Security, 74(??):221–239, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300336.

Shen:2015:IEM

[SPYG15] Chao Shen, Shichao Pei, Zhenyu Yang, and Xiaohong Guan. Input extraction via motion-sensor behavior analysis on smartphones. Computers & Security, 53(??):143–155, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000991.

Salim:2013:BAR

[SRDD13] Farzad Salim, Jason Reid, Uwe Dulleck, and Ed Dawson. Budget-aware role based access control. Computers & Security, 35(??):37–50, June 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481200171X.

Shams:2018:TAS

[SRU18] Erfan A. Shams, Ahmet Rizaner, and Ali Hakan Ulusoy. Trust aware support vector machine intrusion detection and prevention system in vehicular ad hoc networks. Computers & Security, 78(??):245–254, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307569.

Sadan:2011:WUS

[SS11] Zac Sadan and David Schwartz. WhiteScript: Using social network analysis parameters to balance between browser usability and malware exposure. Computers & Security, 30(1):4–12, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000842.

Stanekova:2013:ADM

[SS13] L’ubica Staneková and Martin Stanek. Analysis of dictionary methods for PIN selection. Computers & Security, 39 (part B)(??):289–298, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813001235.

Sadhya:2016:PPS

[SS16] Debanjan Sadhya and Sanjay Kumar Singh. Privacy preservation for soft biometrics based multimodal recognition system. Computers & Security, 58(??):160–179, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816000055.

Sadhya:2017:PRS

[SS17a] Debanjan Sadhya and Sanjay Kumar Singh. Providing robust security measures to Bloom filter based biometric template protection schemes. Computers & Security, 67(??):59– 72, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817300391.

Singh:2017:FIW

[SS17b] Bhupendra Singh and Upasna Singh. A forensic insight into Windows 10 cortana search. Computers & Security, 66(??):142–154, May 2017. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300111.

SanMiguelCarrasco:2018:UID

[SS18a] Rafael San Miguel Carrasco and Miguel-Angel Sicilia. Unsupervised intrusion detection through skip-gram models of network behavior. Computers & Security, 78(??):187–197, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818302700.

Singh:2018:PEA

[SS18b] Bhupendra Singh and Upasna Singh. Program execution analysis in Windows: a study of data sources, their format and comparison of forensic capability. Computers & Security, 74(??):94–114, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300257.

Shameli-Sendi:2016:TIS

[SSABC16] Alireza Shameli-Sendi, Rouzbeh Aghababaei-Barzegar, and Mohamed Cheriet. Taxonomy of information security risk assessment (ISRA). Computers & Security, 57(??):14–30, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815001650.

Shehab:2012:ACO

[SSAK12] Mohamed Shehab, Anna Squicciarini, Gail-Joon Ahn, and Irini Kokkinou. Access control for online social networks third party applications. Computers & Security, 31(8):897–911, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001186.

Shameli-Sendi:2014:TIR

[SSCHL14] Alireza Shameli-Sendi, Mohamed Cheriet, and Abdelwahab Hamou-Lhadj. Taxonomy of intrusion risk assessment and response system. Computers & Security, 45(??):1–16, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000613.

Skopik:2016:PSP

[SSF16] Florian Skopik, Giuseppe Settanni, and Roman Fiedler. A problem shared is a problem halved: a survey on the dimensions of collective cyber defense through security information sharing. Computers & Security, 60(??):154–176, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300347.

Singh:2016:SRI

[SSK16] Karanpreet Singh, Paramvir Singh, and Krishan Kumar. A systematic review of IP traceback schemes for denial of service attacks. Computers & Security, 56(??):111–139, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000930.

Singh:2017:ALH

[SSK17] Karanpreet Singh, Paramvir Singh, and Krishan Kumar. Application layer HTTP-GET flood DDoS attacks: Research landscape and challenges. Computers & Security, 65(??):344–372, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301365.

Singh:2019:ICD

[SSK19] Manmeet Singh, Maninder Singh, and Sanmeet Kaur. Issues and challenges in DNS based botnet detection: a survey. Computers & Security, 86(??):28–52, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301117.

Shahandashti:2015:RUP

[SSNS15] Siamak F. Shahandashti, Reihaneh Safavi-Naini, and Nashad Ahmed Safa. Reconciling user privacy and implicit authentication for mobile devices. Computers & Security, 53(??):215–233, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000838.

Safa:2015:ISC

[SSS^{+}15] Nader Sohrabi Safa, Mehdi Sookhak, Rossouw Von Solms, Steven Furnell, Norjihan Abdul Ghani, and Tutut Herawan. Information security conscious care behaviour formation in organizations. Computers & Security, 53(??):65–78, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000863.

Sohal:2018:CFI

[SSSC18] Amandeep Singh Sohal, Rajinder Sandhu, Sandeep K. Sood, and Victor Chang. A cybersecurity framework to identify malicious edge device in fog computing and cloud-of-things environments. Computers & Security, 74(??):340–354, May 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301827.

Shiravi:2012:TDS

[SSTG12] Ali Shiravi, Hadi Shiravi, Mahbod Tavallaee, and Ali A. Ghorbani. Toward developing a systematic approach to generate benchmark datasets for intrusion detection. Computers & Security, 31(3):357–374, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001672.

Sharma:2013:AAM

[SSVA13] Manisha Sharma, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. AMTRAC: an administrative model for temporal role-based access control. Computers & Security, 39 (part B)(??):201–218, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000965.

Singh:2019:MAB

[SSVA19] Mahendra Pratap Singh, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Managing attribute-based access control policies in a unified framework using data warehousing and in-memory database. Computers & Security, 86(??):183–205, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819301166.

Stefan:2012:RKD

[SSY12] Deian Stefan, Xiaokui Shu, and Danfeng (Daphne) Yao. Robustness of keystroke-dynamics based biometrics against synthetic forgeries. Computers & Security, 31(1):109–121, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811001179.

Senavirathne:2019:IPM

[ST19] Navoda Senavirathne and Vicenç Torra. Integrally private model selection for decision trees. Computers & Security, 83(??):167–181, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302268.

Sayed:2018:ITI

[STA18] Bassam Sayed, Issa Traoré, and Amany Abdelhalim. Iftranspiler: Inlining of hybrid flow-sensitive security monitor for JavaScript. Computers & Security, 75(??):92–117, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818300397.

Shabtai:2014:MMD

[STCM^{+}14] A. Shabtai, L. Tenenboim-Chekina, D. Mimran, L. Rokach, B. Shapira, and Y. Elovici. Mobile malware detection through analysis of deviations in application network behavior. Computers & Security, 43(??):1–18, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000285.

Saidi:2017:AAC

[STSB17] Firas Saidi, Zouheir Trabelsi, Khaled Salah, and Henda Ben Ghezala. Approaches to analyze cyber terrorist communities: Survey and challenges. Computers & Security, 66(??):66–80, May 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817300068.

Su:2010:WWA

[Su10] Ming-Yang Su. WARP: a wormhole-avoidance routing protocol by anomaly detection in mobile ad hoc networks. Computers & Security, 29(2):208–224, March 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001072.

Sultan:2019:PPM

[Sul19] Sari Sultan. Privacy-preserving metering in smart grid for billing, operational metering, and incentive-based schemes: a survey. Computers & Security, 84(??):148–165, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303675.

Shin:2010:AMD

[SVA10] Heechang Shin, Jaideep Vaidya, and Vijayalakshmi Atluri. Anonymization models for directional location based service environments. Computers & Security, 29(1):59–73, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404809000807.

Safa:2016:ISP

[SVF16] Nader Sohrabi Safa, Rossouw Von Solms, and Steven Furnell. Information security policy compliance model in organizations. Computers & Security, 56(??):70–82, February 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001583.

Shmueli:2014:IDE

[SVGE14] Erez Shmueli, Ronen Vaisenberg, Ehud Gudes, and Yuval Elovici. Implementing a database encryption solution, design and implementation issues. Computers & Security, 44(??):33–50, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000509.

Sharma:2019:DRB

[SW19] Shwadhin Sharma and Merrill Warkentin. Do I really belong?: Impact of employment status on information security policy compliance. Computers & Security, 87 (??):Article 101397, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304024.

Sun:2011:IPT

[SWLZ11] Xiaoxun Sun, Hua Wang, Jiuyong Li, and Yanchun Zhang. Injecting purpose and trust into data anonymisation. Computers & Security, 30(5):332–345, July 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000666.

Shropshire:2015:PAI

[SWS15] Jordan Shropshire, Merrill Warkentin, and Shwadhin Sharma. Personality, attitudes, and intentions: Predicting initial adoption of information security behavior. Computers & Security, 49(??):177–191, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000036.

Schryen:2016:DTN

[SWS16] Guido Schryen, Gerit Wagner, and Alexander Schlegel. Development of two novel face-recognition CAPTCHAs: a security and usability study. Computers & Security, 60(??):95– 116, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300293.

Shahriar:2014:EDV

[SWZL14] Hossain Shahriar, Komminist Weldemariam, Mohammad Zulkernine, and Thibaud Lutellier. Effective detection of vulnerable and malicious browser extensions. Computers & Security, 47(??):66–84, November 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000984.

Shih:2011:SMI

[SYCS11] Dong-Her Shih, David C. Yen, Chih-Hung Cheng, and Ming-Hung Shih. A secure multi-item e-auction mechanism with bid privacy. Computers & Security, 30(4):273–287, June 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811000460.

Stoller:2011:SRA

[SYGR11] Scott D. Stoller, Ping Yang, Mikhail I. Gofman, and C. R. Ramakrishnan. Symbolic reachability analysis for parameterized administrative role-based access control. Computers & Security, 30(2–3):148–164, March 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000787.

Shen:2014:AAG

[SYK14] Yue Shen, Zheng Yan, and Raimo Kantola. Analysis on the acceptance of global trust management for unwanted traffic control based on game theory. Computers & Security, 47(??):3–25, November 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000492.

Shuai:2019:AAS

[SYWX19] Mengxia Shuai, Nenghai Yu, Hongxia Wang, and Ling Xiong. Anonymous authentication scheme for smart home environment with provable security. Computers & Security, 86(??):132–146, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818313701.

Shen:2016:UPP

[SYX^{+}16] Chao Shen, Tianwen Yu, Haodi Xu, Gengshan Yang, and Xiaohong Guan. User practice in password security: an empirical study of real-life passwords in the wild. Computers & Security, 61(??):130–141, August 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300657.

Tudor:2018:IDC

[TAP18] Valentin Tudor, Magnus Almgren, and Marina Papatriantafilou. The influence of dataset characteristics on privacy preserving methods in the advanced metering infrastructure. Computers & Security, 76(??):178–196, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481830141X.

Tauber:2011:SCM

[Tau11] Arne Tauber. A survey of certified mail systems provided on the Internet. Computers & Security, 30(6–7): 464–485, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000629.

Thalmann:2014:CDL

[TBDM14] Stefan Thalmann, Daniel Bachlechner, Lukas Demetz, and Markus Manhart. Complexity is dead, long live complexity! How software can help service providers manage security and compliance. Computers & Security, 45(??):172–185, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000935.

Tamjidyamcholo:2014:EMK

[TBSR14] Alireza Tamjidyamcholo, Mohd Sapiyan Bin Baba, Nor Liyana Mohd Shuib, and Vala Ali Rohani. Evaluation model for knowledge sharing in information security professional virtual community. Computers & Security, 43(??):19–34, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814000297.

Tapiador:2011:MMA

[TC11] Juan E. Tapiador and John A. Clark. Masquerade mimicry attack detection: a randomised approach. Computers & Security, 30(5):297–310, July 2011. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000654.

Tran:2016:CRR

[TCNFW16] Hiep Tran, Enrique Campos-Nanez, Pavel Fomin, and James Wasek. Cyber resilience recovery model to combat zero-day malware attacks. Computers & Security, 61(??):19–31, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300505.

Turkmen:2017:FAX

[TdHRZ17] Fatih Turkmen, Jerry den Hartog, Silvio Ranise, and Nicola Zannone. Formal analysis of XACML policies using SMT. Computers & Security, 66(??):185–203, May 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300135.

Tripathy:2018:RBS

[TDJB18] Bata Krishna Tripathy, Debi Prasad Das, Swagat Kumar Jena, and Padmalochan Bera. Risk based security enforcement in software defined network. Computers & Security, 78(??):321–335, September 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301913.

Toland:2010:IPM

[TFE10] Tyrone S. Toland, Csilla Farkas, and Caroline M. Eastman. The inference problem: Maintaining maximal availability in the presence of database updates. Computers & Security, 29(1):88–103, February 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000789.

Tjhai:2010:PTS

[TFPC10] Gina C. Tjhai, Steven M. Furnell, Maria Papadaki, and Nathan L. Clarke. A preliminary two-stage alarm correlation and filtering system using SOM neural network and K-means algorithm. Computers & Security, 29(6):712–723, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000192.

Tsoumas:2012:ICW

[TG12] Bill Tsoumas and Dimitris Gritzalis. Inside cyber warfare: Mapping the cyber underworld. Computers & Security, 31(6):801, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000831.

Tsalis:2014:SCM

[TG14] Nikolaos Tsalis and Dimitris Gritzalis. Securing cloud and mobility: a practitioner’s guide. Computers & Security, 45(??):305, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000212.

Tsalis:2015:BRH

[TG15] Nikolaos Tsalis and Dimitris Gritzalis. Book review: Hacking Web Intelligence: Open Source Intelligence and Web Reconnaissance Concepts and Techniques, Sudhanshu Chauhan, Nutan Kumar Panda, Elsevier Publications, USA (2015). Computers & Security, 55(??):113, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001236.

Tripathi:2018:SRD

[TH18] Nikhil Tripathi and Neminath Hubballi. Slow rate denial of service attacks against HTTP/2 and detection. Computers & Security, 72(??):255–272, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301980.

Toledo:2013:DFS

[THA^{+}13] Nerea Toledo, Marivi Higuero, Jasone Astorga, Marina Aguado, and Jean Marie Bonnin. Design and formal security evaluation of NeMHIP: a new secure and efficient network mobility management protocol based on the host identity protocol. Computers & Security, 32(??):1–18, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001599.

Theoharidou:2010:MLC

[TKG10] Marianthi Theoharidou, Panayiotis Kotzanikolaou, and Dimitris Gritzalis. A multi-layer criticality assessment methodology based on interdependencies. Computers & Security, 29(6):643–658, September 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000210.

Tsohou:2015:ARC

[TKK15] Aggeliki Tsohou, Maria Karyda, and Spyros Kokolakis. Analyzing the role of cognitive and cultural biases in the internalization of information security policies: Recommendations for information security awareness programs. Computers & Security, 52(??):128–141, July 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000565.

Tang:2015:STC

[TL15] Changlong Tang and Jiqiang Liu. Selecting a trusted cloud service provider for your SaaS program. Computers & Security, 50(??):60–73, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000139.

Tondel:2014:ISI

[TLJ14] Inger Anne Tøndel, Maria B. Line, and Martin Gilje Jaatun. Information security incident management: Current practice as reported in the literature. Computers & Security, 45(??):42–57, September 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000819.

Tang:2016:RIH

[TLZL16] Zhenjun Tang, Huan Lao, Xianquan Zhang, and Kai Liu. Robust image hashing via DCT and LLE. Computers & Security, 62(??):133–148, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300852.

Tsalis:2017:EPP

[TMN^{+}17] Nikolaos Tsalis, Alexios Mylonas, Antonia Nisioti, Dimitris Gritzalis, and Vasilios Katos. Exploring the protection of private browsing in desktop browsers. Computers & Security, 67(??):181–197, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300597.

Tormo:2015:TPP

[TMP15] Ginés Dólera Tormo, Félix Gómez Mármol, and Gregorio Martínez Pérez. Towards privacy-preserving reputation management for hybrid broadcast broadband applications. Computers & Security, 49(??):220–238, March 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001564.

Thompson:2017:SBH

[TMW17] Nik Thompson, Tanya Jane McGill, and Xuequn Wang. “Security begins at home”: Determinants of home computer and mobile device security behavior. Computers & Security, 70(??):376–391, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301426.

Tan:2013:CRM

[TOZJ13] Hailun Tan, Diethelm Ostry, John Zic, and Sanjay Jha. A confidential and DoS-resistant multi-hop code dissemination protocol for wireless sensor networks. Computers & Security, 32(??):36–55, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001484.

Tiruvakadu:2018:CWA

[TP18] Divya Sai Keerthi Tiruvakadu and Venkataram Pallapa. Confirmation of wormhole attack in MANETs using honeypot. Computers & Security, 76(??):32–49, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300865.

Tounsi:2018:STT

[TR18] Wiem Tounsi and Helmi Rais. A survey on technical threat intelligence in the age of sophisticated cyber attacks. Computers & Security, 72(??):212–233, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301839.

Torten:2018:ISA

[TRB18] Ron Torten, Carmen Reaiche, and Stephen Boyle. The impact of security awareness on information technology professionals’ behavior. Computers & Security, 79(??):68–79, November 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818304656.

Tsigkritis:2013:AGE

[TS13] Theocharis Tsigkritis and George Spanoudakis. Assessing the genuineness of events in runtime monitoring of cyber systems. Computers & Security, 38(??):76–96, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000564.

Talbot:2018:DRA

[TTCB18] Christopher M. Talbot, Michael A. Temple, Timothy J. Carbino, and J. Addison Betances. Detecting rogue attacks on commercial wireless insteon home automation systems. Computers & Security, 74(??):296–307, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302055.

Tan:2019:MMT

[TTJ19] Hailun Tan, Gene Tsudik, and Sanjay Jha. MTRA: Multi-tier randomized remote attestation in IoT networks. Computers & Security, 81(??):78–93, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307697.

Tas:2016:NSI

[TUB16] Ismail Melih Tas, Bahar Ugurdogan, and Selcuk Baktir. Novel session initiation protocol-based distributed denialof-service attacks and effective defense strategies. Computers & Security, 63(??):29–44, November 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300980.

Tsakalidis:2019:CIA

[TVPV19] George Tsakalidis, Kostas Vergidis, Sophia Petridou, and Maro Vlachopoulou. A cybercrime incident architecture with adaptive response policy. Computers & Security, 83(??):22– 37, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818308150.

Toor:2018:VQA

[TWNC18] Andeep S. Toor, Harry Wechsler, Michele Nappi, and Kim-Kwang Raymond Choo. Visual question authentication protocol (VQAP). Computers & Security, 76(??):285–294, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817302547.

Tang:2019:SAA

[TXM^{+}19] Zhushou Tang, Minhui Xue, Guozhu Meng, Chengguo Ying, Yugeng Liu, Jianan He, Haojin Zhu, and Yang Liu. Securing Android applications via edge assistant third-party library detection. Computers & Security, 80(??):257–272, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818311301.

Trabelsi:2015:SDS

[TZMG15] Zouheir Trabelsi, Safaa Zeidan, Mohammad M. Masud, and Kilani Ghoudi. Statistical dynamic splay tree filters towards multilevel firewall packet filtering enhancement. Computers & Security, 53(??):109–131, September 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481500084X.

Teh:2016:STD

[TZTC16] Pin Shen Teh, Ning Zhang, Andrew Beng Jin Teoh, and Ke Chen. A survey on touch dynamics authentication in mobile devices. Computers & Security, 59(??):210–235, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300256.

Trabelsi:2013:DTA

[TZZG13] Zouheir Trabelsi, Liren Zhang, Safaa Zeidan, and Kilani Ghoudi. Dynamic traffic awareness statistical model for firewall performance enhancement. Computers & Security, 39 (part B)(??):160–172, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000928.

Ucci:2019:SML

[UAB19] Daniele Ucci, Leonardo Aniello, and Roberto Baldoni. Survey of machine learning techniques for malware analysis. Computers & Security, 81(??):123–147, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303808.

Ulas:2016:ARL

[UAK16] Cihan Ulas, Ulas Asik, and Cantürk Karadeniz. Analysis and reconstruction of laser printer information leakages in the media of electromagnetic radiation, power, and signal lines. Computers & Security, 58(??):250–267, May 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300049.

Uzun:2012:SLP

[UB12] Yasin Uzun and Kemal Bicakci. A second look at the performance of neural networks for keystroke dynamics using a publicly available dataset. Computers & Security, 31(5):717– 726, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404812000636.

Haq:2018:MLW

[UCCJ18] Irfan Ul Haq, Sergio Chica, Juan Caballero, and Somesh Jha. Malware lineage in the wild. Computers & Security, 78(??):347–363, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818308605.

Uzunov:2012:SDS

[UFF12] Anton V. Uzunov, Eduardo B. Fernandez, and Katrina Falkner. Securing distributed systems using patterns: a survey. Computers & Security, 31(5):681–703, July 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000661.

Uzunov:2015:SSF

[UFF15] Anton V. Uzunov, Eduardo B. Fernandez, and Katrina Falkner. Security solution frames and security patterns for authorization in distributed, collaborative systems. Computers & Security, 55(??):193–234, November 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815001200.

Ultra:2017:SMS

[UPF17] J. D. Ultra and S. Pancho-Festin. A simple model of separation of duty for access control models. Computers & Security, 68(??):69–80, July 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300652.

Ugarte-Pedrero:2014:AAD

[UPSGF^{+}14] Xabier Ugarte-Pedrero, Igor Santos, Iván García-Ferreira, Sergio Huerta, Borja Sanz, and Pablo G. Bringas. On the adoption of anomaly detection for packed executable filtering. Computers & Security, 43(??):126–144, June 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000522.

Umer:2017:FBI

[USB17] Muhammad Fahad Umer, Muhammad Sher, and Yaxin Bi. Flow-based intrusion detection: Techniques and challenges. Computers & Security, 70(??):238–254, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301165.

Uzunov:2016:SSS

[Uzu16] Anton V. Uzunov. A survey of security solutions for distributed publish/subscribe systems. Computers & Security, 61(??):94–129, August 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816300487.

Vigil:2015:IAN

[VBC^{+}15] Martín Vigil, Johannes Buchmann, Daniel Cabarcas, Christian Weinert, and Alexander Wiesmaier. Integrity, authenticity, non-repudiation, and proof of existence for long-term archiving: a survey. Computers & Security, 50(??):16–32, May 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001849.

VanGundy:2012:NUR

[VC12] Matthew Van Gundy and Hao Chen. Noncespaces: Using randomization to defeat cross-site scripting attacks. Computers & Security, 31(4):612–628, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001477.

vanDeursen:2013:MIS

[vDBD13] Nicole van Deursen, William J. Buchanan, and Alistair Duff. Monitoring information security risks within health care. Computers & Security, 37(??):31–45, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000813.

Vijayanand:2018:IDS

[VDK18] R. Vijayanand, D. Devaraj, and B. Kannapiran. Intrusion detection system for wireless mesh network using multiple support vector machine classifiers with genetic-algorithm-based feature selection. Computers & Security, 77(??):304–314, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818303766.

vanderWalt:2018:CSI

[vdWEG18] Estee van der Walt, J. H. P. Eloff, and Jacomine Grobler. Cyber-security: Identity deception detection on social media platforms. Computers & Security, 78(??):76–89, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818306503.

Venugopal:2018:IDI

[Ven18] A. Vijayalakshmi Venugopal. Inadvertent disclosures of inventions in social media affecting patent rights. Computers & Security, 72(??):136–144, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301967.

Vavilis:2015:AAF

[VEPZ15] Sokratis Vavilis, Alexandru Egner, Milan Petković, and Nicola Zannone. An anomaly analysis framework for database systems. Computers & Security, 53(??):156–173, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000905.

Verkijika:2018:USS

[Ver18] Silas Formunyuy Verkijika. Understanding smartphone security behaviors: an extension of the protection motivation theory with anticipated regret. Computers & Security, 77(??):860–870, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818302505.

Virvilis:2014:ADA

[VG14] Nikos Virvilis and Dimitris Gritzalis. Automatic defense against zero-day polymorphic worms in communication networks. Computers & Security, 42(??):191–192, May 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000182.

Vrakas:2014:OUI

[VGL14] Nikos Vrakas, Dimitris Geneiatakis, and Costas Lambrinoudakis. Obscuring users’ identity in VoIP/IMS environments. Computers & Security, 43(??):145–158, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814000510.

Vitunskaite:2019:SCC

[VHBJ19] Morta Vitunskaite, Ying He, Thomas Brandstetter, and Helge Janicke. Smart cities and cyber security: Are we there yet?a comparative study on the role of standards, third party risk management and security ownership. Computers & Security, 83(??):313–331, June 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818310423.

Varshney:2016:PDU

[VMA16] Gaurav Varshney, Manoj Misra, and Pradeep K. Atrey. A phish detector using lightweight search features. Computers & Security, 62(??):213–228, September 2016. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816300918.

Vennila:2018:DVS

[VMS18] Ganesan Vennila, M. S. K. Manikandan, and M. N. Suresh. Dynamic voice spammers detection using hidden Markov model for voice over Internet protocol network. Computers & Security, 73(??):1–16, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302080.

Virvilis:2015:SBW

[VMTG15] Nikos Virvilis, Alexios Mylonas, Nikolaos Tsalis, and Dimitris Gritzalis. Security busters: Web browser security vs. rogue sites. Computers & Security, 52(??):90–105, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000590.

Varga:2018:DSA

[VPNP18] Mihai Varga, Alina Petrescu-Nita, and Florin Pop. Deadline scheduling algorithm for sustainable computing in Hadoop environment. Computers & Security, 76(??):354–366, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818300038.

Volkamer:2017:UET

[VRRK17] Melanie Volkamer, Karen Renaud, Benjamin Reinheimer, and Alexandra Kunz. User experiences of TORPEDO: TOoltip-poweRed phishing email DetectiOn. Computers & Security, 71(??):100–113, November 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300275.

Voris:2019:AAU

[VSB^{+}19] Jonathan Voris, Yingbo Song, Malek Ben Salem, Shlomo Hershkop, and Salvatore Stolfo. Active authentication using file system decoys and user behavior modeling: results of a large scale study. Computers & Security, 87 (??):Article 101412, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818311258.

Viegas:2018:REE

[VSO^{+}18] Eduardo Viegas, Altair Santin, Luiz Oliveira, André França, Ricardo Jasinski, and Volnei Pedroni. A reliable and energyefficient classifier combination scheme for intrusion detection in embedded systems. Computers & Security, 78(??):16–32, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818306175.

vonSolms:2013:ISC

[vSvN13] Rossouw von Solms and Johan van Niekerk. From information security to cyber security. Computers & Security, 38(??):97–102, October 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000801.

VanNiekerk:2010:ISC

[VV10] J. F. Van Niekerk and R. Von Solms. Information security culture: a management perspective. Computers & Security, 29(4):476–486, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001126.

Vorakulpipat:2012:PSR

[VVS12] Chalee Vorakulpipat, Vasaka Visoottiviseth, and Siwaruk Siwamogsatham. Polite sender: a resource-saving spam email countermeasure based on sender responsibilities and recipient justifications. Computers & Security, 31(3):286–298, May 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000119.

Wijayarathna:2019:WJC

[WA19] Chamila Wijayarathna and Nalin Asanka Gamagedara Arachchilage. Why Johnny can’t develop a secure application? A usability analysis of Java Secure Socket Extension API. Computers & Security, 80(??):54–73, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304887.

Webb:2014:SAM

[WAMS14] Jeb Webb, Atif Ahmad, Sean B. Maynard, and Graeme Shanks. A situation awareness model for information security risk management. Computers & Security, 44(??):1–15, July 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814000571.

West:2016:IFF

[WB16] Jarrod West and Maumita Bhattacharya. Intelligent financial fraud detection: a comprehensive review. Computers & Security, 57(??):47–66, March 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001261.

Wu:2018:CPI

[WCDY18] Zezhi Wu, Xingyuan Chen, Xuehui Du, and Zhi Yang. CDroid: practically implementation a formal-analyzed CIFC model on Android. Computers & Security, 78(??):231–244, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818307776.

Wedge:2012:BDF

[Wed12] Tim Wedge. The basics of digital forensics. Computers & Security, 31(6):800, September 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812000983.

Wessel:2015:IMD

[WHSE15] Sascha Wessel, Manuel Huber, Frederic Stumpf, and Claudia Eckert. Improving mobile device security with operating system-level virtualization. Computers & Security, 52(??):207–220, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000206.

Wei:2018:NBM

[WHZ^{+}18] Hang Wei, Guan-Yu Hu, Zhi-Jie Zhou, Pei-Li Qiao, Zhi-Guo Zhou, and You-Min Zhang. A new BRB model for security-state assessment of cloud computing based on the impact of external and internal environments. Computers & Security, 73(??):207–218, March 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302341.

Wang:2019:RSI

[WHZ^{+}19] Ping Wang, Xing He, Yushu Zhang, Wenying Wen, and Ming Li. A robust and secure image sharing scheme with personal identity information embedded. Computers & Security, 85(??):107–121, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819300914.

Wang:2015:CDF

[WJ15] Ting Wang and Dongyao Ji. Comment on “Design and formal security evaluation of NeMHIP”. Computers & Security, 52(??):159–161, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813000345.

Woo:2011:MVD

[WJAM11] Sung-Whan Woo, HyunChul Joh, Omar H. Alhazmi, and Yashwant K. Malaiya. Modeling vulnerability discovery process in Apache and IIS HTTP servers. Computers & Security, 30(1):50–62, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404810000908.

Jang:2016:ADA

[wJKW^{+}16] Jae wook Jang, Hyunjae Kang, Jiyoung Woo, Aziz Mohaisen, and Huy Kang Kim. Andro-Dumpsys: Anti-malware system based on the similarity of malware creator and malware centric information. Computers & Security, 58(??):125–138, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481600002X.

Weippl:2012:ESE

[WL12] Edgar Weippl and Manuel Leithner. Enterprise security for the executive. Computers & Security, 31(4):636, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481200051X.

Wang:2018:GBM

[WL18] Li-E. Wang and Xianxian Li. A graph-based multifold model for anonymizing data with attributes of multiple types. Computers & Security, 72(??):122–135, January 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481730192X.

Wang:2017:DCD

[WLCC17] Tzy-Shiah Wang, Hui-Tang Lin, Wei-Tsung Cheng, and Chang-Yu Chen. DBod: Clustering and detecting DGAbased botnets using DNS traffic analysis. Computers & Security, 64(??):1–15, January 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301250.

Wang:2019:TPA

[WLF^{+}19] Jinyan Wang, Chen Liu, Xingcheng Fu, Xudong Luo, and Xianxian Li. A three-phase approach to differentially private crucial patterns mining over data streams. Computers & Security, 82(??):30–48, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307417.

Wang:2019:AEI

[WLW^{+}19] Ziyu Wang, Jianwei Liu, Qianhong Wu, Yanting Zhang, Hui Yu, and Ziyu Zhou. An analytic evaluation for the impact of uncle blocks by selfish and stubborn mining in an imperfect Ethereum network. Computers & Security, 87(??):Article 101581, November 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818314172.

Wang:2011:HAB

[WLWG11] Guojun Wang, Qin Liu, Jie Wu, and Minyi Guo. Hierarchical attribute-based encryption and scalable user revocation for sharing data in cloud servers. Computers & Security, 30(5):320–331, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000678.

Wang:2019:DAT

[WLZ^{+}19] Ye Wang, Qingbao Li, Ping Zhang, Zhifeng Chen, and Guimin Zhang. DOPdefender: an approach to thwarting data-oriented programming attacks based on a data-aware automaton. Computers & Security, 81(??):94–106, March 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481830659X.

Wagner:2019:CTI

[WMPA19] Thomas D. Wagner, Khaled Mahbub, Esther Palomar, and Ali E. Abdallah. Cyber threat intelligence sharing: Survey and research directions. Computers & Security, 87 (??):Article 101589, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481830467X.

Wang:2015:SDB

[WN15] Yongge Wang and Tony Nicol. On statistical distance based testing of pseudo random sequences and experiments with PHP and Debian OpenSSL. Computers & Security, 53(??):44–64, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000693.

Woo:2019:DEC

[Woo19] Simon S. Woo. Design and evaluation of 3D CAPTCHAs. Computers & Security, 82(??):49–67, May 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301238.

Workman:2010:BPC

[Wor10] Michael Workman. A behaviorist perspective on corporate harassment online: Validation of a theoretical model of psychological motives. Computers & Security, 29(8):831–839, November 2010. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404810000829.

Wei:2010:MPT

[WP10] Jinpeng Wei and Calton Pu. Modeling and preventing TOCTTOU vulnerabilities in Unix-style file systems. Computers & Security, 29(8):815–830, November 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404810000830.

Wei:2012:TGD

[WP12] Jinpeng Wei and Calton Pu. Toward a general defense against kernel queue hooking attacks. Computers & Security, 31(2):176–191, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811001507.

Wagner:2017:KAV

[WRTA17] Markus Wagner, Alexander Rind, Niklas Thür, and Wolfgang Aigner. A knowledge-assisted visual malware analysis system: Design, validation, and reflection of KA-MAS. Computers & Security, 67(??):1–15, June 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300263.

Wang:2017:PPT

[WS17] Shuo Wang and Richard O. Sinnott. Protecting personal trajectories of social media users through differential privacy. Computers & Security, 67(??):142–163, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817300251.

Wang:2015:TPC

[WW15] Yongzhi Wang and Jinpeng Wei. Toward protecting control flow confidentiality in cloud-based computation. Computers & Security, 52(??):106–127, July 2015. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000553.

Wu:2018:MSF

[WW18] Hao Wu and Zhonghua Wang. Multi-source fusion-based security detection method for heterogeneous networks. Computers & Security, 74(??):55–70, May 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300063.

Wu:2018:TSD

[WWXZ18] Tingmin Wu, Sheng Wen, Yang Xiang, and Wanlei Zhou. Twitter spam detection: Survey of new approaches and comparative study. Computers & Security, 76(??):265–284, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481730250X.

Wei:2018:AAP

[WWY^{+}18] Zhuo Wei, Yongdong Wu, Yanjiang Yang, Zheng Yan, Qingqi Pei, Yajuan Xie, and Jian Weng. AutoPrivacy: Automatic privacy protection and tagging suggestion for mobile social photo. Computers & Security, 76(??):341–353, July 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817302626.

Wang:2019:IID

[WX19] Dong Wang and Zhengquan Xu. Impact of inaccurate data on differential privacy. Computers & Security, 82(??):68– 79, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818304747.

Weishaupl:2018:ISI

[WYS18] Eva Weishäupl, Emrah Yasasin, and Guido Schryen. Information security investments: an exploratory multiple case study on decision-making, evaluation and learning. Computers & Security, 77(??):807–823, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300555.

Wang:2013:EAG

[WZK13] Shuzhen Wang, Zonghua Zhang, and Youki Kadobayashi. Exploring attack graph for cost-benefit security hardening: a probabilistic approach. Computers & Security, 32(??):158–169, February 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404812001496.

Wang:2010:SAS

[WZW^{+}10] Fangwei Wang, Yunkai Zhang, Changguang Wang, Jianfeng Ma, and SangJae Moon. Stability analysis of a SEIQV epidemic model for rapid spreading worms. Computers & Security, 29(4):410–418, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001096.

Wan:2016:PEP

[WZW16] Zhiguo Wan, Wen-Tao Zhu, and Guilin Wang. PRAC: Efficient privacy protection for vehicle-to-grid communications in the smart grid. Computers & Security, 62(??):246–256, September 2016. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404816300761.

Xun:2018:SDL

[XdZMsC18] Peng Xun, Pei dong Zhu, Sabita Maharjan, and Peng shuai Cui. Successive direct load altering attack in smart grid. Computers & Security, 77(??):79–93, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818302554.

Xiao:2010:KSM

[XHC^{+}10] Liang Xiao, Bo Hu, Madalina Croitoru, Paul Lewis, and Srinandan Dasmahapatra. A knowledgeable security model for distributed health information systems. Computers & Security, 29(3):331–349, May 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809000819.

Xiang:2011:EAN

[XHZ11] Yang Xiang, Jiankun Hu, and Wanlei Zhou. Editorial: Advances in network and system security. Computers & Security, 30(5):295–296, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000708.

Xiong:2019:TMS

[XL19] Wenjun Xiong and Robert Lagerström. Threat modeling — a systematic literature review. Computers & Security, 84(??):53–69, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818307478.

Xu:2019:LPI

[XLS^{+}19] Zhou Xu, Yaping Lin, Voundi Koe Arthur Sandor, Zhisheng Huang, and Xinbo Liu. A lightweight privacy and integrity preserving range query scheme for mobile cloud computing. Computers & Security, 84(??):318–333, July 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818309064.

Xu:2017:AEP

[XLT^{+}17] Guowen Xu, Hongwei Li, Chen Tan, Dongxiao Liu, Yuanshun Dai, and Kan Yang. Achieving efficient and privacypreserving truth discovery in crowd sensing systems. Computers & Security, 69(??):114–126, August 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301675.

Xu:2017:EOS

[XMY^{+}17] Rui Xu, Kirill Morozov, Yanjiang Yang, Jianying Zhou, and Tsuyoshi Takagi. Efficient outsourcing of secure k-nearest neighbour query over encrypted database. Computers & Security, 69(??):65–83, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301651.

Xenakis:2014:APT

[XN14] Christos Xenakis and Christoforos Ntantogian. An advanced persistent threat in 3G networks: Attacking the home network from roaming networks. Computers & Security, 40(??):84–94, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404813001685.

Xenakis:2016:USM

[XNP16] Christos Xenakis, Christoforos Ntantogian, and Orestis Panos. (U)SimMonitor: a mobile application for security evaluation of cellular networks. Computers & Security, 60(??):62–78, July 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481630027X.

Xenakis:2011:CEI

[XPS11] Christos Xenakis, Christoforos Panos, and Ioannis Stavrakakis. A comparative evaluation of intrusion detection architectures for mobile ad hoc networks. Computers & Security, 30(1):63–80, January 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481000091X.

Xu:2019:DMD

[XSD19] Congyuan Xu, Jizhong Shen, and Xin Du. Detection method of domain names generated by DGAs based on semantic representation and deep neural network. Computers & Security, 85(??):77–88, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818312938.

Xia:2018:CGB

[XXLZ18] Xiaofang Xia, Yang Xiao, Wei Liang, and Meng Zheng. Coded grouping-based inspection algorithms to detect malicious meters in neighborhood area smart grid. Computers & Security, 77(??):547–564, August 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818304929.

Xiao:2012:MFR

[XYY^{+}12] Da Xiao, Yan Yang, Wenbin Yao, Chunhua Wu, Jianyi Liu, and Yixian Yang. Multiple-file remote data checking for cloud storage. Computers & Security, 31(2):192–205, March 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404811001489.

Yadav:2019:EAM

[Yad19] Ram Mahesh Yadav. Effective analysis of malware detection in cloud computing. Computers & Security, 83(??):14–21, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818304012.

Yang:2019:DBA

[YAXY19] Rupeng Yang, Man Ho Au, Qiuliang Xu, and Zuoxia Yu. Decentralized blacklistable anonymous credentials with reputation. Computers & Security, 85(??):353–371, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818313282.

Yao:2018:KMA

[YCH18] Mei-Ling Yao, Ming-Chuen Chuang, and Chun-Cheng Hsu. The kano model analysis of features for mobile security applications. Computers & Security, 78(??):336–346, September 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818303341.

Yang:2012:RHA

[YCJ12] Gaobo Yang, Ning Chen, and Qin Jiang. A robust hashing algorithm based on SURF for video copy detection. Computers & Security, 31(1):33–39, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001404.

Yilmaz:2018:ADP

[YG18] Ercan Nurcan Yilmaz and Serkan Gönen. Attack detection/ prevention system against cyber attack in industrial control systems. Computers & Security, 77(??):94–105, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303316.

Yang:2019:RCC

[YGX^{+}19] Yu-Guang Yang, Xue-Pei Guo, Gang Xu, Xiu-Bo Chen, Jian Li, Yi-Hua Zhou, and Wei-Min Shi. Reducing the communication complexity of quantum private database queries by subtle classical post-processing with relaxed quantum ability. Computers & Security, 81(??):15–24, March 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311234.

Ye:2019:ESM

[YGZG19] Bei Ye, Yajun Guo, Lei Zhang, and Xiaowei Guo. An empirical study of mnemonic password creation tips. Computers & Security, 85(??):41–50, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819300884.

Yan:2015:LRP

[YHL^{+}15] Qiang Yan, Jin Han, Yingjiu Li, Jianying Zhou, and Robert H. Deng. Leakage-resilient password entry: Challenges, design, and evaluation. Computers & Security, 48(??):196–211, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001540.

Yamada:2017:EPA

[YI17] Asahiko Yamada and Tatsuro Ikeda. Enhanced PKI authentication with trusted product at claimant. Computers & Security, 67(??):324–334, June 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817300044.

Yoon:2010:HSF

[YKH10] Ji Won Yoon, Hyoungshick Kim, and Jun Ho Huh. Hybrid spam filtering for mobile communication. Computers & Security, 29(4):446–459, June 2010. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404809001266.

Yousaf:2017:AIT

[YLB^{+}17] Awais Yousaf, Asim Loan, Radu F. Babiceanu, Leandros Maglaras, and Onaiza Yousaf. Architectural and information theoretic perspectives of physical layer intruders for direct sequence spread spectrum systems. Computers & Security, 70(??):124–143, September 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404817301189.

Yu:2019:PQQ

[YLN^{+}19] Bin Yu, Joseph Liu, Surya Nepal, Jiangshan Yu, and Paul Rimba. Proof-of-QoS: QoS based blockchain consensus protocol. Computers & Security, 87(??):Article 101580, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818313774.

Yu:2019:ABC

[YLW^{+}19] Feng Yu, Qiang Liu, Shu Wu, Liang Wang, and Tieniu Tan. Attention-based convolutional approach for misinformation identification from massive and noisy microblog posts. Computers & Security, 83(??):106–121, June 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818303997.

Yi:2015:PW

[YRA15] Jeong Hyun Yi, Kyung-Hyune Rhee, and Gail-Joon Ahn. Preface — WISA 2014. Computers & Security, 52(??):250, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000723.

Yakura:2019:NMA

[YSN^{+}19] Hiromu Yakura, Shinnosuke Shinozaki, Reon Nishimura, Yoshihiro Oyama, and Jun Sakuma. Neural malware analysis with attention mechanism. Computers & Security, 87 (??):Article 101592, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404819300264.

Yanushkevich:2019:CCE

[YST^{+}19] Svetlana N. Yanushkevich, Kelly W. Sundberg, Nathan W. Twyman, Richard M. Guest, and Vlad P. Shmerko. Cognitive checkpoint: Emerging technologies for biometric-enabled watchlist screening. Computers & Security, 85(??):372–385, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819300951.

Tsai:2016:UOS

[ySTJA^{+}16] Hsin yi Sandy Tsai, Mengtian Jiang, Saleem Alhabash, Robert LaRose, Nora J. Rifon, and Shelia R. Cotten. Understanding online safety behaviors: a protection motivation theory perspective. Computers & Security, 59(??):138–150, June 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816300190.

Yang:2014:EDN

[YtLpZD14] Jie Yang, Hai tao Liu, Zu ping Zhang, and Jian Dong. Extended DMTP: a new protocol for improved graylist categorization. Computers & Security, 40(??):75–83, February 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404813001430.

Yang:2011:CTI

[YW11] Jianhua Yang and David Woolbright. Correlating TCP/IP packet contexts to detect stepping-stone intrusion. Computers & Security, 30(6–7):538–546, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811000721.

Yu:2017:EEG

[YWL^{+}17] Xingjie Yu, Zhan Wang, Yingjiu Li, Liang Li, Wen Tao Zhu, and Li Song. EvoPass: Evolvable graphical password against shoulder-surfing attacks. Computers & Security, 70(??):179–198, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481730113X.

Yan:2014:ESI

[YWND14] Zheng Yan, Guojun Wang, Valtteri Niemi, and Robert H. Deng. Editorial: Special issue on trust in cyber, physical and social computing. Computers & Security, 47(??):1–2, November 2014. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404814001369.

Yi:2017:EIV

[YWS17] Mingxu Yi, Jinxia Wei, and Lingwei Song. Efficient integrity verification of replicated data in cloud computing system. Computers & Security, 65(??):202–212, March 2017. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301560.

Yan:2012:SMA

[YWYZ12] Diqun Yan, Rangding Wang, Xianmin Yu, and Jie Zhu. Steganography for MP3 audio by exploiting the rule of window switching. Computers & Security, 31(5):704–716, July 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000673.

Yang:2018:SSD

[YWZM18] Li Yang, Teng Wei, Fengwei Zhang, and Jianfeng Ma. SADUS: Secure data deletion in user space for mobile devices. Computers & Security, 77(??):612–626, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818300683.

Yuxin:2011:FRS

[YXD^{+}11] Ding Yuxin, Yuan Xuebing, Zhou Di, Dong Li, and An Zhanchao. Feature representation and selection in malicious code detection methods based on static system calls. Computers & Security, 30(6–7):514–524, September/October 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481100068X.

Yang:2012:OMS

[YYJC12] Zhi Yang, Lihua Yin, Shuyuan Jin, and Xingyuan Chen. Optimal mining on security labels for decentralized information flow control. Computers & Security, 31(8):912–924, November 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812001174.

Yin:2015:DSC

[YYL^{+}15] Ting Yin, Gaobo Yang, Leida Li, Dengyong Zhang, and Xingming Sun. Detecting seam carving based image resizing using local binary patterns. Computers & Security, 55(??):130–141, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815001248.

Zhang:2017:BVP

[ZBB17] Xiaolu Zhang, Ibrahim Baggili, and Frank Breitinger. Breaking into the vault: Privacy, security and forensic analysis of Android vault applications. Computers & Security, 70(??):516–531, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817301529.

Zahadat:2015:BSE

[ZBBO15] Nima Zahadat, Paul Blessner, Timothy Blackburn, and Bill A. Olson. BYOD security engineering: a framework and its analysis. Computers & Security, 55(??):81–99, November 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404815000978.

Zhou:2019:SAN

[ZCZ^{+}19] Lu Zhou, Jiageng Chen, Yidan Zhang, Chunhua Su, and Marino Anthony James. Security analysis and new models on the intelligent symmetric key encryption. Computers & Security, 80(??):14–24, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818309647.

Zhou:2011:EDR

[ZDBW11] GuangXu Zhou, Murat Demirer, Coskun Bayrak, and Licheng Wang. Enable delegation for RBAC with secure authorization certificate. Computers & Security, 30(8):780–790, November 2011. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404811000976.

Zeng:2019:ESL

[ZDW19] Jianping Zeng, Jiangjiao Duan, and Chengrong Wu. Empirical study on lexical sentiment in passwords from Chinese websites. Computers & Security, 80(??):200–210, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818307028.

Zhou:2019:FCB

[ZGD19] Luying Zhou, Huaqun Guo, and Gelei Deng. A fog computing based approach to DDoS mitigation in IIoT systems. Computers & Security, 85(??):51–62, August 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311349.

Zhao:2013:CSM

[ZGZ^{+}13] Xinjie Zhao, Shize Guo, Fan Zhang, Tao Wang, Zhijie Shi, Zhe Liu, and Jean-François Gallais. A comprehensive study of multiple deductions-based algebraic trace driven cache attacks on AES. Computers & Security, 39 (part B)(??):173–189, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481300093X.

Zonouz:2013:CPS

[ZH13] Saman Zonouz and Parisa Haghani. Cyber-physical security metric inference in smart grid critical infrastructures based on system administrators’ responsive behavior. Computers & Security, 39 (part B)(??):190–200, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000941.

Zonouz:2013:SCB

[ZHB^{+}13] Saman Zonouz, Amir Houmansadr, Robin Berthier, Nikita Borisov, and William Sanders. Secloud: a cloud-based comprehensive and lightweight security solution for smartphones. Computers & Security, 37(??):215–227, September 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481300031X.

Zhang:2011:PFB

[ZHL^{+}11] Peng Zhang, Jiankun Hu, Cai Li, Mohammed Bennamoun, and Vijayakumar Bhagavatula. A pitfall in fingerprint bio-cryptographic key generation. Computers & Security, 30(5):311–319, July 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404811000459.

Zhuang:2018:PCS

[Zhu18] Yan Zhuang. The performance cost of software obfuscation for Android applications. Computers & Security, 73(??):57– 72, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404817302092.

Zhang:2015:TMP

[ZHZ^{+}15] Yixuan Zhang, Jingsha He, Bin Zhao, Zhiqing Huang, and Ruohong Liu. Towards more pro-active access control in computer systems and networks. Computers & Security, 49(??):132–146, March 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814001813.

Zineddine:2015:VMT

[Zin15] Mhamed Zineddine. Vulnerabilities and mitigation techniques toning in the cloud: a cost and vulnerabilities coverage optimization approach using cuckoo search algorithm with Lévy flights. Computers & Security, 48(??):1–18, February 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404814001333.

Zhi-jun:2012:MBA

[ZjHtMhBs12] Wu Zhi-jun, Zhang Hai-tao, Wang Ming-hua, and Pei Baosong. MSABMS-based approach of detecting LDoS attack. Computers & Security, 31(4):402–417, June 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000521.

Zhao:2017:DEH

[ZJK^{+}17] Rui Zhao, Samantha John, Stacy Karas, Cara Bussell, Jennifer Roberts, Daniel Six, Brandon Gavett, and Chuan Yue. Design and evaluation of the highly insidious extreme phishing attacks. Computers & Security, 70(??):634–647, September 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404817301633.

Zafar:2017:SCC

[ZKM^{+}17] Faheem Zafar, Abid Khan, Saif Ur Rehman Malik, Mansoor Ahmed, Adeel Anjum, Majid Iqbal Khan, Nadeem Javed, Masoom Alam, and Fuzel Jamil. A survey of cloud computing data integrity schemes: Design challenges, taxonomy and future trends. Computers & Security, 65(??):29–49, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404816301377.

Zhou:2013:IAA

[ZLC13] Peng Zhou, Xiapu Luo, and Rocky K. C. Chang. Inference attacks against trust-based onion routing: Trust degree to the rescue. Computers & Security, 39 (part B)(??):431–446, November 2013. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404813001363.

Zhou:2010:SCA

[ZLK10] Chenfeng Vincent Zhou, Christopher Leckie, and Shanika Karunasekera. A survey of coordinated attacks and collaborative intrusion detection. Computers & Security, 29(1):124–140, February 2010. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740480900073X.

Zhong:2018:CSD

[ZLL^{+}18] Chen Zhong, Tao Lin, Peng Liu, John Yen, and Kai Chen. A cyber security data triage operation retrieval system. Computers & Security, 76(??):12–31, July 2018. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818301408.

Zhang:2019:ZCS

[ZLL^{+}19] Yuncong Zhang, Yu Long, Zhen Liu, Zhiqiang Liu, and Dawu Gu. Z-Channel: Scalable and efficient scheme in Zerocash. Computers & Security, 86(??):112–131, September 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818313737.

Zeng:2012:AIB

[ZLLL12] Ying Zeng, FenLin Liu, XiangYang Luo, and ShiGuo Lian. Abstract interpretation-based semantic framework for software birthmark. Computers & Security, 31(4):377–390, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404812000545.

Zhu:2015:DIN

[ZLW^{+}15] Erzhou Zhu, Feng Liu, Zuo Wang, Alei Liang, Yiwen Zhang, Xuejian Li, and Xuejun Li. Dytaint: the implementation of a novel lightweight 3-state dynamic taint analysis framework for x86 binary programs. Computers & Security, 52(??):51– 69, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404815000401.

Zou:2019:BBP

[ZLW19] Renpeng Zou, Xixiang Lv, and Baocang Wang. Blockchainbased photo forensics with permissible transformations. Computers & Security, 87(??):Article 101567, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818313981.

Zaeem:2017:MAI

[ZMYB17] Razieh Nokhbeh Zaeem, Monisha Manoharan, Yongpeng Yang, and K. Suzanne Barber. Modeling and analysis of identity threat behaviors through text mining of identity theft stories. Computers & Security, 65(??):50–63, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404816301559.

Zhang:2011:TCS

[ZNAHK11] Zonghua Zhang, Farid Naït-Abdesselam, Pin-Han Ho, and Youki Kadobayashi. Toward cost-sensitive self-optimizing anomaly detection and response in autonomic networks. Computers & Security, 30(6–7):525–537, September/October 2011. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481100071X.

Zhang:2019:FHM

[ZQY^{+}19] Jixin Zhang, Zheng Qin, Hui Yin, Lu Ou, and Kehuan Zhang. A feature-hybrid malware variants detection using CNN based opcode embedding and BPNN based API embedding. Computers & Security, 84(??):376–392, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818312902.

Zwattendorfer:2015:DSP

[ZS15] Bernd Zwattendorfer and Daniel Slamanig. Design strategies for a privacy-friendly Austrian eID system in the public cloud. Computers & Security, 52(??):178–193, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404815000346.

Zangeneh:2018:CSM

[ZS18] Vahid Zangeneh and Mehdi Shajari. A cost-sensitive move selection strategy for moving target defense. Computers & Security, 75(??):72–91, June 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404817302821.

Zhang:2019:SEF

[ZTC19] Li Zhang, Vrizlynn L. L. Thing, and Yao Cheng. A scalable and extensible framework for Android malware detection and family attribution. Computers & Security, 80(??):120–133, January 2019. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481830419X.

Zhang:2019:FMS

[ZTL^{+}19] Jie Zhang, Zhanyong Tang, Meng Li, Dingyi Fang, Xiaojiang Chen, and Zheng Wang. Find me a safe zone: a countermeasure for channel state information based attacks. Computers & Security, 80(??):273–290, January 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818311337.

Zhao:2013:BDB

[ZTS^{+}13] David Zhao, Issa Traore, Bassam Sayed, Wei Lu, Sherif Saad, Ali Ghorbani, and Dan Garant. Botnet detection based on traffic behavior analysis and flow intervals. Computers & Security, 39 (Part A)(??):2–16, November 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000837.

Zhu:2014:SAB

[ZW14] Feng Zhu and Jinpeng Wei. Static analysis based invariant detection for commodity operating systems. Computers & Security, 43(??):49–63, June 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404814000273.

Zhou:2019:ELA

[ZW19] Yun Zhou and Peichao Wang. An ensemble learning approach for XSS attack detection with domain knowledge and threat intelligence. Computers & Security, 82(??):261–269, May 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404818306370.

Zhu:2019:IED

[ZWNM19] Erzhou Zhu, Peng Wen, Kanqi Ni, and Ruhui Ma. Implementation of an effective dynamic concolic execution framework for analyzing binary programs. Computers & Security, 86(??):1–27, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818300932.

Zhong:2019:SLS

[ZWX^{+}19] Lin Zhong, Qianhong Wu, Jan Xie, Zhenyu Guan, and Bo Qin. A secure large-scale instant payment system based on blockchain. Computers & Security, 84(??):349–364, July 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740481831407X.

Zhang:2011:DCS

[ZY11] Da Zhang and Chai Kiat Yeo. Distributed court system for intrusion detection in mobile ad hoc networks. Computers & Security, 30(8):555–570, November 2011. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001155.

Zhao:2014:TSU

[ZY14] Rui Zhao and Chuan Yue. Toward a secure and usable cloud-based password manager for web browsers. Computers & Security, 46(??):32–47, October 2014. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404814001059.

Zi:2010:IPN

[ZYPL10] Xiaochao Zi, Lihong Yao, Li Pan, and Jianhua Li. Implementing a passive network covert timing channel. Computers & Security, 29(6):686–696, September 2010. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404809001485.

Zhang:2016:CRA

[ZYRZ16] Hao Zhang, Danfeng (Daphne) Yao, Naren Ramakrishnan, and Zhibin Zhang. Causality reasoning about network events for detecting stealthy malware activities. Computers & Security, 58(??):180–198, May 2016. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816000043.

Zhang:2014:TON

[ZYWZ14] Hao Zhang, Nenghai Yu, Yonggang Wen, and Weiming Zhang. Towards optimal noise distribution for privacy preserving in data aggregation. Computers & Security, 45(??):210–230, September 2014. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S016740481400090X.

Zhao:2012:TRS

[ZZ12] Xingwen Zhao and Fangguo Zhang. Tracing and revoking scheme for dynamic privileges against pirate rebroadcast. Computers & Security, 31(1):59–69, February 2012. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404811001386.

Zhanwei:2019:ADM

[ZZ19] Song Zhanwei and Liu Zenghui. Abnormal detection method of industrial control system based on behavior model. Computers & Security, 84(??):166–178, July 2019. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818307405.

Zhou:2017:ENQ

[ZZC17] Lu Zhou, Youwen Zhu, and Aniello Castiglione. Efficient k-NN query over encrypted data in cloud with limited key-disclosure and offline data owner. Computers & Security, 69(??):84–96, August 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301663.

Zhu:2013:NIC

[ZZZY13] Hegui Zhu, Cheng Zhao, Xiangde Zhang, and Lianping Yang. A novel iris and chaos-based random number generator. Computers & Security, 36(??):40–48, July 2013. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404813000321.